A Bibliography of Papers in Lecture Notes in Computer Science (1995), Part 1 of 2

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

11 March 2017 Version 1.26

Title word cross-reference

(2) [1066]. 2 [480, 615, 284, 474, 2018]. 3 [495, 483, 482, 499, 491, 502, 472, 480, 464, 459, 494, 511, 448, 1014, 129, 460, 506, 493, 500, 918, 513, 501, 445, 128, 2015, 937, 461, 508]. 4 [484]. A^{2} [1082, 1997]. β [1271]. f [398]. k [1715, 1021, 405]. λ [1639]. μ [1656, 1658]. p [1065]. π [1568, 296, 768]. r [851]. → [748, 720]. y^{2} + axy ≡ x^{3} (mod n) [1057].

-Aided [1063]. -Calculus [296, 1568, 1658, 768, 1656, 1639]. -Cell [1271]. -Codes [1082, 1997]. -Colored [1715]. -CUT [1021]. -D [128]. -Dependent [398]. -Dimensional [615]. -Divergence [1787]. -DNF [437]. -fold [851]. -partition [405].

0-1 [1029, 1449]. 01H [476]. 02 [1538].

1 [869, 1010, 961]. 10rO0s [720].

2 [602, 985, 1708, 990, 992, 926]. 2-adic [1969]. 2-Spiral [602]. 2-Structures [1708].

3-valued [1163]. 30 [493].

5 [996, 909].

7 [1905]. 7-Designs [1905].

8-bit [1091]. 80 [2045].

9 [1499].

= [725, 329, 556, 1656].

A-codes [851]. Abdominal [1494]. Abduction [1163, 1792, 1152]. Abductive [1805, 1184, 1162, 1161]. Abelian [836]. Able [616]. Abstract [1097, 806, 643, 296, 1100, 527, 543, 595, 518, 1817, 536, 642, 549, 2024, 539, 359, 377]. Abstracting [1120]. Abstraction [1652, 772, 1655, 1488]. Abstractions [559, 75]. Abundance [1225]. Abundance-Distributions [1225]. Academia [186]. Academic [200]. Accelerated [1291]. acceleration [854]. Accelerator [1350]. accelerators [1091]. Acceptability [1189]. Access [1738, 88, 2009]. Accessibility [631]. Accessing [2008]. According [1824]. Accumulation [1278]. Achieve [572]. Acid [1510]. Acid-Base [1510]. Acoustical [1808]. Acquisition [253, 1508, 255, 1386, 252]. ACRI [961]. ACRI-1 [961]. Across [2043]. Action [1020, 1864, 39, 1880]. Actions [1845, 23]. Activation [1315, 485]. Active [512, 112, 1123, 1221]. Activity [1527, 1278]. Actor [1412]. Actor-based [1412]. Acute [1501, 1494]. Acyclic [170, 389, 1629, 1887]. Ada [1143]. Adaptable [453, 458]. Adaptation [1488, 1250, 1213]. Adaptative [1382]. Adapting [668]. Adaption [468]. Adaptive [1205, 1270, 1245, 169, 548, 1233, 1259, 1262, 1280, 642, 1394, 1377, 1939, 1503, 24, 449, 995]. Adaptivity [1239]. Addition [1987, 1986]. Additional [2009]. Additive [1355]. adic [1969]. Adjacency [605]. Adjusting [1517]. Administering [823]. Admissible [1469]. Adult [1274]. Advance [224]. Advanced [1096, 562]. Advantages [874]. adversarial [1967]. advice [416]. Adviser [1488]. aerodynamic [1015]. aerospace [958]. Affine [585, 68, 920, 930].

Affinity [66]. AFL [1711]. after [543]. against [1956, 834]. Age [1225, 1511, 1557]. Agency [30]. Agent [26, 34, 18, 19, 21, 28, 35, 29, 36, 31, 32, 1764, 23, 38, 40, 25]. Agent-Based [35]. Agent-Oriented [34, 40, 25]. Agent-Passing [1764]. Agents [1248, 20, 27, 1485, 1231, 33, 642, 1237, 24, 39]. Aggressive [1245]. AGM [1930]. Agreement [279, 1978]. AI [1798]. Aided [1653, 1063, 451, 471, 1876]. Air [249, 94]. Air-Traffic [249]. Airport [534]. Alerts [1484]. Alethic [1797]. Algebra [797, 516, 65, 1531, 472, 1649, 109, 2014, 545, 1593, 1887]. Algebraic [743, 1591, 1535, 1537, 1738, 1897, 521, 117, 1574, 1768, 527, 1586, 530, 1888, 1914, 1566, 1462, 1898, 341, 1711, 1579, 813, 535, 1923, 1925, 325, 1553, 771, 783, 519]. Algebraic-Geometric [1914]. Algebras [806, 608, 524, 590, 1132, 1562, 540, 1439]. Algol [2057]. Algol-Like [2057]. Algorithm [2025, 70, 1291, 1367, 1330, 646, 1539, 1993, 169, 1294, 1729, 1611, 1612, 465, 560, 158, 1692, 1326, 1066, 1052, 1028, 1099, 1624, 1927, 2035, 731, 167, 1312, 863, 434, 963, 1873, 437, 432, 967, 989, 950, 982]. Algorithmic [385, 1464, 1665, 168, 76, 999, 412]. Algorithms [1727, 780, 410, 270, 1369, 580, 1023, 1730, 1773, 1716, 1732, 292, 1302, 1258, 618, 1021, 1051, 409, 1717, 650, 2014, 146, 297, 1651, 1584, 1133, 1532, 289, 1323, 1359, 273, 1822, 172, 131, 1623, 1216, 679, 140, 90, 1243, 272, 405, 682, 1877, 1878, 1641]. Alias [77]. Alignment [65, 1610, 1616, 630, 1430, 1623, 965]. All-Optical [1365]. Allocation [795, 1939]. Almost [1043]. Alphabet [615]. Alphabet-Independent [615]. Alternation [1734]. Ambiguity [626]. Ambiguous [1827, 1809]. Ames [1949]. among [1084, 2040, 1985]. amplification [1975]. Amplitude [510]. AN2 [974]. Analog [1356, 1369]. Analogical [655]. analogue [862]. Analyses [697, 793]. Analysis [230, 498, 1820, 77, 1527, 996, 1543, 480, 2028, 1514, 661, 1409, 1310, 830, 1307, 1278, 1577, 487, 648, 313, 475, 1665, 1417, 1676, 1679, 1575, 1292, 1852, 493, 1548, 1618, 277, 1211, 500, 1619, 8, 1342, 112, 273, 683, 1408, 264, 1381, 1522, 1496, 81, 1474, 476, 1940, 1273, 799, 481, 1814, 10, 394, 1475, 789, 1455, 561, 1203, 848, 912, 942, 437, 981, 909, 984, 983]. Analytic [888]. Analytical [1266, 673]. Analyzer [797]. Analyzing [503, 150]. Anatomic [502]. Anatomical [459, 468]. Anatomy [466, 460, 762]. Anatomy-Based [466]. AND/OR [1869]. Angelic [1889]. Angiography [480, 500]. Angle [135]. ANGN [1267]. Animal [24]. Animals [1209]. Anisotropic [511, 927]. ANL [1946]. ANL/IBM [1946]. ANN [1267]. Annotated [1804, 1171, 244]. anonymity [1061]. Anonymous [1092, 1981]. Answer [1187]. Answers [1455]. Ant [1246, 1247]. Anterior [505]. Anti [1470]. Anti-Foundation [1470]. Antibiotics [1488]. Anticipatory [27]. Antidirected [410]. Any [616]. applicability [1060]. Application [1311, 1428, 1497, 1148, 805, 1538, 502, 512, 1540, 1401, 1716, 1477, 250, 1307, 1278, 1807, 1382, 503, 688, 583, 1262, 1052, 1381, 1561, 2011, 783, 1312, 929, 923, 836, 413, 992, 1000]. Applications [1479, 1726, 1138, 1401, 396, 1458, 235, 1717, 1657, 1667, 1733, 157, 1787, 1123, 3, 1785, 799, 389, 5, 895, 903, 1008, 852, 951, 1005, 1995, 907, 1633]. Applicative [2052]. Applied [498, 740, 732, 1494, 1016, 934]. Applying [1415, 182, 1585, 234, 1326, 1511]. Approach [196, 1828, 1830, 1103, 1423, 1157, 1406, 1481, 524, 1372, 19, 1810, 206, 1514, 87, 1845, 383, 1671, 703, 830, 210, 649, 208, 1786, 180, 1801, 1610, 139, 1208, 341, 1855, 1135, 277, 173, 30, 130, 8, 1323, 673, 63, 194, 1744, 1355, 1522, 635, 1626, 807, 1866, 680, 1503, 72, 1668, 771, 39, 638, 991, 1016, 1644, 912, 999, 908, 996, 1002, 1887, 864, 997]. Approaches [470, 1933, 168, 1492]. Approximability [621]. Approximable [1448]. Approximate [1802, 1604]. Approximating [1022, 1731]. Approximation [1698, 1023, 1678, 1021, 427, 1853, 407, 124, 1994, 1623, 1699, 1001]. Approximations [1778, 1740, 1919, 421]. April [36]. Arbiter [1997]. Arbitrary [1449, 300]. arbitration [1080]. Arborescences [1051]. Arboricity [138]. Arc [1107, 1105]. Arc-Consistency [1107]. Architectural [1936, 1630]. Architecture [1878, 1263, 18, 1518, 1807, 1299, 29, 1471, 1384, 599, 1373, 1927, 962, 1363, 1695, 863, 911, 941, 961]. Architectures [1426, 93, 1308, 32, 1239, 336, 1637, 910, 893, 1010]. archiving [955]. Area [1225, 162, 1515]. Areas [691]. Argumentation [1151, 1183, 1767]. arising [915]. Arithmetic [1647, 741]. Arithmetical [1364]. Arjuna [1631]. ARPEGE [917]. ARPEGE/IFS [917]. Arrangement [638]. Array [67, 1614, 78, 92, 1927]. Arrays [319, 969, 1884]. Arterial [492]. Artifacts [1425]. Artificial [1225, 1335, 1388, 269, 1210, 1253, 1264, 1303, 1266, 1292, 1254, 1197, 1202, 1215, 1218, 1251, 1191, 37, 1192, 1334, 1190, 1232, 1195, 1340, 1194]. Ascends [215]. ASF [815]. ASIC [348]. Asking [1811]. ASL [520]. Asociative [1382]. Aspects [1464, 757, 266, 235, 1814, 850]. Assemblies [1324]. Assertions [1659]. Assess [484]. Assessment [1150, 685, 1137, 1136, 1141, 1003]. Assignment [1032]. Assist [812]. Assistant [891]. assistants [1888]. Assisted [1587, 505, 475]. Assisting [231, 813]. Associated [1787]. Association [2004]. Associative [1389, 749, 1283, 1352]. Associative-Commutative [749]. Assumption [1652, 773, 1155]. Assumption-Commitment [1652].

Assumption/Guarantee [773]. Assumptions [1671]. Assurance [200, 1500]. Asteroidal [1732]. ASTRE [747]. Asymmetric [1960]. Asynchronous [1662, 1552, 1880]. Asynchronously [1357, 1882]. Asynchrony [1673, 2040]. ATEMPT [925]. Atlas [473, 461]. ATM [974, 904, 900]. ATMS [1832]. Atomic [1672, 2032, 949]. Attack [1998, 1996, 1956, 834, 840]. Attacking [859, 1058]. Attacks [1997, 1972, 1077, 1989, 1974]. Attentional [1325]. Attitude [1302]. Attitudes [17]. Attraction [10]. Attractor [1307, 1327]. Attractors [1811]. Attribute [678]. Attributed [1096]. Attributes [2009, 675]. Audio [1676]. Auditory [1276]. Augment [1943]. Augmentation [138, 1052]. Augmenting [1035, 1672]. Authentication [1070, 1081, 1974, 1973, 1080]. Author [1093, 2000, 2046]. authority [1072]. Automata [315, 2049, 1720, 1663, 316, 1317, 579, 616, 284, 1746, 1671, 303, 1735, 626, 1230, 304, 1667, 288, 627, 629, 633, 1625, 636, 765, 1228, 1250, 1668, 440]. Automata-Theoretic [1671, 1668]. Automated [747, 750, 1839, 1409, 830, 443, 1676, 1584, 594, 513, 891, 1654, 2016]. Automatic [1753, 66, 1595, 1909, 459, 1309, 581, 1299, 93, 1655, 474, 341, 1499, 672, 822, 882, 336, 801, 718, 1872, 1016, 899]. Automatical [468]. Automatically [1419]. Automation [1471, 1147, 689]. Automaton [616, 1620, 1229, 1204]. Automorphism [1905]. Autonomous [1235, 1298, 1237, 1242]. Autonomy [18]. Autoradiography [474]. Autoscheduling [91]. AV [1828]. AV-Qualified [1828]. avalanche [844]. Average [114, 437, 330, 319, 318]. AVL [162]. Avoidance [1322, 1234]. Avoiding [1648, 3]. Axiom [354, 1470]. Axiomatic [1835, 533, 1855]. Axiomatizability [777, 1710]. Axis [678, 434]. Axis-Parallel [678, 434].

B [1538, 816, 501]. B-Deformable [501]. B-Toolkit [816]. Back [1391, 1316]. Back-Propagation [1316]. Background [2]. Backmarking [1109]. Backpropagation [1286, 1292, 1390, 989]. Balance [1510]. Balancing [1333, 2013, 952, 940, 951, 949, 948, 950]. Banking [1138]. Base [1840, 1510]. Based [1151, 1497, 499, 348, 1502, 870, 644, 1505, 509, 610, 685, 1481, 1837, 1518, 1314, 1840, 1670, 206, 1338, 1804, 255, 1514, 1489, 1519, 1844, 1264, 383, 1378, 527, 1029, 1596, 1847, 35, 648, 1120, 1803, 1102, 1564, 1385, 650, 2014, 466, 1230, 665, 1498, 1493, 1854, 566, 2042, 467, 462, 157, 8, 1323, 1471, 1492, 473, 878, 793, 1565, 1812, 1376, 1521, 1593, 1516, 801, 1488, 1868, 1528, 538, 1524, 1937, 2030, 1304, 74, 497, 554, 1507, 444, 1520, 240, 1421, 1819, 1312, 853, 1114, 1156, 1536, 1805, 1149, 1406, 1992, 1962, 1240, 1380, 1964]. based [1412, 1544, 1969, 852, 1956, 933, 247, 1634, 1057, 954, 1859, 1818, 1966, 1083, 767, 1955, 679, 828, 862, 868, 187, 338, 933, 1552]. Bases [1165, 1830, 611, 1849, 1913, 1041, 261, 1920, 583, 1504, 1276, 405]. Basic [1702, 1094, 1674, 369, 731, 982]. Basis [1361, 1319, 1328, 256, 1040, 1360, 1368, 1817, 1456]. Batch [1945]. Bats [1232]. Bauhaus [1118]. Bayesian [1481, 1836, 1847, 1786, 1852, 1853, 1861, 1351, 1313, 1775]. BDD [1649, 1552, 338]. BDD-Based [338, 1552]. BDDs [607]. Be [1951, 761, 1911, 1217, 1290, 1958, 1267]. Beams [458]. Bedfellows [258]. Behavior [1271, 646, 1240, 1235, 266, 537, 1124]. Behavior-based [1240]. Behavioral [111, 1277]. Behaviors [1480, 440]. Behaviour [1290, 28, 529, 1246, 226, 1298, 32, 1405, 1239, 3, 1300, 39]. Behavioural [1659, 522, 1567, 1245, 662, 705, 772, 1276]. Behaviours [1546]. Being [1443]. Belief [19, 1846, 1849, 1853, 1771, 1769]. Beliefs [1509]. Benchmark [332, 910, 911]. Benchmark-Circuits [332]. Benchmarking [985]. Benchmarks [2022, 983]. Bend [132]. Bend-Minimum [132]. Bends [133]. Benefits [1148, 1149, 1942]. bent [843]. Berlekamp [1927]. Berlin [943]. Best [188]. Better [662]. Between [2049, 806, 607, 238, 186, 628, 271, 1202, 636, 1439, 1411, 884, 1601, 1983, 1414, 1612, 1614, 1984, 377, 1939, 2019, 1629]. Beyond [329]. Bias [1333]. Bibliographical [577]. Bicriteria [1749]. Bidirectional [1332]. Bidirectionality [1107]. Binary [1904, 1445, 1918, 1037, 1865, 840]. Binding [793, 1250]. Binding-Time [793]. Binomial [1904]. Biocomputation [274]. Biological [1197, 630, 1619, 1345]. Biology [1190]. Biomedical [1479, 452, 8]. Biomolecular [1254]. Biopsy [1502]. Bipartite [143]. Bipartition [1024]. bisection [950, 1021]. Bisimilarity [779, 2048]. Bisimulation [768]. Bisimulations [1735]. bit [1091]. Bits [1722, 1087, 1086, 1088, 1085]. Bivariate [1900]. Black [1955]. Blackmailing [1991]. Blind [1992, 1074, 852]. Blinding [1076, 1957]. Blob [8]. Block [133, 1066, 1994, 904, 853, 1989, 982, 1015]. block-sparse [982]. blocking [1643]. Blood [491, 1489]. board [962]. Bodies [1039]. Boehm [736]. Boltzmann [1357, 1279, 1396, 1395]. Bone [498]. Book [400]. Boolean [588, 102]. boosting [413]. booth [1090]. Bootstrap [666, 1136]. Bottleneck [1032]. Bottom [766, 1157]. Bottom-Up [1157, 766]. Bound [1971, 1287, 381, 140, 1970, 837]. Boundaries [490]. Boundary [394, 988]. Bounded [780, 1832, 1730, 20, 327, 281, 626, 297, 1028, 1470, 434, 1894]. Bounded-Degree [297]. Boundedness [123, 122]. Bounding [133, 436]. Bounds [605, 606, 1738, 1747, 609, 306, 1723, 1783, 282, 321, 1549, 631, 131, 1899, 408, 416, 1080]. Box [1531, 1955]. boxes [1984]. BP [1344, 1291]. BP-gammaL1 [1291]. BPP [320]. Brain [1296, 1260, 496, 474, 467, 469, 473, 485, 461, 486]. Branching [1674, 1672, 408]. breadth [1892]. breadth-first [1892]. Break [1720, 1067, 1978]. Breaking [1981, 1084]. Breast [1514, 1491, 1523]. Bridge [1202]. Bridging [186, 1638]. bright [1011]. broadband [901]. Broadcast [1977]. Broadcasting [1589, 397]. Bronchus [513]. Browsing [825]. Bruijn [1078, 391, 739]. Brzozowski [586]. Bubbles [497]. Buffer [1688]. Buffer-requirement [1688]. buffers [1894]. Bug [762]. Builders [224]. Building [880, 1478, 1495, 1636, 1631]. Built [1742]. Built-in [1742]. Bulk [517]. Burnside [586]. Business [1417]. buying [847]. Bypass [404].

C [86, 967, 971, 565]. CAB [1540]. Cactus [1614]. Calculating [792]. calculation [973]. calculations [1012]. Calculi [736, 1764, 1798, 554, 540, 372]. Calculus [787, 725, 2056, 1434, 296, 1663, 1477, 1804, 1568, 1658, 2059, 1589, 1575, 1438, 768, 1176, 564, 1862, 552, 1443, 1656, 767, 351, 1879, 1883, 364, 361, 1437, 1889, 366, 367, 1639, 1891, 376, 377, 378]. Calibration [504]. Call [76, 376]. call-by-value [376]. Caltech [1880]. Cambridge [254]. Camera [504]. Can [1339, 761, 1911, 1958, 794, 1217, 868, 1290]. Cancer [1491, 1523]. Canonical [1461]. cap [1740, 598, 104]. Capabilities [1943, 1254]. Capability [1137, 1298]. Capable [456, 1229]. Capacitated [1051]. Capacities [1838]. Capstone [202]. Card [1358, 1995]. Cardiac [1525, 927]. Cardiology [1515]. CARDS [205]. Care [29, 1496, 1516]. Cares [1617]. CARMI [1944]. Cartesian [2025, 391, 1463, 1194]. CAS [471]. Cascades [1996]. Case [1098, 516, 1039, 1505, 19, 1909, 206, 28, 648, 650, 511, 1588, 330, 798, 1046, 319, 335, 690, 1237, 1488, 1455, 1504, 784, 437, 416]. Case-Based [1505, 648, 650, 1488]. Case-Databases [1504]. Cases [643, 1508]. Cash [610]. Cataloging [830]. Categorical [2058, 2056, 733, 2061, 732, 2059, 376, 378, 356]. Categories [1451, 1463]. Category [2057, 1755, 1564, 2048, 778, 2054]. Category-Based [1564]. Category-Theoretic [2057]. Causal [1778, 1483, 1546, 1777, 1861, 1865, 1765, 1643]. Causality [296, 1763, 1575]. Causally [1532, 1634]. Causes [1809]. CAVEAT [1653]. Cayley [636]. CC [591, 601]. CEA [1009]. CEA/CEL [1009]. CEA/CEL-V [1009]. Cell [1271, 1550]. Cells [1272]. Cellular [1356, 4, 616, 284, 1230, 636, 1228, 1250, 1229, 1350]. Center [215, 1511, 860, 1947]. Centered [715, 708]. Central [1280]. centres [942]. CERN [990]. Certificates [610, 1076]. CFD [1012, 1014]. CG [156, 982]. CG-algorithm [982]. Chain [1050, 1834, 609, 1752, 1667, 1300]. Chained [752]. Chains [1791, 1510, 1987, 1986]. Challenge [1897, 952]. Challenges [1415, 1682, 964]. Chances [139]. Change [1811, 713, 715]. changes [420]. Changing [17]. Channel [1981, 848, 900]. Chaos [1193, 1809]. Chaotic [1291, 1811, 1810, 1808, 1305, 3]. Character [1602, 1380, 1382, 1324]. Characterisation [26, 690]. Characteristic [349]. Characteristics [1949, 844]. Characterization [1049, 1756, 3, 151, 720, 908]. Characterizations [1726, 1168, 1167]. Characterizing [905, 431]. checked [1459]. Checking [1659, 1578, 1652, 1658, 1671, 1674, 1108, 1170, 1109, 1654, 597, 1552, 1664, 1112]. checkpoint [1637]. Chemical [1897]. Chemistry [1197, 963]. CHI [228]. Chicken [1602]. Children [1511, 2030].

Chip [1366, 1111]. Choice

[354, 101, 1548, 274, 1224]. Choose [1824]. Choosing [1403]. Chor [1058]. Chordal [1718]. chosen [1974]. chosen-content [1974]. Church [302]. ChyPro [2045]. Cid [86]. Ciphers [1994]. ciphertexts [1069]. Circuit [1647, 330, 337, 419]. Circuits [606, 1369, 282, 278, 341, 1724, 283, 332, 1662, 1552]. Circumscribing [1864]. Circumscription [293]. Circumscriptive [1185]. Class [785, 1706, 1705, 1537, 1538, 2028, 2005, 1679, 1651, 675, 2030, 558, 1411, 555]. Class-and [2030]. Class-Driven [675]. Classes [2029, 1082, 1910, 1717, 281, 1186, 326, 373]. Classical [1831, 1802, 886, 1182]. Classification [1383, 1837, 1318, 677, 1237, 678, 1352, 1351, 1313, 682]. Classifier [1257, 659, 1334, 1304]. Classroom [189]. Clausal [243]. Clean [1130]. climate [914]. Climbing [1112]. Clinical [1497, 1506, 1490, 505, 1515, 1507]. Clique [1024]. Clock [1996, 1972]. clock-controlled [1972]. clocked [1077]. Clocks [1743]. Clones [662]. Closed [1463, 1277]. Closure [1670, 1714, 1737, 1195]. Closure-Condition [1670]. Closures [389]. Clouds [1200]. CLP [1103, 591, 593, 598, 589]. CLP-Languages [593]. club [961]. Cluster [1868, 980, 949]. Cluster-Based [1868]. Clustering [464, 1214, 1328]. Clusters [64, 1948, 975]. CM [996, 909]. CM-5 [996, 909]. CMM [187]. CMM-Based [187]. CMU [183]. CNN [1374]. Co [263, 1205, 1208, 474, 671, 1209]. Co-Evolution [1208, 1209]. Co-Evolutionary [1205]. Co-Occurrences [263]. Co-operative [671]. Co-Registration [474]. Coalgebras [531, 1573]. coarse [939]. coarse-mesh [939]. coaxial [956]. COBRA [1504].

Cochlear [1321, 1276]. CodA [2033]. Code [331, 1896, 805, 796, 1016, 985, 939, 1014, 1018, 938, 948, 1015]. Coded [1603]. Codes [1902, 1904, 1082, 88, 1907, 1910, 1729, 1914, 1915, 1997, 1757, 1918, 1921, 853, 969, 1962, 1080, 1010, 1974, 851, 868]. Codewords [1902]. Coding [308, 695, 5, 1276, 428]. Codings [309]. Coevolution [1248, 1207, 1206]. Cofree [1573]. Cog [1217]. Cognitive [18, 1277, 1222]. Cognodynamics [1402]. Coherence [1769]. Coherent [695]. Coins [1979, 1075]. Collaboration [238]. Collaboratory [14, 15, 16]. Collapse [1724]. Collection [727, 560, 2035]. Collections [822, 1593]. Collective [1300]. Collinearity [585]. Collision [1234]. Collisions [858]. Colonies [1246]. Colony [1247]. Color [1375, 1406]. COLOR-X [1406]. Colored [603, 1715, 1544]. Coloring [311]. Coloured [1543, 1547]. Colouring [620]. Colourings [1050]. Columnar [1265]. Combination [743, 243, 672, 1377]. Combinations [781]. Combinatorial [1398, 1080, 839, 1037, 851]. Combinatorics [1897]. Combinators [563, 361]. Combined [1510, 2017]. Combining [516, 463, 1508, 1055, 1683, 687, 1030, 1112]. combustion [937]. Comfortable [1428]. Comments [1993]. commerce [945]. Commitment [1652, 20, 107]. Common [1605, 624, 310, 256, 239, 1624]. Communicating [523, 1117, 129, 24, 1882]. Communication [80, 307, 279, 390, 1414, 306, 82, 705, 79, 68, 631, 1218, 2043, 892, 1174, 955, 1634, 1894, 906]. Communication-Free [68]. communications [901, 979]. Communities [228]. Commutations [1714]. Commutative [516, 749]. Commutativity [569]. Comonad [732]. Compact [293, 1961]. Compactness [1745]. Comparability [382]. Comparative [1301, 648, 1588, 42]. Comparing [351, 1613, 1360, 793]. Comparison [2022, 470, 467, 1492, 1626, 691, 909]. Comparisons [1607, 1007]. Compartimentation [1198]. Compartmental [1266, 1480]. Compass [1233]. Compatibility [1602]. Compatible [1044]. Competition [1301]. Compilation [80, 72]. Compiler [64, 817, 1459, 919, 895, 969]. Compilers [71]. Compiling [63]. Complete [1903, 1583, 1040, 1247, 582, 768, 540, 720, 371, 284, 1446]. Completeness [278, 1452, 781, 717, 1582, 782, 1684, 1455, 378]. Completion [1043, 753, 722, 583]. Complex [670, 1817, 698, 1421, 952, 957]. Complexity [1697, 1725, 279, 99, 1042, 390, 306, 1161, 327, 160, 624, 1971, 625, 1613, 397, 330, 149, 387, 288, 1197, 1467, 326, 631, 1958, 1899, 1926, 681, 1203, 1699, 442, 422, 841, 1970, 423, 425]. Component [562, 1471]. Component-Based [1471]. Components [1498]. Composable [1538]. Composed [1087, 1088]. Composing [1165, 1893]. Composition [1122, 1251, 10, 337, 1411]. Compositional [1661, 1580, 1670, 21, 337, 1882]. Compositive [1789]. comprehension [361]. Comprehensions [2060]. Comprehensive [206]. Compressing [1716, 2023]. Compression [827]. Computability [323]. Computation [1903, 1296, 1722, 1157, 1539, 956, 1908, 1158, 1159, 87, 1723, 1071, 1260, 1914, 935, 1565, 3, 544, 1228, 1229, 1931, 980]. Computation-Universal [1228]. Computational [1073, 1027, 1042, 160, 732, 753, 1272, 1467, 849, 953, 1084, 927, 988]. Computationally [1247]. Computations [803, 395, 316, 1923, 778, 926, 1646, 998]. Compute [1865, 1002, 1065]. Computer [219, 209, 1897, 505, 1258, 475, 1653, 466, 890, 184, 451, 179, 471, 642, 1511, 476, 1876, 1942, 910, 965, 939, 847, 906].

Computer-Aided [451, 1876]. Computer-Integrated [466]. Computerized [461]. Computers [560, 1941, 957, 1017, 937, 1015, 958, 934]. Computing [15, 1601, 1169, 233, 82, 624, 1612, 559, 1254, 1714, 194, 1160, 1189, 775, 1774, 976, 901, 896, 964, 910, 987, 946, 935, 914, 1876, 1633]. Concatenation [1454]. Concept [650, 1466, 1858, 1472, 1429, 1513, 175, 653]. Concepts [1702, 1094, 732, 1476, 952]. Conceptual [1499, 1687, 1405, 1407, 1420]. concerning [433]. Concurrency [803, 1691, 1575, 596]. Concurrent [727, 316, 1540, 1100, 1580, 1709, 1592, 1577, 246, 550, 305, 596, 795, 552, 544, 547, 248, 23, 554, 1124, 561, 558, 40, 563]. Condition [1670]. Conditional [585, 1783, 730]. Conditioning [1322]. Conditions [1223, 1511, 783]. Cone [1040]. conference [1976]. Configuration [706, 393]. Confirmation [1867]. Confirmer [1959]. Conflict [1859]. Conflicts [1117]. Confluence [1592, 770, 731, 730]. Confluent [394]. Congenital [1504]. Conjecture [328, 586, 2051]. Conjunction [789]. Connected [1053]. Connection [870, 1054, 874, 133, 271, 878, 909]. Connectionist [1387, 1317, 1346]. Connectionist-Symbolic [1346]. Connectionists [1320]. Connections [1293]. Connective [1833]. Connectives [1799]. Connectivity [1262, 1052, 1721, 1742]. conquer [1878, 997]. Consciousness [1336]. Consequence [1178, 1837, 1844]. Consequences [297, 533, 1724, 1277, 381]. Conserving [1940]. Consistency [1107, 589, 1783, 1337, 1105, 1645]. Consistent [1281, 1532]. Constant [1095, 1607, 799]. Constant-Space [1607]. Constrained [508, 950, 433]. Constraint [743, 118, 883, 587, 589, 1100, 1670, 1113, 1592, 1108, 592, 244, 593, 247, 596, 597, 125, 635, 1110, 599, 1106, 601]. Constraint-Generating [118]. Constraint-Solving [589]. Constraints [1821, 1165, 658, 482, 588, 165, 590, 1791, 1101, 62, 1102, 753, 594, 759, 595, 1164, 721, 1656, 121]. Constructible [1928]. Constructing [1849, 1729, 1130]. Construction [1591, 1953, 1425, 1717, 1914, 2053, 1341, 723, 1229, 1875, 1886, 1952]. Constructions [315, 1891]. Constructive [1367, 591, 1593]. constructors [379]. Contained [1450]. containing [969]. Contemporary [1255, 724]. Content [826, 1974]. Contest [141]. Context [1571, 302, 612, 1519, 760, 1447, 314, 442]. Context-Free [1571, 1447, 314, 442]. Context-Sensitive [302, 612]. Contexts [19]. Contextual [1822, 1216]. continental [959]. Continuation [558]. Continuations [804]. Continuous [613, 1810, 1576, 1315, 1279, 668, 675]. Contour [1375]. Contradictory [1839]. Contrast [1268, 932]. Contributions [716]. Control [1669, 1537, 1811, 1805, 1399, 1235, 1238, 1302, 28, 82, 1823, 648, 1676, 1691, 1943, 820, 1259, 2047, 672, 73, 1280, 1496, 1486, 1273, 1484, 679, 76]. Control-Rule [648]. Controlled [1996, 1972]. Controller [1544]. Controllers [1347, 295, 1236, 1400]. Controlling [1534, 745, 1298, 547]. Convection [991]. Convenience [1784]. Convenient [2048]. Convergence [1449, 1059]. Convergent [1570]. Conversion [822, 1390]. Convex [1039, 399, 323, 2054, 427]. Convexly [134]. Convicted [1443]. COO [712]. Cooperating [1263, 599]. Cooperation [14, 12, 13, 1238, 1244, 1744, 1304]. Cooperative [1414, 235, 1485, 1262, 1503]. Coordinated [28]. Coordinates [159, 1918]. Coordinating [1472]. Coordination [1114, 1115, 1119, 710, 705, 1247, 1121, 1123, 1237, 1124, 709]. Coprocessor [1358]. coproducts [362].

Coq [359, 371, 1572]. Core [183, 1086, 1085]. Cornell [1947]. Corpus [826]. Correcting [1754, 853, 1962, 868]. Correction [1757, 510]. Correctness [1567, 1116, 341, 1459, 1641, 801, 334]. Correctness-Preserving [1116]. correlation [1972, 1077, 1984, 840]. Correspondence [1713]. Corresponding [1495]. Cortex [1274]. Cortical [462, 1261, 1265, 954]. Coscheduling [1937]. Cost [1832, 1577, 1046, 1254, 1496, 1350]. Cost-Effectiveness [1496]. Costs [1729]. Count [93]. Countable [1589]. Counter [872]. counterexamples [440]. Counting [606, 281, 1064]. Couple [2059]. Coupled [98, 1646]. Course [1, 193, 202, 214]. Cover [390, 1056, 1383]. Covering [1898, 1918, 387, 1921]. Coverings [1038]. CPO [785, 1581]. CPOs [2054]. CPUs [549]. CPWM [1368]. CRAFT [982]. CRASH [1018]. crashworthiness [1018]. CRAY [982, 988, 963, 993, 939, 909, 1009, 943, 998]. Creation [168, 33]. creativity [412]. Criteria [9, 1985]. Criterion [1329]. Critic [1521]. Critical [1675, 984]. Cross [99, 1614, 973]. Crossing [400, 140]. Cruciate [505]. Cryptanalysis [856, 1982, 1983, 1060, 864, 1955, 857]. crypto [1091]. crypto-accelerators [1091]. cryptographic [866, 1964, 845, 844]. Cryptography [1906, 1999, 1950, 836]. Cryptosystem [1720, 1058, 862]. cryptosystems [1962, 1084]. CS [990, 992, 926]. CS-2 [990, 992, 926]. CSCW [11, 700, 228, 1634, 239]. CSL [43]. CSP [1877, 1881]. CT [509, 467, 513]. Cubes [512]. cubic [1057]. Cues [825]. Current [707, 1366]. Curricula [182]. Curriculum [185, 183, 184, 198]. curve [834]. Curved [492, 490]. Curves [604, 322, 1916, 1057, 1064, 431]. Curvilinear [489]. Customising [692]. Cut [1442, 1040, 1021]. Cuts [1055, 1022, 1731].

Cycles [410]. Cyclic [629].

D [500, 513, 495, 483, 482, 499, 491, 502, 472, 480, 464, 459, 284, 494, 511, 448, 1014, 129, 460, 506, 474, 493, 918, 2018, 501, 445, 128, 1927, 2015, 484, 937, 461, 508]. D.S.A. [1958]. D0P [101]. D2PC [107]. D3 [1495]. D4 [1930]. Dags [161]. DAI [22, 25]. DAISY [37]. Damgaard [858]. Darboux [1932]. Dark [1232]. DARPA [911]. Darwin [1639]. Data [495, 1702, 1704, 1703, 1165, 2022, 604, 66, 483, 658, 491, 67, 113, 1348, 1506, 1652, 527, 2003, 2023, 543, 2014, 79, 511, 448, 94, 1575, 506, 474, 452, 2001, 92, 2020, 813, 518, 1687, 536, 73, 564, 1685, 794, 1461, 1522, 1593, 485, 567, 450, 2032, 508, 72, 539, 486, 150, 449, 75, 565, 542, 975, 919, 995, 869, 996, 369, 902]. Data-Flow [79, 1575]. Data-Parallel [67, 92, 72, 565, 919]. Data-Specifications [1461]. Database [818, 1428, 1694, 2022, 1829, 2008, 110, 1683, 2007, 1425, 1682, 1686, 94, 1693, 1517, 1696, 258, 1599, 1687, 1684, 461, 2002, 96, 121, 831, 921, 895]. Databases [658, 100, 1681, 114, 2005, 2006, 1597, 2004, 124, 262, 112, 95, 125, 2011, 108, 1504, 1507, 1775, 922]. Dataflow [566, 799]. Dataflow-Based [566]. Datalog [123, 1166, 116, 125, 101, 120]. Datapath [1655]. Dataspace [829]. Datatypes [806, 1885]. Dates [1044]. Davies [1998]. DBMS [2021]. Deadlock [317]. Deaf [1321]. dealer [1954]. Debate [758]. Debugger [557]. Debugging [925, 790]. Decade [757]. Decentralised [704]. Decentralized [1238]. Decidability [1436, 1568, 765]. Decidable [370, 628]. Deciding [723]. Decision [390, 1491, 1501, 663, 321, 665, 651, 652, 1776, 1771, 305, 7, 1495, 1496, 1482, 1515, 413, 435, 1526]. Decision-Making [390]. Decision-Support [1491, 1501, 1526]. decision-theoretic [413]. Decision-Tree [665]. Declarations [1563, 788].

Declarative [383, 1599, 168]. Declustering [2013]. Decomposable [397]. Decomposing [2059]. Decomposition [1045, 1842, 640]. decoupled [961]. Dedicated [1352]. Deduction [746, 1803, 594, 247]. Deductive [1165, 111, 56, 1597]. Deep [99]. Deeper [1410]. Default [1156, 1178, 1180, 1465, 1173, 1848, 1851, 1857, 1859, 1860, 1864, 1866, 1856]. Defaults [1871]. Defeasible [1870]. Definability [1760, 1446, 1470]. Defined [564, 1925]. definedness [1438]. Defining [531]. Definite [1455]. Definition [1706, 1705, 509, 600]. Definitions [1670]. Deformable [495, 501]. Deformations [483, 508]. Deformed [1378, 1999]. Deformity [503]. Degree [1754, 297, 1742, 775]. Degrees [605]. Delay [1563, 1385]. Deletion [881]. Delineation [451]. delivery [1643]. Delphi [208]. Delta [1038, 1740, 598, 104]. Delta-Coverings [1038]. delta-Languages [104]. delta-Levelability [1740]. Demand [1937]. Demand-Based [1937]. demands [420, 905]. Demonstration [816, 155, 157]. Dempster [1768, 1868]. Dempster-Shafer [1768, 1868]. Denial [1155]. Dense [482, 1961]. Density [328, 415, 957]. Dentritic [1296]. Deontic [250, 1870]. Dependability [1150]. Dependence [316, 74, 75]. Dependencies [118, 658, 117, 114, 1066, 73]. Dependent [398]. Depends [1839, 418]. Depth [282, 283, 1227]. Depth-Three [282]. Derivation [1877, 1881]. Derivatives [315]. Derived [270]. Deriving [1637]. Describe [1912, 1687]. Describing [797, 523, 556, 226]. Description [41, 252, 428]. Descriptions [697, 1615, 1487]. Descriptive [263]. Design [1344, 1428, 570, 569, 808, 604, 1536, 230, 517, 348, 848, 1116, 197, 231, 1675, 234, 235, 1544, 220, 511, 1127, 1140, 1749, 217, 345, 694, 335, 1251, 1407, 2043, 601, 1213, 40, 894, 1873, 1635, 941, 868, 1015]. Designated [1959]. Designed [1217]. Designing [171, 212, 1420]. Designs [1905]. Desirability [2002]. desktop [962]. DESV [869]. DESV-1 [869]. Detail [1685]. Detailed [1099]. Detecting [66, 503, 1569, 74]. Detection [470, 317, 511, 1384, 485]. Determination [1137, 1929]. Determine [208]. Determining [1310, 1750]. Determinism [572]. Deterministic [572, 1642, 303, 1739, 120, 765]. Detour [1314]. Develop [1348, 1215]. Developing [1540, 204, 1611, 1579, 1283, 832]. Development [1428, 1347, 699, 1580, 924, 1585, 757, 758, 1586, 760, 1422, 193, 1432, 271, 198, 1217, 1346, 892, 1524, 1141, 689, 1500, 241, 929, 371, 693]. Developmental [1216]. Device [1338]. Devices [1590, 1389, 1734]. Dexterity [455]. Diabetes [1486, 1516]. Diabetic [1528]. Diagnosis [1491, 1501, 1510, 1494]. Diagnostic [1526]. Diagnostics [1854, 1495]. Diagonal [290]. Diagrams [133, 1779, 350, 57, 584, 1776, 1569, 1482]. Diagraphs [170]. Dictionaries [251]. Dictionary [1628]. difference [1001]. Differences [76]. Different [884, 1861, 1798, 1439]. Differential [482, 1983, 1059, 1925, 1926, 1929, 1932, 864]. Differentiation [1214, 693]. Difficult [1106]. Difficulty [153]. Diffusion [497, 893, 939]. Digital [819, 818, 833, 823, 480, 322, 830, 826, 506, 1359, 828, 821, 824, 832, 827, 831, 853, 1964, 862, 1958]. Digraph [1053, 1025]. Digraphs [1728]. Dijkstra [1889]. Dilemma [1224]. Dimension [145, 436]. Dimensional [127, 615, 291, 1608, 1912, 584, 492, 1618, 969]. dimensionality [973]. Dimensions [474, 324]. Diophantine [742]. Direct [1931]. Directed [411, 1022, 132, 1606, 817, 1471, 145, 389, 1105, 1203, 1887]. Direction [819]. Directions [144, 2017]. Discipline [1935]. Discontinuity [1861]. DISCOUNT [746]. Discounted [7]. Discovery [658, 1905, 880, 2005, 1208, 2004, 412]. Discrete [1576, 944, 1758, 295, 2019, 1992, 852, 1966, 1084, 862, 1065]. Discretization [675]. Disguising [752]. Disjunctive [100, 1157, 659, 1186, 1152, 1153, 1160, 1185]. disk [986]. Dispatch [2036]. Displacements [1924]. Display [493]. Distance [1612, 401, 1867, 1629]. Distributed [2021, 80, 746, 803, 1357, 646, 1117, 390, 1031, 526, 701, 115, 235, 29, 559, 1651, 78, 1692, 1532, 91, 86, 1123, 194, 37, 547, 98, 567, 2040, 1304, 975, 1637, 899, 1642, 924, 970, 1017, 1645, 997, 1641, 1006, 1631]. Distributed-Memory [1357, 78, 86]. Distribution [66, 1904, 67, 1744, 1980, 1976, 860]. Distributions [1225, 330, 92, 1059, 1938, 415, 969]. Disturbed [1823]. Divergence [1787, 1762]. Diverse [192, 2043]. Diversification [274]. Diversity [1784]. Divertibility [1963]. divide [1878, 997]. divide-and-conquer [1878, 997]. Dividers [1647]. Divisible [1979]. Division [1901]. Divisor [624]. DL [833]. DL-Raid [833]. DM [1989]. DMR [213]. DNA [867, 1289]. DNA-Like [1289]. DNF [437, 1340]. Do [2037, 2038, 1197]. Doctrine [2055]. Document [822, 825, 824]. Documentation [815]. Documents [823, 826]. Domain [1148, 613, 1709, 1523, 1424, 734, 595, 1110, 1814, 108, 1504]. Domain-Independent [108]. Domains [1571, 1101, 668]. Dominating [1732]. Domination [386]. Domino [380]. Dominoes [388]. Don’t [1617, 1201]. Dosage [1517]. double [1989]. Dragging [180]. Dragging-Through [180]. Drawability [164]. Drawable [166]. Drawing [127, 170, 134, 141, 132, 135, 163, 158, 154, 156, 174, 175, 167, 148]. Drawings [171, 137, 162, 136, 130, 131]. Drift [668].

Driven [209, 675, 1250, 991, 1536, 981, 926]. Drug [1517]. DSPL [978]. DSS [1003, 854]. Dual [659, 1915]. Dualities [2049]. Duration [1663]. During [1424, 514]. Dynamic [1727, 952, 117, 523, 1406, 292, 527, 1577, 313, 1849, 1729, 1483, 460, 1214, 1328, 149, 1046, 1326, 795, 1721, 1341, 1944, 107, 1942, 949, 790, 950, 995, 957, 948]. Dynamical [9, 537, 10, 1242]. Dynamics [1811, 4, 1199, 1856, 991, 940, 966, 953].

E-Collaboratory [15]. E-Unification [745, 871]. E6 [1930]. E8 [1930]. E=MC02 [901]. Each [133]. Eager [728, 782]. Early [1501, 1511]. easily [1443]. easy [419, 1113]. Echocardiographic [475]. Ecology [1245]. Economics [1342]. Edge [470, 1193, 136, 620, 1052, 312]. Edge-Colouring [620]. Edge-Connectivity [1052]. Editing [450, 240, 1629]. Education [219, 180, 215, 186, 211, 192, 194, 1561, 1516, 188, 178]. Educational [212]. Educators [177]. Effect [654, 1528]. Effective [1821, 2052, 343, 1914]. Effectiveness [71, 1496]. Effects [660, 1457, 1343, 1845, 1864]. Efficiency [1500, 1061]. Efficient [1659, 580, 1603, 1539, 660, 1806, 2005, 2014, 1690, 1610, 1611, 974, 297, 1679, 1072, 560, 1981, 1987, 723, 1554, 1130, 405, 846, 1986, 1976, 1884, 1995, 937, 989]. Efficiently [170, 245]. Efforts [257]. EFTTP [752]. Egocentric [1234]. Eiffel [217]. eigenvalue [997]. Eigenvectors [159]. EL-SIM [1347]. Elastic [508]. Elasticity [512]. elections [847]. electric [980, 946]. electromagnetic [936]. electromagnetics [988]. Electromyographic [1527]. Electronic [610, 1979, 473, 825, 1236, 945, 846]. element [1016, 998, 988]. Elementary [65, 884]. Elements [264, 1226]. Elevation [2023]. Elimination [883, 386, 881, 782, 1899, 916]. Elliptic [968, 834, 1064]. EM-X [567]. Embeddable [310]. embedded [1005]. Embedding [153, 1972, 1048, 1185]. Embeddings [400]. Embodied [1231]. Emergence [268]. Emergent [1218]. Emphasis [172]. Empirical [1778, 673]. Emulates [1253]. Emulation [1358]. enabling [945]. Enact [686]. eNCE [394]. Encoding [1289, 1499, 172, 1572]. Encryption [1960, 1977, 1993, 835, 869]. End [1597]. Endangered [201]. Endoscopy [477]. Endosymbiosis [1210]. Ends [572]. Energies [1756]. Energy [1254, 1384, 1788, 433]. Engineer [193, 178]. Engineering [185, 181, 222, 706, 204, 224, 225, 206, 177, 195, 183, 1424, 180, 228, 184, 218, 798, 1410, 211, 192, 194, 202, 239, 198, 1418, 229, 1874]. Engineers [238, 240, 1560]. Enhanced [443, 504, 960]. enhancement [932]. Enhancing [589, 1239]. Enough [709]. Enquiry [1686]. Entailment [1837, 1192]. Entire [1901]. Entities [2016]. Entropy [1865, 1002]. Enumerating [324]. Enveloping [708]. Environment [1694, 2045, 808, 13, 1257, 1536, 833, 260, 1540, 347, 1348, 1347, 1499, 91, 1346, 1143, 809, 1243, 708, 815, 923, 951]. environmental [1003]. Environments [233, 1223, 715, 192, 229, 241, 986, 1640]. Epistemic [1845, 1807, 1188]. Epistemology [1191]. Epsilon [1316]. Epsilon-Back-Propagation [1316]. EQ [87]. Equal [1042]. Equality [1788, 362]. Equals [282]. Equation [1911]. Equational [746, 750, 1670, 1710, 1564, 1559, 532, 752, 584, 744]. Equations [742, 1759, 523, 1745, 267, 531, 1468, 1925, 1926, 1929, 930, 373, 1001]. Equilibrium [1252]. Equivalence [766, 1541, 285, 765, 1761, 373, 435]. equivalence-queries [435]. Equivalences [1568, 768, 305, 1555]. Equivalent [1711]. Erasing [628]. Error [234, 1757, 277, 1332, 853, 1962, 868]. error-correcting [853, 1962, 868]. Errors [1222, 1628]. Escaping [1194]. escrow [1069]. ESPRIT [1003, 1020]. Essay [1796]. Essential [1242, 178]. Establishing [181, 826]. ESTEREL [44, 1657]. Estimates [1688]. Estimating [1784]. Estimation [1387, 482, 1138]. Estimators [1319, 428]. eta [362, 357]. eta-equality [362]. Euclidean [1909]. Europe [13]. European [14, 901, 15, 16, 257, 1020, 942, 959]. Europort [1010]. Europort-1 [1010]. EUROPORT2 [1020]. Evaluate [1529]. Evaluating [89, 221, 1491, 443, 1699]. Evaluation [223, 348, 1145, 1167, 225, 1338, 1490, 123, 1525, 251, 1254, 1140, 1231, 1771, 1431, 1524, 601, 638, 987, 984, 1009, 911, 941]. Evaluations [1489, 1824]. Even [1915, 318]. Event [1571, 1536, 1406, 1477, 1176, 81]. Event-driven [1536]. Eventual [572]. Every [1154]. Evidence [1779, 1767]. Evidential [921]. Evolute [1308]. Evolution [275, 185, 1249, 1208, 1244, 271, 1209, 1212, 1255, 1200, 1204]. Evolutionary [1205, 1596, 1199, 1609, 1613, 1241, 1211, 1623, 1190, 1224, 272, 1312]. Evolutive [1369]. Evolve [2043]. Evolving [1215, 1216, 1236]. Examination [1103]. Example [1505, 170, 732]. Examples [573, 872, 655, 429]. exception [364]. Exceptionally [1106]. Exceptions [647, 1890]. Exchange [1075, 848]. Executable [791, 245]. Executing [35]. Execution [64, 78, 548, 896, 936]. Exemplar [644]. Exemplar-Based [644]. Existence [1756]. Existential [1672, 1454]. existing [865]. Expander [312]. Expanding [360]. Expected [1605, 1721]. Experience [1148, 231, 206, 1319, 990, 1140, 1431, 202, 2043, 2039, 1147, 952]. Experiences [1146, 533, 1018, 2024]. Experiential [190]. Experiment [1685, 894]. Experimental [498, 1329, 1279]. Experiments [697, 1245, 1693, 688, 982]. Expert [1490, 1509, 1530, 1507, 416]. Expertise [1505]. Explained [289]. Explaining [1405]. Explanations [1480]. Explicit [739, 357, 372]. Exploit [1236]. Exploitation [1700]. Exploiting [893, 953, 298, 564, 338]. Exploration [1542, 452]. Exploring [1879, 225, 150]. exponent [834]. Exponential [1865, 394, 866, 431]. Exponentials [1919]. Exponentiation [1906, 622, 1987, 1986]. Expressing [110]. Expressions [315, 1804, 401, 419, 441, 437]. Expressive [1743, 102, 783, 1871]. Expressiveness [1834]. Extended [1534, 1435, 1184, 1163, 1859, 1862, 1160, 2024, 438]. Extending [350, 2009, 73, 1684]. Extension [1751, 1796, 1769, 1177]. Extensional [1855, 360]. Extensions [1156, 352, 1029, 576, 1179, 1923, 1969]. External [108]. Extra [730]. Extracting [358, 506, 1890, 1285, 1340]. Extraction [251, 513]. Extrapolation [1311]. Extreme [324]. eye [933]. Eyes [1602].

Facet [1408]. Facets [1042]. Facilitate [1493]. Facilities [558]. Factoring [867]. Factorisation [1988]. factorization [993]. Factorized [1913]. Factors [1931]. Fair [1661, 570, 1075, 1762, 1074, 1668]. Fairness [1671, 1460]. Fall [1814]. False [89]. Falsity [1151]. FAM [1823]. Families [1904, 526, 317]. Family [301, 1027]. Fans [1036]. Fast [1097, 1719, 1906, 1358, 1604, 169, 622, 1911, 1299, 1057, 469, 633, 1927, 955, 1077]. Faster [1169, 1052]. fault [1642]. Faults [312]. Favour [695]. FCI [963]. FCSR [1078]. FD [589, 601, 1111]. FDDs [607]. Feasibility [1416]. Feasible [1456, 936]. Feature [464, 256, 467, 1812, 1099]. Feature-Based [467, 1812]. Features [463, 487, 1384, 654, 1864, 1819]. Federated [1694, 1423]. Feedback [1338, 1238, 1022, 1969, 1432]. Feedforward [1309, 1500]. FENDER [680]. Few [575]. fibre [900]. Field [1489, 1193, 1065]. Fields [1754, 622, 1281, 1914, 1262, 1265, 980, 1064]. File [1588]. Files [2022, 556]. Filling [604]. Filtered [1238, 841]. Filters [1268, 671, 1689]. Final [367, 95]. Find [404, 1609]. Finding [605, 489, 1716, 310, 69, 277, 1066]. Findings [1429]. Fine [1700, 64, 62, 93, 63, 774, 549]. Fine-Grain [1700, 64, 62, 93, 63, 549]. Finite [315, 1669, 1720, 1754, 1317, 1444, 1712, 1760, 1101, 622, 1735, 872, 595, 1666, 627, 629, 764, 1625, 1001, 1110, 1016, 1064, 998]. Finite-State [1666, 1625]. Finitely [1442]. FIRE [985]. Firing [1290]. First [516, 187, 117, 505, 1445, 1849, 764, 555, 248, 1866, 755, 1450, 558, 1932, 1871, 1892]. First-class [555]. First-Order [516, 1445, 1849, 764, 248, 1871]. Fisher [1434]. Fisher-Honsell-Mitchell [1434]. Fit [483]. Fitness [1200]. Five [264]. Fixed [1659, 2056, 1623]. Fixed-Point [1659]. Fixpoint [1670]. Flat [1691]. Flavour [1133]. Fleng [557]. Flexibility [630]. Flexible [233, 1849, 951, 1550, 1624, 809, 508, 1416, 1884]. Flow [77, 1033, 79, 130, 793, 76, 956, 1575, 913]. Flow-Based [793]. Flow-Insensitive [77]. Flows [1035, 1046, 948]. Fluents [1864]. fluid [953]. fluids [957]. Focus [50, 892]. Focusing [2005]. Foes [1320]. fold [851]. Folding [1253, 1199, 1252]. Folding-Landscapes [1199]. Folds [1131]. Following [474]. Foraging [1246]. forgery [1967]. Form [1440, 483, 740, 1670, 1468]. Formal [349, 808, 2034, 222, 1881, 348, 800, 1908, 20, 232, 21, 1675, 1458, 234, 350, 195, 761, 760, 344, 845, 3, 2010, 1561, 714, 1926, 807, 578, 1929, 1893, 1980, 72, 1421, 393, 353]. Formalising [1438]. Formalism [1096, 236, 1850, 553]. Formalisms [226, 256]. Formalization [231, 1442, 23]. Formation [1511, 1220]. Formed [1539]. Forms [659, 1461, 103]. Formulae [1804]. Formulating [1043]. Formulations [1023]. Fortran [1008, 953]. Forward [1404, 1109, 1365, 1112]. Foundation [1470, 1513]. Foundations [12, 1580, 1765]. Founded [1591, 1169, 1170, 1158, 1159]. Four [1918]. Fourier [508]. FPGA [1367]. Fractal [1808, 1305]. Fragments [265]. Frame [264, 809]. Frameless [448]. Frames [1846, 541]. Framework [2034, 1935, 446, 1178, 1184, 1587, 1794, 1412, 1409, 1029, 1792, 235, 79, 1598, 344, 1478, 1579, 1153, 484, 1871, 542, 942]. Frameworks [518]. France [917]. Free [2058, 1571, 1039, 483, 608, 1732, 1729, 1442, 1447, 1548, 68, 1814, 815, 442, 314, 1090, 433]. Free-Form [483]. Freedom [200]. Frequency [1343, 280]. Friends [1320]. Front [1597]. Front-End [1597]. Frontier [628, 95]. Fudgets [1129]. Fugitive [406]. Fugitive-Search [406]. Full [827, 917]. Full-Text [827]. Fully [747, 296, 1721, 377]. Function [426, 1330, 1754, 1319, 1914, 1286, 1517, 1360, 484, 840, 858, 862]. Functional [126, 1098, 1873, 1240, 1683, 1490, 114, 2041, 1125, 2012, 1786, 1128, 1133, 1459, 1131, 1227, 1251, 1555, 1684, 1130, 1877, 957, 1126]. functionals [374]. Functioning [1263]. Functions [1901, 1361, 1912, 1566, 1853, 1328, 1081, 1922, 1771, 1368, 1086, 1928, 108, 1079, 421, 866, 843, 1956, 1989, 1085, 844, 417]. Fundamental [1701]. Further [1245]. Fusing [446]. Fusion [491]. Future [15, 820, 179, 1011]. Fuzzy [1815, 1330, 1789, 491, 1820, 1347, 1804, 1319, 1786, 1544, 1509, 1328, 1943, 1776, 1785].

GA [1113]. GA-easy [1113]. GA-hard [1113]. Gain [1823]. Galen [1497, 1513]. Gallery [176]. Galois [1908, 1931]. Game [1168, 1670, 199]. Game-Theoretic [1670]. Games [1453, 406, 1667, 1750, 1666, 1244, 276]. Gamma [1309, 1917]. gamma-Planes [1917]. gammaL1 [1291]. Gap [1241, 186, 630, 1638]. Garbage [727, 560, 2035]. gas [940]. Gates [606, 282]. Gauss [622]. Gaussian [1330, 1313]. GCD [1923]. General [742, 1039, 1073, 1157, 1406, 1794, 1971, 1287, 1478, 1468, 1185, 1557, 1016, 896, 1964, 1970, 1892, 1065, 417]. Generalisation [1332]. Generalization [1706, 1705, 646, 1709, 336, 174, 413, 1060]. Generalizations [1440]. Generalized [1602, 1747, 1100, 703, 290, 303, 1035, 1916, 1736, 1550, 996]. Generated [286, 1521, 1419, 784]. Generating [118, 88, 1473, 99, 1649, 751, 1480, 872, 2020, 676]. Generation [603, 1369, 115, 817, 93, 78, 1493, 899, 923]. Generator [1968, 1280]. Generators [280, 792, 842, 840]. Generic [259, 805, 1593, 790]. Genericity [737]. Genesis [1273]. Genetic [331, 646, 1302, 1258, 465, 1326, 273, 1216, 679, 1243, 272]. Genetics [266, 267]. Genome [275, 1619]. Genomes [1612]. Genotype [275]. Genotype-Phenotype [275]. Gentle [798]. Gentzen [1437]. Gentzen-style [1437]. Geographic [2022, 2004]. Geographical [2008]. Geology [1401]. Geometric [152, 1388, 604, 1914, 154]. Geometrical [1376, 175]. Geometry [585, 1909]. Getting [461]. GF} [1066]. Gigabit [900]. Gini [1784]. Gini-Simpson [1784]. Girard [1435, 368]. GIS [903, 2012, 2013]. Given [1922, 1796]. GLB [389]. GLB-Closures [389]. Global [70, 483, 1648, 1115, 560, 1247, 1327]. Goal [33]. Goedel [2]. GOLDRUSH [1695]. Gomory [1055]. Good [1609]. Goppa [1921]. Gossiping [1719]. Governed [1121]. Government [1426]. Gradient [1323]. Grafting [663]. Grain [1700, 64, 62, 93, 63, 549]. Grammar [1249, 156, 1744]. Grammars [171, 612, 1339, 384, 392, 783, 442]. Grammatical [1744]. grand [964]. Granularity [1506, 548]. Graph [152, 1097, 1094, 1096, 411, 171, 127, 577, 1514, 134, 141, 617, 383, 527, 159, 155, 623, 1545, 146, 407, 387, 1040, 1133, 462, 156, 157, 631, 73, 633, 384, 1718, 174, 147, 104, 392, 1742, 394, 783, 393, 563, 1892, 950, 1098, 1095]. Graph-Based [1514]. Graph-Grammar [156]. Graph-Visualization [155]. Graphbots [1758]. GraphEd [146]. Grapheme [1390]. Grapheme-Phoneme [1390]. Graphic [1685, 2016]. Graphical [1098, 230, 146, 1421, 150, 1883]. Graphs [142, 602, 1049, 391, 1725, 1715, 403, 1834, 1773, 396, 398, 137, 1732, 170, 386, 143, 406, 166, 1022, 618, 132, 169, 1606, 1051, 382, 135, 153, 136, 620, 397, 129, 1048, 158, 149, 144, 154, 629, 312, 1721, 172, 636, 145, 140, 167, 389, 405, 638, 1629, 351, 1887]. Gray [1375]. Greatest [624]. Grey [1374]. Grid [137, 133, 163, 1014, 1001]. GRIN [1041]. Gröbner [1913, 1041]. Ground [1452, 583, 239, 723, 1195]. Grounding [1341, 1192]. Group [222, 1905, 1965, 221, 220, 582, 213, 754, 1644, 1061, 1634, 1633]. Groups [1908, 581, 1259, 547, 1931]. Growing [302, 612]. Growth [275]. GSM [1148]. GSOS [785]. GSPN [1542]. Guarantees [18]. Guarded [546]. Guards [1099]. GUI [224]. Guidance [515, 448, 447]. Guide [577, 1431, 412]. Guided [673, 656, 457]. Guidelines [1143]. Guiding [754, 1195].

h [1787]. Hadamard [1915]. Halfspace [151]. Halting [628]. Hamiltonian [410]. HAMLET [1005]. Hand [730, 933]. hand-eye [933]. Handle [224, 1801]. Handles [1548]. Handling [662, 592, 1857, 364]. Handwritten [1380, 1381]. Hard [177, 1086, 1085, 617, 1113]. Hardware [1650, 343, 1458, 350, 2037, 1655, 2038, 332, 1353, 1359, 346, 1354, 1236, 863, 865]. Hardware-Oriented [1353]. Hardware-Verification [332]. Harmony [558]. Harmony/2 [558]. Hash [1082, 1081, 1817, 1086, 1956, 1989, 1085, 858, 1955]. Hash-Key [1817]. hashing [859]. Having [1724]. Hazards [333]. HCl [222, 224, 225]. HDL [740]. Head [738]. Head- [738]. Health [1496]. Heart [450]. heat [935]. Hebb [1265]. Hebbian [6, 1293]. Hedonic [1248]. Heidegger [1194]. Her [178]. Heritage [716]. Heterogeneous [1419, 980, 903]. Heuristic [1773, 1327, 175]. Heuristics [1805]. Hidden [1034, 1110]. hiding [906]. Hierarchical [1535, 1235, 1270, 781, 1212, 1556, 908]. Hierarchy [2028, 1247, 536, 774, 408]. High [1538, 964, 986, 1238, 1101, 1279, 779, 344, 1268, 914, 1938, 1486, 455, 2013, 968, 971, 1200, 946, 953, 935, 902, 1636, 906, 1015, 1632]. high-integrity [1636]. high-latency [906]. High-Level [1101]. High-Pass [1238]. High-Performance [968, 914]. High-Variability [1938]. Higher [863, 769, 740, 113, 1436, 359, 591, 1566, 772, 1128, 550, 889, 1459, 719, 795, 735, 324, 1570]. Higher-Order [769, 740, 113, 1436, 591, 1566, 772, 1128, 550, 889, 795, 735, 1570, 359, 1459]. Highly [1823, 1696]. Hilbert [1040]. Hill [1112]. HiLog [113]. Hints [1854]. HIPPI [900]. Hippocampus [1283]. Historical [2]. History [1481]. HOL [811, 54, 1888]. HOL-UNITY [811]. Home [828]. Home-based [828]. Homeostatic [1811]. Homogeneous [1932]. homomorphic [835]. Homomorphism [1466]. Honsell [1434]. HOOD [1004]. Hopfield [7, 433, 1300]. Horizontal [463]. Hospital [1498, 1499, 1484]. hours [936]. HPC [936]. HPC-Reducing [936]. HPCN [942, 945, 1005]. HPF [987, 969, 982].

HTTDs [54]. Human [234, 238, 1510, 462, 1089]. Humans [1343]. Hybrid [1815, 1663, 1680, 1380, 1518, 1772, 1665, 1735, 1346, 678, 682, 1014]. hydrodynamic [959]. Hydroelectric [1537]. hydrogen [938]. Hypercube [297, 973]. Hypercubes [1988]. Hypermedia [2045, 1556]. Hyperplane [652]. Hyperstack [486]. Hypertext [1412, 822, 1622, 241]. Hypertree [385]. Hypotheses [1787]. Hypothesis [269, 650, 656, 681]. Hypothetical [1175].

I/O [895, 561]. IBM [985, 960]. ICD [1499]. ICD-9 [1499]. ICL [1695]. Ideal [1912, 1047]. Ideals [1901, 1903]. Identical [1043]. Identification [1399, 1068, 2005, 848, 424, 954, 1083, 1089]. Identities [1925, 755, 540]. Identity [1063, 1386]. Identity-mapping [1386]. Ignorant [1482]. II [1807, 1467]. III [756]. Ill [1308]. Ill-Posed [1308]. Illustrations [269]. ILP [656, 681]. Image [1311, 464, 830, 475, 488, 1374, 453, 8, 481, 1503, 497, 457, 444, 1371, 911]. Image-Guided [457]. Imaged [492]. Images [470, 482, 446, 1502, 463, 1372, 1712, 459, 487, 1808, 494, 465, 496, 493, 467, 513, 1818, 1375, 445, 476, 1373, 484, 490, 447, 955, 954, 932]. Imaging [499, 1808, 2027]. Imagining [598]. Imai [1089]. Imbalanced [1811]. Immune [1219]. Impact [1426, 1134, 1142, 211, 1003]. Impaired [1517]. Imperative [787, 542]. Implant [1321]. Implement [2032]. Implementation [1356, 1097, 1098, 1338, 957, 2014, 146, 1041, 341, 566, 1353, 1359, 1355, 1495, 1006, 1927, 601, 1363, 1284, 565, 542, 960, 1884, 974, 1643, 917, 1090, 989, 962, 1065, 883]. Implementations [725, 1567, 1917, 1360]. Implemented [752]. Implementing [117, 1361, 1033, 1294, 906]. Implication [1799]. Implications [15, 1810]. Implicit [1760, 918]. Implicitization [1916]. implies [1967]. Import [2022, 788]. Import/ Export [788]. impredicative [1890]. Improve [1516]. Improved [1023, 1021, 1523, 131, 1739, 1958, 1058]. Improvement [187, 1998, 1137, 199, 1135, 1136, 1404, 189]. Improving [71, 646, 1316]. ImpUNITY [1895]. In-the-Many [702]. Inaccuracies [153]. Incidence [861]. Including [1997, 1989]. Inclusion [281]. Inclusions [540]. Inconsistency [1830, 1175]. Inconsistent [1831]. Incorporates [1345]. Increased [1500]. Increment [123]. Incremental [1537, 1329, 660, 123, 1208, 78, 1359, 1181, 2035]. Indecomposable [1910]. Independence [1835, 1793, 744, 340]. Independent [224, 615, 1056, 108, 1878]. Index [1093, 2000, 2046, 78, 1667, 775]. Indexing [726, 827]. Indifference [620]. Individual [1312]. Induced [2049]. Induction [660, 1885, 1462, 666, 1825, 674, 676, 679, 1507]. Induction-based [679]. Inductive [1595, 1670, 661, 662, 649, 796, 8, 654, 336, 1470, 418, 1891]. Inductively [1523]. Industrial [1146, 1401, 1342, 1561, 1143, 977, 1012, 1018, 1010, 1015]. Industry [1426, 182, 210, 186, 178, 945, 958, 934]. Inequalities [777, 1024, 1056]. inequational [369]. Inequations [742]. Inertia [1843]. inevitable [1091]. Infarction [1501]. Infection [1484]. Inference [805, 1801, 768, 882, 1187, 1172, 1785, 1469, 418]. Inferences [2017]. Inferencing [2016]. Infinitary [1760, 736]. Infinite [1581, 1652, 285, 286, 1651, 276, 775]. Infinite-State [1651]. Infinitesimal [1871]. Influence [166, 1779, 1286, 1776, 1482]. InfoHarness [1419]. InfoMall [945]. Information [1827, 1295, 1789, 1831, 1473, 1426, 706, 201, 1481, 1839, 1506, 1586, 1851, 533, 850, 1498, 2004, 1771, 1787, 1822, 1562, 1770, 1407, 544, 128, 1419, 821, 1416, 1533, 1975, 1061, 1634]. Information-Based [1481]. information-theoretic [1061]. Informational [1788, 1790]. Infrastructure [15]. Inheritance [1257, 111, 1683, 1183, 532, 873, 2030, 1411]. Inhomogeneities [510]. Initial [1450, 1887]. initial-algebra [1887]. Initialization [1331]. Initiated [1115]. Initiative [179]. injection [1642]. Ink [163]. Input [1271, 1457, 1310, 402]. Input-Output [402]. Input/Output [1457]. Inputs [1297, 1917, 1448]. insecure [854]. Insensitive [77]. Insertions [1731]. Insights [1647]. Inspection [1230, 1550]. Inspiration [1263, 2039]. Instance [327]. Instantiated [1483]. Instantiation [1181]. Institution [806, 538]. Institutional [541]. Institutions [529, 536]. Instruction [347]. Insufficient [1522]. Integer [1027, 1030, 1041, 335, 741]. Integrals [1932]. Integrated [196, 260, 28, 1422, 466, 901, 912]. Integrating [230, 700, 1102, 1544, 829]. Integration [707, 1502, 1423, 664, 184, 168, 1922, 673, 1654, 1504, 1520, 924, 693]. Integrity [1694, 121, 1636]. Intel [982, 907]. Intelligence [1511, 37]. Intelligent [1347, 1943, 451, 30]. intense [938]. intensive [910, 895]. Interacting [553, 39]. Interaction [36, 2047, 1239, 1939]. Interactions [270, 1289, 1120, 2040]. Interactive [880, 1348, 155, 712, 226, 451, 157, 929]. Interactively [1677]. Interactivistm [1240]. Interconnected [1300]. Interface [1497, 223, 230, 562, 1587, 697, 221, 235, 238, 237, 1696, 2042, 1687, 1420, 227, 229, 1275]. Interface-Based [2042]. Interfaces [1536, 231, 221]. Internal [1386, 1764, 767]. International [1137]. Internet [902]. Interobserver [1523]. Interoperability [1599]. interpolation [426, 373]. Interpretable [699]. Interpretation [1502, 1100, 1162, 1489, 1806, 1463, 595, 1685, 353, 354, 839]. Interprocedural [80, 77, 799]. Intersection [1903, 1747, 1436, 225, 241, 370]. Interspecies [1218]. Interval [1821, 587, 299, 382, 247, 1025, 2018]. Interval-based [247]. Intervalizing [1715]. Intervals [1782]. Intra [448]. Intra-Operative [448]. Intravascular [493, 492]. Intrinsic [842, 422, 423]. Introducing [1822, 1345]. Introduction [11, 568, 1564, 798]. Introductory [193]. Intuitionistic [1458, 878, 877, 879, 356, 1875]. Invariance [1853]. Invariances [1285]. Invasive [456]. Invention [110]. Inverse [1931]. Inversion [1659, 284]. inversions [858]. Investigation [1257, 1527]. Investigations [250]. Investment [188]. Invited [1619, 1627]. Invocation [804]. Involvement [182]. ionization [938]. iPSC [1949]. iPSC/860 [1949]. IRIX [1935]. Irregular [494, 969]. irregularly [1077]. Isabelle [887]. ISHs [1282]. Isomerism [1897]. Isometry [1910]. Isomorphic [404, 1437]. Isomorphism [411]. Isomorphisms [1569]. Issues [818, 707, 223, 870, 1401, 224, 1933, 1396, 694, 241, 1630, 900]. Iterated [283, 1224, 1961]. Iteration [93]. Iterative [666, 967, 998]. ITU [1003]. ITU-Land [1003].

Jargon [1474]. JASMINUM [231]. jet [956]. JIGSAW [657]. Job [1949, 1933, 1945, 1948]. Jobs [1042, 1043, 1044, 1937]. Join [2021, 115]. Joins [1689]. Juggling [1814]. Junction [1772]. just [693]. just-in-time [693].

K6 [1040]. Kernel [600, 1006]. Kernels [1313]. Key [818, 1720, 1820, 610, 1076, 279, 1401, 861, 1817, 1472, 1980, 863, 848, 1976, 1069, 860, 693, 1995, 862, 1978]. Key-Picture [1820]. Keys [114, 1990]. keystream [842]. Kidney [445]. KIV [52]. Knapsack [1028, 1956]. knapsacks [1961]. Knee [505, 1530]. Knowledge [26, 1827, 1828, 1830, 1508, 1963, 1518, 1087, 1804, 1489, 2005, 1596, 1108, 1102, 1424, 1849, 251, 1498, 261, 1493, 1492, 1565, 1521, 1496, 1516, 1528, 1524, 2017, 1475, 205, 1869, 461, 1504, 1520, 653, 178, 1092, 849, 1088, 1091]. Knowledge-Based [1518, 1489, 1498, 1493, 1492, 1565, 1516, 1528, 1524, 1520]. Knowledge-Bases [1504]. Kohonen [1393, 1375]. Kolmogorov [419, 1002, 424, 102]. Kolmogorov-easy [419]. Konrad [943]. Konrad-Zuse-Zentrum [943]. KP [1029]. KSE [1777]. KZ [2055]. KZ-Doctrine [2055].

L0k [1449]. Labelings [136]. Labelled [404, 402, 875]. Laboratory [1493, 832]. Lamarckian [1257]. Lambda [1434, 361, 2059, 736, 1438, 1439, 351, 725, 1437, 366, 367, 372, 376, 377, 378]. lambda-calculi [372]. Lambda-calculus [361, 725, 351, 1437, 366, 367, 376, 377, 378]. Lambek [886]. LAN [900]. Lanczos [1066]. Land [1003]. Landscape [1200]. Landscapes [1199]. Language [1697, 517, 521, 608, 895, 2057, 2026, 257, 1119, 1506, 1683, 87, 826, 1175, 530, 551, 566, 36, 157, 84, 1687, 695, 804, 1122, 254, 547, 597, 2016, 829, 557, 38, 240, 601, 1456, 1761, 558, 681, 40, 1349, 775, 565, 894]. Language-Based [240]. Languages [126, 785, 1098, 111, 302, 113, 1100, 696, 591, 2012, 625, 2059, 2037, 2038, 97, 1447, 1459, 1736, 694, 287, 795, 264, 314, 1551, 104, 774, 578, 394, 1684, 72, 102, 519, 429, 439, 593]. Large [1415, 1344, 1951, 2005, 533, 158, 1078, 829, 1147, 1362, 931, 936, 1007]. Largest [310]. Laser [454, 938]. LASSAP [1007]. latency [906]. Lattice [1039, 1385, 940, 1058]. Lattice-Based [1385]. Lattices [1858, 677, 175]. Law [1121, 1825]. Law-Governed [1121]. Laws [1449]. Layered [32]. Layout [171, 169, 129, 168, 147]. Layouts [133]. Lazy [524, 728, 782]. LCM [58]. Leadership [204]. Learn [1339, 1217]. Learnability [419, 439]. Learnable [1242]. Learned [1523, 238, 1140, 1631]. learners [430]. Learning [1311, 931, 1291, 415, 1330, 440, 643, 644, 441, 421, 1306, 1329, 1314, 182, 647, 1309, 659, 1847, 830, 1319, 1755, 648, 664, 649, 639, 650, 1318, 1286, 321, 1266, 666, 1298, 1308, 1328, 669, 670, 1231, 640, 671, 641, 1323, 672, 673, 1386, 1494, 1218, 633, 1861, 654, 1221, 1239, 1222, 677, 1325, 438, 435, 1285, 1400, 653, 1312, 434, 426, 414, 442, 422, 413, 437, 427, 423, 429, 425, 431, 989, 213]. Lecture [1619, 1627]. Left [99]. Left-Deep [99]. Leftmost [737]. Legal [69]. Leighton [1978]. Lemma [737, 1857, 1060, 750]. Length [1042, 1605, 114, 1287, 1757, 1617, 1989, 428]. Lengths [1443]. Lesions [1514]. Less [163, 456, 1091]. Lessons [1219, 238, 1631]. Letter [766, 1729]. Letter-to-Letter [766]. Level [26, 1538, 1101, 1278, 344, 1375, 1486, 730, 790, 2033]. Level-Confluence [730]. Levelability [1740]. Levels [739, 1522, 415]. Lexemes [252]. Lexical [259, 263, 255]. Lexicon [260, 265]. Libraries [819, 823, 830, 821, 827, 831]. Library [818, 833, 826, 1127, 1812, 828, 824, 832]. Lie [1443]. Life [1225, 269, 1253, 1196, 1197, 1202, 1191, 1192, 1190, 1195, 1194]. Lifschitz [775]. Liftable [1453]. Lifting [2055, 1056]. Ligament [505]. Light [1233, 530, 60]. lightweight [1643]. Like [2057, 1289]. Likelihood [485, 417]. Likelihood-Ratio [485]. Limit [1761]. Limitations [1942, 2002]. limited [863]. Limits [1, 1646]. Linda [1118, 1121]. Line [610, 322, 399, 1922, 1380, 292, 413, 409, 1733, 417]. Linear [742, 1902, 65, 1441, 1904, 1982, 1663, 1680, 472, 1983, 1732, 1444, 615, 291, 887, 1953, 404, 618, 1910, 1971, 593, 773, 1048, 546, 550, 277, 888, 1994, 1032, 1926, 857, 2002, 1932, 853, 356, 841, 1970, 1060, 864, 998, 1952]. Linear-Time [615, 618, 773]. Linguistic [1336, 256]. Linguistically [1406]. Linguistically-based [1406]. Linguistics [826]. Link [881]. Linking [1975, 1840]. Links [1983, 1411]. LINPACK [983]. List [1454]. Lists [321, 435]. Liveness [1661, 1534]. Liver [1502, 509]. Living [1251]. Load [940, 2013, 951, 949, 948, 950]. Load-Balancing [2013]. Local [1661, 1387, 1830, 1838, 1578, 1398, 526, 1278, 1266, 1327, 1664, 1515, 1895]. Localisation [261]. Locality [70, 559]. Locally [1463]. Locating [502]. Locking [105]. Locomotion [1280]. LOFT [813]. Log [1448, 1721, 326, 1084]. Log-Approximable [1448]. log-Space [326]. logarithm [1992, 852, 1966]. logarithms [862, 1065]. Logic [1797, 729, 1563, 1660, 1650, 1829, 587, 1441, 521, 1831, 1168, 1178, 1833, 1184, 242, 110, 613, 113, 1444, 1804, 1163, 1760, 887, 250, 1162, 647, 661, 1183, 662, 1161, 1806, 1458, 1451, 244, 1175, 1564, 1442, 1752, 1559, 1850, 1800, 772, 886, 1851, 773, 546, 788, 550, 247, 243, 1857, 1164, 716, 1859, 245, 1176, 1227, 878, 764, 1153, 1593, 1188, 635, 1160, 655, 676, 1182, 1870, 557, 561, 879, 771, 637, 1284, 25, 1155, 1177, 1872, 356]. Logical [1178, 1580, 1446, 22, 1164, 518, 23, 1698]. Logics [1799, 100, 1464, 294, 874, 1674, 1803, 875, 890, 1672, 1447, 1582, 1469, 876]. Logistic [1289]. LOKI [1990, 857]. Long [1904, 1921]. Longest [1624]. Look [1628]. Look-Up [1628]. Loop [89, 1170, 75, 920]. Loosely [98]. loss [416, 904, 417]. Lossless [2001]. Lost [716]. LotoTis [807]. Low [1356, 328, 834, 1268, 1350]. Low-Cost [1350]. Low-Power [1356]. Lower [606, 1738, 609, 306, 1723, 1971, 282, 321, 631, 837, 1899, 381, 140, 408, 1769, 1970].

LPG [521]. LSI [863, 1917]. LTP [1318]. LU [993]. Lucas [862]. Lukaszewicz [1860]. Lung [513]. Lustre [45]. LVQ [1351]. Lymph [1492].

Machine

[1097, 1361, 1587, 830, 251, 258, 566, 1459, 909, 641, 1494, 549, 1124, 973, 898, 943]. Machine-Assisted [1587]. Machine-checked [1459]. Machine-Readable [251]. Machinery [728]. Machines [1669, 1738, 1357, 1317, 1279, 1207, 78, 1696, 1396, 628, 1360, 632, 86, 1251, 1044, 1395]. MacKE [891]. Maclane [1909, 2051]. Macromolecules [1619]. made [936]. Magic [1159, 855]. Magnetic [469, 167]. Magnetic-Spring [167]. magnetohydrodynamics [918]. Maintaining [1579]. Maintenance [1550, 1431, 689]. Major [1550]. Majorization [1054]. Make [1053]. Making [390, 1610, 1771, 1482]. Malformation [1504]. Malign [330]. Mammalian [1283]. Mammograms [489]. Mammographic [488]. Management [1694, 260, 1477, 1412, 1782, 1425, 1682, 702, 1530, 1485, 1733, 211, 202, 1947, 1948, 1524, 1814, 809, 393, 946, 897]. Management-Oriented [202]. Managing [1506, 1813, 192, 214]. Mandelbrot [634]. Manipulation [1421]. MANTRA [1361]. Manufacturing [1550]. Many [749, 696, 1104, 1807, 1442, 702, 838]. Many-to-One [749]. Many-Valued [1807]. Map [275, 1305, 174]. Mapped [1357, 1376]. Mapping [1367, 920, 1292, 1493, 7, 1386, 1006]. Mappings [69]. Maps [1306, 1315, 1353, 130, 1375, 1222, 2010, 2013]. Marginal [1865, 1798, 1774]. Marginals [1774]. Mark [752]. Markov [1050, 1791, 1510, 7, 864, 1557, 1774]. Markovian [491]. Marriage [2030]. Masks [468]. Massey [1927]. massive [1878]. Massively [1257, 560, 566, 1364, 893, 965, 939, 947]. Master [183]. Matcher [749]. Matching [2031, 1095, 1603, 1380, 1604, 615, 291, 1031, 1606, 1607, 1608, 407, 1615, 1617, 1618, 467, 1620, 1622, 1626, 441, 954, 379]. Matchings [1611]. Mate [274]. Mathematical [1211, 500, 716]. Mathematics [1874, 1, 763, 1560]. Matrices [1577, 1717, 1915, 1787, 1047]. Matrix [605, 609, 1293, 1924, 877]. Matroid [1036]. Matroids [142]. Matsui [1982, 1060]. Matsumoto [1089]. Matter [1254]. Mature [2035]. Maturity [215]. Maude [534]. Max [1035, 1021]. Maxima [1648]. Maximal [1039, 1907, 398, 1716]. Maximum [385, 1033, 621, 139, 277, 1865, 417]. May [1217, 372]. Mazurkiewicz [1709]. MCM [58]. MDL [429, 674]. Mean [1193]. Means [572, 745, 1172, 1496, 1322, 589, 1342]. Measure [1789, 1755, 674, 1867]. Measurement [216, 686, 1825, 1147, 981]. Measurements [1205, 982]. Measures [1855, 1449, 1787, 1770, 1769]. Measuring [803, 401, 481]. Mechanical [1477, 430]. Mechanics [763]. Mechanism [729]. Mechanisms [1273, 5]. Mechanized [796, 340, 342]. Mechanizing [1583]. Medical [483, 470, 446, 1505, 1473, 463, 1477, 464, 459, 1519, 487, 1808, 494, 1509, 29, 1476, 453, 1495, 1482, 1474, 1475, 1503, 1513, 1500, 478, 955, 932]. Medicine [1493, 1520]. Medium [1130]. Meet [1899]. Meets [1245, 452]. MegaSERVER [1695]. MEIKO [926]. MEM [1379]. Membership [1330]. Memory [80, 1357, 1648, 255, 1382, 78, 298, 91, 86, 1721, 1939, 1995, 1234, 1637, 977, 1017]. Memory-Based [255]. Merge [318]. Merged [1839]. Merging [652, 1131]. Merits [1644]. Mermaid [941]. Mesh [1739, 939]. Meshes [395, 494]. Message [2036, 1120, 1081, 1966, 926, 1939, 852, 1018, 1643]. Message-driven [926]. Message-Passing [1939, 1018]. Messages [307, 1719]. Meta [713, 1417, 852, 2033, 554, 815, 901, 809]. meta-blind [852]. Meta-Environment [815]. META-Frame [809]. Meta-level [2033]. Meta-message [852]. Meta-Model [1417]. Meta-process [713]. Metabolism [1204]. Metaclasses [1702, 1706, 1707, 1704]. Metadata [1419]. Metafiles [2008]. METAGEN [1600]. Metalevel [1108]. Metalogical [1457]. Metaobject [2032]. METATEM [246, 248]. Metatheory [879]. Meteo [917]. Meteo-France [917]. Method [870, 1033, 1380, 294, 874, 1752, 1690, 503, 642, 1405, 878, 336, 1418, 167, 338, 1500, 918, 988]. Methodology [1964, 208, 1136, 929]. Methods [219, 222, 1820, 589, 232, 1675, 1658, 234, 1803, 760, 220, 1898, 535, 1288, 672, 1181, 1561, 203, 1814, 1352, 1507, 935, 998]. Metric [242, 1582]. Metrics [1146, 1815, 687, 683, 1418]. MHz [493]. Micali [1978]. Mickey [1045]. Mickey-Mouse [1045]. Micro [1269]. Micro-Structures [1269]. Microcirculation [481]. microcontrollers [1091]. Microeconomic [1941]. microprocessor [1880]. Microprocessors [334]. Microscope [515]. Microsurgery [455]. Migrating [16]. migration [1019]. Milord [1807]. Milord-II [1807]. MIMD [1988, 941, 934]. MIMD-Factorisation [1988]. Mimicry [1248]. Mind [1194, 420]. Mind-Set [1194]. Minimal [1902, 1550, 1718, 1028, 75, 1697, 424]. Minimising [1227]. Minimization [649, 1448]. Minimizing [1609]. Minimum [137, 390, 1022, 618, 132, 1046, 1731, 428]. Minimum-Width [137]. mining [921]. Mitchell [1434]. mix [1090]. mix-type [1090]. Mixed [1441, 1027, 1918, 1362, 405].

Mixed-Integer [1027]. ML [2061, 377]. MLP [1291, 1312]. Mobile [1763]. Mobility [1758, 1764, 767]. Möbius [1838]. Mod [282, 1065]. Mod-Gates [282]. Modal [1659, 874, 1458, 1803, 875, 1469, 294, 876]. Modality [491, 464]. Mode [1366, 1221, 561]. Model [1497, 1702, 259, 1591, 1387, 483, 883, 805, 509, 1149, 111, 880, 1253, 1578, 1240, 1506, 206, 1760, 1652, 20, 1658, 1671, 1674, 1199, 1544, 1417, 1266, 1282, 1374, 1493, 1854, 1272, 1673, 881, 2047, 1121, 1337, 8, 1550, 1274, 1280, 1123, 1381, 1420, 2010, 1654, 597, 198, 1552, 714, 120, 1556, 1379, 1283, 1664, 1141, 968, 1421, 1276, 106, 334, 1154, 1322, 337, 353, 356, 995, 849, 1638, 365, 981, 959, 927]. Model-Based [1497, 206, 1854, 8, 1149]. model-driven [981]. Model-Landscapes [1199]. Modeling [1661, 1828, 502, 1406, 1477, 1810, 383, 1498, 1268, 1261, 271, 1273, 907, 691]. Modelled [1290]. Modelling [196, 1411, 1704, 1096, 1703, 491, 512, 1219, 696, 1336, 1414, 701, 1246, 236, 1476, 1673, 711, 1639, 462, 32, 704, 1407, 1413, 1513, 698, 1416, 1533, 941]. Models [785, 1225, 495, 1479, 1441, 1508, 706, 1139, 1542, 699, 1581, 733, 687, 1649, 232, 1307, 28, 1297, 1480, 736, 1353, 1269, 1818, 1685, 1198, 642, 1405, 1408, 692, 1496, 1160, 508, 2002, 784, 1638]. Moderate [1742]. modes [856]. Modula [59]. Modula-3 [59]. Modular [606, 1537, 1348, 1543, 528, 788, 631, 1569, 1363, 1213, 863]. Modularity [717]. Modulation [1264]. Modules [2061, 556]. Modulo [2048]. Molecular [159, 1190, 2015, 991, 940, 966]. Molecule [1252]. Monadic [1650, 1444, 1446, 1742]. Monads [1131, 1126]. money [1964, 846]. Mongruences [1573]. Monitoring [1478]. Monoids [308, 586]. Monolingual [265]. Monotone [1838]. Monotonic [1837, 647]. monotonicity [420]. Monotonous [632]. months [936]. Morphism [286].

Morphisms [776]. Morphology [500]. Morse [506]. Most [2029]. Motion [482, 480, 1278, 476]. Motivated [33]. Motor [1222, 1237]. Motorola [215]. Mouse [1602, 1045]. MPP [937]. MR [496, 467, 500, 476, 486]. MR-Angiography [500]. MRF [462]. MRI [498, 474, 510, 485, 450, 508, 449]. MSC [1017]. MSC/NASTRAN [1017]. MTA [1934]. Multi [26, 1311, 1935, 470, 1505, 491, 19, 464, 21, 1258, 703, 1022, 1608, 875, 548, 1599, 898, 1770, 2011, 23, 1870, 969, 999, 1015]. Multi-Agent [26, 19, 21, 23]. multi-algorithmic [999]. multi-block [1015]. Multi-Cuts [1022]. Multi-Database [1599]. Multi-Dimensional [1608, 969]. Multi-Discipline [1935]. Multi-Expertise [1505]. multi-machine [898]. Multi-Modal [875]. Multi-Modality [491, 464]. Multi-Objective [1770]. Multi-operation [898]. Multi-Parent [1258]. Multi-Scale [470, 2011]. Multi-threaded [548]. Multi-Valued [1311]. Multi-View [703]. Multiacuity [1387]. Multialgebras [540]. multicast [1632]. Multiclass [1377]. Multicomputers [1939]. Multicut [1023]. Multidatabase [98]. Multidisciplinary [1496]. Multifeature [835]. Multigraphs [1052, 76]. Multihead [627]. Multilayer [1377, 1365]. Multilevel [384]. Multimedia [1682, 1123, 955]. Multimodal [1282, 447]. Multiparty [306]. multipermutations [1955]. Multiple [1935, 1829, 2029, 89, 1527, 1306, 1806, 1422, 1610, 722, 630, 1623, 1626, 1379, 653, 122, 856, 1961, 916]. multiple-iterated [1961]. Multiple-Knowledge [653]. Multiple-Processor [1935]. Multiple-valued [1806]. Multiple-Writer [89]. Multiplication [283, 1900, 863]. Multiplicative [836]. Multipliers [1647]. Multiprocessor [1938, 567, 977].

Multiprocessors [1357, 1937]. Multiprogrammed [1937]. Multiresolution [2010]. Multiround [1070]. Multiscale [487, 954]. Multisensor [1822]. multiset [371]. multisignature [1967]. Multislices [1529]. Multivariate [652, 1922]. Multiword [252]. MUSiC [1145]. Must [1951]. Mutable [547]. Mutation [1326, 1206, 273, 63]. Mutual [1068]. Mutualism [268]. mutually [1072]. MVP [686]. MVP-S [686]. Myocardial [1501, 484]. MyWorld [25].

N3S [1012]. Naive [1686]. Name [1796]. Named [637]. Naming [1432]. Narrowing [721, 782, 1570]. NAS [1948, 983]. NASA [1949]. National [832]. Natural [1506, 826, 1432, 763, 829, 1572]. Natural-Language [1506]. Navier [1015]. Navier-Stokes [1015]. Navigating [1233]. Navigation [262]. NC [410, 398, 1716]. NCX [565]. Near [1031, 1973]. Near-Optimal [1031]. Nearest [617, 682, 1312]. Nearly [1154, 1078]. Nearly-Stable [1154]. Need [2037, 2038, 1865, 958]. Needs [208, 1191, 1870]. Negation [729, 591, 101, 1177]. negative [354]. Neighbor [682, 1312]. Neighborhood [2015]. Neighbour [617]. Neighbourhood [385]. Neocortex [468]. NEPHARM [1517]. Nervous [1277]. NESIS [1506]. nests [920]. Net [1536, 1759, 1544, 1403, 778, 1556]. Nets [2049, 780, 1534, 1535, 1537, 1538, 1539, 523, 1541, 1578, 1543, 317, 1679, 304, 779, 1546, 1547, 1548, 1549, 1550, 1342, 873, 1552, 1795, 1553, 1554, 1555, 1557, 1558, 1533, 996]. NETTOOL [1346]. Network [1034, 15, 1383, 1291, 1401, 1219, 1358, 1372, 1491, 1501, 1378, 1847, 1333, 1294, 1385, 665, 1943, 1749, 1337, 1274, 1277, 754, 1376, 1365, 1283, 1265, 1515, 1331, 1345, 1393, 974, 933]. Networked [821, 940, 966]. networking [914]. Networks [1778, 1344, 1311, 1263, 1356, 1388, 1367, 1330, 1369, 1502, 1361, 646, 1389, 1301, 1527, 1339, 1836, 1380, 4, 1756, 1338, 1336, 1302, 1264, 1309, 1289, 1303, 1199, 639, 1849, 1299, 1318, 1483, 1370, 1287, 1266, 297, 1292, 1852, 1293, 1321, 1777, 1326, 1360, 629, 1261, 300, 7, 1332, 1215, 1861, 1355, 1397, 1354, 676, 1765, 1304, 148, 1350, 1362, 1520, 1400, 1363, 1340, 1371, 1500, 1390, 1284, 1775, 931, 952, 903, 436, 932, 1955, 904, 906, 433]. Neural [1344, 1263, 1356, 1388, 1383, 1367, 1330, 1369, 1502, 1361, 646, 1402, 1389, 1301, 1527, 1401, 1339, 1358, 1372, 1380, 4, 1756, 1338, 1336, 1491, 1501, 1302, 1264, 1378, 1309, 1525, 1303, 1294, 1385, 1281, 1318, 1370, 1287, 1266, 1282, 1403, 304, 665, 1298, 1308, 933, 1321, 1326, 1259, 1360, 932, 1269, 1337, 1342, 1274, 1277, 7, 1332, 1215, 754, 1355, 1381, 1376, 1334, 1354, 1373, 1273, 676, 1365, 1283, 1265, 1331, 1352, 1351, 1350, 1520, 1275, 5, 1400, 1345, 1363, 1324, 1340, 1371, 1500, 1284, 931, 436, 433, 1349]. Neural-Network [1401]. Neuro [1347]. Neuro-Fuzzy [1347]. Neurobiological [1263]. Neurocomputer [1364]. Neuroconsciousness [1335]. Neurology [1521]. Neuromathematics [1288]. Neuronal [1297]. Neurons [1311, 1290, 1315, 1305]. Neurosurgery [448, 471]. Neurosurgical [515, 447]. Neutral [1199]. neutron [939]. NEXP [329]. next [923]. next-generation [923]. NIZK [1092]. No [1020, 1247]. Node [173, 1492]. Noise [660, 1241]. Noisy [1339]. Non [1797, 1151, 1291, 884, 482, 1441, 1831, 1837, 1453, 647, 465, 243, 628, 804, 1621, 335, 1940, 825, 120, 1766, 1534, 1643, 836]. non-Abelian [836]. Non-Alethic [1797]. non-blocking [1643]. Non-Chaotic [1291]. Non-Deterministic [120]. Non-elementary [884]. Non-Erasing [628]. Non-Liftable [1453]. Non-Linear [1441]. Non-Monotonic [1837, 647].

Non-Refutable [1151]. Non-Restoring [335]. Non-Rigid [482]. Non-Segmented [465]. Non-Speculative [804]. Non-Standard [1621, 1766]. Non-Textual [825]. Non-Trivializable [1831]. Non-Work-Conserving [1940]. Nondeterminism [626]. Nondeterministic [784]. Nonfinite [777, 1710]. Nonlinear [1023, 4, 1823, 1665, 1079, 918]. Nonlinearity [1985]. nonlinearly [841]. Nonmonotonic [100, 1183, 1846]. Nonmonotonicity [1187]. nonperfect [837]. Nonrecursive [123]. nonrestoring [863]. Nonself [1219]. Nonsensing [1746]. Nonstandard [611]. Nonstationary [1044]. Normal [659, 1920, 1153, 320, 1154]. Normalization [2058, 1427, 738]. Normalized [583]. Northwest [959]. Notation [563]. Notations [1422, 418]. Note [1920, 1770, 430]. Notes [545]. Notion [1541, 538, 1195]. Notions [1867]. Novel [460, 1014]. NP [284, 617, 1446, 1724, 1742]. NP-Complete [284, 1446]. NP-hard [617]. NSL [1349]. NSL-Neural [1349]. Number [1722, 133, 1607, 1609, 144, 325, 1522, 1064, 1065]. numberings [424]. Numbers [611, 662, 400, 140, 320, 1969]. Numeric [1825, 654, 1871]. Numerical [1792]. Numerically [910].

o [725, 895, 561]. O0p101 [720]. OBDD [1648]. OBDDs [1748]. Object [787, 1344, 1694, 1114, 1702, 1701, 1706, 1707, 1705, 2022, 2034, 230, 499, 1115, 2026, 111, 1681, 2028, 525, 1519, 2041, 1813, 1102, 530, 817, 2037, 2038, 1597, 2044, 551, 1547, 534, 1140, 451, 186, 217, 1337, 1123, 537, 37, 2030, 490, 2027, 2024, 444, 1819, 1411]. Object-Based [499, 1102, 2030]. Object-Oriented [1702, 1701, 1707, 2022, 2034, 230, 2026, 2028, 525, 1519, 2037, 2038, 1597, 551, 534, 1140, 186, 37, 1694].

Objective [1770]. Objectory [206]. Objects [1415, 2049, 2029, 1434, 502, 472, 526, 506, 1476, 770, 1818, 325, 552, 547, 2040, 1379, 508, 1411]. Oblique [678]. Oblivious [1068]. OBS [689]. Observations [1532]. Obstacle [1322]. Obtaining [902]. Occurrences [263]. OCore [551]. Odd [318]. Odd-Even [318]. OFDDs [1748]. Off [610]. Off-Line [610]. offline [414]. offs [1958]. oil [915]. OKFDDs [1748]. omega [1680, 1466, 288, 287, 774, 784, 1449]. omega-Automata [288]. omega-Languages [287, 774]. omega-Regular [1680]. omega-Regularity [1466]. omega-Rule [784]. On-Line [399, 1380, 292, 409, 1733, 417, 413]. Onboard [689]. One [749, 696, 616, 1445, 619, 1927, 964, 868]. One-D [1927]. Ongoing [97]. Online [414]. onto [1367, 1357, 1019, 1493, 7]. Ontological [1513]. Open [34, 2012, 2042, 694, 1621, 1860, 1640]. Operant [1322]. operation [856, 898]. Operational [521, 1139, 1119, 363]. Operations [1295, 512, 555, 106, 1439]. Operative [448, 514, 671]. Operator [1165, 1186]. Operators [242, 511, 272, 1879]. Optical [1365]. Optimal [1740, 1271, 725, 1726, 1367, 1960, 395, 1730, 99, 946, 162, 1031, 306, 2007, 1729, 304, 319, 1750, 1973, 1637, 432, 428]. Optimal-Area [162]. Optimality [266]. Optimisation [1692, 1689, 999, 944]. optimising [972]. optimistic [896]. Optimization [70, 1398, 1054, 2028, 591, 1199, 119, 267, 1208, 274, 112, 335, 1397, 98, 1371, 355]. Optimizations [80]. Optimize [1752]. Optimizer [796]. Optimizing [725, 79, 83, 92, 1181, 1243, 963]. Optimum [1377]. Opus [84]. Oracle [632]. Order [769, 516, 1650, 740, 1339, 113, 1436, 1444, 1445, 591, 1849, 1279, 1566, 532, 772, 1594, 1128, 550, 889, 719, 1025, 795, 735, 764, 1570, 248, 1866, 1450, 1871, 359, 1459, 1984, 1643, 379]. Order-Sorted [1566, 532]. ordered [1634, 1632]. Ordering [609, 1054, 719, 637, 371]. Orderings [385, 1444, 386, 1453, 751, 722, 1751, 718]. Orders [316, 115, 382, 151]. ordinal [418]. Organisational [13]. Organising [1303, 39]. Organisms [1208, 1218, 1251, 1213]. Organization [1262, 1218, 1265]. Organizing [270, 1306, 1315, 1353, 1375, 957, 1203]. Orgy [1258]. Orientation [1343]. Orientations [138]. Oriented [142, 1344, 1702, 1701, 1707, 2022, 2034, 230, 2026, 111, 34, 2028, 525, 1519, 817, 2037, 2038, 1597, 1036, 551, 534, 1140, 186, 1353, 217, 202, 37, 807, 40, 25, 1694]. Origin [1197]. Orthogonal [1383, 132, 1382, 131]. Orthopaedics [456]. Oscillatory [1271, 1297, 1293]. ossifrage [855]. Other [616, 1707]. Outerplanar [161]. Outline [1]. Output [402, 840]. Overhead [90]. Overlapping [979]. Overlay [444]. Overload [1386]. Overloading [1128]. Overview [87, 639, 84, 454, 519, 978]. Oz [600].

P [1740, 1535, 101, 545, 2060]. P-Algebra [545]. P-Levelability [1740]. P/T [1535]. P0N0P [329]. PAC [438]. Package [1007, 924, 913]. Packet [300]. Packing [1051]. Packlib [1348]. Page [619]. Paging [619]. PaiLisp [548]. Pain [1494]. Paintbrush [451]. Pairs [1732, 402, 1610, 1186]. Pairwise [1616]. PAM [1018]. PAM-CRASH [1018]. Pancreatic [1271]. Pandore [919]. Paper [822, 214]. Paperfolding [301]. PAPERS [64]. Paradigm [1298, 415, 1000]. paradox [368]. Paragon [982, 907]. Parallel [570, 571, 569, 1257, 1357, 980, 1730, 2057, 1116, 748, 1542, 67, 1963, 903, 322, 1012, 1674, 1949, 1933, 82, 556, 265, 93, 559, 560, 1584, 548, 1696, 1046, 551, 83, 92, 566, 669, 1360, 1017, 804, 86, 984, 194, 3, 336, 1944, 1044, 81, 994, 1011, 1927, 1948, 678, 1937, 1942, 1941, 1364, 915, 557, 1362, 72, 998, 1015, 988, 96, 934, 565, 1877, 976, 434, 921, 919, 1003, 894, 985, 995, 910, 912, 895, 908, 929, 956, 996, 1002, 899, 924, 1013, 923, 893, 1019, 972, 965, 939, 928, 957, 955, 925]. parallel [981, 951, 970, 932, 918, 1010, 997, 1000, 1006, 926, 1646, 938, 983, 922, 989, 948, 947, 1989]. parallel-DM [1989]. Parallel-Sequential [1362]. Parallelisation [1014, 959, 916, 896, 913]. Parallelism [70, 1700, 1007, 564, 63, 76, 90, 1004]. Parallelity [401]. Parallelization [973, 116, 68, 967, 1878, 1872, 1016, 930, 917]. Parallelizing [71, 1008, 90]. Parameterized [1916, 536, 1553]. Parameters [263, 406, 1310, 1286, 1345]. Parametric [1836, 1574, 1683, 618, 353]. PARCS [955]. PaReDuX [748]. Parent [1258]. PARIX [1006]. ParMod [967]. ParMod-C [967]. Parsers [2041, 1125]. Parsing [156, 1627]. Part [1442, 924]. Partial [315, 1167, 524, 1824, 1545, 1594, 1917, 1714, 1181, 1925, 1160, 1926, 731, 1929, 105, 637]. Partial-Ordering [637]. Partially [1801, 1305]. Particularly [1440]. partition [897, 405]. Partitioning [1034, 616, 73, 633, 1940, 1939, 1942, 950]. Partitions [2011]. Partners [1224]. parts [1886]. Pass [1238]. Passing [1764, 1939, 1018]. Past [242, 179]. Patches [502]. Patching [667]. Path [411, 1728, 589, 292, 1675, 382, 492, 401, 371, 984]. Path-Consistency [589]. Path-width [382]. Pathology [1492]. Pathophysiological [1480]. Paths [410, 1716, 1104, 1610, 1035]. Pathway [1262]. Patient [1473, 1485, 1478]. Pattern [1369, 1380, 615, 291, 966, 1606, 1333, 1608, 1318, 1615, 1272, 1510, 1618, 1619, 1280, 1377, 1273, 1819, 441, 429, 439].

Pattern-Matching [1615]. Patterns [88, 705, 1262, 1620, 1809, 2043, 1900]. payment [1067]. Payoff [671]. PCF [929]. PDE [995]. Peak [1200]. Pedagogic [891]. 2 [558]. 860 [1949]. Analysis [1675]. CEL-V [1009]. Export [788]. Guarantee [773]. IBM [1946]. IFS [917]. IP [905]. NASTRAN [1017]. OR [1869]. Output [1457]. Schaltenbrand-Wahren [473]. Signer [1063]. T [1535]. unstructured [1014]. Pentium [762]. people [976]. Peptide [1626]. Perception [1370, 1321, 264]. Perceptrons [1392, 1377, 1083]. Perfect [1087, 1953, 1088, 961, 1952]. Performance [987, 85, 1549, 909, 1231, 807, 2013, 557, 961, 1243, 968, 638, 964, 912, 946, 986, 972, 953, 935, 981, 902, 914, 984, 911, 982, 1015, 1632, 971]. Performance-Oriented [807]. Performances [661, 1064]. Periodic [1272]. Periodic-Pattern-Selective [1272]. Periods [622, 1078]. Peripheral [1489]. PERMAS [1016]. Permutation [297]. Permutations [402]. Persistent [1278, 708]. Person [954]. Personal [218, 190]. Personalised [1473]. Perspective [706, 237, 1809]. Perspectives [1682, 1276]. Petri [1536, 1538, 1539, 1541, 996, 1578, 1543, 1544, 1679, 779, 1547, 1549, 1550, 1552, 778, 1553, 1555, 1556, 1557, 1558, 1533]. petroleum [934]. Pharmacokinetic [1517]. Phase [107, 105]. Phenomena [1337, 937]. Phenotype [275]. Philosophical [1765]. Phoneme [1390]. Photocoagulation [454]. Phylogenetic [1609]. Physical [1336]. physics [917]. Pi [1575]. Pi-Calculus [1575]. PIAF [990]. Piano [661]. Picture [1820]. Piecewise [325]. Piecewise-Smooth [325]. piling [1060]. piling-up [1060]. Pipelined [2036, 334]. Pipelines [333]. Pipes [1373]. Pixel [1818]. Pixel-based [1818]. PL [1905]. PLACA [38]. placement [967]. Placing [1727]. plaintext [1974]. Planar [398, 170, 618, 2003, 132, 153, 136, 397, 139]. Planarity [399, 160, 161]. Plane [137, 134, 163, 1750]. Planes [1917]. Planning [1529, 479, 643, 1103, 1487, 457, 447]. Plans [686, 1688]. Plant [1537]. Plasticity [1274]. Platform [224, 146, 922]. Platforms [2043, 985, 1019]. PLATO [812]. Plotkin [613]. PM [694, 690]. PMC [797]. PN [1538, 841]. PN-sequences [841]. Pocket [432]. Point [1659, 743, 2056, 1144, 1484, 151]. Point-Halfspace [151]. Pointers [77]. Points [324, 1064]. Poisson [994]. Polaris [71]. Policies [1940]. Poly [294, 1467, 1721]. Poly-Log [1721]. Poly-space [1467]. Polyflow [1013]. Polygonal [2013]. Polygons [602]. Polyhedra [1054]. Polyhedral [1030, 139, 1036]. Polymorphic [2026]. Polymorphism [1683, 1128, 353, 360]. Polynomial [1740, 1901, 1903, 780, 1024, 751, 1911, 1612, 278, 97, 1666, 1923, 1737, 1900, 731, 775]. Polynomial-Size [278, 1666]. Polynomial-Time [1612, 97]. Polynomials [624, 1916, 1932]. PolyTOIL [2026]. Polytope [1026, 1025]. Polytopes [1036]. Polytrees [1778]. Pomsets [109]. Population [509, 266, 267, 1200]. Population-Based [509]. Portable [1945, 919, 1018, 922, 948]. Porting [963, 1020, 939, 1010]. Posed [1308]. Posets [2054]. Position [445]. Positional [1043]. Positioning [173, 933]. Positive [1238, 122, 429]. Possibilistic [1833, 1794, 1835, 1791, 1792, 639, 1795]. Possibilities [1832]. Possibility [1793]. Possible [331, 249, 1713]. Post [1713]. Poster [176]. Posteriori [1827]. Postulates [1843]. Potential [503]. Potentially [1990]. Power [1356, 1537, 110, 589, 1743, 651, 102, 783, 540, 418, 946, 849]. Powerdomain [2054]. Powerdomains [1460]. Powerlists [1584]. powers [1084]. Powerstructures [1460]. Practical [216, 173, 1143, 240, 1313, 1956, 1090]. Practice [687, 757, 758, 1610]. Practising [178]. Practitioners [1432]. Pragmatics [253, 1686]. PRAM [1739]. Preattentive [1373]. Prebiotic [1197]. Precise [799]. Precomputation [1987, 1986]. preconditioners [934]. Predicate [1804, 1445, 885, 10, 879]. Predicates [645, 591]. Predict [1528]. Predicting [1253, 416]. Prediction [1863, 417]. Preference [1805, 1870]. Prefix [396, 1729]. Prefix-Free [1729]. Pregnancy [1528]. Preliminary [1527, 754, 917, 1429, 911]. PREPARE [969]. Preprocessing [1105, 1092]. Presence [77, 1222, 379]. Present [179]. Presentation [172]. Presentations [532, 582]. Preserving [1116, 1469]. Pretty [1127]. Pretty-printing [1127]. prevent [847]. Priming [1334]. Primitive [1713]. Principle [1859, 1251]. Principled [335]. Principles [1701, 216, 182, 1260]. printing [1127]. Prioritized [1841, 1859]. Priority [1175, 402]. Prisoner [1224]. privacy [1975]. Private [1722]. Probabilistic [572, 571, 614, 1842, 639, 1783, 1849, 1483, 1282, 116, 627, 326, 1861, 1766, 486, 1644, 1972]. Probabilities [1836, 277, 1785, 1865, 1769, 1871, 430]. Probability [1782, 1783, 1855, 1326, 1780, 412]. Problem [1034, 766, 609, 197, 398, 1033, 1302, 284, 1674, 285, 1043, 139, 1308, 311, 574, 1713, 628, 575, 640, 672, 1227, 1052, 1408, 1032, 1028, 1341, 634, 586, 1624, 174, 1798, 1250, 405, 1992, 852, 1966, 997, 1083, 950]. Problems [1023, 1027, 1747, 249, 1026, 1144, 756, 292, 1113, 1029, 618, 621, 623, 732, 409, 1446, 1370, 1898, 1292, 139, 407, 387, 1448, 1749, 631, 305, 7, 1621, 1511, 3, 1110, 1106, 755, 1931, 1629, 999, 1336, 893, 944, 936, 1015].

Procedure [1156, 1376]. procedures [1895]. Process [797, 209, 187, 1399, 706, 685, 1253, 697, 699, 1144, 1137, 696, 1568, 687, 700, 1649, 210, 1813, 702, 710, 199, 545, 218, 191, 715, 190, 1135, 1136, 711, 1404, 1211, 36, 305, 274, 695, 189, 704, 693, 692, 552, 1947, 714, 684, 689, 554, 691, 698, 708, 709, 1644, 713]. Process-Based [685]. Process-Centered [708]. Processes [769, 1763, 1674, 1823, 712, 220, 129, 715, 770, 2048, 553, 1558, 1877, 1882]. Processing [1388, 1295, 1107, 969, 1348, 99, 255, 1385, 488, 1374, 1944, 1354, 975, 903]. Processor [1935, 795, 1940, 1352]. Processors [2034, 2036]. procrastination [418]. ProCSuS [554]. Producing [1282]. Product [2025, 1146, 1971, 1737, 1879, 1970, 693]. Production [1828, 1949, 1014, 943]. Productivity [179]. Products [391, 99, 1924, 1147]. Profession [181]. Professionals [209]. Profoundly [1321]. Prognostic [1525]. Program [569, 1816, 747, 1168, 800, 204, 1875, 1158, 1886, 556, 128, 655, 801, 790, 1154, 355]. Programmable [4, 1544]. Programmed [1096, 393]. Programming [2045, 587, 521, 1027, 1184, 242, 34, 1162, 1519, 661, 1183, 662, 1161, 1592, 928, 1451, 244, 1175, 1729, 1129, 2059, 1041, 1594, 1598, 1128, 788, 1132, 550, 566, 1164, 2033, 217, 1131, 596, 635, 812, 567, 203, 2039, 557, 38, 561, 1456, 558, 121, 1155, 1177, 894, 1000, 1631, 412, 1126]. Programs [729, 570, 571, 1563, 1178, 1116, 110, 67, 750, 1580, 1163, 2028, 647, 1806, 1029, 82, 1577, 1171, 1752, 1166, 1030, 1657, 1186, 546, 243, 83, 92, 1666, 1152, 1922, 245, 1176, 672, 654, 1153, 1188, 81, 1160, 1185, 408, 1182, 1130, 1872, 1881, 925, 981, 1890, 1891, 949]. Progress [1205, 757, 436]. Progression [1528]. Project [1137, 686, 978, 1217, 901, 1003, 1004]. Projection [2018]. Projects [203, 1500].

Proliferative [1514]. Prolog [1182]. promises [1011]. Promising [819]. Proof [2058, 570, 1583, 347, 887, 796, 341, 878, 336, 891, 1654, 2051, 1882, 1888, 371]. Proofs [1441, 1595, 1583, 1963, 751, 734, 667, 1063, 721, 1443, 877, 718, 1177, 358, 1092, 849, 1891, 1641, 374]. Propagation [1772, 1779, 1483, 1391, 799, 1365, 1316, 893, 927]. Properties [349, 1168, 1680, 1111, 796, 1651, 1917, 2044, 536, 312, 1037, 1796, 1644, 866, 369, 370]. Property [1657]. Proposed [2061]. Propositional [293, 1464, 1806, 649, 890, 885, 1862, 1443, 1182, 1284, 876]. Propositions [1672]. Prospects [757, 436]. Protein [1618, 1863]. Proteins [1253]. Protocol [1676, 2042, 300, 1471, 107, 892, 1632, 1978]. Protocol-Directed [1471]. Protocols [2041, 1070, 2032, 1980, 1637, 848, 845]. Prototype [1501, 817, 682]. Prototypes [1318]. Prototypicality [644]. Prototyping [800, 87, 236, 753, 227]. provably [1067]. Prover [752, 1063, 1859, 1860, 892, 810]. Prover/Signer [1063]. Provers [1857]. Providing [1422, 1061]. Proving [516, 870, 1567, 1802, 1909, 343, 1652, 737, 812, 801, 724]. Proximity [165, 164, 1234]. Pruning [887, 663, 591, 664, 1292, 1692, 652]. Pseudo [588]. Pseudo-Boolean [588]. Pseudorandom [280]. PSL [2042]. Psychiatry [1505]. Psychological [1337, 1429]. PTIME [104, 1450]. PTTP [883]. PTTP-Implementation [883]. Public [1720, 863, 1995, 862]. public-key [863, 1995, 862]. Publication [820]. Pulser [339]. pulses [938]. Purchaser [1144]. Pure [1029, 1459, 352]. purpose [896]. Purposes [265]. Pursuit [1750]. Push [1033]. Push-Relabel [1033]. Pushdown [626, 1752]. Puzzling [657]. PVM [975, 977, 960, 899, 974, 949]. PVMe [960]. PVS [1585, 346].

Q [1999, 1218]. Q-Deformed [1999]. Q-Learning [1218]. Quad [2003]. Quadratic [1027, 502]. Quadrics [989, 911]. Qualified [1828]. Qualitative [1479, 1480, 1487, 1323, 1781, 1863, 1522, 1780]. Quality [1145, 1134, 1138, 1144, 1142, 1410, 1140, 1430, 1496, 1516, 1147, 200, 214, 1500]. Quantification [1672]. Quantifying [1148]. Quantitative [348, 1135, 500, 1781]. Quantization [1522]. Quantum [1068, 1999, 963, 973]. quantum-chemistry [963]. Quasi [1831, 1377, 1228, 1001]. Quasi-Classical [1831]. Quasi-Optimum [1377]. Quasi-Uniform [1228, 1001]. Quasigroup [755]. Queen [1205]. Queries [1728, 123, 101, 97, 125, 120, 2015, 108, 419, 438, 435]. Query [126, 517, 1692, 98, 1688, 102, 442]. Querying [100]. Quest [97]. Questions [177]. Queues [402]. Queuing [1943]. quotient [365].

R [353]. Rabin [1667, 775]. Rachmaninoff [661]. RADD [1428]. Radial [1361, 1328, 1360, 1368, 1313]. Radical [1903]. Radiological [1526]. Radiosurgery [458, 457]. Radiotherapic [1529]. Radiotherapy [479]. Radius [1898, 1921]. radix [863]. Raid [833]. Raise [55]. Ramified [1467, 1969]. Rampart [1636]. Random [603, 1738, 1722, 1087, 114, 284, 115, 1448, 462, 312, 1032, 1284, 1088]. Randomized [1023, 1723, 298]. Randomness [1917, 1954]. Range [1271, 1372]. Ranges [279]. Ranked [1326]. Ranking [2006]. Rankings [403]. Rapid [1428]. Rat [474]. rate [1989]. Rates [1206]. Ratio [485]. Rational [759, 1922, 1179, 763, 776, 431]. Ray [499, 490]. RBF [1330, 1299]. Re [1472]. Re-Usable [1472]. Reachability [1545, 129, 1548, 1553]. reacting [948]. Reaction [497]. Reaction-Diffusion [497]. Reactive [1661, 802, 670, 32, 1239, 973].

Readable [251]. Real [1753, 808, 1383, 1290, 1678, 662, 1677, 193, 1292, 1579, 1191, 1239, 2040, 1664, 444, 637, 992, 944, 1005]. Real-Time [808, 1678, 1677, 193, 1579, 2040, 1664, 444, 637, 992, 1005]. Real-World [1239, 944]. Realistic [1297, 507]. Reality [443, 1241, 460, 504, 514, 478]. Realizability [617, 1668]. realization [354]. Realized [2054]. Realizing [283]. really [1091]. Realm [2014]. Realm-Based [2014]. Reals [599]. Realtime [504]. Reasoning [333, 769, 571, 1830, 791, 1831, 1805, 1802, 1477, 1180, 1839, 19, 1173, 1576, 1652, 1409, 1846, 350, 648, 639, 1544, 1509, 339, 1584, 1452, 584, 1858, 1864, 2016, 1866, 1856, 1871, 1697]. Rebuilding [1648]. Receipt [1090]. Receipt-free [1090]. Receiver [1115]. Receiver-Initiated [1115]. Receiving [1245]. Receptive [1281, 1262, 1265]. RECIT [1474]. Recognising [1685]. Recognition [1311, 502, 1380, 1378, 1382, 1510, 1377, 1379, 824, 151, 1324, 1819, 966]. Recognizable [1571, 611, 1551]. Recognizing [166, 1381]. Recombination [1256, 273]. Recombinative [669]. reconciliation [1975]. Reconfigurable [1364]. Reconfiguration [947]. Reconstructing [495, 490]. Reconstruction [2058, 491, 505, 474, 492, 501]. Record [1473]. Recovery [1813, 852, 1966]. Rectangle [143, 166]. Rectangle-Visibility [143]. rectangles [434]. Rectilinear [160]. Recurrence [614, 1467, 930]. Recurrent [1339, 1287]. Recursion [1435, 666, 801, 103, 1885]. Recursion-Based [801]. Recursive [1759, 645, 750, 623, 1294, 1468, 122, 421]. Recursively [752, 564]. Recycling [1087, 1088]. Red [1205]. Redrawing [152]. Reduce [1108, 1332]. Reduced [347, 973]. Reducibility [723]. Reducing [89, 936]. Reduction [2058, 580, 1333, 93, 149, 885, 737, 722, 754, 1867, 1058, 904].

Redundancy [2007, 298]. Redundant [1110]. ReDux [748]. Reengineering [213]. reference [377]. Refined [1531]. Refinement [788, 802, 534, 342, 1879, 1894]. refinements [369]. Reflections [1495]. Reflex [1298]. Reflexive [1807, 714]. Refusal [1224]. Refutable [1151]. Refutation [876, 721]. Regenerative [1557, 1275]. Regenerative-Type [1275]. Region [1384]. Regions [2019]. Register [1971, 1970]. registers [1972, 1077, 1969]. Registration [472, 464, 443, 466, 448, 465, 474, 445, 444]. Regular [315, 803, 1680, 745, 1339, 138, 136, 312, 774, 775, 441]. Regularity [1466]. Rehabilitation [1511]. Rehearsal [507]. Reinforcement [669, 671]. Reinforcing [182]. Relabel [1033]. Related [1768, 406, 139, 407, 1621, 151, 1629, 916]. Relating [733, 265]. Relation [607, 438, 1894]. Relational [2022, 1828, 658, 2050, 1775]. Relations [1659, 806, 1178, 1837, 614, 1844, 313, 281, 1452, 575, 1742, 2019]. Relationship [1429, 1411]. Relationships [2018, 1084, 1985]. Relative [1908, 366]. Relaxations [1030]. Release [1044]. Relevance [1338]. Relevant [691]. Reliability [1794, 1839]. Remotely [553]. Renal [1517]. Rendering [492]. Reordering [69]. Reorganization [22]. repair [1978]. Replacement [1530]. Replete [2053]. Replicated [336]. Replicator [1198]. Replicators [1201]. Report [1148, 1802, 141, 190, 1493, 1431, 1475, 176]. Representable [165, 145]. Representation [253, 1240, 1804, 487, 2003, 1102, 1786, 488, 2001, 1396, 2018, 1386, 2015, 1475, 1869, 1421, 1775, 1774, 1439]. Representations [293, 143, 144, 1212, 1472, 1625, 1766, 653]. Represented [1725]. Representing [1519, 1585, 35]. Reproducibility [1523]. Reproducing [1229]. Reproduction [1258, 1230]. required [1954]. Requirement [1648, 1688]. Requirements [196, 1977, 1139, 1409, 195, 1424, 1410, 1122, 1947, 1354, 1948, 1980, 923]. Resampling [1507]. Research [819, 219, 204, 239, 691, 241, 914, 941, 958]. reservoir [915]. Reshaping [88]. Resilient [1079]. Resistance [1271]. Resolution [1156, 1453, 1859, 1926, 724, 1455]. Resolution-based [1156]. Resolvent [1931]. Resolvents [1908]. Resonance [469, 1394]. Resource [20, 62, 327, 1733, 1947, 905]. Resource-Bounded [327]. Resources [1727, 257, 1944, 1236, 863]. respect [417]. Respiratory [1273]. Responsibilities [178]. Restoring [335]. Restricted [740, 408]. restriction [1697]. Restrictive [1076]. Restructuring [680]. Results [1329, 1042, 1161, 663, 1279, 407, 85, 1582, 1621, 744, 754, 408, 1224, 755, 122, 1699, 920, 917, 983, 378]. Retinal [1268, 454]. Retinopathy [1528]. Retrieval [459, 1851, 1818, 1419]. Retrospective [510]. Reusability [1815, 1141]. Reusable [1425]. Reuse [1816, 1424, 1142, 667, 1812, 205, 1143]. Reversals [1809]. Reversible [616]. Review [1290, 226]. Revised [1802]. Revision [1840, 1171, 1862, 121, 1174]. Revisited [1225, 1219, 614, 528, 1693, 717, 681]. Rewrite [729, 1096, 752, 719, 735, 635, 730, 738, 741, 357]. Rewriting [727, 1097, 1098, 1094, 1095, 585, 580, 740, 1581, 733, 577, 579, 756, 528, 732, 581, 728, 576, 1452, 781, 574, 583, 716, 112, 812, 578, 755, 724, 393, 371]. Rich [106]. Right [1799, 730]. Right-Hand [730]. Rigid [482, 871]. Rings [516]. Ripping [1075]. Risk [685]. Rivest [1058]. RNA [1601, 270, 965, 1199, 1252, 1202]. Road [2024]. Robot [455, 1236, 1243]. Robotic [456, 458, 457]. Robotics [1241]. Robots [1192, 1242]. Robust [674]. ROI [1359].

Role [569, 644, 195, 193, 274, 1825, 1561]. Roles [686, 39]. Root [335]. ROSE [2014]. Rosser [302]. Rough [1507]. Routing [299, 1733, 300, 638]. RSA [1057, 834]. RSA-type [1057]. Rule [1114, 1502, 1670, 1519, 648, 1523, 674, 1285, 1796, 784, 357]. Rule-Based [1670, 1519, 1114]. Rule-Given [1796]. Rules [570, 1828, 592, 1801, 1166, 1417, 752, 2004, 882, 1172, 1216, 1469, 677, 1507, 1340, 122]. Running [318, 990]. Runtime [84, 919]. Russian [215]. RUTH [657].

S [2053, 686, 1984]. S-boxes [1984].

S-Replete [2053]. S. [2051]. S.A.M.C.E.F. [1019]. s02DES [857]. S4 [1469]. Safe [2026]. Safety [1657, 1651]. SAMMI [1019]. Sampling [1607, 511]. SAR [975]. Satisfaction [1113, 1104, 772, 1106]. Satisfiability [1842, 409]. Satisfying [923]. Saturated [1307]. SBTA [625]. Scalability [970, 654]. Scalable [1935, 1696, 985, 955]. Scalar [74]. Scale [470, 1007, 1374, 2011]. Scaling [1309]. Scanning [1325]. Scans [290]. scattering [973]. Scheduler [1677, 1941]. Scheduling [1934, 1821, 1935, 1042, 62, 1933, 1823, 1945, 1043, 1936, 1943, 1733, 1946, 63, 1938, 1044, 1947, 1688, 549, 898]. Schema [1314]. Schemata [1427]. Scheme [3, 853, 859, 361, 1083, 1892, 1090, 868, 428, 1089]. Schemes [1073, 1965, 1953, 299, 1954, 836, 852, 1057, 839, 860, 1967, 1966, 837, 1995, 1952]. Schizophrenic [2049]. Scholarly [820]. Schrödinger [994]. Science [219, 1897, 184]. Scientific [14, 12, 16, 13, 87, 1949, 832, 987, 935]. Scientific-Technical [832]. Scoliosis [514]. Scoring [1616]. Screening [461]. Screens [158]. SDF [815]. SDL [49]. SEAMLESS [1596]. Search [1398, 887, 406, 623, 1693, 1619, 274, 1419, 1106]. Searching [1725, 1747, 830]. Second [1650, 1339, 1444, 1091]. Second-Order [1650, 1339, 1444]. Secondary [1199]. Secondary-Structure [1199]. Secret [1073, 610, 1076, 1953, 1071, 1954, 1072, 839, 837, 1952]. Secret-Key [610, 1076]. secrets [838]. sections [973]. Secure [854, 1070, 1071, 1973, 848, 1976, 1069, 1067]. Securing [1069]. Security [860, 835, 1962, 850, 851]. sedimentation [991]. Seeded [2020]. Seeing [1232]. Segment [1450]. Segmentation [1372, 322, 496, 453, 469, 1375, 1384, 1373, 1503, 508, 497, 1371, 486, 449]. Segmented [465]. SEI [215]. Select [1493]. Selection [1820, 1299, 1259, 813, 274, 273, 203]. Selective [1272]. Self [270, 1534, 1754, 1219, 1306, 1303, 1315, 1281, 556, 1915, 1230, 1353, 1968, 1375, 2039, 1265, 1229, 1203, 957]. Self-Consistent [1281]. Self-Controlling [1534]. Self-Correcting [1754]. Self-Describing [556]. Self-Dual [1915]. Self-Inspection [1230]. Self-Organising [1303]. Self-Organization [1265]. Self-Organizing [270, 1306, 1315, 1353, 1203, 957]. Self-Reproducing [1229]. Self-shrinking [1968]. Selforganizing [1305]. Semantic [1704, 1703, 1574, 1670, 1111, 872, 873, 1795, 1195]. Semantically [106]. Semantics [1151, 1535, 1169, 521, 1538, 296, 1834, 740, 1167, 1157, 111, 1119, 1100, 1158, 1457, 1173, 1159, 246, 2061, 530, 817, 788, 1170, 1153, 1188, 120, 1870, 1572, 1189, 363, 367, 376]. Semi [1383, 1440, 1479, 843, 781, 1499, 731, 776, 918]. Semi-Automatic [1499]. Semi-bent [843]. Semi-Completeness [781]. Semi-cover [1383]. semi-implicit [918]. Semi-Qualitative [1479]. Semi-Thue [731]. Semi-Trace [776]. Semi-Unification [1440]. Semidefinite [1030]. Semigroups [1745]. Sender [1115]. Sender-Initiated [1115]. SENN [1500]. Sensing [1746]. Sensitive [302, 612]. Sensitivity [1343, 1310, 1852, 1268]. Sensor [445]. Sensory [1222, 1237].

Sensory-Motor [1222, 1237]. Separated [536]. Separating [1024]. Separation [1026, 1376]. Separators [286, 1718]. Sequence [1007, 1056, 1610, 1623, 1626, 840]. Sequences [482, 301, 1741, 475, 1971, 630, 8, 1970, 1078, 841]. Sequent [1437, 877]. Sequential [1508, 526, 1607, 341, 277, 1645, 544, 1362]. Series [1311, 1046, 401, 1925, 681, 682, 1002]. Series-Parallel [1046]. Series-Parallelity [401]. Server [1063, 854, 955]. Service [1140, 1430, 1938, 1640]. Services [833, 1636]. Session [713, 701, 683]. Set [1034, 1789, 328, 347, 398, 1716, 1026, 294, 1333, 1912, 2050, 474, 759, 1452, 1617, 1187, 125, 1470, 1507, 1194]. Set-Relations [1452]. Set-Theoretic [294]. SETH [1490]. SETHEO [892]. Sets [1539, 611, 328, 745, 1022, 82, 1120, 78, 323, 2020, 813, 1714, 104, 723, 490]. Setting [818]. Several [624]. Sewage [1373]. Sexual [274]. Seymour [1049]. Shafer [1768, 1868]. shallow [1001]. Shape [509, 503, 500]. Shapes [1388, 458]. Sharable [1513]. Share [1951]. Shared [1223, 298, 91, 86, 1637, 977, 849]. Shared-Memory [91, 86, 977]. shareholders [1967]. shares [837]. Sharing [1073, 805, 89, 1953, 1062, 1071, 1936, 2044, 861, 1496, 547, 351, 1954, 836, 1072, 838, 951, 1634, 839, 837, 1952]. shelf [959]. Shell [1495]. Shift [1971, 1970, 1972, 1077]. Shift-Register [1971, 1970]. SHNN [1362]. Shock [497]. Shock-Based [497]. Shooting [1726]. Short [1073, 1719, 1615, 1721, 938]. Shortest [1728, 292, 1610, 440]. Shortest-Paths [1610]. Should [1267]. SHPC [897]. shrinking [1968]. Shuffle [608, 777, 1710]. Shunting [1355]. Shuttle [689]. Sides [730]. sieve [1065]. Sigmoid [1286]. Signable [1036]. Signal [46, 797, 808]. Signaling [1245]. Signals [1388]. Signature [1965, 1062, 1958, 853, 852, 1966, 862, 868]. Signatures [1959, 1957, 1991, 1063, 1074, 854, 1992, 1061]. Significance [1310]. SIM [1347]. Similarities [338]. Similarity [1601, 1844]. Similarity-Based [1844]. Simple [1388, 1440, 1905, 1677, 796, 1282, 158, 737, 1269, 1734, 167, 364, 1884, 365, 429, 1091, 85]. Simplicity [280, 9]. Simplification [494, 1259, 368]. Simplified [871]. Simplifying [663]. simply [378]. simply-typed [378]. Simpson [1784]. Simulate [616, 1227]. Simulated [512]. Simulating [546, 24, 938]. Simulation [1479, 1257, 479, 61, 477, 297, 304, 1480, 1241, 1739, 1226, 1554, 948, 1350, 1349, 1018, 913]. Simulations [1205, 298, 2048, 636, 940, 928, 915]. Simulator [1344, 454, 1345]. Simultaneous [307, 1379]. Single [201, 1979, 339, 1200, 962]. single-board [962]. Single-Skilled [201]. Single-Term [1979]. Singleton [1433]. Singly [1915]. Singly-Even [1915]. singular [1057]. Situated [1231, 39]. Situation [1176]. Situational [706]. Size [1951, 1343, 278, 1666, 837]. Sized [1130]. Sizes [313, 297]. Skeptical [1179]. Skilled [201]. Skills [209, 204, 640, 39]. Skills-Driven [209]. SL2 [859]. Slender [314]. Slices [492]. Slicing [790]. Small [279, 1728, 291, 158, 1724, 283, 1628]. Smaller [1607, 1625]. Smalltalk [2045]. Smalltalk-80 [2045]. Smart [556, 1995]. SMoLCS [814]. Smooth [325]. Snakes [453]. Social [760, 1223]. Societies [24]. Society [1243]. Soft [1301, 272]. Softpar [1004]. Software [1529, 185, 1816, 1146, 1815, 181, 222, 1148, 187, 1150, 1145, 685, 1139, 1134, 197, 204, 1138, 216, 1, 1348, 224, 699, 1144, 1137, 225, 206, 1585, 757, 177, 758, 210, 183, 760, 180, 710, 228, 184, 199, 193, 218, 238, 1657, 261, 1142, 190, 792, 1135, 1136, 1432, 1485, 215, 211, 192, 1817, 1431, 194, 1430, 202, 692, 239, 198, 714, 2043, 207, 1147, 689, 229, 691, 698, 240, 1560, 241, 178, 1874, 1069, 951, 1893]. Solids [512]. Solution [742, 1912, 1266, 792, 936, 634, 867, 996, 893, 994, 1090, 915, 998]. Solutions [1713, 1929]. Solvable [766]. Solve [670, 1015]. Solved [1911]. Solvers [599, 968, 995]. Solving [743, 65, 1759, 588, 249, 589, 593, 672]. Some [818, 580, 866, 874, 177, 621, 1279, 1370, 1693, 576, 716, 305, 1288, 744, 754, 1273, 1429, 691, 433, 122, 1780, 1962, 442]. SOPHOCLE [454]. Sort [532, 318]. Sorted [113, 1566, 532, 1582]. Sorting [319]. Sorts [113]. SOS [554]. Sound [1457, 1801]. Source [1376, 790]. Source-Level [790]. Sources [320]. Soviet [1993]. SP [1946]. SP1 [985]. SP1-2 [985]. SP2 [960]. Space [126, 604, 1115, 1977, 1542, 1543, 4, 464, 291, 1607, 650, 1610, 1936, 2044, 326, 2060, 678, 450, 689, 434, 1467]. space-bounded [434]. Spaces [1758, 759, 1037, 656, 681]. Spanning [391, 618, 1051, 621]. Sparse [993, 395, 88, 328, 490, 982]. Spatial [2021, 1343, 2005, 2007, 2014, 2009, 2006, 2004, 2001, 1244, 2018, 95, 2010, 2011, 2016, 2002, 1689]. Spatiotemporal [481]. Special [2037, 2038]. Specialist [253]. Specialization [645, 1223]. Specializer [800]. Species [201]. Specific [2029, 1000]. Specification [1537, 805, 800, 562, 21, 526, 529, 1171, 195, 712, 530, 1588, 533, 798, 256, 802, 2042, 534, 535, 1868, 392, 393, 519, 1880]. Specifications [791, 1574, 525, 1566, 773, 792, 813, 1176, 882, 1569, 1461, 807, 1877, 1881]. Specificity [1848]. Specify [806]. Specifying [2012, 405]. SPECT [445]. Spectra [1445]. spectral [950]. SPECTRE [657]. Spectrum [51, 850]. Speculative [804]. Specware [1893]. Speech [1820, 1385, 1321]. Speed [1387, 1107, 1286, 340, 298].

Speed-independence [340]. Speeds [1044]. Speedup [1730, 1646, 907]. Speedups [884]. Sphere [1173]. Spherical [1907, 1001]. SPICE [1137]. Spike [1354, 1595]. Spike-Processing [1354]. Spinal [503]. Spiral [602]. Splitn [1761]. Splitn-Language [1761]. Splitting [1052]. SPMD [83, 907]. spread [850]. Spring [167]. SQL [1696]. Square [1717, 335]. squeamish [855]. SS [1029]. Stable [1167, 111, 1026, 1186, 1160, 120, 1154]. StackThreads [549]. Stage [1866]. Stages [758]. Staging [214]. Standard [1137, 2009, 583, 1621, 1427, 877, 1766, 869, 377, 1958]. Standard-Sequent [877]. Standardizing [257]. Standards [1681]. Stanford [810]. STAR [1011]. Starting [1484]. State [1669, 1542, 1317, 1543, 1712, 1652, 1846, 625, 1651, 1666, 1341, 1625]. Statecharts [47, 786]. Statement [78]. States [1807, 1510]. Static [1727, 1158, 2028, 173, 795]. Statistical [1318, 1785, 675, 2011, 983, 842, 438]. Statisticians [1320]. Step [117, 662, 1692, 810]. Steps [1585]. Stepwise [712]. Stereo [446, 515]. Stereotactic [448, 471]. Stimulus [1343]. Stochastic [1225, 1660, 1541, 1290, 1850, 1800, 1297, 1549, 1550, 1721, 425, 1557, 996, 433]. Stock [566, 549]. Stokes [1015]. Stopping [1329]. Storage [2008, 298]. STORE [1381]. STORM [749]. StormCast [1630]. Straight [1922]. Straight-Line [1922]. Strange [1811]. Strategies [2021, 1425, 1208, 1750, 1486, 1255, 276, 565, 973, 1064, 915]. Stratification [1784]. Stratified [1830, 110, 1180, 1465, 101]. Stream [1556]. strength [1984]. strengthen [865]. Strengthening [1055]. Striate [1274]. Strict [374]. Strictness [105, 789]. String [1603, 1607, 581, 1622, 1627, 849]. Stringology [1621]. Strings [1601, 1339, 1716, 1615, 1617]. Strong [308, 404, 782, 2048, 844]. Strongly [1053]. Structural [333, 1514, 1462, 288, 103, 338]. Structure [1034, 1039, 385, 1253, 1836, 2003, 1199, 2009, 1437, 545, 1618, 326, 1192, 567, 1242, 965, 423, 851, 949]. Structured [1497, 823, 913, 1243, 1500, 1014]. structured/unstructured [1014]. Structures [604, 2052, 502, 489, 1444, 1270, 1768, 459, 404, 861, 2047, 1269, 786, 844, 1265, 1220, 1421, 542, 1708]. Stubborn [1539]. Student [203]. Students [204]. Studies [339, 690]. Study [1778, 1098, 498, 1027, 1301, 19, 1588, 798, 335, 1237, 1429, 937, 927]. Studying [267, 755]. Sturmian [1736]. Style [1652, 1437]. Sub [104]. Sub-PTIME [104]. Subclass [1680]. Subclasses [1551, 442]. Subdivisions [2003, 678]. Subgraph [404, 139]. Subgraphs [391, 1095]. Subjective [1222]. Subnetworks [1300]. Suboptimal [1313]. Subrecursion [1456]. Subsequence [1624]. Subsequences [1605]. Substitution [745, 726, 1467]. Substitutions [739, 372]. Substructural [1833]. Subsumption [119]. Subtraction [480]. Subtrees [310]. Subtypes [520]. Subtyping [2031, 1433, 1434, 1436, 355]. Subword [285]. Successful [1401, 238]. Successive [1678]. Succinctly [1725]. Suffix [1604, 1717, 1614]. Sum [1610]. Sum-of-Pairs [1610]. Summarising [894]. summation [1879]. Super [781, 958]. Super-Hierarchical [781]. supercomputer [926, 989, 962]. supercomputers [966, 970]. superpositions [427]. Superquadrics [483, 501]. Supersequences [1605]. Supervised [1306]. Supervisory [1669]. Support [1491, 1501, 1525, 350, 235, 1422, 2009, 2037, 2038, 1936, 1634, 1495, 1524, 1515, 1416, 709, 895, 1893, 1526]. Supporting [1097, 14, 833, 1407, 2015, 461, 1894].

Supports [1896]. SuRE [1598]. Surface [494, 511, 448, 2015, 508, 916]. surface-related [916]. Surfaces [502]. Surgery [1525, 443, 466, 471, 456, 514, 507]. Surveillance [1484]. Survey [208, 237, 582, 42, 254, 691, 845]. suspected [1967]. SVM [1008]. SVM-Fortran [1008]. Symbiosis [1209]. Symbol [1341]. Symbolic [1295, 1590, 1545, 1249, 57, 595, 2018, 1346]. Symbols [1381]. Symmetric [1912, 767, 997]. Symmetries [1545, 289]. Symmetrized [582]. Symmetry [1671, 503, 154]. Synapse [1366]. Synapsis [1368]. Synaptic [1264, 1278, 1297]. Synchronization [81]. Synchronized [1544, 345, 765]. Synchronous [341, 1673, 555]. synchrony [1638]. Syntactic [885]. Syntax [1837, 359]. Syntax-Based [1837]. Synthesis [1753, 780, 348, 56, 750, 1596, 344, 1326, 295, 672, 655, 276, 1668]. Synthesizing [1891]. Synthetic [1245]. System [811, 746, 1816, 747, 805, 2008, 1595, 1583, 1527, 750, 1804, 1490, 1489, 1585, 1378, 1675, 659, 1596, 1847, 155, 1102, 1786, 236, 443, 897, 1588, 1945, 1530, 1247, 1599, 802, 1946, 1499, 154, 471, 156, 84, 1277, 473, 37, 1474, 1600, 1528, 1524, 1929, 1147, 1513, 461, 1351, 554, 1507, 393, 1203, 1526, 831, 478, 1880, 1976, 1882, 1069, 1635, 1630, 933, 371, 1890, 1067, 994, 1631, 862, 353]. Systematic [1996]. Systems [1415, 742, 818, 26, 1661, 1753, 1701, 1098, 1094, 1707, 1096, 808, 1095, 1660, 1257, 1537, 301, 580, 1502, 1680, 706, 523, 34, 1117, 1240, 1518, 1414, 1678, 1652, 21, 528, 290, 35, 648, 2007, 1425, 1911, 1586, 1913, 22, 1509, 1745, 1249, 1665, 1417, 1655, 533, 1480, 1589, 339, 1298, 1498, 1693, 1943, 753, 1767, 781, 9, 574, 1478, 1197, 2042, 1321, 1579, 768, 770, 719, 295, 670, 575, 1492, 735, 537, 1407, 1744, 1239, 1570, 1495, 98, 1948, 1556, 731, 10, 821, 1664, 23, 1942, 730, 698, 1520, 738, 1533, 741, 39, 96, 1932]. systems [352, 1644, 912, 908, 956, 946, 899, 986, 1642, 923, 897, 850, 970, 370, 876, 1645, 915, 947]. Systolic [1361, 287, 1927].

T3D

[963, 993, 939, 909, 1009, 982, 943, 998, 988]. Tableau [879]. Tableaus [873]. Tableaux [884, 875, 872, 889, 885, 888]. Tables [651]. Tag [301]. Tagged [476, 450]. Take [1865]. Taking [1481]. Talairach [473]. Talairach-Tournoux [473]. Talairach-Tournoux/ Schaltenbrand-Wahren [473]. Tapes [1207]. TAPSOFT [757]. Task [41]. Tasks [1821, 670, 1496, 547]. Tatzelwurm [53]. Tau [1451]. Taxonomies [1472]. TCP [905]. TCP/IP [905]. teacher [415]. Teaching [196, 197, 216, 206, 191, 460, 217, 202, 207, 1560]. Team [203]. Teams [233]. Technical [13, 832]. Technique [1586, 1653, 1318, 267, 451, 112, 1627]. Techniques [743, 1802, 225, 64, 1519, 2005, 195, 1333, 1425, 716, 1494, 1868, 140, 755, 227, 790, 5, 921, 1699, 896, 979]. Technologies [1511]. Technology [700, 710, 711, 186, 673, 229, 1005]. Teeth [1602]. Telecardiology [1512]. Telecollaboration [452]. Telecommunication [148]. Telecommunications [1657, 798]. Teleology [1341]. Tell [1197]. Template [1203]. Template-Directed [1203]. Temporal [1311, 1821, 1660, 242, 244, 773, 247, 243, 1672, 810, 245, 1582, 1285, 771, 637, 1881]. Tense [1464]. Tentative [690]. Tera [1934]. Term [585, 590, 1581, 733, 1979, 528, 781, 754, 812, 724, 1439, 371]. Terminal [531, 1327]. terminate [372]. Termination [573, 1709, 528, 751, 734, 371, 719, 718, 720, 1889, 374]. Terminological [1804, 1162, 261, 262]. Terminology [253, 260, 258]. Terms [253, 1441, 1649]. Ternary [1918]. Tertiary [1253]. Test [1369, 1917, 1493, 813, 1788, 485, 723, 899]. Testbed [25]. Testing [1784, 1148, 1150, 1149, 399, 160, 761, 1787, 1762, 161, 1925, 1099, 731, 912]. Tests [505]. Text [1378, 1693, 827, 358]. Texts [1603]. Textual [825]. Texture [1383, 498, 1384]. Their [604, 396, 1100, 1605, 1777, 211, 882, 389, 415, 836, 281, 852, 1917, 439, 1785]. Theorem [516, 870, 1802, 1045, 1909, 343, 1652, 1857, 1859, 1860, 812, 892, 724, 2]. Theorem-Proving [1802]. Theorems [752, 433]. Theoretic [1151, 2057, 1670, 294, 1671, 1662, 1668, 1061, 413]. Theoretical [5]. Theories [1165, 1156, 522, 750, 1180, 1465, 1185, 1454]. Theorist [94]. Theorizing [642]. Theory [880, 1794, 1768, 1760, 1708, 757, 1710, 266, 1793, 758, 1755, 710, 1193, 1463, 734, 1214, 506, 2050, 1767, 1711, 1197, 1259, 1860, 754, 1899, 1394, 1226, 1947, 1470, 1867, 1868, 538, 680, 1507, 1698, 1875, 957]. Theory-Based [538]. Therapy [1487, 1471, 1486, 1488, 1524]. There [1260, 31]. Thermodynamics [1252]. Third [379]. Third-order [379]. THOR [1554]. threaded [548]. Threads [549]. Three [269, 1301, 127, 1307, 282, 474, 492, 1618, 1360, 1269]. Three-Dimensional [127, 492, 1618]. Three-Dimensions [474]. Threshold [283, 1967, 676, 836]. Threshold-multisignature [1967]. Thresholding [1262]. Thue [574, 575, 731]. Tiered [1691]. Tight [605, 664, 416]. Tightening [1029]. tightly [1646]. tightly-coupled [1646]. Tile [1204]. Time [1740, 1753, 808, 1095, 1383, 395, 1024, 1506, 1732, 615, 291, 1576, 1678, 404, 1674, 618, 1677, 1723, 1385, 1612, 193, 1936, 773, 1048, 1476, 97, 1579, 1721, 1215, 793, 1938, 544, 1865, 318, 2040, 1556, 394, 1664, 444, 1790, 637, 775, 682, 1002, 992, 951, 1005, 840, 693, 1506]. Time-Delay [1385]. TIME-NESIS [1506]. Time-Objects [1476]. Time-Optimal [395]. time-sharing [951]. time-varying [840]. Timed [1549, 295, 1558]. Times [1043, 936]. Timing [1590, 1679, 57]. Tissue [496, 927]. TKR [1530]. TKR-tool [1530]. TLT [48]. Today [13]. Together [657]. Tolerance [152]. Tolerant [234]. Tomograms [469, 461]. Tomography [1928]. Too [197, 761]. Tool [1428, 66, 1423, 1491, 1501, 1653, 1559, 460, 813, 514, 891, 1516, 812, 147, 809, 150, 896, 899, 923, 981, 937, 1530]. toolbox [930]. Toolkit [816, 1777, 972, 1636]. Tools [223, 1149, 221, 237, 227, 708, 815, 1964, 897]. ToolSet [814]. Top [1420]. Topics [1839]. Topography [462]. Topological [1039, 2001, 535, 538, 2019]. Topologically [1611, 453]. Topology [1333, 1863, 2017]. Toroidal [130]. Torus [1048]. TOSCANA [150]. Total [1530]. Totality [789]. Totally [1722, 1634, 1632]. Toupie [1656, 597]. Tour [656]. Tournaments [410]. Tournoux/ Schaltenbrand [473]. Towers [1923]. Toxicology [1490]. Trabecular [498]. Trace [580, 308, 1662, 1551, 776]. traceability [1069, 1967]. Traces [1578, 309, 1709]. Track [483]. Tracking [1205, 1822, 476]. trade [1958]. trade-offs [1958]. Trading [420]. Traffic [249]. Train [2035]. Trained [1265]. Trainer [1521]. Training [209, 187, 210, 208, 213, 186, 1561, 205, 478, 213]. Trans [1188]. Trans-Epistemic [1188]. Transaction [1638, 711, 1407, 106]. Transactions [525, 1691, 637]. Transcendental [1754]. Transducers [766, 1625]. Transductions [776]. transfer [935]. Transferring [205]. Transformation [1816, 747, 1426, 1158, 1653, 1182, 718]. Transformations [89, 1116, 110, 1712, 527, 69, 68, 786, 1427, 801, 75]. Transformed [1318]. Transformers [104, 10]. Transforming [877]. Transforms [1838]. Transient [1690, 1287]. Transients [715]. Transis [1635]. Transition [1840, 1791, 786, 1894]. Transition-Based [1840]. Transitions [1841, 345, 1557]. Translation [1435, 294, 1803, 265, 258, 882, 377]. Translation-Based [1803]. translations [351]. Translators [392]. TransLexis [260]. Translocation [1612]. transmission [902]. Transputer [933]. Transputer-based [933]. transversal [1892]. trapdoor [1961]. Treatment [1529, 2057, 1835, 457]. Tree [766, 1727, 395, 1024, 579, 887, 618, 621, 382, 726, 93, 665, 1614, 287, 1623, 1865, 765, 90]. Tree-width [382]. Trees [603, 165, 99, 1604, 162, 170, 1772, 663, 1051, 1717, 1609, 402, 321, 1613, 652, 2020, 1099, 764, 676, 1553, 1869, 1774, 435]. Treewidth [380, 1730, 1728, 381]. Trend [1487]. Trends [211]. Tri [290]. Tri-Diagonal [290]. Triangular [1913, 915]. Triangulation [1773, 1718]. tridiagonal [997]. Trigonometric [1916]. Triple [1732, 392]. Triple-Free [1732]. Trivializable [1831]. TRIWAQ [913]. Troll [60, 530]. True [1575, 412]. Trust [794]. trusted [1072]. Tuple [119]. Tuples [1616]. turbulent [956, 937]. Turing [628]. Tutorial [4, 345, 346]. Two [470, 89, 733, 291, 1053, 1692, 1918, 584, 627, 1582, 1492, 1861, 107, 1866, 105, 103]. Two-Connected [1053]. Two-Dimensional [291, 584]. Two-Phase [105]. Two-Phase-Commitment [107]. Two-Sorted [1582]. Two-Stage [1866]. Two-Way [627]. Type [2026, 543, 281, 1463, 1557, 1275, 352, 1875, 370, 1890, 1090, 379, 1057]. Type-Safe [2026]. Typeable [738]. typechecker [375]. Typed [363, 2059, 439, 256, 1438, 372, 378]. Types [1433, 520, 517, 263, 1436, 527, 2014, 518, 536, 1861, 2032, 539, 365, 369, 377]. Typical [1826, 1845, 1307]. Typing [1574, 539, 366].

UBS [1098, 1095]. UBS-Graph [1098, 1095]. Ultrasonic [1388]. Ultrasound [493, 492]. Unambiguous [1737]. Unbound [120]. Unbounded [623, 145, 1894]. Uncertain [1828, 1858]. Uncertainty [1782, 1813, 409, 1825, 1781, 1863, 1814, 1766]. Unconditionally [1973]. Uncoupled [1124]. Undecidability [317, 779]. Undecidable [3]. Undefined [1827]. Undeniable [1991]. Undergraduate [204, 182]. Understanding [195, 212, 1410, 911]. Undirected [169, 1629]. Undiscounted [7]. Uneasy [258]. Unequal [1729]. Unfoldings [1662]. Unidirectionally [1300]. Unification [1440, 745, 871, 744]. Unified [79, 30, 63, 484, 167, 106, 864]. Uniform [2044, 1228, 1177, 1001]. Uniformly [115]. Uniformly-Distributed [115]. Unifying [1153]. unions [429]. Unique [210]. Unit [1044, 676, 1364]. Units [1279]. Unity [1583, 811, 568, 1580, 1895]. Univariate [624]. Universal [516, 1082, 1260, 1085, 1086, 1228]. Universality [628]. University [178, 215]. Unix [1588]. Unknown [1253, 93]. Unlearning [6]. Unstructured [495]. Unsupervised [1321]. Untyped [366, 367]. Unwanted [2030]. Update [1335, 1840, 1843, 1166, 1172, 1684]. Updates [1841, 1175, 121]. Updating [1791]. Upper [1747, 1287, 1899]. UPT [1140]. Upward [162, 132, 160, 163, 130, 804, 161]. Usability [1438]. Usable [1472]. Usage [126]. Use [1034, 604, 1145, 385, 206, 1266, 1241, 261, 594, 686, 1499, 1718, 1419, 1429, 975, 1301, 430]. used [418]. Useful [1235, 1741]. Usefulness [1256]. User [1497, 223, 230, 1587, 231, 221, 1144, 235, 237, 1685, 1420, 828, 227, 229, 923]. Uses [212]. Using [1659, 572, 1699, 1388, 1383, 66, 896, 1650, 65, 482, 355, 1369, 1107, 1805, 1473, 491, 509, 512, 1802, 1149, 1139, 472, 1527, 231, 464, 1782, 2028, 874, 64, 1652, 1519, 661, 1409, 616, 515, 1675, 1108, 1912, 119, 212, 1043, 1403, 1851, 465, 1991, 163, 506, 1693, 69, 753, 341, 798, 1854, 451, 500, 156, 453, 1662, 1771, 1487, 1687, 1818, 1386, 633, 264, 346, 1863, 445, 1482, 1865, 1987, 825, 677, 481, 2043, 678, 892, 2035, 1379, 10, 2032, 508, 1400, 1500, 393, 1234, 991, 1537, 854, 865, 929, 1604, 1585, 1309, 1310, 925]. using [1584, 2042, 1018, 1550, 967, 1631, 948, 1689, 988, 1986]. Usually [1660]. Utilities [1776]. Utility [648]. Utilizing [1671, 901]. Utopia [2024].

V [404, 566, 2060, 1009]. V-Comprehensions [2060]. V-Structures [404]. Valid [1611]. Validating [1139, 443]. Validation [207]. Value [110, 1779, 74, 376]. Value-Based [74]. Valued [1311, 1807, 1163, 1806]. Variability [1938]. Variable [1757, 1617, 782]. Variable-Length [1757]. Variables [1804, 1442, 1483, 1110, 730, 1895]. Variance [1333]. Variances [1788]. variants [1995]. Variate [1784]. variation [869]. Variations [1902]. Varieties [608, 1780]. Various [339, 631]. Varying [1315, 840]. Vascular [500]. VC [436]. VC-dimension [436]. VCG [147]. Vector [1383, 1337, 1987, 966, 1986]. Vehicle [1235]. Velocity [1337]. Ventilation [1477]. Ventilator [1524]. ventricular [927]. Verifiable [1062, 1071]. Verification [349, 811, 1563, 2034, 1650, 1647, 1587, 1580, 343, 1576, 1678, 1675, 1458, 1653, 1657, 340, 332, 1063, 342, 1662, 345, 346, 1781, 1552, 207, 338, 337, 845]. Verification/Analysis [1675]. verified [375]. Verifying [1680, 1585, 1677, 1651, 892]. Version [1412, 1013]. Versioning [1690]. Versions [884, 1411]. Versus [1327, 1748, 414, 1746, 420, 431, 200]. Vertex [390]. Vertical [463]. Vertically [32]. Vessels [491]. via [419, 848, 946, 966, 1510, 68]. Video [1373, 831]. View [743, 1531, 1144, 703, 2003, 1462, 116, 1313]. Viewpoint [1295]. Viewpoints [684]. Views [705, 1196, 325]. Vinci [155]. Virtual [479, 460, 514, 450, 478, 1638]. Viruses [1202]. Visibility [142, 602, 143, 144]. Vision [444]. Vision-Based [444]. Vista [5]. Visual [1295, 1820, 236, 149, 157, 1262, 1950]. Visualisation [447]. Visualization [155, 443, 556, 448, 504, 1817, 513, 507, 128, 829]. Visualizations [697]. VLSI [1353, 1366, 1355, 1351, 968, 1363]. Vocabulary [264]. Volume [511, 506, 492, 508]. Volumetric [446]. votes [847]. voting [1090]. VPF [2022]. VR [145]. VR-Representable [145]. vs [1740, 307, 1638, 90, 1203].

W [1696]. W-SQL [1696]. Wadge [775]. Wahren [473]. water [1001]. wave [927]. Way [1404, 820, 627]. Ways [1413]. Ways-of-Working [1413]. Weak [1821, 612, 1054, 1957, 779, 278, 1990, 768]. Weakening [1799]. Weakly [1741]. weakness [842]. Web [212, 641]. Weight [1904]. Weighted [1368]. Well [1591, 1169, 1539, 1158, 1159, 1170]. Well-Founded [1591, 1169, 1170, 1158, 1159]. Where [31, 1967]. Which [1904, 820]. whole [858]. whose [1839, 1804]. Wicked [197]. Wide [641, 212]. Width [137, 382]. Wildcards [1608]. Will [31]. Winning [1666]. within [152, 1583]. Without [2030, 415, 1811, 1843, 1072, 1443, 405]. Word [286, 574, 575, 586, 578]. Word-Rewriting [578]. Words [285, 1713, 1714, 775, 855]. Work [1940]. Workbench [1428]. Worker [201]. Workflow [699]. Working [222, 269, 221, 220, 1413, 214]. Workload [1949, 952, 908]. Works [1660].

Workstation [64, 1515, 975, 952, 980, 986]. workstations [940, 966]. World [1173, 1292, 1239, 188, 944, 212, 641]. World-Wide [212]. Worlds [249]. worst [416]. worst-case [416]. WPVM [976]. Writer [89].

X [499, 1406, 567, 490]. X-Ray [499, 490]. XHIVE [929]. XSim [1345]. Xt [311]. Xt-Coloring [311].

Year [187]. Yekutieli [634]. Yprel [1304].

Zentrum [943]. Zero [1087, 849, 1912, 1088, 1091]. Zero-Knowledge [1087, 849, 1088, 1091]. ZPL [85]. Zuse [943].

References

Chaitin:1995:LMC

[1] G. J. Chaitin. The limits of mathematics — course outline and software. Lecture Notes in Computer Science, 888:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tarski:1995:HBG

[2] J. Tarski. Historical background of Goedel’s theorem. Lecture Notes in Computer Science, 888:3–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perrone:1995:FSA

[3] A. L. Perrone. A formal scheme for avoiding undecidable problems — applications to chaotic behaviour, characterization and parallel computation. Lecture Notes in Computer Science, 888:9–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chua:1995:CNN

[4] L. O. Chua, T. Roska, and T. Kozek. Cellular neural networks — A tutorial on programmable nonlinear dynamics in space. Lecture Notes in Computer Science, 888:53–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanHemmen:1995:NCT

[5] J. L. Van Hemmen and R. Ritz. Neural coding — A theoretical vista of mechanisms, techniques, and applications. Lecture Notes in Computer Science, 888:75–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wimbauer:1995:HU

[6] S. Wimbauer and J. L. Van Hemmen. Hebbian unlearning. Lecture Notes in Computer Science, 888:121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Murgu:1995:MDU

[7] A. Murgu. Mapping discounted and undiscounted Markov decision problems onto Hopfield neural networks. Lecture Notes in Computer Science, 888:137–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Molander:1995:BAB

[8] S. Molander. “blob” analysis of biomedical image sequences — A model-based and an inductive approach. Lecture Notes in Computer Science, 888:169–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurka:1995:SCD

[9] P. Kůrka. Simplicity criteria for dynamical systems. Lecture Notes in Computer Science, 888:189–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sintzoff:1995:ADS

[10] M. Sintzoff and F. Geurts. Analysis of dynamical systems using predicate transformers — attraction and composition. Lecture Notes in Computer Science, 888:227–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:IC

[11] Anonymous. An introduction to CSCW. Lecture Notes in Computer Science, 889:7–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:FSC

[12] Anonymous. Foundations of scientific cooperation. Lecture Notes in Computer Science, 889:60–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:TOT

[13] Anonymous. Today’s organisational and technical environment for scientific cooperation in Europe. Lecture Notes in Computer Science, 889:74–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:ECS

[14] Anonymous. An European “collaboratory” for supporting scientific cooperation. Lecture Notes in Computer Science, 889:102–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:ICF

[15] Anonymous. Implications of the “Ecollaboratory for a future European computing and network infrastructure”. Lecture Notes in Computer Science, 889:140–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:MTE

[16] Anonymous. Migrating towards an European Scientific Collaboratory. Lecture Notes in Computer Science, 889: 190–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bell:1995:CA

[17] J. Bell. Changing attitudes. Lecture Notes in Computer Science, 890:40–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castelfranchi:1995:GAC

[18] C. Castelfranchi. Guarantees for autonomy in cognitive agent architecture. Lecture Notes in Computer Science, 890:56–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cimatti:1995:MAR

[19] A. Cimatti and L. Serafini. Multi-agent reasoning with belief contexts: the approach and a case study. Lecture Notes in Computer Science, 890:71–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongha:1995:TFM

[20] P. Dongha. Toward a formal model of commitment for resource bounded agents. Lecture Notes in Computer Science, 890:86–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dunin-Keplicz:1995:CFS

[21] B. Dunin-Keplicz and J. Treur. Compositional formal specification of multiagent systems. Lecture Notes in Computer Science, 890:102–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guichard:1995:LRD

[22] F. Guichard and J. Ayel. Logical reorganization of DAI systems. Lecture Notes in Computer Science, 890:118– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soutchanski:1995:LFC

[23] M. Soutchanski and E. Ternovskaia. Logical formalization of concurrent actions for multi-agent systems. Lecture Notes in Computer Science, 890:129– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Staniford:1995:SAS

[24] G. Staniford and R. Paton. Simulating animal societies with adaptive communicating agents. Lecture Notes in Computer Science, 890:145–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wooldridge:1995:MLA

[25] M. Wooldridge. This is MyWorld: The logic of an agent-oriented DAI testbed. Lecture Notes in Computer Science, 890:160–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aitken:1995:KLC

[26] J. S. Aitken, F. Schmalhofer, and N. Shadbolt. A knowledge level characterisation of multi-agent systems. Lecture Notes in Computer Science, 890: 179–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ekdahl:1995:TAA

[27] B. Ekdahl, E. Astor, and P. Davidsson. Towards anticipatory agents. Lecture Notes in Computer Science, 890: 191–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferguson:1995:ICC

[28] I. A. Ferguson. Integrated control and coordinated behaviour: a case for agent models. Lecture Notes in Computer Science, 890:203–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1995:AAD

[29] J. Huang, N. R. Jennings, and J. Fox. An agent architecture for distributed medical care. Lecture Notes in Computer Science, 890:219–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malec:1995:UAI

[30] J. Malec. A unified approach to intelligent agency. Lecture Notes in Computer Science, 890:233–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moffat:1995:WTW

[31] D. Moffat and N. Frijda. Where there’s a will there’s an agent. Lecture Notes in Computer Science, 890:245–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1995:MRB

[32] J. P. Mueller, M. Pischel, and M. Thiel. Modelling reactive behaviour in vertically layered agent architectures. Lecture Notes in Computer Science, 890: 261–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Norman:1995:GCM

[33] T. J. Norman and D. Long. Goal creation in motivated agents. Lecture Notes in Computer Science, 890:277– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burkhard:1995:AOP

[34] H.-D. Burkhard. Agent-oriented programming for open systems. Lecture Notes in Computer Science, 890:291– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisher:1995:REA

[35] M. Fisher. Representing and executing agent-based systems. Lecture Notes in Computer Science, 890:307–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCabe:1995:AAP

[36] F. G. McCabe and K. L. Clark. April — agent process interaction language. Lecture Notes in Computer Science, 890:324–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Poggi:1995:DOO

[37] A. Poggi. DAISY: An object-oriented system for distributed artificial intelligence. Lecture Notes in Computer Science, 890:341–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1995:PAP

[38] S. R. Thomas. The PLACA agent programming language. Lecture Notes in Computer Science, 890:355–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wavish:1995:RSB

[39] P. Wavish and M. Graham. Roles, skills and behaviour: a situated action approach to organising systems of interacting agents. Lecture Notes in Computer Science, 890:371–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weerasooriya:1995:DCA

[40] D. Weerasooriya, A. Rao, and K. Ramamohanarao. Design of a concurrent agent-oriented language. Lecture Notes in Computer Science, 890:386– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lindner:1995:TD

[41] T. Lindner. Task description. Lecture Notes in Computer Science, 891:7–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lewerentz:1995:CS

[42] C. Lewerentz and T. Lindner. Comparative survey. Lecture Notes in Computer Science, 891:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noekel:1995:C

[43] K. Noekel and K. Winkelmann. CSL. Lecture Notes in Computer Science, 891:55–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Budde:1995:E

[44] R. Budde. ESTEREL. Lecture Notes in Computer Science, 891:75–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holenderski:1995:L

[45] L. Holenderski. Lustre. Lecture Notes in Computer Science, 891:101– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amagbegnon:1995:S

[46] T. Pascalin Amagbegnon, P. Le Guernic, H. Marchand, and E. Rutten. Signal. Lecture Notes in Computer Science, 891:113–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damm:1995:S

[47] W. Damm, H. Hungar, P. Kelb, and R. Schloer. Statecharts. Lecture Notes in Computer Science, 891:131– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cuellar:1995:T

[48] J. Cuellar and M. Huber. TLT. Lecture Notes in Computer Science, 891: 151–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heinkel:1995:S

[49] S. Heinkel and T. Lindner. SDL. Lecture Notes in Computer Science, 891: 171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuchs:1995:F

[50] M. Fuchs and J. Philipps. Focus. Lecture Notes in Computer Science, 891: 185–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dranidis:1995:S

[51] D. Dranidis and S. Gastinger. Spectrum. Lecture Notes in Computer Science, 891:199–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schellhorn:1995:K

[52] G. Schellhorn and A. Burandt. KIV. Lecture Notes in Computer Science, 891:229–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klingenbeck:1995:T

[53] S. Klingenbeck and T. Kaeuft. Tatzelwurm. Lecture Notes in Computer Science, 891:247–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cardell-Oliver:1995:HH

[54] R. Cardell-Oliver. HTTDs and HOL. Lecture Notes in Computer Science, 891:261–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erasmy:1995:R

[55] F. Erasmy and E. Sekerinski. Raise. Lecture Notes in Computer Science, 891:277–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burghardt:1995:DS

[56] J. Burghardt. Deductive synthesis. Lecture Notes in Computer Science, 891:295–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Korf:1995:STD

[57] F. Korf and R. Schloer. Symbolic timing diagrams. Lecture Notes in Computer Science, 891:311–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wieringa:1995:LM

[58] R. Wieringa. LCM and MCM. Lecture Notes in Computer Science, 891: 333–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rueping:1995:M

[59] A. Rueping and E. Sekerinski. Modula-3. Lecture Notes in Computer Science, 891:357–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herzig:1995:TL

[60] R. Herzig and N. Vlachantonis. Troll light. Lecture Notes in Computer Science, 891:373–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brauer:1995:S

[61] A. Brauer and T. Lindner. Simulation. Lecture Notes in Computer Science, 891:383–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feautrier:1995:FGS

[62] P. Feautrier. Fine-grain scheduling under resource constraints. Lecture Notes in Computer Science, 892:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Novack:1995:MSU

[63] S. Novack and A. Nicolau. Mutation scheduling: A unified approach to compiling for fine-grain parallelism. Lecture Notes in Computer Science, 892:16–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dietz:1995:CTF

[64] H. G. Dietz, W. E. Cohen, T. Muhammad, and T. I. Mattox. Compiler techniques for fine-grain execution on workstation clusters using PAPERS. Lecture Notes in Computer Science, 892:31–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bau:1995:SAU

[65] D. Bau, I. Kodukula, V. Kotlyar, and K. Pingali. Solving alignment using elementary linear algebra. Lecture Notes in Computer Science, 892:46–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ayguade:1995:DUA

[66] E. Ayguade, J. Garcia, M. Girones, and J. Labarta. Detecting and using affinity in an automatic data distribution tool. Lecture Notes in Computer Science, 892:61–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chatterjee:1995:ADD

[67] S. Chatterjee, R. Schreiber, T. J. Sheffler, and J. R. Gilbert. Array distribution in data-parallel programs. Lecture Notes in Computer Science, 892:76–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lim:1995:CFP

[68] A. W. Lim and M. S. Lam. Communication-free parallelization via affine transformations. Lecture Notes in Computer Science, 892:92–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kelly:1995:FLR

[69] W. Kelly and W. Pugh. Finding legal reordering transformations using mappings. Lecture Notes in Computer Science, 892:107–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Appelbe:1995:NAG

[70] B. Appelbe, S. Doddapaneni, and C. Hardnett. A new algorithm for global optimization for parallelism and locality. Lecture Notes in Computer Science, 892:125–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blume:1995:PIE

[71] W. Blume, R. Eigenmann, K. Faigin, and J. Grout. Polaris: Improving the effectiveness of parallelizing compilers. Lecture Notes in Computer Science, 892:141–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trescher:1995:FAC

[72] J. A. Trescher, L. C. Breebaart, P. F. G. Dechering, and A. B. Poelman. A formal approach to the compilation of data-parallel languages. Lecture Notes in Computer Science, 892: 155–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nakanishi:1995:DPG

[73] T. Nakanishi, K. Joe, A. Fukuda, and K. Araki. The data partitioning graph: Extending data and control dependencies for data partitioning. Lecture Notes in Computer Science, 892:170–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoltz:1995:DVB

[74] E. Stoltz and M. Wolfe. Detecting value-based scalar dependence. Lecture Notes in Computer Science, 892:186– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1995:MDD

[75] Y.-Q. Yang, C. Ancourt, and F. Irigoin. Minimal data dependence abstractions for loop transformations. Lecture Notes in Computer Science, 892:201– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sgro:1995:DAP

[76] V. Sgro and B. G. Ryder. Differences in algorithmic parallelism in control flow and call multigraphs. Lecture Notes in Computer Science, 892:217–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burke:1995:FII

[77] M. Burke, P. Carini, J.-D. Choi, and M. Hind. Flow-insensitive interprocedural alias analysis in the presence of pointers. Lecture Notes in Computer Science, 892:234–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaushik:1995:IGI

[78] S. D. Kaushik, C.-H. Huang, and P. Sadayappan. Incremental generation of index sets for array statement execution on distributed-memory machines. Lecture Notes in Computer Science, 892:251–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1995:UDF

[79] M. Gupta, E. Schonberg, and H. Srinivasan. A unified data-flow framework for optimizing communication. Lecture Notes in Computer Science, 892:266– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Agrawal:1995:ICO

[80] G. Agrawal and J. Saltz. Interprocedural communication optimizations for distributed memory compilation. Lecture Notes in Computer Science, 892:283–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramanujam:1995:AES

[81] J. Ramanujam and A. Mathew. Analysis of event synchronization in parallel programs. Lecture Notes in Computer Science, 892:300–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrante:1995:CCS

[82] J. Ferrante. Computing communication sets for control parallel programs. Lecture Notes in Computer Science, 892:316–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krishnamurthy:1995:OPS

[83] A. Krishnamurthy and K. Yelick. Optimizing parallel SPMD programs. Lecture Notes in Computer Science, 892: 331–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mehrotra:1995:OOL

[84] P. Mehrotra and M. Haines. An overview of the Opus language and runtime system. Lecture Notes in Computer Science, 892:346–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1995:SPR

[85] C. Lin and L. Snyder. SIMPLE performance results in ZPL. Lecture Notes in Computer Science, 892:361–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikhil:1995:CPS

[86] R. S. Nikhil. Cid: A parallel, “sharedmemory” C for distributed-memory machines. Lecture Notes in Computer Science, 892:376–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Derby:1995:EON

[87] T. Derby, R. Schnabel, and B. Zorn. EQ: Overview of a new language approach for prototyping scientific computation. Lecture Notes in Computer Science, 892:391–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bik:1995:RAP

[88] A. J. C. Bik, P. M. W. Knijnenburg, and H. A. G. Wijshoff. Reshaping access patterns for generating sparse codes. Lecture Notes in Computer Science, 892:406–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodin:1995:ETL

[89] F. Bodin, T. Montaut, and E. D. Granston. Evaluating two loop transformations for reducing multiple-writer false sharing. Lecture Notes in Computer Science, 892:423–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Solworth:1995:PTA

[90] J. A. Solworth and B. B. Reagan. Parallelizing tree algorithms: Overhead vs. parallelism. Lecture Notes in Computer Science, 892:440–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreira:1995:ADS

[91] J. E. Moreira and C. D. Polychronopoulos. Autoscheduling in a distributed shared-memory environment. Lecture Notes in Computer Science, 892:455–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kunchithapadam:1995:OAD

[92] K. Kunchithapadam and B. P. Miller. Optimizing array distributions in dataparallel programs. Lecture Notes in Computer Science, 892:472–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hiraki:1995:ART

[93] K. Hiraki. Automatic reduction tree generation for fine-grain parallel architectures when iteration count is unknown. Lecture Notes in Computer Science, 892:487–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imielinski:1995:DAW

[94] T. Imielinski. Data on air — what’s in it for the database theorist. Lecture Notes in Computer Science, 893:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paredaens:1995:SDF

[95] J. Paredaens. Spatial databases, the final frontier. Lecture Notes in Computer Science, 893:14–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weikum:1995:PDS

[96] G. Weikum. Parallel database systems. Lecture Notes in Computer Science, 893:33–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolaitis:1995:LPT

[97] P. G. Kolaitis. Languages for polynomial-time queries — an ongoing quest. Lecture Notes in Computer Science, 893:38–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salza:1995:DQO

[98] S. Salza, G. Barone, and T. Morzy. Distributed query optimization in loosely coupled multidatabase systems. Lecture Notes in Computer Science, 893: 40–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cluet:1995:CGO

[99] S. Cluet and G. Moerkotte. On the complexity of generating optimal leftdeep processing trees with cross products. Lecture Notes in Computer Science, 893:54–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bonatti:1995:QDD

[100] P. A. Bonatti and T. Eiter. Querying disjunctive databases through nonmonotonic logics. Lecture Notes in Computer Science, 893:68–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greco:1995:DQS

[101] S. Greco, D. Sacca, and C. Zaniolo. DATALOG queries with stratified negation and choice: from P to D0P. Lecture Notes in Computer Science, 893:82–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tyszkiewicz:1995:KEP

[102] J. Tyszkiewicz. On the Kolmogorov expressive power of Boolean query languages. Lecture Notes in Computer Science, 893:97–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suciu:1995:TFS

[103] D. Suciu and L. Wong. On two forms of structural recursion. Lecture Notes in Computer Science, 893:111–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sazonov:1995:CDL

[104] V. Sazonov and A. Lisitsa. cap deltalanguages for sets and Sub-PTIME graph transformers. Lecture Notes in Computer Science, 893:125–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soisalon-Soininen:1995:PST

[105] E. Soisalon-Soininen and T. Yloenen. Partial strictness in two-phase locking. Lecture Notes in Computer Science, 893:139–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vingralek:1995:UTM

[106] R. Vingralek, H. Ye, Y. Breitbart, and H.-J. Schek. Unified transaction model for semantically rich operations. Lecture Notes in Computer Science, 893:148–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raz:1995:DTP

[107] Y. Raz. The dynamic two-phase-commitment (D2PC) protocol. Lecture Notes in Computer Science, 893:162– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suciu:1995:DIQ

[108] D. Suciu. Domain-independent queries on databases with external functions. Lecture Notes in Computer Science, 893:177–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grumbach:1995:AP

[109] S. Grumbach and T. Milo. An algebra for pomsets. Lecture Notes in Computer Science, 893:191–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cabibbo:1995:PSL

[110] L. Cabibbo. On the power of stratified logic programs with value invention for expressing database transformations. Lecture Notes in Computer Science, 893:208–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bugliesi:1995:SMS

[111] M. Bugliesi and H. M. Jamil. A stable model semantics for behavioral inheritance in deductive object oriented languages. Lecture Notes in Computer Science, 893:222–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montesi:1995:RTA

[112] D. Montesi and R. Torlone. A rewriting technique for the analysis and the optimization of active databases. Lecture Notes in Computer Science, 893: 238–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1995:SHS

[113] W. Chen and M. Kifer. Sorted HiLog: Sorts in higher-order logic data languages. Lecture Notes in Computer Science, 893:252–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demetrovics:1995:ALK

[114] J. Demetrovics, G. O. H. Katona, D. Miklos, and O. Seleznjev. The average length of keys and functional dependencies in (random) databases. Lecture Notes in Computer Science, 893: 266–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galindo-Legaria:1995:UDR

[115] C. A. Galindo-Legaria, A. Pellenkoft, and M. L. Kersten. Uniformlydistributed random generation of join orders. Lecture Notes in Computer Science, 893:280–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lifschitz:1995:PVD

[116] S. Lifschitz and V. Vianu. A probabilistic view of datalog parallelization. Lecture Notes in Computer Science, 893:294–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1995:FST

[117] N. Bidoit and S. De Amo. A first step towards implementing dynamic algebraic dependencies. Lecture Notes in Computer Science, 893:308–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baudinet:1995:CGD

[118] M. Baudinet, J. Chomicki, and P. Wolper. Constraint-generating dependencies. Lecture Notes in Computer Science, 893:322–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harinarayan:1995:OUT

[119] V. Harinarayan and A. Gupta. Optimization using tuple subsumption. Lecture Notes in Computer Science, 893: 338–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sacca:1995:DND

[120] D. Sacca. Deterministic and nondeterministic stable model semantics for unbound DATALOG queries. Lecture Notes in Computer Science, 893: 353–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

WiktorMarek:1995:RPD

[121] V. Wiktor Marek and M. Trusczynski. Revision programming, database updates and integrity constraints. Lecture Notes in Computer Science, 893: 368–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1995:SPR

[122] K. Wang. Some positive results for boundedness of multiple recursive rules. Lecture Notes in Computer Science, 893:383–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dong:1995:IBN

[123] G. Dong and J. Su. Increment boundedness and nonrecursive incremental evaluation of datalog queries. Lecture Notes in Computer Science, 893:397– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Libkin:1995:AD

[124] L. Libkin. Approximation in databases. Lecture Notes in Computer Science, 893:411–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Revesz:1995:DQS

[125] P. Z. Revesz. Datalog queries of set constraint databases. Lecture Notes in Computer Science, 893:425–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abiteboul:1995:SUF

[126] S. Abiteboul and G. Hillebrand. Space usage in functional query languages. Lecture Notes in Computer Science, 893:439–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1995:TDG

[127] R. F. Cohen, P. Eades, T. Lin, and F. Ruskey. Three-dimensional graph drawing. Lecture Notes in Computer Science, 894:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reiss:1995:DVP

[128] S. P. Reiss. 3-D visualization of program information. Lecture Notes in Computer Science, 894:12–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jeron:1995:LRG

[129] T. Jeron and C. Jard. 3D layout of reachability graphs of communicating processes. Lecture Notes in Computer Science, 894:25–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohar:1995:FAU

[130] B. Mohar and P. Rosenstiehl. A flow approach to upward drawings of toroidal maps. Lecture Notes in Computer Science, 894:33–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papakostas:1995:IAB

[131] A. Papakostas and I. G. Tollis. Improved algorithms and bounds for orthogonal drawings. Lecture Notes in Computer Science, 894:40–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Foessmeier:1995:BMO

[132] U. Foessmeier and M. Kaufmann. On bend-minimum orthogonal upward drawing of directed planar graphs. Lecture Notes in Computer Science, 894: 52–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Even:1995:GLB

[133] S. Even and G. Granot. Grid layouts of block diagrams — bounding the number of bends in each connection. Lecture Notes in Computer Science, 894:64–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Djidjev:1995:DGC

[134] H. N. Djidjev. On drawing a graph convexly in the plane. Lecture Notes in Computer Science, 894:76–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garg:1995:DAG

[135] A. Garg. On drawing angle graphs. Lecture Notes in Computer Science, 894:84–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

He:1995:REL

[136] X. He and M.-Y. Kao. Regular edge labelings and drawings of planar graphs. Lecture Notes in Computer Science, 894:96–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chrobak:1995:MWG

[137] M. Chrobak and S. Nakano. Minimumwidth grid drawings of plane graphs. Lecture Notes in Computer Science, 894:104–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeFraysseix:1995:ROA

[138] H. De Fraysseix and P. O. De Mendez. Regular orientations, arboricity, and augmentation. Lecture Notes in Computer Science, 894:111–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juenger:1995:PAM

[139] M. Juenger and P. Mutzel. The polyhedral approach to the maximum planar subgraph problem: New chances for related problems. Lecture Notes in Computer Science, 894:119–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shahrokhi:1995:CNG

[140] F. Shahrokhi, L. A. Szekely, and I. Vrt’o. Crossing numbers of graphs, lower bound techniques and algorithms. Lecture Notes in Computer Science, 894:131–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1995:GDC

[141] P. Eades and J. Marks. Graph drawing contest report. Lecture Notes in Computer Science, 894:143–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abello:1995:VGO

[142] J. Abello and K. Kumar. Visibility graphs and oriented matroids. Lecture Notes in Computer Science, 894:147– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dean:1995:RVR

[143] A. M. Dean and J. P. Hutchinson. Rectangle-visibility representations of bipartite graphs. Lecture Notes in Computer Science, 894:159–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kranakis:1995:NDV

[144] E. Kranakis, D. Krizanc, and J. Urrutia. On the number of directions in visibility representations of graphs. Lecture Notes in Computer Science, 894:167–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Romanik:1995:DVR

[145] K. Romanik. Directed VR-Representable graphs have unbounded dimension. Lecture Notes in Computer Science, 894:177–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Himsolt:1995:GGP

[146] M. Himsolt. GraphEd: a graphical platform for the implementation of graph algorithms. Lecture Notes in Computer Science, 894:182–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sander:1995:GLT

[147] G. Sander. Graph layout through the VCG tool. Lecture Notes in Computer Science, 894:194–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tollis:1995:DTN

[148] I. G. Tollis and C. Xia. Drawing telecommunication networks. Lecture Notes in Computer Science, 894:206– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kimelman:1995:RVC

[149] D. Kimelman, B. Leban, T. Roth, and D. Zernik. Reduction of visual complexity in dynamic graphs. Lecture Notes in Computer Science, 894:218– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogt:1995:TGT

[150] F. Vogt and R. Wille. TOSCANA — a graphical tool for analyzing and exploring data. Lecture Notes in Computer Science, 894:226–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tanenbaum:1995:CRP

[151] P. J. Tanenbaum, M. T. Goodrich, and E. R. Scheinerman. Characterization and recognition of point-halfspace and related orders. Lecture Notes in Computer Science, 894:234–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abellanas:1995:RGW

[152] M. Abellanas, F. Hurtado, and P. A. Ramos. Redrawing a graph within a geometric tolerance. Lecture Notes in Computer Science, 894:246–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godau:1995:DEP

[153] M. Godau. On the difficulty of embedding planar graphs with inaccuracies. Lecture Notes in Computer Science, 894:254–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manning:1995:SDG

[154] J. Manning, M. Atallah, K. Cudjoe, and J. Lozito. A system for drawing graphs with geometric symmetry. Lecture Notes in Computer Science, 894:262–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Froehlich:1995:DIG

[155] M. Froehlich and M. Werner. Demonstration of the interactive graphvisualization system da Vinci. Lecture Notes in Computer Science, 894:266– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McCreary:1995:CGD

[156] C. McCreary, F.-S. Shieh, and H. Gill. CG: a graph drawing system using graph-grammar parsing. Lecture Notes in Computer Science, 894:270– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McWhirter:1995:DIG

[157] J. D. McWhirter. A demonstration of interactive graph based visual language applications. Lecture Notes in Computer Science, 894:274–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaugars:1995:SAD

[158] K. Kaugars, J. Reinfelds, and A. Brazma. A simple algorithm for drawing large graphs on small screens. Lecture Notes in Computer Science, 894:278–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fowler:1995:MGE

[159] P. W. Fowler, T. Pisanski, and J. Shawe-Taylor. Molecular graph eigenvectors for molecular coordinates. Lecture Notes in Computer Science, 894:282–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garg:1995:CCU

[160] A. Garg and R. Tamassia. On the computational complexity of upward and rectilinear planarity testing. Lecture Notes in Computer Science, 894:286– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papakostas:1995:UPT

[161] A. Papakostas. Upward planarity testing of outerplanar dags. Lecture Notes in Computer Science, 894:298– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crescenzi:1995:OAU

[162] P. Crescenzi and A. Piperno. Optimalarea upward drawings of AVL trees. Lecture Notes in Computer Science, 894:307–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jourdan:1995:UDP

[163] G.-V. Jourdan, I. Rival, and N. Zaguia. Upward drawing on the plane grid using less ink. Lecture Notes in Computer Science, 894:318–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiBattista:1995:PD

[164] G. Di Battista, W. Lenhart, and G. Liotta. Proximity drawability. Lecture Notes in Computer Science, 894:328– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bose:1995:PCR

[165] P. Bose, G. Di Battista, W. Lenhart, and G. Liotta. Proximity constraints and representable trees. Lecture Notes in Computer Science, 894:340– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

ElGindy:1995:RRI

[166] H. ElGindy, G. Liotta, A. Lubiw, and H. Meijer. Recognizing rectangle of influence drawable graphs. Lecture Notes in Computer Science, 894:352– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sugiyama:1995:SUM

[167] K. Sugiyama and K. Misue. A simple and unified method for drawing graphs: Magnetic-spring algorithm. Lecture Notes in Computer Science, 894:364– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1995:IDA

[168] T. Lin and P. Eades. Integration of declarative and algorithmic approaches for layout creation. Lecture Notes in Computer Science, 894:376–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frick:1995:FAL

[169] A. Frick, A. Ludwig, and H. Mehldau. A fast adaptive layout algorithm for undirected graphs. Lecture Notes in Computer Science, 894:388–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cruz:1995:DGE

[170] I. F. Cruz and A. Garg. Drawing graphs by example efficiently: Trees and planar acyclic diagraphs. Lecture Notes in Computer Science, 894:404– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandenburg:1995:DGD

[171] F. J. Brandenburg. Designing graph drawings by layout graph grammars. Lecture Notes in Computer Science, 894:416–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noik:1995:EPE

[172] E. G. Noik. Encoding presentation emphasis algorithms for graphs. Lecture Notes in Computer Science, 894:428– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Luo:1995:PAS

[173] J. Luo and K. Miriyala. A practical approach to static node positioning. Lecture Notes in Computer Science, 894:436–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saalfeld:1995:MGG

[174] A. Saalfeld. Map generalization as a graph drawing problem. Lecture Notes in Computer Science, 894:444– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stumme:1995:GHD

[175] G. Stumme and R. Wille. A geometrical heuristic for drawing concept lattices. Lecture Notes in Computer Science, 894:452–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whitesides:1995:PGR

[176] S. H. Whitesides. Poster gallery report. Lecture Notes in Computer Science, 894:461–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fairley:1995:SHQ

[177] R. E. Fairley. Some hard questions for software engineering educators. Lecture Notes in Computer Science, 895:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zucconi:1995:EKP

[178] L. Zucconi. Essential knowledge for the practising software engineer and the responsibilities of university and industry for her education. Lecture Notes in Computer Science, 895:5–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maly:1995:CPI

[179] K. J. Maly, D. E. Ray, J. C. Wild, and I. B. Levinstein. Computer productivity initiative: Past, present, and future. Lecture Notes in Computer Science, 895:15–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gorski:1995:SEE

[180] J. Gorski. Software engineering education: A dragging-through approach. Lecture Notes in Computer Science, 895:31–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barbacci:1995:ESE

[181] M. Barbacci, P. Douglas, D. Carver, and J. Werth. Establishing software engineering as a profession. Lecture Notes in Computer Science, 895:47–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dick:1995:IIU

[182] G. N. Dick and S. F. Jones. Industry involvement in undergraduate curricula: Reinforcing learning by applying the principles. Lecture Notes in Computer Science, 895:51–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garlan:1995:CMS

[183] D. Garlan, A. Brown, D. Jackson, and J. Tomayko. The CMU master of software engineering core curriculum. Lecture Notes in Computer Science, 895:65–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hilburn:1995:ISE

[184] T. B. Hilburn, I. Hirmanpour, and A. Kornecki. The integration of software engineering into a computer science curriculum. Lecture Notes in Computer Science, 895:87–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atkinson:1995:ESE

[185] C. Atkinson, D. Eichmann, and C. McKay. An evolution of a software engineering curriculum. Lecture Notes in Computer Science, 895:99–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manns:1995:OOT

[186] M. L. Manns and G. I. Puhr. Objectoriented technology education and training: Bridging the gap between academia and industry. Lecture Notes in Computer Science, 895:115– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beckman:1995:CBS

[187] K. Beckman and R. Burd. CMM-Based software process improvement training: The first year. Lecture Notes in Computer Science, 895:117–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waddoups:1995:EWB

[188] R. O. Waddoups. Education — the world’s best investment. Lecture Notes in Computer Science, 895:119– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moore:1995:PIC

[189] M. M. Moore and T. Brennan. Process improvement in the classroom. Lecture Notes in Computer Science, 895:123– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khajenoori:1995:PSP

[190] S. Khajenoori and I. Hirmanpour. Personal software process: An experiential report. Lecture Notes in Computer Science, 895:131–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacquot:1995:PTP

[191] J.-P. Jacquot. The process of teaching process. Lecture Notes in Computer Science, 895:147–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mead:1995:MSE

[192] N. R. Mead, M. F. Browning, D. Carter, and N. Gibbs. Managing software engineering education in diverse environments. Lecture Notes in Computer Science, 895:159–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoover:1995:RSE

[193] C. L. Hoover. The role of the software engineer in real-time software development: An introductory course. Lecture Notes in Computer Science, 895: 167–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paprzycki:1995:PDC

[194] M. Paprzycki, R. Wasniowski, and J. Zalewski. Parallel and distributed computing education: A software engineering approach. Lecture Notes in Computer Science, 895:187–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

France:1995:URF

[195] R. B. France and M. M. Larrondo-Petrie. Understanding the role of formal specification techniques in requirements engineering. Lecture Notes in Computer Science, 895:207–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allen:1995:IAT

[196] P. Allen and L. Semmens. An integrated approach to teaching requirements modelling. Lecture Notes in Computer Science, 895:223–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Budgen:1995:TSD

[197] D. Budgen. Is teaching software design a ‘wicked’ problem too? Lecture Notes in Computer Science, 895:239– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robinson:1995:SEC

[198] J. S. Robinson. Software engineering curriculum development model. Lecture Notes in Computer Science, 895: 257–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holeman:1995:SPI

[199] R. Holeman. The software process improvement game. Lecture Notes in Computer Science, 895:261–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stinson:1995:AFV

[200] M. C. Stinson. Academic freedom versus quality assurance. Lecture Notes in Computer Science, 895:263–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Browning:1995:ESS

[201] M. F. Browning. Endangered species? the single-skilled information worker. Lecture Notes in Computer Science, 895:267–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parrish:1995:ETM

[202] A. Parrish, D. Brown, and D. Cordes. Experience in teaching a managementoriented capstone software engineering course. Lecture Notes in Computer Science, 895:279–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scott:1995:TSM

[203] T. J. Scott and J. H. Cross. Team selection methods for student programming projects. Lecture Notes in Computer Science, 895:295–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burnstein:1995:DLS

[204] I. Burnstein and C. R. Carlson. Developing leadership skills in software engineering students through an undergraduate research program. Lecture Notes in Computer Science, 895:305– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1995:CTT

[205] K. E. Smith. CARDS training: Transferring reuse knowledge. Lecture Notes in Computer Science, 895:325– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coyne:1995:TMC

[206] R. F. Coyne, A. H. Dutoit, B. Bruegge, and D. Rothenberger. Teaching more comprehensive model-based software engineering: Experience with objectory’s use case approach. Lecture Notes in Computer Science, 895:339– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shepard:1995:TSV

[207] T. Shepard. On teaching software verification and validation. Lecture Notes in Computer Science, 895:375– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geier:1995:DSM

[208] J. D. Geier. The Delphi survey methodology: An approach to determine training needs. Lecture Notes in Computer Science, 895:389–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bayman:1995:SDP

[209] S. Bayman, M. H. Blumberg, H. Cumm, and C. Diallo. A skills-driven process for training computer professionals. Lecture Notes in Computer Science, 895:403–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fortin:1995:TUA

[210] P. D. Fortin, C. A. Jeske, J. R. Lakey, and K. B. Urquhart. Is this training? A unique approach to software process training in industry. Lecture Notes in Computer Science, 895:409–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mead:1995:ETT

[211] N. R. Mead. Education trends and their impact on management of software engineering education. Lecture Notes in Computer Science, 895:419– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hefley:1995:UUD

[212] W. E. Hefley. Understanding, using and designing for educational uses of the World-Wide Web. Lecture Notes in Computer Science, 895:431–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mahy:1995:TLR

[213] I. Mahy. From TRAINING to LEARN-ING: The reengineering of training at DMR Group Inc. Lecture Notes in Computer Science, 895:433–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Talbert:1995:MQC

[214] S. Talbert and E. Gyalokay. Managing quality in course staging — A working paper. Lecture Notes in Computer Science, 895:437–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Livingston:1995:RSC

[215] J. Livingston, K. Wright, and L. Shafer. A Russian software center ascends the SEI maturity: Education by Motorola University. Lecture Notes in Computer Science, 895:439–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Card:1995:TPP

[216] D. N. Card. Teaching practical principles of software measurement. Lecture Notes in Computer Science, 895: 443–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McKim:1995:TOO

[217] J. C. McKim. Teaching object oriented programming and design with eiffel. Lecture Notes in Computer Science, 895:445–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Humphrey:1995:PPS

[218] W. S. Humphrey. The personal process in software engineering. Lecture Notes in Computer Science, 895:447– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Almstrum:1995:RMC

[219] V. L. Almstrum, D. Burton, C. Christi, and A. Fleury. Research methods in computer science education. Lecture Notes in Computer Science, 895:449– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grudin:1995:WGD

[220] J. Grudin and R. T. Fielding. Working group on design methods and processes. Lecture Notes in Computer Science, 896:4–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chimera:1995:WGE

[221] R. Chimera. Working group on evaluating user interfaces and user interface tools. Lecture Notes in Computer Science, 896:9–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bass:1995:WGF

[222] L. Bass. Working group on formal methods in HCl and software engineering. Lecture Notes in Computer Science, 896:14–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bass:1995:IEU

[223] L. Bass, G. Abowd, and R. Kazman. Issues in the evaluation of user interface tools. Lecture Notes in Computer Science, 896:17–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chimera:1995:PIG

[224] R. Chimera. Platform independent GUI builders advance software engineering to handle HCl issues. Lecture Notes in Computer Science, 896:28–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coutaz:1995:ETE

[225] J. Coutaz. Evaluation techniques: Exploring the intersection of HCl and software engineering. Lecture Notes in Computer Science, 896:35–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harrison:1995:RFD

[226] M. D. Harrison and D. J. Duke. A review of formalisms for describing interactive behaviour. Lecture Notes in Computer Science, 896:49–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szekely:1995:UIP

[227] P. Szekely. User interface prototyping: Tools and techniques. Lecture Notes in Computer Science, 896:76–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grudin:1995:SEC

[228] J. Grudin and S. Poltrock. Software engineering and the CHI and CSCW communities. Lecture Notes in Computer Science, 896:93–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taylor:1995:UIT

[229] R. N. Taylor. User interface technology and software engineering environments. Lecture Notes in Computer Science, 896:113–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beck:1995:IOO

[230] A. Beck, C. Janssen, A. Weisbecker, and J. Ziegler. Integrating objectoriented analysis and graphical user interface design. Lecture Notes in Computer Science, 896:127–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carneiro-Coffin:1995:EUJ

[231] L. M. F. Carneiro-Coffin, D. D. Cowan, C. J. P. Lucena, and D. Smith. An experience using JASMINUM — formalization assisting with the design of user interfaces. Lecture Notes in Computer Science, 896:141–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duke:1995:FMF

[232] D. J. Duke and M. D. Harrison. From formal models to formal methods. Lecture Notes in Computer Science, 896: 159–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fafchamps:1995:CEF

[233] D. Fafchamps and P. Garg. Computing environments for flexible teams. Lecture Notes in Computer Science, 896: 174–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fields:1995:AFM

[234] B. Fields, P. Wright, and M. Harrison. Applying formal methods for human error tolerant design. Lecture Notes in Computer Science, 896:185–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gellersen:1995:SUI

[235] H.-W. Gellersen. Support of user interface design aspects in a framework for distributed cooperative applications. Lecture Notes in Computer Science, 896:196–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godfrey:1995:PVF

[236] M. W. Godfrey, R. C. Holt, and S. Mancoridis. Prototyping a visual formalism for system modelling. Lecture Notes in Computer Science, 896: 211–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keller:1995:UIT

[237] R. K. Keller. User interface tools: a survey and perspective. Lecture Notes in Computer Science, 896:225– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Isaacs:1995:LLS

[238] E. A. Isaacs, T. Morris, and T. K. Rodriguez. Lessons learned from a successful collaboration between software engineers and human interface engineers. Lecture Notes in Computer Science, 896:232–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pickering:1995:SEC

[239] J. M. Pickering and R. E. Grinter. Software engineering and CSCW: A common research ground. Lecture Notes in Computer Science, 896:241–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDeVanter:1995:PLB

[240] M. L. Van De Vanter. Practical language-based editing for software engineers. Lecture Notes in Computer Science, 896:251–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ziv:1995:RII

[241] H. Ziv and L. J. Osterweil. Research issues in the intersection of hypertext and software development environments. Lecture Notes in Computer Science, 896:268–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brzoska:1995:TLP

[242] C. Brzoska. Temporal logic programming with metric and past operators. Lecture Notes in Computer Science, 897:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kono:1995:CCN

[243] S. Kono. A combination of clausal and non clausal temporal logic programs. Lecture Notes in Computer Science, 897:40–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fruehwirth:1995:TLA

[244] T. Fruehwirth. Temporal logic and annotated constraint logic programming. Lecture Notes in Computer Science, 897:58–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Merz:1995:EET

[245] S. Merz. Efficiently executable temporal logic programs. Lecture Notes in Computer Science, 897:69–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisher:1995:TSC

[246] M. Fisher. Towards a semantics for concurrent METATEM. Lecture Notes in Computer Science, 897:86–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koehler:1995:CDI

[247] J. Koehler and R. Treinen. Constraint deduction in an interval-based temporal logic. Lecture Notes in Computer Science, 897:103–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reynolds:1995:TFO

[248] M. Reynolds. Towards first-order concurrent METATEM. Lecture Notes in Computer Science, 897:118–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cavalcanti:1995:SAT

[249] M. Cavalcanti. Solving air-traffic problems with “possible worlds”. Lecture Notes in Computer Science, 897:144– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DenHaan:1995:IAD

[250] N. Den Haan. Investigations into the application of Deontic logic. Lecture Notes in Computer Science, 897:157– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ide:1995:KEM

[251] N. Ide and J. Veronis. Knowledge extraction from machine-readable dictionaries: An evaluation. Lecture Notes in Computer Science, 898:19–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Storrer:1995:DAM

[252] A. Storrer and U. Schwall. Description and acquisition of multiword lexemes. Lecture Notes in Computer Science, 898:35–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ahmad:1995:PST

[253] K. Ahmad. Pragmatics of specialist terms: The acquisition and representation of terminology. Lecture Notes in Computer Science, 898:51–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Procter:1995:CLS

[254] P. Procter. The Cambridge language survey. Lecture Notes in Computer Science, 898:77–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daelemans:1995:MBL

[255] W. Daelemans. Memory-based lexical acquisition and processing. Lecture Notes in Computer Science, 898:85–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krieger:1995:TFF

[256] H.-U. Krieger. Typed feature formalisms as a common basis for linguistic specification. Lecture Notes in Computer Science, 898:101–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calzolari:1995:EET

[257] N. Calzolari. European efforts towards standardizing language resources. Lecture Notes in Computer Science, 898: 121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koch:1995:MTT

[258] K. Koch. Machine translation and terminology database — uneasy bedfellows? Lecture Notes in Computer Science, 898:131–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachut:1995:GLM

[259] D. Bachut, I. Duquennoy, L. Humphreys, and T. Kyriakopoulou. A generic lexical model. Lecture Notes in Computer Science, 898:141–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blaeser:1995:TIE

[260] B. Blaeser. TransLexis: An integrated environment for lexicon and terminology management. Lecture Notes in Computer Science, 898:159–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karkaletsis:1995:UTK

[261] E. Karkaletsis, C. Spyropoulos, and G. Vouros. The use of terminological knowledge bases in software localisation. Lecture Notes in Computer Science, 898:175–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayer:1995:NTT

[262] R. Mayer. Navigation through terminological databases. Lecture Notes in Computer Science, 898:189–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caroli:1995:TLC

[263] F. Caroli. Types of lexical cooccurrences: Descriptive parameters. Lecture Notes in Computer Science, 898:203–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ostler:1995:PVF

[264] N. Ostler. Perception vocabulary in five languages — towards an analysis using frame elements. Lecture Notes in Computer Science, 898:219–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heid:1995:RPM

[265] U. Heid. Relating parallel monolingual lexicon fragments for translation purposes. Lecture Notes in Computer Science, 898:231–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ewens:1995:AOB

[266] W. J. Ewens and A. Hastings. Aspects of optimality behavior in population genetics theory. Lecture Notes in Computer Science, 899:7–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hastings:1995:OTS

[267] A. Hastings and G. A. Fox. Optimization as a technique for studying population genetics equations. Lecture Notes in Computer Science, 899:18–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duchateau-Nguyen:1995:EM

[268] G. Duchateau-Nguyen, G. Weisbuch, and L. Peliti. Emergence of mutualism. Lecture Notes in Computer Science, 899:27–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bedau:1995:TIA

[269] M. A. Bedau. Three illustrations of artificial life’s working hypothesis. Lecture Notes in Computer Science, 899: 53–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Banzhaf:1995:SOA

[270] W. Banzhaf. Self-organizing algorithms derived from RNA interactions. Lecture Notes in Computer Science, 899:69–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mjolsness:1995:MCB

[271] E. Mjolsness, C. D. Garrett, J. Reinitz, and D. H. Sharp. Modeling the connection between development and evolution. Lecture Notes in Computer Science, 899:103–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voigt:1995:SGO

[272] H.-M. Voigt. Soft genetic operators in evolutionary algorithms. Lecture Notes in Computer Science, 899:123– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muehlenbein:1995:ASM

[273] H. Muehlenbein and D. Schlierkamp-Voosen. Analysis of selection, mutation and recombination in genetic algorithms. Lecture Notes in Computer Science, 899:142–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1995:RMC

[274] G. F. Miller and P. M. Todd. The role of mate choice in biocomputation: Sexual selection as a process of search, optimization and diversification. Lecture Notes in Computer Science, 899:169–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Altenberg:1995:GGE

[275] L. Altenberg. Genome growth and the evolution of the genotype-phenotype map. Lecture Notes in Computer Science, 899:205–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1995:SSI

[276] W. Thomas. On the synthesis of strategies in infinite games. Lecture Notes in Computer Science, 900:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Louchard:1995:FML

[277] G. Louchard. Finding the maximum with linear error probabilities: A sequential analysis approach. Lecture Notes in Computer Science, 900:14–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juedes:1995:CWC

[278] D. W. Juedes and J. H. Lutz. Completeness and weak completeness under polynomial-size circuits. Lecture Notes in Computer Science, 900:26–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:1995:CCK

[279] J.-Y. Cai, R. J. Lipton, L. Longpre, and M. Ogihara. Communication complexity of key agreement on small ranges. Lecture Notes in Computer Science, 900:38–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1995:PGF

[280] Y. Han and L. A. Hemaspaandra. Pseudorandom generators and the frequency of simplicity. Lecture Notes in Computer Science, 900:50–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hertrampf:1995:CBC

[281] U. Hertrampf. Classes of bounded counting type and their inclusion relations. Lecture Notes in Computer Science, 900:60–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Green:1995:LBD

[282] F. Green. Lower bounds for depththree circuits with equals and modgates. Lecture Notes in Computer Science, 900:71–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krause:1995:RIM

[283] M. Krause. On realizing iterated multiplication by small depth threshold circuits. Lecture Notes in Computer Science, 900:83–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durand:1995:RNC

[284] B. Durand. A random NP-Complete problem for inversion of 2D cellular automata. Lecture Notes in Computer Science, 900:95–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fagnot:1995:SEP

[285] I. Fagnot. On the subword equivalence problem for infinite words. Lecture Notes in Computer Science, 900: 107–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garel:1995:SIW

[286] E. Garel. On the separators on an infinite word generated by a morphism. Lecture Notes in Computer Science, 900:119–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monti:1995:STO

[287] A. Monti and A. Peron. Systolic tree omega-languages. Lecture Notes in Computer Science, 900:131–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krishnan:1995:SCO

[288] S. C. Krishnan, A. Puri, and R. K. Brayton. Structural complexity of omega-automata. Lecture Notes in Computer Science, 900:143–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Minkwitz:1995:AES

[289] T. Minkwitz. Algorithms explained by symmetries. Lecture Notes in Computer Science, 900:157–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1995:GST

[290] P. F. Fischer, F. P. Preparata, and J. E. Savage. Generalized scans and tri-diagonal systems. Lecture Notes in Computer Science, 900:168–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crochemore:1995:TDP

[291] M. Crochemore, L. Gasieniec, W. Plandowski, and W. Rytter. Twodimensional pattern matching in linear time and small space. Lecture Notes in Computer Science, 900:181–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Djidjev:1995:LDA

[292] H. N. Djidjev, G. E. Pantziou, and C. D. Zaroliagis. On-line and dynamic algorithms for shortest path problems. Lecture Notes in Computer Science, 900:193–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cadoli:1995:CRP

[293] M. Cadoli, F. M. Donini, and M. Schaerf. On compact representations of propositional circumscription. Lecture Notes in Computer Science, 900:205–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DAgostino:1995:STT

[294] G. D’Agostino, A. Montanari, and A. Policriti. A set-theoretic translation method for (poly)modal logics. Lecture Notes in Computer Science, 900:217–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maler:1995:SDC

[295] O. Maler, A. Pnueli, and J. Sifakis. On the synthesis of discrete controllers for timed systems. Lecture Notes in Computer Science, 900:229–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boreale:1995:FAS

[296] M. Boreale and D. Sangiorgi. A fully abstract semantics for causality in the π-calculus. Lecture Notes in Computer Science, 900:243–254, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hromkovic:1995:SPN

[297] J. Hromkovic, K. Lorys, P. Kanarek, and R. Klasing. On the sizes of permutation networks and consequences for efficient simulation of hypercube algorithms on bounded-degree networks. Lecture Notes in Computer Science, 900:255–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MeyeraufderHeide:1995:ESR

[298] F. Meyer auf der Heide, C. Scheideler, and V. Stemann. Exploiting storage redundancy to speed up randomized shared memory simulations. Lecture Notes in Computer Science, 900:267– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flammini:1995:IRS

[299] M. Flammini, G. Gambosi, and S. Salomone. Interval routing schemes. Lecture Notes in Computer Science, 900: 279–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MeyeraufderHeide:1995:PRP

[300] F. Meyer auf der Heide and B. Voecking. A packet routing protocol for arbitrary networks. Lecture Notes in Computer Science, 900:291–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bercoff:1995:FTS

[301] C. Bercoff. A family of tag systems for paperfolding sequences. Lecture Notes in Computer Science, 900:303– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buntrock:1995:GCS

[302] G. Buntrock and F. Otto. Growing context-sensitive languages and Church–Rosser languages. Lecture Notes in Computer Science, 900:313– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giammarresi:1995:DGA

[303] D. Giammarresi and R. Montalbano. Deterministic generalized automata.

Lecture Notes in Computer Science, 900:325–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Indyk:1995:OSA

[304] P. Indyk. Optimal simulation of automata by neural nets. Lecture Notes in Computer Science, 900:337–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1995:CPE

[305] A. R. Meyer. Concurrent process equivalences: Some decision problems. Lecture Notes in Computer Science, 900: 349–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duris:1995:OLB

[306] P. Duris and J. D. P. Rolim. Optimal lower bounds on the multiparty communication complexity. Lecture Notes in Computer Science, 900:350– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Babai:1995:SMV

[307] L. Babai, P. G. Kimmel, and S. V. Lokam. Simultaneous messages vs. communication. Lecture Notes in Computer Science, 900:361–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruyere:1995:CSC

[308] V. Bruyere and C. De Felice. Coding and strong coding in trace monoids. Lecture Notes in Computer Science, 900:373–??, 1995. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Diekert:1995:CT

[309] V. Diekert, A. Muscholl, and K. Reinhardt. On codings of traces. Lecture Notes in Computer Science, 900:385– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1995:FLC

[310] A. Gupta and N. Nishimura. Finding largest common embeddable subtrees. Lecture Notes in Computer Science, 900:397–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaller:1995:XCP

[311] D. Kaller, A. Gupta, and T. Shermer. The Xt-coloring problem. Lecture Notes in Computer Science, 900: 409–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikoletseas:1995:EPR

[312] S. E. Nikoletseas and P. G. Spirakis. Expander properties in random regular graphs with edge faults. Lecture Notes in Computer Science, 900:421– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gardy:1995:DAS

[313] D. Gardy and G. Louchard. Dynamic analysis of the sizes of relations. Lecture Notes in Computer Science, 900:433–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Raz:1995:SCF

[314] D. Raz. On slender context-free languages. Lecture Notes in Computer Science, 900:445–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Antimirov:1995:PDR

[315] V. Antimirov. Partial derivatives of regular expressions and finite automata constructions. Lecture Notes in Computer Science, 900:455–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bracho:1995:DOC

[316] F. Bracho, M. Droste, and D. Kuske. Dependence orders for computations of concurrent automata. Lecture Notes in Computer Science, 900:467–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fabret:1995:UDD

[317] A.-C. Fabret and A. Petit. On the undecidability of deadlock detection in families of nets. Lecture Notes in Computer Science, 900:479–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rueb:1995:ART

[318] C. Rueb. On the average running time of odd-even merge sort. Lecture Notes in Computer Science, 900:491– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kunde:1995:OAC

[319] M. Kunde, R. Niedermeier, K. Reinhardt, and P. Rossmanith. Optimal average case sorting on arrays. Lecture Notes in Computer Science, 900: 503–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strauss:1995:NNS

[320] M. Strauss. Normal numbers and sources for BPP. Lecture Notes in Computer Science, 900:515–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hancock:1995:LBL

[321] T. Hancock, T. Jiang, M. Li, and J. Tromp. Lower bounds on learning decision lists and trees. Lecture Notes in Computer Science, 900:527– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damaschke:1995:LSD

[322] P. Damaschke. Line segmentation of digital curves in parallel. Lecture Notes in Computer Science, 900:539– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kummer:1995:CCS

[323] M. Kummer and M. Schaefer. Computability of convex sets. Lecture Notes in Computer Science, 900:550– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ottmann:1995:EEP

[324] T. Ottmann, S. Schuierer, and S. Soundaralakshmi. Enumerating extreme points in higher dimensions. Lecture Notes in Computer Science, 900:

562–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petitjean:1995:NVP

[325] S. Petitjean. The number of views of piecewise-smooth algebraic objects. Lecture Notes in Computer Science, 900:571–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Macarie:1995:SLS

[326] I. I. Macarie. On the structure of log-space probabilistic complexity classes. Lecture Notes in Computer Science, 900:583–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fortnow:1995:RBI

[327] L. Fortnow and M. Kummer. Resourcebounded instance complexity. Lecture Notes in Computer Science, 900:597– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buhrman:1995:SSC

[328] H. Buhrman and M. Hermo. On the sparse set conjecture for sets with low density. Lecture Notes in Computer Science, 900:609–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fenner:1995:BPN

[329] S. A. Fenner and L. Fortnow. Beyond P0N0P = NEXP. Lecture Notes in Computer Science, 900:619–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jakoby:1995:MDA

[330] A. Jakoby, R. Reischuk, and C. Schindelhauer. Malign distributions for average case circuit complexity. Lecture Notes in Computer Science, 900:628– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arques:1995:PCG

[331] D. G. Arques and C. J. Michel. A possible code in the genetic code. Lecture Notes in Computer Science, 900: 640–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kropf:1995:BCH

[332] T. Kropf. Benchmark-circuits for hardware-verification. Lecture Notes in Computer Science, 901:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aagaard:1995:RAP

[333] M. Aagaard and M. Leeser. Reasoning about pipelines with structural hazards. Lecture Notes in Computer Science, 901:13–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Windley:1995:CMP

[334] P. J. Windley and M. L. Coe. A correctness model for pipelined microprocessors. Lecture Notes in Computer Science, 901:33–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OLeary:1995:NRI

[335] J. O’Leary, M. Leeser, J. Hickey, and M. Aagaard. Non-restoring integer square root: A case study in design by principled optimization. Lecture Notes in Computer Science, 901:52–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pierre:1995:AGM

[336] L. Pierre. An automatic generalization method for the inductive proof of replicated and parallel architectures. Lecture Notes in Computer Science, 901:72–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhu:1995:CCM

[337] Z. Zhu. A compositional circuit model and verification by composition. Lecture Notes in Computer Science, 901: 92–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanEijk:1995:ESS

[338] C. A. J. Van Eijk and G. L. J. M. Janssen. Exploiting structural similarities in a BDD-Based verification method. Lecture Notes in Computer Science, 901:110–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johnson:1995:SSP

[339] S. D. Johnson, P. S. Miner, and A. Camilleri. Studies of the single pulser in various reasoning systems. Lecture Notes in Computer Science, 901:126–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kishinevsky:1995:MVS

[340] M. Kishinevsky and J. Staunstrup. Mechanized verification of speedindependence. Lecture Notes in Computer Science, 901:146–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kitamichi:1995:ACP

[341] J. Kitamichi, S. Morioka, T. Higashino, and K. Taniguchi. Automatic correctness proof of the implementation of synchronous sequential circuits using an algebraic approach. Lecture Notes in Computer Science, 901:165–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maretti:1995:MVR

[342] N. Maretti. Mechanized verification of refinement. Lecture Notes in Computer Science, 901:185–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cyrluk:1995:ETP

[343] D. Cyrluk, S. Rajan, N. Shankar, and M. K. Srivas. Effective theorem proving for hardware verification. Lecture Notes in Computer Science, 901:203– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kropf:1995:FFH

[344] T. Kropf, K. Schneider, and R. Kumar. A formal framework for high level synthesis. Lecture Notes in Computer Science, 901:223–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mellergaard:1995:TDV

[345] N. Mellergaard and J. Staunstrup. Tutorial on design verification with synchronized transitions. Lecture Notes in Computer Science, 901:239–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Owre:1995:TUP

[346] S. Owre, J. M. Rushby, N. Shankar, and M. K. Srivas. A tutorial on using PVS for hardware verification. Lecture Notes in Computer Science, 901: 258–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Busch:1995:RIS

[347] H. Busch. A reduced instruction set proof environment. Lecture Notes in Computer Science, 901:280–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bezzi:1995:QEF

[348] G. Bezzi, M. Bombana, P. Cavalloro, and S. Conigliaro. Quantitative evaluation of formal based synthesis in ASIC design. Lecture Notes in Computer Science, 901:286–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Allemand:1995:FVC

[349] M. Allemand. Formal verification of characteristic properties. Lecture Notes in Computer Science, 901:292– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fisler:1995:EFR

[350] K. Fisler. Extending formal reasoning with support for hardware diagrams. Lecture Notes in Computer Science, 901:298–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asperti:1995:CLC

[351] A. Asperti and C. Laneve. Comparing lambda-calculus translations in sharing graphs. Lecture Notes in Computer Science, 902:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barthe:1995:EPT

[352] G. Barthe. Extensions of pure type systems. Lecture Notes in Computer Science, 902:16–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellucci:1995:MFP

[353] R. Bellucci, M. Abadi, and P.-L. Curien. A model for formal parametric polymorphism: a PER interpretation for System R. Lecture Notes in Computer Science, 902:32–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berardi:1995:RNI

[354] S. Berardi, M. Bezem, and T. Coquand. A realization of the negative interpretation of the axiom of choice. Lecture Notes in Computer Science, 902:47–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berardi:1995:USP

[355] S. Berardi and L. Boerio. Using subtyping in program optimization. Lecture Notes in Computer Science, 902:63–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bierman:1995:WCM

[356] G. M. Bierman. What is a categorical model of intuitionistic linear logic? Lecture Notes in Computer Science, 902:78–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Briaud:1995:EER

[357] D. Briaud. An explicit Eta rewrite rule. Lecture Notes in Computer Science, 902:94–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coscoy:1995:ETP

[358] Y. Coscoy, G. Kahn, and L. Thery. Extracting text from proofs. Lecture Notes in Computer Science, 902:109– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Despeyroux:1995:HOA

[359] J. Despeyroux, A. Felty, and A. Hirschowitz. Higher-order abstract syntax in coq. Lecture Notes in Computer Science, 902:124–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiCosmo:1995:EEP

[360] R. Di Cosmo and A. Piperno. Expanding extensional polymorphism. Lecture Notes in Computer Science, 902:139– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dowek:1995:LCC

[361] G. Dowek. Lambda-calculus, combinators and the comprehension scheme. Lecture Notes in Computer Science, 902:154–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghani:1995:EEC

[362] N. Ghani. eta-equality for coproducts. Lecture Notes in Computer Science, 902:171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goguen:1995:TOS

[363] H. Goguen. Typed operational semantics. Lecture Notes in Computer Science, 902:186–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeGroote:1995:SCE

[364] P. De Groote. A simple calculus of exception handling. Lecture Notes in Computer Science, 902:201–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofman:1995:SMQ

[365] M. Hofman. A simple model for quotient types. Lecture Notes in Computer Science, 902:216–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holmes:1995:ULC

[366] M. R. Holmes. Untyped lambdacalculus with relative typing. Lecture Notes in Computer Science, 902:235– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honsell:1995:FSU

[367] F. Honsell and M. Lenisa. Final semantics for untyped lambda-calculus. Lecture Notes in Computer Science, 902: 249–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hurkens:1995:SGP

[368] A. J. C. Hurkens. A simplification of Girard’s paradox. Lecture Notes in Computer Science, 902:266–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kondoh:1995:BPD

[369] H. Kondoh. Basic properties of data types with inequational refinements. Lecture Notes in Computer Science, 902:279–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurata:1995:DPI

[370] T. Kurata and M. Takahashi. Decidable properties of intersection type systems. Lecture Notes in Computer Science, 902:297–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leclerc:1995:TPT

[371] F. Leclerc. Termination proof of term rewriting system with the multiset path ordering. A complete development in the system coq. Lecture Notes in Computer Science, 902:312–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mellies:1995:TLC

[372] P. A. Mellies. Typed lambda-calculi with explicit substitutions may not terminate. Lecture Notes in Computer Science, 902:328–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padovani:1995:ECI

[373] V. Padovani. On equivalence classes of interpolation equations. Lecture Notes in Computer Science, 902:335– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandePol:1995:SFT

[374] J. Van de Pol and H. Schwichtenberg. Strict functionals for termination proofs. Lecture Notes in Computer Science, 902:350–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pollack:1995:VT

[375] R. Pollack. A verified typechecker. Lecture Notes in Computer Science, 902: 365–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pravato:1995:CSC

[376] A. Pravato, S. Ronchi della Rocca, and L. Roversi. Categorical semantics of the call-by-value lambda-calculus. Lecture Notes in Computer Science, 902: 381–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ritter:1995:FAT

[377] E. Ritter and A. M. Pitts. A fully abstract translation between a lambdacalculus with reference types and standard ML. Lecture Notes in Computer Science, 902:397–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simpson:1995:CCR

[378] A. K. Simpson. Categorical completeness results for the simply-typed lambda-calculus. Lecture Notes in Computer Science, 902:414–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Springintveld:1995:TOM

[379] J. Springintveld. Third-order matching in the presence of type constructors. Lecture Notes in Computer Science, 902:428–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1995:DT

[380] H. L. Bodlaender and J. Engelfriet. Domino treewidth. Lecture Notes in Computer Science, 903:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramachandramurthi:1995:LBT

[381] S. Ramachandramurthi. A lower bound for treewidth and its consequences. Lecture Notes in Computer Science, 903:14–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Garbe:1995:TWP

[382] R. Garbe. Tree-width and path-width of comparability graphs of interval orders. Lecture Notes in Computer Science, 903:26–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ebert:1995:DAG

[383] J. Ebert and A. Franzke. A declarative approach to graph based modeling. Lecture Notes in Computer Science, 903:38–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parisi-Presicce:1995:MGG

[384] F. Parisi-Presicce and G. Piersanti. Multilevel graph grammars. Lecture Notes in Computer Science, 903:51–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brandstaedt:1995:AUH

[385] A. Brandstaedt, V. D. Chepoi, and F. F. Dragan. The algorithmic use of hypertree structure and maximum neighbourhood orderings. Lecture Notes in Computer Science, 903:65–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahlhaus:1995:DEO

[386] E. Dahlhaus, P. Hammer, F. Maffray, and S. Olariu. On domination elimination orderings and domination graphs. Lecture Notes in Computer Science, 903:81–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kratochvil:1995:CGC

[387] J. Kratochvil, A. Proskurowski, and J. A. Telle. Complexity of graph covering problems. Lecture Notes in Computer Science, 903:93–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kloks:1995:D

[388] T. Kloks, D. Kratsch, and H. Mueller. Dominoes. Lecture Notes in Computer Science, 903:106–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Turau:1995:GCD

[389] V. Turau and W. Chen. GLB-Closures in directed acyclic graphs and their applications. Lecture Notes in Computer Science, 903:121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crescenzi:1995:MVC

[390] P. Crescenzi and L. Trevisan. Minimum vertex cover, distributed decisionmaking, and communication complexity. Lecture Notes in Computer Science, 903:130–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreae:1995:CPG

[391] T. Andreae, M. Noelle, and G. Schreiber. Cartesian products of graphs as spanning subgraphs of De Bruijn graphs. Lecture Notes in Computer Science, 903:140–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuerr:1995:SGT

[392] A. Schuerr. Specification of graph translators with triple graph grammars. Lecture Notes in Computer Science, 903:151–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Westfechtel:1995:UPG

[393] B. Westfechtel. Using programmed graph rewriting for the formal specification of a configuration management system. Lecture Notes in Computer Science, 903:164–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Skodinis:1995:ETA

[394] K. Skodinis and E. Wanke. Exponential time analysis of confluent and boundary eNCE graph languages. Lecture Notes in Computer Science, 903:180– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhagavathi:1995:TOT

[395] D. Bhagavathi, V. Bokka, H. Gurla, and S. Olariu. Time-optimal tree computations on sparse meshes. Lecture Notes in Computer Science, 903:193– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaudhuri:1995:PGT

[396] S. Chaudhuri and T. Hagerup. Prefix graphs and their applications. Lecture Notes in Computer Science, 903: 206–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jakoby:1995:CBP

[397] A. Jakoby, R. Reischuk, and C. Schindelhauer. The complexity of broadcasting in planar and decomposable graphs. Lecture Notes in Computer Science, 903:219–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1995:MDS

[398] Z.-Z. Chen. The maximal f-dependent set problem for planar graphs is in NC. Lecture Notes in Computer Science, 903:232–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiBattista:1995:LCP

[399] G. Di Battista, R. Tamassia, and L. Vismara. On-line convex planarity testing. Lecture Notes in Computer Science, 903:242–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shahrokhi:1995:BEC

[400] F. Shahrokhi, O. Sykora, L. A. Szekely, and I. Vrt’o. Book embeddings and crossing numbers. Lecture Notes in Computer Science, 903:256–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naumann:1995:MDS

[401] V. Naumann. Measuring the distance to series-parallelity by path expressions. Lecture Notes in Computer Science, 903:269–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golin:1995:LTP

[402] M. Golin and S. Zaks. Labelled trees and pairs of input-output permutations in priority queues. Lecture Notes in Computer Science, 903:282–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1995:RG

[403] H. L. Bodlaender, J. S. Deogun, K. Jansen, and T. Kloks. Rankings of graphs. Lecture Notes in Computer Science, 903:292–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doerr:1995:BSV

[404] H. Doerr. Bypass strong V-structures and find an isomorphic labelled subgraph in linear time. Lecture Notes in Computer Science, 903:305–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wada:1995:EAM

[405] K. Wada, A. Takaki, and K. Kawaguchi. Efficient algorithms for a mixed k-partition problem of graphs without specifying bases. Lecture Notes in Computer Science, 903:319–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dendris:1995:FSG

[406] N. D. Dendris, L. M. Kirousis, and D. M. Thilikos. Fugitive-search games on graphs and related parameters. Lecture Notes in Computer Science, 903: 331–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kajitani:1995:NAR

[407] Y. Kajitani, J. D. Cho, and M. Sarrafzadeh. New approximation results on graph matching and related problems. Lecture Notes in Computer Science, 903:343–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sieling:1995:NLB

[408] D. Sieling and I. Wegener. New lower bounds and hierarchy results for restricted branching programs. Lecture Notes in Computer Science, 903:359– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giaccio:1995:LAS

[409] R. Giaccio. On-line algorithms for satisfiability problems with uncertainty. Lecture Notes in Computer Science, 903:371–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bampis:1995:NAA

[410] E. Bampis, Y. Manoussakis, and I. Milis. NC algorithms for antidirected Hamiltonian paths and cycles in tournaments. Lecture Notes in Computer Science, 903:387–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Babel:1995:DPG

[411] L. Babel, I. Ponomarenko, and G. Tinhofer. Directed path graph isomorphism. Lecture Notes in Computer Science, 903:395–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Solomonoff:1995:DAP

[412] R. J. Solomonoff. The discovery of algorithmic probability: A guide for the programming of true creativity. Lecture Notes in Computer Science, 904:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freund:1995:DTG

[413] Y. Freund and R. E. Schapire. A decision-theoretic generalization of online learning and an application to boosting. Lecture Notes in Computer Science, 904:23–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ben-David:1995:OLV

[414] S. Ben-David, E. Kushilevitz, and Y. Mansour. Online learning versus offline learning. Lecture Notes in Computer Science, 904:38–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ben-David:1995:LDT

[415] S. Ben-David and M. Lindenbaum. Learning distributions by their density levels — a paradigm for learning without a teacher. Lecture Notes in Computer Science, 904:53–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haussler:1995:TWC

[416] D. Haussler, J. Kivinen, and M. K. Warmuth. Tight worst-case loss bounds for predicting with expert advice. Lecture Notes in Computer Science, 904:69–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamanishi:1995:LML

[417] K. Yamanishi. On-line maximum likelihood prediction with respect to general loss functions. Lecture Notes in Computer Science, 904:84–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambainis:1995:PPI

[418] A. Ambainis. The power of procrastination in inductive inference: How it depends on used ordinal notations. Lecture Notes in Computer Science, 904: 99–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balcazar:1995:LKE

[419] J. L. Balcazar, H. Buhrman, and M. Hermo. Learnability of Kolmogoroveasy circuit expressions via queries. Lecture Notes in Computer Science, 904:112–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lange:1995:TMD

[420] S. Lange and T. Zeugmann. Trading monotonicity demands versus mind changes. Lecture Notes in Computer Science, 904:125–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Case:1995:LRF

[421] J. Case, S. Kaufmann, E. Kinber, and M. Kummer. Learning recursive functions from approximations. Lecture Notes in Computer Science, 904:140– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1995:ICL

[422] R. Freivalds, E. Kinber, and C. H. Smith. On the intrinsic complexity of learning. Lecture Notes in Computer Science, 904:154–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jain:1995:SIC

[423] S. Jain and A. Sharma. The structure of intrinsic complexity of learning. Lecture Notes in Computer Science, 904:169–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1995:KNM

[424] R. Freivalds and S. Jain. Kolmogorov numberings and minimal identification. Lecture Notes in Computer Science, 904:182–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rissanen:1995:SCL

[425] J. Rissanen. Stochastic complexity in learning. Lecture Notes in Computer Science, 904:196–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anthony:1995:FLI

[426] M. Anthony and P. Bartlett. Function learning from interpolation. Lecture Notes in Computer Science, 904: 211–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gurvits:1995:ALC

[427] L. Gurvits and P. Koiran. Approximation and learning of convex superpositions. Lecture Notes in Computer Science, 904:222–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vovk:1995:MDL

[428] V. G. Vovk. Minimum description length estimators under the optimal coding scheme. Lecture Notes in Computer Science, 904:237–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kilpelaeinen:1995:MLU

[429] P. Kilpelaeinen, H. Mannila, and E. Ukkonen. MDL learning of unions of simple pattern languages from positive examples. Lecture Notes in Computer Science, 904:252–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1995:NUP

[430] E. Martin and D. Osherson. A note on the use of probabilities by mechanical learners. Lecture Notes in Computer Science, 904:261–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuurmans:1995:CRV

[431] D. Schuurmans. Characterizing rational versus exponential learning curves. Lecture Notes in Computer Science, 904:272–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muselli:1995:PAO

[432] M. Muselli. Is pocket algorithm optimal? Lecture Notes in Computer Science, 904:287–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandenBerg:1995:STC

[433] J. Van den Berg and J. C. Bioch. Some theorems concerning the free energy of (un)constrained stochastic Hopfield neural networks. Lecture Notes in Computer Science, 904:298–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ameur:1995:SBL

[434] F. Ameur. A space-bounded learning algorithm for axis-parallel rectangles. Lecture Notes in Computer Science, 904:313–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simon:1995:LDL

[435] H. U. Simon. Learning decision lists and trees with equivalence-queries. Lecture Notes in Computer Science, 904: 322–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karpinski:1995:BVD

[436] M. Karpinski and A. Macintyre. Bounding VC-dimension of neural networks: Progress and prospects. Lecture Notes in Computer Science, 904:337– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golea:1995:ACA

[437] M. Golea. Average case analysis of a learning algorithm for -DNF expressions. Lecture Notes in Computer Science, 904:342–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shamir:1995:LES

[438] E. Shamir and C. Shwartzman. Learning by extended statistical queries and its relation to PAC learning. Lecture Notes in Computer Science, 904:357– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koshiba:1995:TPL

[439] T. Koshiba. Typed pattern languages and their learnability. Lecture Notes in Computer Science, 904:367–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergadano:1995:LBA

[440] F. Bergadano and S. Varricchio. Learning behaviors of automata from shortest counterexamples. Lecture Notes in Computer Science, 904:380–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brazma:1995:LRE

[441] A. Brazma. Learning of regular expressions by pattern matching. Lecture Notes in Computer Science, 904: 392–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Domingo:1995:QCL

[442] C. Domingo and V. Lavin. The query complexity of learning some subclasses of context-free grammars. Lecture Notes in Computer Science, 904:404– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grimson:1995:EVA

[443] W. E. L. Grimson, G. J. Ettinger, S. J. White, and P. L. Gleason. Evaluating and validating an automated registration system for enhanced reality visualization in surgery. Lecture Notes in Computer Science, 905:3–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Uenohara:1995:VBO

[444] M. Uenohara and T. Kanade. Visionbased object registration for real-time image overlay. Lecture Notes in Computer Science, 905:13–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peria:1995:UPS

[445] O. Peria, L. Chevalier, A. Francois-Joubert, and J.-P. Caravel. Using a 3D position sensor for registration of SPECT and US images of the kidney. Lecture Notes in Computer Science, 905:23–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Betting:1995:NFF

[446] F. Betting, J. Feldmar, N. Ayache, and F. Devernay. A new framework for fusing stereo images with volumetric medical images. Lecture Notes in Computer Science, 905:30–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhao:1995:VMI

[447] J. Zhao, A. C. F. Colchester, C. J. Henri, and D. J. Hawkes. Visualisation of multimodal images for neurosurgical planning and guidance. Lecture Notes in Computer Science, 905:40–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henri:1995:RSD

[448] C. J. Henri, A. C. F. Colchester, J. Zhao, and D. J. Hawkes. Registration of 3D surface data for intraoperative guidance and visualization in frameless stereotactic neurosurgery. Lecture Notes in Computer Science, 905:47–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wells:1995:ASM

[449] W. M. Wells, W. E. L. Grimson, R. Kikinis, and F. A. Jolesz. Adaptive segmentation of MRI data. Lecture Notes in Computer Science, 905:59–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Serra:1995:VSE

[450] L. Serra, T. Poston, H. Ng, and P. A. Heng. Virtual space editing of tagged MRI heart data. Lecture Notes in Computer Science, 905:70–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maes:1995:CAI

[451] F. Maes, D. Vandermeulen, P. Suetens, and G. Marchal. Computer-aided interactive object delineation using an intelligent paintbrush technique. Lecture Notes in Computer Science, 905:77–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klinker:1995:BDE

[452] G. Klinker, I. Carlbom, W. Hsu, and D. Terzopoulos. Biomedical data exploration meets telecollaboration. Lecture Notes in Computer Science, 905:84–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McInerney:1995:MIS

[453] T. McInerney and D. Terzopoulos. Medical image segmentation using topologically adaptable snakes. Lecture Notes in Computer Science, 905:92–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meseure:1995:SRL

[454] P. Meseure, J.-F. Rouland, P. Dubois, and S. Karpf. SOPHOCLE: A retinal laser photocoagulation simulator: Overview. Lecture Notes in Computer Science, 905:105–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schenker:1995:NRH

[455] P. S. Schenker, H. Das, and T. R. Ohm. A new robot for high dexterity microsurgery. Lecture Notes in Computer Science, 905:115–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OToole:1995:TMC

[456] R. V. O’Toole, D. A. Simon, B. Jaramaz, and O. Ghattas. Towards more capable and less invasive robotic surgery in orthopaedics. Lecture Notes in Computer Science, 905:123– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tombropoulos:1995:TPI

[457] R. Tombropoulos, A. Schweikard, J.-C. Latombe, and J. R. Adler. Treatment planning for image-guided robotic radiosurgery. Lecture Notes in Computer Science, 905:131–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schweikard:1995:RRB

[458] A. Schweikard, R. Tombropoulos, and J. R. Adler. Robotic radiosurgery with beams of adaptable shapes. Lecture Notes in Computer Science, 905:138– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Declerck:1995:ARA

[459] J. Declerck, G. Subsol, J.-P. Thirion, and N. Ayache. Automatic retrieval of anatomical structures in 3D medical images. Lecture Notes in Computer Science, 905:153–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kancherla:1995:NVR

[460] A. R. Kancherla, J. P. Rolland, D. L. Wright, and G. Burdea. A novel virtual reality tool for teaching dynamic 3D anatomy. Lecture Notes in Computer Science, 905:163–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suzuki:1995:SSG

[461] H. Suzuki, K. Yoshizaki, M. Matsuo, and J. Kashio. A supporting system for getting tomograms and screening with a computerized 3D brain atlas and a knowledge database. Lecture Notes in Computer Science, 905:170–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mangin:1995:MBR

[462] J.-F. Mangin, J. Regis, I. Bloch, and V. Frouin. A MRF based random graph modelling the human cortical topography. Lecture Notes in Computer Science, 905:177–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bookstein:1995:CVH

[463] F. L. Bookstein. Combining “vertical” and “horizontal” features from medical images. Lecture Notes in Computer Science, 905:184–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collignon:1995:MMM

[464] A. Collignon, D. Vandermeulen, P. Suetens, and G. Marchal. 3D multimodality medical image registration using feature space clustering. Lecture Notes in Computer Science, 905:195– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacq:1995:RNS

[465] J.-J. Jacq and C. Roux. Registration of non-segmented images using a genetic algorithm. Lecture Notes in Computer Science, 905:205–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamadeh:1995:ABR

[466] A. Hamadeh, S. Lavallee, R. Szeliski, and P. Cinquin. Anatomy-based registration for computer-integrated surgery. Lecture Notes in Computer Science, 905:212–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maintz:1995:CFB

[467] J. B. A. Maintz, P. A. Van den Elsen, and M. A. Viergever. Comparison of feature-based matching of CT and MR brain images. Lecture Notes in Computer Science, 905:219–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kruggel:1995:AAA

[468] F. Kruggel. Automatical adaption of anatomical masks to the neocortex. Lecture Notes in Computer Science, 905:231–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mittelhaeusser:1995:FSB

[469] G. Mittelhaeusser and F. Kruggel. Fast segmentation of brain magnetic resonance tomograms. Lecture Notes in Computer Science, 905:237–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beil:1995:CTM

[470] W. Beil and H. S. Stiehl. Comparison of two multi-scale approaches to edge detection in medical images. Lecture Notes in Computer Science, 905:242– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Masutani:1995:CAS

[471] Y. Masutani, K. Masamune, M. Suzuki, and T. Dohi. Computer aided surgery (CAS) system for stereotactic neurosurgery. Lecture Notes in Computer Science, 905:247–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burel:1995:ROU

[472] G. Burel, H. Henocq, and J.-Y. Catros. Registration of 3D objects using linear algebra. Lecture Notes in Computer Science, 905:252–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nowinski:1995:TTS

[473] W. L. Nowinski, A. Fang, B. T. Nguyen, and R. Raghavan. Talairach-Tournoux/Schaltenbrand-Wahren based electronic brain atlas system. Lecture Notes in Computer Science, 905:257– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1995:CRM

[474] B. Kim, K. A. Frey, S. Mukhopadhyay, and B. D. Ross. Co-registration of MRI and autoradiography of rat brain in three-dimensions following automatic reconstruction of 2D data set. Lecture Notes in Computer Science, 905:262– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giachetti:1995:CAA

[475] A. Giachetti, G. Gigli, and V. Torre. Computer assisted analysis of echocardiographic image sequences. Lecture Notes in Computer Science, 905:267– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reynard:1995:CTT

[476] D. Reynard, A. Blake, A. Azzawi, and P. Styles. Computer tracking of tagged 01H MR images for motion analysis. Lecture Notes in Computer Science, 905:272–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geiger:1995:SE

[477] B. Geiger and R. Kikinis. Simulation of endoscopy. Lecture Notes in Computer Science, 905:277–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ziegler:1995:VRM

[478] R. Ziegler, W. Mueller, G. Fischer, and M. Goebel. A virtual reality medical training system. Lecture Notes in Computer Science, 905:282–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bendl:1995:VSR

[479] R. Bendl, A. Hoess, and W. Schlegel. Virtual simulation in radiotherapy planning. Lecture Notes in Computer Science, 905:287–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coatrieux:1995:MAD

[480] J.-L. Coatrieux, F. Mao, C. Toumoulin, and R. Collorec. 2D and 3D motion analysis in digital subtraction angiography. Lecture Notes in Computer Science, 905:295–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1995:MMU

[481] Y. Sato, J. Chen, S. Yamamoto, and S. Tamura. Measuring microcirculation using spatiotemporal image analysis. Lecture Notes in Computer Science, 905:302–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benayoun:1995:DNR

[482] S. Benayoun, C. Nastar, and N. Ayache. Dense non-rigid motion estimation in sequences of 3D images using differential constraints. Lecture Notes in Computer Science, 905:309– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bardinet:1995:SFF

[483] E. Bardinet, L. D. Cohen, and N. Ayache. Superquadrics and free-form deformations: A global model to fit and track 3D medical data. Lecture Notes in Computer Science, 905:319– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shi:1995:UFA

[484] P. Shi, G. Robinson, A. Chakraborty, and L. Staib. A unified framework to assess myocardial function from 4D images. Lecture Notes in Computer Science, 905:327–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruan:1995:DBA

[485] S. Ruan, C. Jaggi, J. M. Constans, and D. Bloyet. Detection of brain activation from MRI data by likelihoodratio test. Lecture Notes in Computer Science, 905:341–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vincken:1995:PHS

[486] K. L. Vincken, A. S. E. Koster, and M. A. Viergever. Probabilistic hyperstack segmentation of MR brain data. Lecture Notes in Computer Science, 905:351–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fidrich:1995:MRA

[487] M. Fidrich and J.-P. Thirion. Multiscale representation and analysis of features from medical images. Lecture Notes in Computer Science, 905:358– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Highnam:1995:RMI

[488] R. Highnam, M. Brady, and B. Shepstone. A representation for mammographic image processing. Lecture Notes in Computer Science, 905:365– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerneaz:1995:FCS

[489] N. Cerneaz and M. Brady. Finding curvilinear structures in mammograms. Lecture Notes in Computer Science, 905:372–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sullivan:1995:RCO

[490] S. Sullivan, A. Noble, and J. Ponce. On reconstructing curved object boundaries from sparse sets of X-ray images. Lecture Notes in Computer Science, 905:385–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloch:1995:RBV

[491] I. Bloch, C. Pellot, F. Sureda, and A. Herment. 3D reconstruction of blood vessels by multi-modality data fusion using fuzzy and Markovian modelling. Lecture Notes in Computer Science, 905:392–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lengyel:1995:TDR

[492] J. Lengyel, D. P. Greenberg, A. Yeung, and E. Alderman. Three-dimensional reconstruction and volume rendering of intravascular ultrasound slices imaged on a curved arterial path. Lecture Notes in Computer Science, 905: 399–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kluytmans:1995:ADM

[493] M. Kluytmans, C. J. Bouma, B. M. Ter Haar Romeny, and G. Pasterkamp. Analysis and 3D display of 30 MHz intravascular ultrasound images. Lecture Notes in Computer Science, 905:406– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gourdon:1995:SIS

[494] A. Gourdon. Simplification of irregular surface meshes in 3D medical images. Lecture Notes in Computer Science, 905:413–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Algorri:1995:DMR

[495] M.-E. Algorri and F. Schmitt. Deformable models for reconstructing unstructured 3D data. Lecture Notes in Computer Science, 905:420–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1995:SBT

[496] T. Kapur, W. E. L. Grimson, and R. Kikinis. Segmentation of brain tissue from MR images. Lecture Notes in Computer Science, 905:429–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tek:1995:SBR

[497] H. Tek and B. B. Kimia. Shockbased reaction-diffusion bubbles for image segmentation. Lecture Notes in Computer Science, 905:434–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bezy-Wendling:1995:MTA

[498] J. Bezy-Wendling, A. Bruno, and P. Reuze. MRI texture analysis applied to trabecular bone: An experimental study. Lecture Notes in Computer Science, 905:439–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benjamin:1995:OBX

[499] R. Benjamin. Object-based 3D X-ray imaging. Lecture Notes in Computer Science, 905:444–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Masutani:1995:QVS

[500] Y. Masutani, T. Kurihara, M. Suzuki, and T. Dohi. Quantitative vascular shape analysis for 3D MR-Angiography using mathematical morphology. Lecture Notes in Computer Science, 905: 449–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neveu:1995:BDS

[501] M. Neveu, D. Faudot, and B. Derdouri. B-deformable superquadrics for 3D reconstruction. Lecture Notes in Computer Science, 905:455–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bricault:1995:MOP

[502] I. Bricault and O. Monga. Modeling 3D objects with patches of quadratic surfaces: Application to the recognition and locating of anatomic structures. Lecture Notes in Computer Science, 905:460–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ishikawa:1995:MAS

[503] S. Ishikawa, H. Kosaka, K. Kato, and Y. Otsuka. A method of analyzing a shape with potential symmetry and its application to detecting spinal deformity. Lecture Notes in Computer Science, 905:465–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mellor:1995:RCC

[504] J. P. Mellor. Realtime camera calibration for enhanced reality visualization. Lecture Notes in Computer Science, 905:471–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dessenne:1995:CAK

[505] V. Dessenne, S. Lavallee, R. Julliard, and P. Cinquin. Computer assisted knee anterior cruciate ligament reconstruction: First clinical tests. Lecture Notes in Computer Science, 905:476– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karron:1995:EOV

[506] D. B. Karron and J. Cox. Extracting 3D objects from volume data using digital Morse theory. Lecture Notes in Computer Science, 905:481–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pflesser:1995:TRV

[507] B. Pflesser, U. Tiede, and K. H. Hoehne. Towards realistic visualization for surgery rehearsal. Lecture Notes in Computer Science, 905:487–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Szekely:1995:SOM

[508] G. Szekely, A. Kelemen, C. Brechbuehler, and G. Gerig. Segmentation of 3D objects from MRI volume data using constrained elastic deformations of flexible Fourier surface models. Lecture Notes in Computer Science, 905:495–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boes:1995:LDC

[509] J. L. Boes, C. R. Meyer, and T. E. Weymouth. Liver definition in CT using a population-based shape model. Lecture Notes in Computer Science, 905: 506–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1995:RCM

[510] C. R. Meyer, P. H. Bland, and J. Pipe. Retrospective correction of MRI amplitude inhomogeneities. Lecture Notes in Computer Science, 905:513–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hamitouche:1995:DNS

[511] C. Hamitouche, C. Roux, and J.- L. Coatrieux. Design of new surface detection operators in the case of an anisotropic sampling of 3D volume data. Lecture Notes in Computer Science, 905:523–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bro-Nielsen:1995:MES

[512] M. Bro-Nielsen. Modelling elasticity in solids using active cubes — application to simulated operations. Lecture Notes in Computer Science, 905: 535–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mori:1995:AEV

[513] K. Mori, J. Hasegawa, J. Toriwaki, and H. Anno. Automated extraction and visualization of bronchus from 3D CT images of lung. Lecture Notes in Computer Science, 905:542–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peuchot:1995:VRO

[514] B. Peuchot, A. Tanguy, and M. Eude. Virtual reality as an operative tool during scoliosis surgery. Lecture Notes in Computer Science, 905:549–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Edwards:1995:NGU

[515] P. J. Edwards, D. L. G. Hill, D. J. Hawkes, and R. Spink. Neurosurgical guidance using the stereo microscope. Lecture Notes in Computer Science, 905:555–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachmair:1995:CAU

[516] L. Bachmair, H. Ganzinger, and J. Stuber. Combining algebra and universal algebra in first-order theorem proving: The case of commutative rings. Lecture Notes in Computer Science, 906:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beeri:1995:BTQ

[517] C. Beeri. Bulk types and query language design. Lecture Notes in Computer Science, 906:30–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meseguer:1995:ADT

[518] J. Meseguer and N. Marti-Oliet. From abstract data types to logical frameworks. Lecture Notes in Computer Science, 906:48–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wirsing:1995:ASL

[519] M. Wirsing. Algebraic specification languages: An overview. Lecture Notes in Computer Science, 906:81–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aspinall:1995:TSA

[520] D. Aspinall. Types, subtypes, and ASL. Lecture Notes in Computer Science, 906:116–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bert:1995:OSA

[521] D. Bert and R. Echahed. On the operational semantics of the algebraic and logic programming language LPG. Lecture Notes in Computer Science, 906: 132–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1995:BT

[522] M. Bidoit and R. Hennicker. Behavioural theories. Lecture Notes in Computer Science, 906:153–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broy:1995:EDD

[523] M. Broy. Equations for describing dynamic nets of communicating systems. Lecture Notes in Computer Science, 906:170–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cerioli:1995:LAP

[524] M. Cerioli. A lazy approach to partial algebras. Lecture Notes in Computer Science, 906:188–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denker:1995:TOO

[525] G. Denker. Transactions in objectoriented specifications. Lecture Notes in Computer Science, 906:203–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrich:1995:LSD

[526] H.-D. Ehrich and A. Sernadas. Local specification of distributed families of sequential objects. Lecture Notes in Computer Science, 906:219–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1995:DAD

[527] H. Ehrig, M. Loewe, and F. Orejas. Dynamic abstract data types based on algebraic graph transformations. Lecture Notes in Computer Science, 906:236–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1995:MTT

[528] M. Fernandez and J. P. Jouannaud. Modular termination of term rewriting systems revisited. Lecture Notes in Computer Science, 906:255–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiadeiro:1995:IBS

[529] J. L. Fiadeiro and J. F. Costa. Institutions for behaviour specification. Lecture Notes in Computer Science, 906: 273–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gogolla:1995:ASO

[530] M. Gogolla and R. Herzig. An algebraic semantics for the object specification language TROLL light. Lecture Notes in Computer Science, 906:290– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hensel:1995:DET

[531] U. Hensel and H. Reichel. Defining equations in terminal coalgebras. Lecture Notes in Computer Science, 906: 307–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hintermeier:1995:SIO

[532] C. Hintermeier, C. Kirchner, and H. Kirchner. Sort inheritance for ordersorted equational presentations. Lecture Notes in Computer Science, 906: 319–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hussmann:1995:ASL

[533] H. Hussmann. Axiomatic specification of large information systems: Experiences and consequences. Lecture Notes in Computer Science, 906:336– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lechner:1995:OOA

[534] U. Lechner, C. Lengauer, and M. Wirsing. An object-oriented airport: Specification and refinement in maude. Lecture Notes in Computer Science, 906: 351–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meinke:1995:TMA

[535] K. Meinke. Topological methods for algebraic specification. Lecture Notes in Computer Science, 906:368–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mossakowski:1995:HIS

[536] T. Mossakowski. A hierarchy of institutions separated by properties of parameterized abstract data types. Lecture Notes in Computer Science, 906: 389–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parisi-Presicce:1995:DBO

[537] F. Parisi-Presicce and A. Pierantonio. Dynamical behavior of object systems. Lecture Notes in Computer Science, 906:406–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sernadas:1995:TBT

[538] A. Sernadas, C. Sernadas, and J. M. Valenca. A theory-based topological notion of institution. Lecture Notes in Computer Science, 906:420–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Underwood:1995:TAD

[539] J. Underwood. Typing abstract data types. Lecture Notes in Computer Science, 906:437–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walicki:1995:MPA

[540] M. Walicki and S. Meldal. Multialgebras, power algebras and complete calculi of identities and inclusions. Lecture Notes in Computer Science, 906:453–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolter:1995:IF

[541] U. Wolter. Institutional frames. Lecture Notes in Computer Science, 906: 469–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zucca:1995:IDS

[542] E. Zucca. Implementation of data structures in an imperative framework. Lecture Notes in Computer Science, 906:483–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gogolla:1995:WAD

[543] M. Gogolla and M. Cerioli. What is an abstract data type after all? Lecture Notes in Computer Science, 906: 499–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pratt:1995:TIS

[544] V. Pratt. Time and information in sequential and concurrent computation. Lecture Notes in Computer Science, 907:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honda:1995:NPA

[545] K. Honda. Notes on P-algebra 1: Process structure. Lecture Notes in Computer Science, 907:25–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanovich:1995:SGP

[546] M. I. Kanovich. Simulating guarded programs in linear logic. Lecture Notes in Computer Science, 907:45–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Queinnec:1995:SMO

[547] C. Queinnec. Sharing mutable objects and controlling groups of tasks in a concurrent and distributed language. Lecture Notes in Computer Science, 907:70–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawamoto:1995:MTP

[548] S.-I. Kawamoto and T. Ito. Multithreaded PaiLisp with granularity adaptive parallel execution. Lecture Notes in Computer Science, 907:94–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taura:1995:SAM

[549] K. Taura, S. Matsuoka, and A. Yonezawa. StackThreads: An abstract machine for scheduling fine-grain threads on stock CPUs. Lecture Notes in Computer Science, 907:121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kobayashi:1995:HOC

[550] N. Kobayashi and A. Yonezawa. Higher-order concurrent linear logic programming. Lecture Notes in Computer Science, 907:137–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Konaka:1995:POO

[551] H. Konaka, T. Tomokiyo, M. Maeda, and Y. Ishikawa. A parallel objectoriented language OCore. Lecture Notes in Computer Science, 907:167– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pierce:1995:COP

[552] B. C. Pierce and D. N. Turner. Concurrent objects in a process calculus. Lecture Notes in Computer Science, 907:187–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Satoh:1995:FRI

[553] I. Satoh and M. Tokoro. A formalism for remotely interacting processes. Lecture Notes in Computer Science, 907:216–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Togashi:1995:PMS

[554] A. Togashi, S. Yoshida, S. Kimura, and N. Shiratori. ProCSuS: A meta system for concurrent process calculi based on SOS. Lecture Notes in Computer Science, 907:229–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reppy:1995:FCS

[555] J. H. Reppy. First-class synchronous operations. Lecture Notes in Computer Science, 907:235–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Halstead:1995:SDF

[556] R. H. Halstead. Self-describing files + smart modules = parallel program visualization. Lecture Notes in Computer Science, 907:253–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tatemura:1995:PDP

[557] J. Tatemura, H. Koike, and H. Tanaka. A performance debugger for a parallel logic programming language fleng. Lecture Notes in Computer Science, 907: 284–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wakita:1995:FCC

[558] K. Wakita. First class continuation facilities in concurrent programming language harmony/2. Lecture Notes in Computer Science, 907:300–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jagannathan:1995:LAP

[559] S. Jagannathan. Locality abstractions for parallel and distributed computing. Lecture Notes in Computer Science, 907:320–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamada:1995:AEG

[560] T. Kamada, S. Matsuoka, and A. Yonezawa. An algorithm for efficient global garbage collection on massively parallel computers. Lecture Notes in Computer Science, 907:346–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ueda:1995:MAC

[561] K. Ueda. I/O mode analysis in concurrent logic programming. Lecture Notes in Computer Science, 907:356– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broy:1995:ACI

[562] M. Broy. Advanced component interface specification. Lecture Notes in Computer Science, 907:369–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yoshida:1995:GNC

[563] N. Yoshida. Graph notation for concurrent combinators. Lecture Notes in Computer Science, 907:393–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishimura:1995:CED

[564] S. Nishimura and A. Ohori. A calculus for exploiting data parallelism on recursively defined data. Lecture Notes in Computer Science, 907:413–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuasa:1995:DPC

[565] T. Yuasa, T. Kijima, and Y. Konishi. The data-parallel C language NCX and its implementation strategies. Lecture Notes in Computer Science, 907:433–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kusakabe:1995:DBM

[566] S. Kusakabe and M. Amamiya. A dataflow-based massively parallel programming language “V” and its implementation on a stock parallel machine. Lecture Notes in Computer Science, 907:457–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sato:1995:PDD

[567] M. Sato, Y. Kodama, H. Sakane, and Y. Yamaguchi. Programming with distributed data structure for EM-X multiprocessor. Lecture Notes in Computer Science, 907:472–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:IU

[568] Anonymous. An introduction to UNITY. Lecture Notes in Computer Science, 908:17–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:RCP

[569] Anonymous. The role of commutativity in parallel program design. Lecture Notes in Computer Science, 908:33–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:DPR

[570] Anonymous. On the design of proof rules for fair parallel programs. Lecture Notes in Computer Science, 908:59–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:RAP

[571] Anonymous. Reasoning about probabilistic parallel programs. Lecture Notes in Computer Science, 908:99–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:EDU

[572] Anonymous. Eventual determinism: Using probabilistic means to achieve deterministic ends. Lecture Notes in Computer Science, 908:149–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dershowitz:1995:ET

[573] N. Dershowitz. 33 examples of termination. Lecture Notes in Computer Science, 909:16–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lallement:1995:WPT

[574] G. Lallement. The word problem for Thue rewriting systems. Lecture Notes in Computer Science, 909:27–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matiyasevich:1995:WPT

[575] Y. Matiyasevich. Word problem for Thue systems with a few relations. Lecture Notes in Computer Science, 909: 39–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirchner:1995:SER

[576] H. Kirchner. Some extensions of rewriting. Lecture Notes in Computer Science, 909:54–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courcelle:1995:GRB

[577] B. Courcelle. Graph rewriting: A bibliographical guide. Lecture Notes in Computer Science, 909:74–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Senizergues:1995:FLW

[578] G. Senizergues. Formal languages and word-rewriting. Lecture Notes in Computer Science, 909:75–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dauchet:1995:RTA

[579] M. Dauchet. Rewriting and tree automata. Lecture Notes in Computer Science, 909:95–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertol:1995:ERA

[580] M. Bertol and V. Diekert. On efficient reduction algorithms for some trace rewriting systems. Lecture Notes in Computer Science, 909:114–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gilman:1995:AGS

[581] R. H. Gilman. Automatic groups and string rewriting. Lecture Notes in Computer Science, 909:127–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeChenadec:1995:SSC

[582] P. Le Chenadec. A survey of symmetrized and complete group presentations. Lecture Notes in Computer Science, 909:135–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marche:1995:NRA

[583] C. Marche. Normalized rewriting — application to ground completion and standard bases. Lecture Notes in Computer Science, 909:154–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lafont:1995:ERT

[584] Y. Lafont. Equational reasoning with two-dimensional diagrams. Lecture Notes in Computer Science, 909:170– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balbiani:1995:AGC

[585] P. Balbiani and L. F. Del Cerro. Affine geometry of collinearity and conditional term rewriting. Lecture Notes in Computer Science, 909:196–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rey:1995:BMW

[586] J.-F. Rey. Burnside monoids, word problem and the conjecture of brzozowski. Lecture Notes in Computer Science, 909:214–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benhamou:1995:ICL

[587] F. Benhamou. Interval constraint logic programming. Lecture Notes in Computer Science, 910:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bockmayr:1995:SPB

[588] A. Bockmayr. Solving pseudo-Boolean constraints. Lecture Notes in Computer Science, 910:22–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codognet:1995:ECS

[589] P. Codognet and G. Nardiello. Enhancing the constraint-solving power of clp(FD) by means of path-consistency methods. Lecture Notes in Computer Science, 910:39–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Comon:1995:CTA

[590] H. Comon. Constraints in term algebras. Lecture Notes in Computer Science, 910:62–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fages:1995:CNP

[591] F. Fages. Constructive negation by pruning and optimization higher-order predicates for CLP and CC languages. Lecture Notes in Computer Science, 910:68–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fruehwirth:1995:CHR

[592] T. Fruehwirth. Constraint handling rules. Lecture Notes in Computer Science, 910:90–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imbert:1995:LCS

[593] J.-L. J. Imbert. Linear constraint solving in CLP-Languages. Lecture Notes in Computer Science, 910:108– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirchner:1995:UCA

[594] H. Kirchner. On the use of constraints in automated deduction. Lecture Notes in Computer Science, 910: 128–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

LeCharlier:1995:AIF

[595] B. Le Charlier. Abstract interpretation and finite domain symbolic constraints. Lecture Notes in Computer Science, 910:147–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montanari:1995:CCC

[596] U. Montanari and F. Rossi. Concurrency and concurrent constraint programming. Lecture Notes in Computer Science, 910:171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauzy:1995:TCL

[597] A. Rauzy. Toupie: A constraint language for model checking. Lecture Notes in Computer Science, 910:193– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ridoux:1995:ICC

[598] O. Ridoux. Imagining CLP (cap delta, ). Lecture Notes in Computer Science, 910:209–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rueher:1995:ACC

[599] M. Rueher. An architecture for cooperating constraint solvers on reals. Lecture Notes in Computer Science, 910: 231–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smolka:1995:DKO

[600] G. Smolka. The definition of kernel oz. Lecture Notes in Computer Science, 910:251–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanHentenryck:1995:DIE

[601] P. Van Hentenryck, V. Saraswat, and Y. Deville. Design, implementation, and evaluation of the constraint language cc(FD). Lecture Notes in Computer Science, 910:293–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abello:1995:VGS

[602] J. Abello and K. Kumar. Visibility graphs of 2-Spiral polygons. Lecture Notes in Computer Science, 911:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alonso:1995:RGC

[603] L. Alonso and R. Schott. Random generation of colored trees. Lecture Notes in Computer Science, 911:16–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asano:1995:SFC

[604] T. Asano, D. Ranjan, T. Roos, and E. Welzl. Space filling curves and their use in the design of geometric data structures. Lecture Notes in Computer Science, 911:36–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balasubramanian:1995:TBF

[605] R. Balasubramanian, V. Raman, and G. Srinivasaraghavan. Tight bounds for finding degrees from the adjacency matrix. Lecture Notes in Computer Science, 911:49–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barrington:1995:LBM

[606] D. M. Barrington and H. Straubing. Lower bounds for modular counting by circuits with modular gates. Lecture Notes in Computer Science, 911:60–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becker:1995:RBB

[607] B. Becker, R. Drechsler, and R. Werchner. On the relation between BDDs and FDDs. Lecture Notes in Computer Science, 911:72–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloom:1995:FSA

[608] S. L. Bloom and Z. Esik. Free shuffle algebras in language varieties. Lecture Notes in Computer Science, 911:99–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bradford:1995:LBM

[609] P. G. Bradford, V. Choppella, and G. J. E. Rawlins. Lower bounds for the matrix chain ordering problem. Lecture Notes in Computer Science, 911:112–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brands:1995:LEC

[610] S. Brands. Off-line electronic cash based on secret-key certificates. Lecture Notes in Computer Science, 911: 131–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruyere:1995:RSN

[611] V. Bruyere and G. Hansel. Recognizable sets of numbers in nonstandard bases. Lecture Notes in Computer Science, 911:167–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buntrock:1995:WGC

[612] G. Buntrock and G. Niemann. On weak growing context-sensitive grammars. Lecture Notes in Computer Science, 911:180–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

CallejasBedregal:1995:LPC

[613] B. R. Callejas Bedregal and B. M. Acioly. Logic of Plotkin continuous domain. Lecture Notes in Computer Science, 911:195–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaudhuri:1995:PRR

[614] S. Chaudhuri and D. Dubhashi. (probabilistic) recurrence relations revisited. Lecture Notes in Computer Science, 911:207–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crochemore:1995:LTA

[615] M. Crochemore and W. Rytter. On linear-time alphabet-independent 2dimensional pattern matching. Lecture Notes in Computer Science, 911:220– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durand-Lose:1995:RCA

[616] J. Durand-Lose. Reversible cellular automaton able to simulate any other reversible one using partitioning automata. Lecture Notes in Computer Science, 911:230–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eades:1995:NNG

[617] P. Eades and S. Whitesides. Nearest neighbour graph realizability is NPhard. Lecture Notes in Computer Science, 911:245–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez-Baca:1995:LTA

[618] D. Fernandez-Baca and G. Slutzki. Linear-time algorithms for parametric minimum spanning tree problems on planar graphs. Lecture Notes in Computer Science, 911:257–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feuerstein:1995:PMT

[619] E. Feuerstein. Paging more than one page. Lecture Notes in Computer Science, 911:272–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

HerreradeFigueiredo:1995:ECI

[620] C. M. Herrera de Figueiredo, J. Meidanis, and C. P. De Mello. On edgecolouring indifference graphs. Lecture Notes in Computer Science, 911:286– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Galbiati:1995:ASM

[621] G. Galbiati, A. Morzenti, and F. Maffioli. On the approximability of some maximum spanning tree problems. Lecture Notes in Computer Science, 911: 300–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gao:1995:GPF

[622] S. Gao, J. Von zur Gathen, and D. Panario. Gauss periods and fast exponentiation in finite fields. Lecture Notes in Computer Science, 911:311–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasarch:1995:USR

[623] W. I. Gasarch and K. S. Guimaraes. Unbounded search and recursive graph problems. Lecture Notes in Computer Science, 911:323–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez-Vega:1995:CCG

[624] L. Gonzalez-Vega. On the complexity of computing the greatest common divisor of several univariate polynomials. Lecture Notes in Computer Science, 911:332–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gruska:1995:SCS

[625] J. Gruska, A. Monti, M. Napoli, and D. Parente. State complexity of SBTA languages. Lecture Notes in Computer Science, 911:346–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herzog:1995:PAB

[626] C. Herzog. Pushdown automata with bounded nondeterminism and bounded ambiguity. Lecture Notes in Computer Science, 911:358–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Macarie:1995:MTW

[627] I. I. Macarie. Multihead two-way probabilistic finite automata. Lecture Notes in Computer Science, 911:371– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Margenstern:1995:NET

[628] M. Margenstern. Non-erasing Turing machines: a new frontier between a decidable halting problem and universality. Lecture Notes in Computer Science, 911:386–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matamala:1995:CAN

[629] M. Matamala and E. Goles. Cyclic automata networks on finite graphs. Lecture Notes in Computer Science, 911: 398–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meidanis:1995:MAB

[630] J. Meidanis and J. C. Setubal. Multiple alignment of biological sequences with gap flexibility. Lecture Notes in Computer Science, 911:411–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meinel:1995:LBM

[631] C. Meinel and S. Waack. Lower bounds for the modular communication complexity of various graph accessibility problems. Lecture Notes in Computer Science, 911:427–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mundhenk:1995:MOM

[632] M. Mundhenk. On monotonous oracle machines. Lecture Notes in Computer Science, 911:436–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oommen:1995:ULA

[633] B. J. Oommen and E. V. De St. Croix. On using learning automata for fast graph partitioning. Lecture Notes in Computer Science, 911:449–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prodinger:1995:SPY

[634] H. Prodinger. Solution of a problem of Yekutieli and Mandelbrot. Lecture Notes in Computer Science, 911:461– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Richard:1995:RAC

[635] G. Richard and F. Saubion. A rewrite approach for constraint logic programming. Lecture Notes in Computer Science, 911:469–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roka:1995:SBC

[636] Z. Roka. Simulations between cellular automata on Cayley graphs. Lecture Notes in Computer Science, 911: 483–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1995:TLR

[637] F. Wang. A temporal logic for real-time partial-ordering with named transactions. Lecture Notes in Computer Science, 911:494–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamakawa:1995:NAR

[638] P. M. Yamakawa, H. Ebara, and H. Nakano. A new approach for routing in arrangement graphs and its performance evaluation. Lecture Notes in Computer Science, 911:509–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gebhardt:1995:RLP

[639] J. Gebhardt and R. Kruse. Reasoning and learning in probabilistic and possibilistic networks: An overview. Lecture Notes in Computer Science, 912:3–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Michie:1995:PDL

[640] D. Michie. Problem decomposition and the learning of skills. Lecture Notes in Computer Science, 912:17–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitchell:1995:MLW

[641] T. M. Mitchell. Machine learning in the World Wide Web. Lecture Notes in Computer Science, 912:32–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohlsson:1995:ACM

[642] S. Ohlsson and J. J. Jewett. Abstract computer models: Towards a new method for theorizing about adaptive agents. Lecture Notes in Computer Science, 912:33–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergmann:1995:LAP

[643] R. Bergmann and W. Wilke. Learning abstract planning cases. Lecture Notes in Computer Science, 912:55–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biberman:1995:RPE

[644] Y. Biberman. The role of prototypicality in exemplar-based learning. Lecture Notes in Computer Science, 912:77–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bostroem:1995:SRP

[645] H. Bostroem. Specialization of recursive predicates. Lecture Notes in Computer Science, 912:92–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Branke:1995:DGA

[646] J. Branke, U. Kohlmorgen, and H. Schmeck. A distributed genetic algorithm improving the generalization behavior of neural networks. Lecture Notes in Computer Science, 912:107– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dimopoulos:1995:LNM

[647] Y. Dimopoulos and A. Kakas. Learning non-monotonic logic programs: Learning exceptions. Lecture Notes in Computer Science, 912:122–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Francis:1995:CUA

[648] A. G. Francis and A. Ram. A comparative utility analysis of case-based reasoning and control-rule learning systems. Lecture Notes in Computer Science, 912:138–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gamberger:1995:MAP

[649] D. Gamberger. A minimization approach to propositional inductive learning. Lecture Notes in Computer Science, 912:151–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Griffiths:1995:CSH

[650] A. D. Griffiths and D. G. Bridge. On concept space and hypothesis space in case-based learning algorithms. Lecture Notes in Computer Science, 912: 161–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohavi:1995:PDT

[651] R. Kohavi. The power of decision tables. Lecture Notes in Computer Science, 912:174–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kubat:1995:PMD

[652] M. Kubat and D. Flotzinger. Pruning multivariate decision trees by hyperplane merging. Lecture Notes in Computer Science, 912:190–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandeMerckt:1995:MKR

[653] T. Van de Merckt and C. Decaestecker. Multiple-knowledge representations in concept learning. Lecture Notes in Computer Science, 912:200–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paliouras:1995:ENF

[654] G. Paliouras and D. S. Bree. The effect of numeric features on the scalability of inductive learning programs. Lecture Notes in Computer Science, 912: 218–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sadohara:1995:ALP

[655] K. Sadohara and M. Haraguchi. Analogical logic program synthesis from examples. Lecture Notes in Computer Science, 912:232–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tausend:1995:GTT

[656] B. Tausend. A guided tour through hypothesis spaces in ILP. Lecture Notes in Computer Science, 912:245– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ade:1995:JPT

[657] H. Ade and H. Bostroem. JIGSAW: Puzzling together RUTH and SPEC-TRE. Lecture Notes in Computer Science, 912:263–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bell:1995:DCD

[658] S. Bell and P. Brockhausen. Discovery of constraints and data dependencies in relational databases. Lecture Notes in Computer Science, 912:267–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Escazut:1995:LDN

[659] C. Escazut and P. Collard. Learning disjunctive normal forms in a dual classifier system. Lecture Notes in Computer Science, 912:271–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conroy:1995:ENE

[660] G. V. Conroy and D. M. Dutton. The effects of noise on efficient incremental induction. Lecture Notes in Computer Science, 912:275–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dovey:1995:ARP

[661] M. J. Dovey. Analysis of Rachmaninoff’s piano performances using inductive logic programming. Lecture Notes in Computer Science, 912:279– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dzeroski:1995:HRN

[662] S. Dzeroski, L. Todorovski, and T. Urbancic. Handling real numbers in inductive logic programming: A step towards better behavioural clones. Lecture Notes in Computer Science, 912: 283–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esposito:1995:SDT

[663] F. Esposito, D. Malerba, and G. Semeraro. Simplifying decision trees by pruning and grafting: New results. Lecture Notes in Computer Science, 912:287–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuernkranz:1995:TIP

[664] J. Fuernkranz. A tight integration of pruning and learning. Lecture Notes in Computer Science, 912:291–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ivanova:1995:DTB

[665] I. Ivanova and M. Kubat. Decision-tree based neural network. Lecture Notes in Computer Science, 912:295–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jorge:1995:LRI

[666] A. Jorge and P. Brazdil. Learning recursion with iterative bootstrap induction. Lecture Notes in Computer Science, 912:299–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolbe:1995:PPR

[667] T. Kolbe and C. Walther. Patching proofs for reuse. Lecture Notes in Computer Science, 912:303–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kubat:1995:ADC

[668] M. Kubat and G. Widmer. Adapting to drift in continuous domains. Lecture Notes in Computer Science, 912: 307–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Likas:1995:PRR

[669] A. Likas, K. Blekas, and A. Stafylopatis. Parallel recombinative reinforcement learning. Lecture Notes in Computer Science, 912:311–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin:1995:LSC

[670] M. Martin and U. Cortes. Learning to solve complex tasks for reactive systems. Lecture Notes in Computer Science, 912:315–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mikami:1995:COR

[671] S. Mikami, Y. Kakazu, and T. C. Fogarty. Co-operative reinforcement learning by payoff filters. Lecture Notes in Computer Science, 912:319–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1995:ASC

[672] W. Mueller and F. Wysotzki. Automatic synthesis of control programs by combination of learning and problem solving methods. Lecture Notes in Computer Science, 912:323–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikolaev:1995:ALG

[673] N. J. Nikolaev and E. N. Smirnov. Analytical learning guided by empirical technology: An approach to integration. Lecture Notes in Computer Science, 912:327–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pfahringer:1995:NMM

[674] B. Pfahringer. A new MDL measure for robust rule induction. Lecture Notes in Computer Science, 912:331–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Richeldi:1995:CDS

[675] M. Richeldi and M. Rossotto. Classdriven statistical discretization of continuous attributes. Lecture Notes in Computer Science, 912:335–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sahami:1995:GNN

[676] M. Sahami. Generating neural networks through the induction of threshold logic unit trees. Lecture Notes in Computer Science, 912:339–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sahami:1995:LCR

[677] M. Sahami. Learning classification rules using lattices. Lecture Notes in Computer Science, 912:343–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulmeister:1995:HCU

[678] B. Schulmeister and M. Bleich. Hybrid classification: Using axis-parallel and oblique subdivisions of the attribute space. Lecture Notes in Computer Science, 912:347–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sebag:1995:IBC

[679] M. Sebag, M. Schoenauer, and C. Ravise. An induction-based control for genetic algorithms. Lecture Notes in Computer Science, 912:351–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sommer:1995:FAT

[680] E. Sommer. FENDER: An approach to theory restructuring. Lecture Notes in Computer Science, 912:356–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weber:1995:LSR

[681] I. Weber, B. Tausend, and I. Stahl. Language series revisited: The complexity of hypothesis spaces in ILP. Lecture Notes in Computer Science, 912:360–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wisotzki:1995:PNN

[682] C. Wisotzki and F. Wysotzki. Prototype, nearest neighbor and hybrid algorithms for time series classification. Lecture Notes in Computer Science, 912:364–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Obbink:1995:MAS

[683] J. H. Obbink. Metrics and analysis session. Lecture Notes in Computer Science, 913:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sommerville:1995:PV

[684] I. Sommerville, G. Kotonya, S. Viller, and P. Sawyer. Process viewpoints. Lecture Notes in Computer Science, 913:2–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broeckers:1995:PBS

[685] A. Broeckers. Process-based software risk assessment. Lecture Notes in Computer Science, 913:9–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lott:1995:URM

[686] C. Lott, B. Hoisl, and H. D. Rombach. The use of roles and measurement to enact project plans in MVP-S. Lecture Notes in Computer Science, 913:30–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBunje:1995:CPM

[687] T. De Bunje and A. Saunders. Combining process models and metrics in practice. Lecture Notes in Computer Science, 913:49–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Madhavji:1995:AE

[688] N. H. Madhavji. Application experiments. Lecture Notes in Computer Science, 913:54–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strader:1995:SSO

[689] L. B. Strader, M. K. Aune, J. A. Rodgers, and M. A. Beims. Space shuttle onboard software (OBS) development and maintenance process automation. Lecture Notes in Computer Science, 913:55–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robertson:1995:PCS

[690] I. Robertson. PM case studies: A tentative characterisation. Lecture Notes in Computer Science, 913:60–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Totland:1995:SCS

[691] T. Totland and R. Conradi. A survey and comparison of some research areas relevant to software process modeling. Lecture Notes in Computer Science, 913:65–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perez:1995:CSP

[692] G. Perez, K. El Emam, and N. H. Madhavji. Customising software process models. Lecture Notes in Computer Science, 913:70–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Obbink:1995:PDI

[693] J. H. Obbink. Process differentiation and integration: the key to just-in-time in product development. Lecture Notes in Computer Science, 913:79–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montangero:1995:OID

[694] C. Montangero. Open issues in the design of PM languages. Lecture Notes in Computer Science, 913:93–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montangero:1995:FCP

[695] C. Montangero. In favour of a coherent process coding language. Lecture Notes in Computer Science, 913:94–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conradi:1995:PML

[696] R. Conradi and C. Liu. Process modelling languages: One or many? Lecture Notes in Computer Science, 913: 98–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carr:1995:EPI

[697] D. C. Carr, A. Dandekar, and D. E. Perry. Experiments in process interface descriptions, visualizations and analyses. Lecture Notes in Computer Science, 913:119–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tully:1995:SPM

[698] C. Tully. The software process and the modelling of complex systems. Lecture Notes in Computer Science, 913:138– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chroust:1995:IPM

[699] G. Chroust. Interpretable process models for software development and workflow. Lecture Notes in Computer Science, 913:144–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DiNitto:1995:IPT

[700] E. Di Nitto and A. Fuggetta. Integrating process technology and CSCW. Lecture Notes in Computer Science, 913:154–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engels:1995:DMS

[701] G. Engels. Distributed modelling session. Lecture Notes in Computer Science, 913:162–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graw:1995:PMM

[702] G. Graw and V. Gruhn. Process management in-the-many. Lecture Notes in Computer Science, 913:163–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Estublier:1995:GMV

[703] J. Estublier and N. Belkhatir. A generalized multi-view approach. Lecture Notes in Computer Science, 913:179– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nuseibeh:1995:DPM

[704] B. Nuseibeh, J. Kramer, A. Finkelstein, and U. Leonhardt. Decentralised process modelling. Lecture Notes in Computer Science, 913:185–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groenewegen:1995:CBV

[705] L. Groenewegen and G. Engels. Coordination by behavioural views and communication patterns. Lecture Notes in Computer Science, 913:189– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brinkkemper:1995:CSP

[706] S. Brinkkemper, F. Harmsen, and H. Oei. Configuration of situational process models: an information systems engineering perspective. Lecture Notes in Computer Science, 913:193– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambriola:1995:CII

[707] V. Ambriola, G. A. Cignoni, and C. Fernstroem. Current issues on integration. Lecture Notes in Computer Science, 913:197–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valetto:1995:EPT

[708] G. Valetto and G. Kaiser. Enveloping “persistent” tools for a processcentered environment. Lecture Notes in Computer Science, 913:200–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yang:1995:CPS

[709] Y. Yang. Coordination for process support is not enough! Lecture Notes in Computer Science, 913:205–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greenwood:1995:CTS

[710] R. M. Greenwood. Coordination theory and software process technology. Lecture Notes in Computer Science, 913: 209–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larsen:1995:TTP

[711] J.-O. Larsen and P. Lago. Transaction technology for process modelling. Lecture Notes in Computer Science, 913:214–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godart:1995:SSI

[712] C. Godart and D. Dietrich. Stepwise specification of interactive processes in COO. Lecture Notes in Computer Science, 913:220–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conradi:1995:SCM

[713] R. Conradi. Session on change and meta-process. Lecture Notes in Computer Science, 913:240–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sa:1995:RFS

[714] J. Sa and B. C. Warboys. A reflexive formal software process model. Lecture Notes in Computer Science, 913:

241–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaba:1995:TCP

[715] A. B. Kaba and J.-C. Derniame. Transients change processes in process centered environments. Lecture Notes in Computer Science, 913:255–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matiyasevich:1995:SML

[716] Y. Matiyasevich. On some mathematical logic contributions to rewriting techniques: Lost heritage. Lecture Notes in Computer Science, 914:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marchiori:1995:MCR

[717] M. Marchiori. Modularity of completeness revisited. Lecture Notes in Computer Science, 914:2–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steinbach:1995:ATP

[718] J. Steinbach. Automatic termination proofs with transformation orderings. Lecture Notes in Computer Science, 914:11–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lysne:1995:TOH

[719] O. Lysne and J. Piris. A termination ordering for higher order rewrite systems. Lecture Notes in Computer Science, 914:26–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zantema:1995:CCT

[720] H. Zantema and A. Geser. A complete characterization of termination O0p101 → 10rO0s. Lecture Notes in Computer Science, 914:41–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nieuwenhuis:1995:NRP

[721] R. Nieuwenhuis. On narrowing, refutation proofs and constraints. Lecture Notes in Computer Science, 914:56–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurihara:1995:CMR

[722] M. Kurihara, H. Kondo, and A. Ohuchi. Completion for multiple reduction orderings. Lecture Notes in Computer Science, 914:71–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmid:1995:TEC

[723] K. Schmid and R. Fettig. Towards an efficient construction of test sets for deciding ground reducibility. Lecture Notes in Computer Science, 914:86–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stickel:1995:TRC

[724] M. E. Stickel. Term rewriting in contemporary resolution theorem proving. Lecture Notes in Computer Science, 914:101–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asperti:1995:OOL

[725] A. Asperti. o! = 1: Optimizing optimal lambda-calculus implementations. Lecture Notes in Computer Science, 914:102–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graf:1995:STI

[726] P. Graf. Substitution tree indexing. Lecture Notes in Computer Science, 914:117–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alouini:1995:CGC

[727] I. Alouini. Concurrent garbage collection for concurrent rewriting. Lecture Notes in Computer Science, 914:132– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamperman:1995:LRE

[728] J. F. T. Kamperman and H. R. Walters. Lazy rewriting and eager machinery. Lecture Notes in Computer Science, 914:147–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anantharaman:1995:RML

[729] S. Anantharaman and G. Richard. A rewrite mechanism for logic programs with negation. Lecture Notes in Computer Science, 914:163–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Suzuki:1995:LCC

[730] T. Suzuki, A. Middeldorp, and T. Ida. Level-confluence of conditional rewrite systems with extra variables in righthand sides. Lecture Notes in Computer Science, 914:179–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Senizergues:1995:PAT

[731] G. Senizergues. A polynomial algorithm testing partial confluence of basic semi-Thue systems. Lecture Notes in Computer Science, 914:194–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gehrke:1995:PRA

[732] W. Gehrke. Problems in rewriting applied to categorical concepts by the example of a computational comonad. Lecture Notes in Computer Science, 914:210–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1995:RTC

[733] A. Corradini, F. Gadducci, and U. Montanari. Relating two categorical models of term rewriting. Lecture Notes in Computer Science, 914:225– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kahrs:1995:TDT

[734] S. Kahrs. Towards a domain theory for termination proofs. Lecture Notes in Computer Science, 914:241–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nipkow:1995:HOR

[735] T. Nipkow. Higher-order rewrite systems. Lecture Notes in Computer Science, 914:256–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kennaway:1995:ILC

[736] R. Kennaway, J. W. Klop, R. Sleep, and F.-J. De Vries. Infinitary lambda calculi and Boehm models. Lecture Notes in Computer Science, 914:257– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuper:1995:PGL

[737] J. Kuper. Proving the genericity lemma by leftmost reduction is simple. Lecture Notes in Computer Science, 914: 271–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanBakel:1995:HNT

[738] S. Van Bakel and M. Fernandez. (head- )normalization of typeable rewrite systems. Lecture Notes in Computer Science, 914:279–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lescanne:1995:ESB

[739] P. Lescanne and J. Rouyer-Degli. Explicit substitutions with de Bruijn’s levels. Lecture Notes in Computer Science, 914:294–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boulton:1995:RFH

[740] R. J. Boulton. A restricted form of higher-order rewriting applied to an HDL semantics. Lecture Notes in Computer Science, 914:309–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walters:1995:RSI

[741] H. R. Walters and H. Zantema. Rewrite systems for integer arithmetic. Lecture Notes in Computer Science, 914:324– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdulrab:1995:GSS

[742] H. Abdulrab and M. Maksimenko. General solution of systems of linear Diophantine equations and inequations. Lecture Notes in Computer Science, 914:339–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baader:1995:CCS

[743] F. Baader and K. U. Schulz. Combination of constraint solving techniques: An algebraic point of view. Lecture Notes in Computer Science, 914:352– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otto:1995:SIR

[744] F. Otto, P. Narendran, and D. J. Dougherty. Some independence results for equational unification. Lecture Notes in Computer Science, 914: 367–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burghardt:1995:RSS

[745] J. Burghardt. Regular substitution sets: A means of controlling Eunification. Lecture Notes in Computer Science, 914:382–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Avenhaus:1995:DSD

[746] J. Avenhaus, J. Denzinger, and M. Fuchs. DISCOUNT: A system for distributed equational deduction. Lecture Notes in Computer Science, 914: 397–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellegarde:1995:ATF

[747] F. Bellegarde. ASTRE: Towards a fully automated program transformation system. Lecture Notes in Computer Science, 914:403–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Buendgen:1995:PRP

[748] R. Buendgen, M. Goebel, and W. Kuechlin. Parallel ReDux → PaRe-DuX. Lecture Notes in Computer Science, 914:408–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1995:SMO

[749] T. Chen and S. Anantharaman. STORM: A many-to-one associativecommutative matcher. Lecture Notes in Computer Science, 914:414–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chazarain:1995:LSA

[750] J. Chazarain and S. Muller. LEMMA: a system for automated synthesis of recursive programs in equational theories. Lecture Notes in Computer Science, 914:420–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giesl:1995:GPO

[751] J. Giesl. Generating polynomial orderings for termination proofs. Lecture Notes in Computer Science, 914:426– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holmes:1995:DRC

[752] M. R. Holmes. Disguising recursively chained rewrite rules as equational theorems, as implemented in the prover EFTTP mark 2. Lecture Notes in Computer Science, 914:432–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kirchner:1995:PCC

[753] H. Kirchner and P.-E. Moreau. Prototyping completion with constraints using computational systems. Lecture Notes in Computer Science, 914:438– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paccanaro:1995:GTR

[754] A. Paccanaro. Guiding term reduction through A neural network: Some preliminary results for the group theory. Lecture Notes in Computer Science, 914:444–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stickel:1995:SQI

[755] M. E. Stickel and H. Zhang. Studying quasigroup identities by rewriting techniques: Problems and first results. Lecture Notes in Computer Science, 914:450–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dershowitz:1995:PRI

[756] N. Dershowitz, J.-P. Jouannaud, and J. W. Klop. Problems in rewriting III. Lecture Notes in Computer Science, 914:457–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrig:1995:DTA

[757] H. Ehrig and B. Mahr. A decade of TAPSOFT: Aspects of progress and prospects in theory and practice of software development. Lecture Notes in Computer Science, 915:3–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Floyd:1995:TPS

[758] C. Floyd. Theory and practice of software development — stages in a debate. Lecture Notes in Computer Science, 915:25–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kozen:1995:RSS

[759] D. Kozen. Rational spaces and set constraints. Lecture Notes in Computer Science, 915:42–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goguen:1995:FMS

[760] J. A. Goguen and Luqi. Formal methods and social context in software development. Lecture Notes in Computer Science, 915:62–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaudel:1995:TCF

[761] M.-C Gaudel. Testing can be formal, too. Lecture Notes in Computer Science, 915:82–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pratt:1995:APB

[762] V. Pratt. Anatomy of the Pentium bug. Lecture Notes in Computer Science, 915:97–107, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL ftp://boole.stanford.edu/
pub/FDIV/anapent.ps.gz.

Pratt:1995:RMN

[763] V. Pratt. Rational mechanics and natural mathematics. Lecture Notes in Computer Science, 915:108–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Potthoff:1995:FOL

[764] A. Potthoff. First-order logic on finite trees. Lecture Notes in Computer Science, 915:125–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salomaa:1995:DED

[765] K. Salomaa. Decidability of equivalence for deterministic synchronized tree automata. Lecture Notes in Computer Science, 915:140–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andre:1995:EPL

[766] Y. Andre and F. Bossut. The equivalence problem for letter-to-letter bottom-up tree transducers in solvable. Lecture Notes in Computer Science, 915:155–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sangiorgi:1995:SCB

[767] D. Sangiorgi. I: A symmetric calculus based on internal mobility. Lecture Notes in Computer Science, 915: 172–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lin:1995:CIS

[768] H. Lin. Complete inference systems for weak bisimulation equivalences in the π-calculus. Lecture Notes in Computer Science, 915:187–201, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Amadio:1995:RAH

[769] R. M. Amadio and M. Dam. Reasoning about higher-order processes. Lecture Notes in Computer Science, 915: 202–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1995:CPS

[770] X. Liu and D. Walker. Confluence of processes and systems of objects. Lecture Notes in Computer Science, 915:217–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VonKarger:1995:AAT

[771] B. Von Karger. An algebraic approach to temporal logic. Lecture Notes in Computer Science, 915:232–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmann:1995:BAB

[772] M. Hofmann and D. Sannella. On behavioural abstraction and behavioural satisfaction in higher-order logic. Lecture Notes in Computer Science, 915: 247–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonsson:1995:AGS

[773] B. Jonsson and Y.-K. Tsay. Assumption/guarantee specifications in lineartime temporal logic. Lecture Notes in Computer Science, 915:262–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Selivanov:1995:FHR

[774] V. Selivanov. Fine hierarchy of regular omega-languages. Lecture Notes in Computer Science, 915:277–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wilke:1995:CWD

[775] T. Wilke and H. Yoo. Computing the Wadge degree, the Lifschitz degree, and the Rabin index of a regular language of infinite words in polynomial time. Lecture Notes in Computer Science, 915:288–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wacrenier:1995:STM

[776] P.-A. Wacrenier. Semi-trace morphisms and rational transductions. Lecture Notes in Computer Science, 915: 303–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bloom:1995:NAS

[777] S. L. Bloom and Z. Esik. Nonfinite axiomatizability of shuffle inequalities. Lecture Notes in Computer Science, 915:318–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sassone:1995:CPN

[778] V. Sassone. On the category of Petri net computations. Lecture Notes in Computer Science, 915:334–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jancar:1995:HUW

[779] P. Jancar. High undecidability of weak bisimilarity for Petri nets. Lecture Notes in Computer Science, 915:349– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badouel:1995:PAS

[780] E. Badouel, L. Bernardinello, and P. Darondeau. Polynomial algorithms for the synthesis of bounded nets. Lecture Notes in Computer Science, 915: 364–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

KrishnaRao:1995:SCH

[781] M. R. K. Krishna Rao. Semicompleteness of hierarchical and superhierarchical combinations of term rewriting systems. Lecture Notes in Computer Science, 915:379–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okui:1995:LNS

[782] S. Okui, A. Middeldorp, and T. Ida. Lazy narrowing: Strong completeness and eager variable elimination. Lecture Notes in Computer Science, 915: 394–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1995:EPA

[783] A. Wagner. On the expressive power of algebraic graph grammars with application conditions. Lecture Notes in Computer Science, 915:409–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Walicki:1995:GMO

[784] M. Walicki and S. Meldal. Generated models and the omega-rule: The nondeterministic case. Lecture Notes in Computer Science, 915:424–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aceto:1995:CMC

[785] L. Aceto and A. Ingolfsdottir. CPO models for a class of GSOS languages. Lecture Notes in Computer Science, 915:439–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peron:1995:STS

[786] A. Peron. Statecharts, transition structures and transformations. Lecture Notes in Computer Science, 915:454– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abadi:1995:IOC

[787] M. Abadi and L. Cardelli. An imperative object calculus. Lecture Notes in Computer Science, 915:471–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karali:1995:RIE

[788] I. Karali and C. Halatsis. A refinement of import/export declarations in modular logic programming and its semantics. Lecture Notes in Computer Science, 915:486–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Solberg:1995:STA

[789] K. L. Solberg. Strictness and totality analysis with conjunction. Lecture Notes in Computer Science, 915:501– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tip:1995:GTS

[790] F. Tip. Generic techniques for sourcelevel debugging and dynamic program slicing. Lecture Notes in Computer Science, 915:516–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertot:1995:RES

[791] Y. Bertot and R. Fraer. Reasoning with executable specifications. Lecture Notes in Computer Science, 915:531– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kieburtz:1995:CSG

[792] R. B. Kieburtz, F. Bellegarde, J. Bell, and J. Hook. Calculating software generators from solution specifications. Lecture Notes in Computer Science, 915:546–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palsberg:1995:CFB

[793] J. Palsberg. Comparing flow-based binding-time analyses. Lecture Notes in Computer Science, 915:561–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oerbaek:1995:CYT

[794] P. Oerbaek. Can you trust your data? Lecture Notes in Computer Science, 915:575–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nielson:1995:SDP

[795] H. R. Nielson and F. Nielson. Static and dynamic processor allocation for higher-order concurrent languages. Lecture Notes in Computer Science, 915:590–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geser:1995:MIP

[796] A. Geser. Mechanized inductive proof of properties of a simple code optimizer. Lecture Notes in Computer Science, 915:605–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1995:DSA

[797] H. R. Andersen and M. Mendler. Describing a signal analyzer in the process algebra PMC. Lecture Notes in Computer Science, 915:620–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kleuker:1995:GIS

[798] S. Kleuker. A gentle introduction to specification engineering using a case study in telecommunications. Lecture Notes in Computer Science, 915:636– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sagiv:1995:PID

[799] M. Sagiv, T. Reps, and S. Horwitz. Precise interprocedural dataflow analysis with applications to constant propagation. Lecture Notes in Computer Science, 915:651–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blazy:1995:FSP

[800] S. Blazy and P. Facon. Formal specification and prototyping of a program specializer. Lecture Notes in Computer Science, 915:666–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sands:1995:PCR

[801] D. Sands. Proving the correctness of recursion-based automatic program transformations. Lecture Notes in Computer Science, 915:681–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lano:1995:RSS

[802] K. Lano. Reactive system specification and refinement. Lecture Notes in Computer Science, 915:696–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bareau:1995:MCR

[803] C. Bareau, B. Caillaud, C. Jard, and R. Thoraval. Measuring concurrency of regular distributed computations. Lecture Notes in Computer Science, 915:711–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreau:1995:NSU

[804] L. Moreau. Non-speculative and upward invocation of continuations in a parallel language. Lecture Notes in Computer Science, 915:726–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bert:1995:MIS

[805] D. Bert and C. Oriat. A model inference system for generic specification with application to code sharing. Lecture Notes in Computer Science, 915:741–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumeister:1995:RAD

[806] H. Baumeister. Relations as abstract datatypes: An institution to specify relations between algebras. Lecture Notes in Computer Science, 915:756– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schieferdecker:1995:POF

[807] I. Schieferdecker. Performance-oriented formal specifications — the LotoTis approach. Lecture Notes in Computer Science, 915:772–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:SFD

[808] Anonymous. Signal: A formal design environment for real-time systems. Lecture Notes in Computer Science, 915:789–790, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steffen:1995:MFE

[809] B. Steffen, T. Margaria, and A. Classen. The META-Frame: An environment for flexible tool management. Lecture Notes in Computer Science, 915:791– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manna:1995:SST

[810] Z. Manna, N. Bjoerner, A. Browne, and E. Chang. STeP: The Stanford Temporal Prover. Lecture Notes in Computer Science, 915:793–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1995:HUV

[811] F. Andersen, U. Binau, K. Nyblad, and K. D. Petersen. The HOL-UNITY verification system. Lecture Notes in Computer Science, 915:795–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sampaio:1995:PTA

[812] A. J. Sampaio, A. M. Haeberer, C. T. Prates, and C. D. Ururahy. PLATO: A tool to assist programming as term rewriting and theorem proving. Lecture Notes in Computer Science, 915: 797–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marre:1995:LTA

[813] B. Marre. LOFT: A tool for assisting selection of test data sets from algebraic specifications. Lecture Notes in Computer Science, 915:799–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Astesiano:1995:ST

[814] E. Astesiano, G. Reggio, and F. Morando. The SMoLCS ToolSet. Lecture Notes in Computer Science, 915: 801–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VandenBrand:1995:AME

[815] M. G. J. Van den Brand and E. Visser. The ASF+SDF meta-environment: Documentation tools for free! Lecture Notes in Computer Science, 915: 803–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Core:1995:BTD

[816] B. Core. The B-toolkit demonstration. Lecture Notes in Computer Science, 915:805–806, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guedes:1995:OOS

[817] L. C. C. Guedes, E. H. Haeusler, and J. L. Rangel. Object oriented semantics directed compiler generation: A prototype. Lecture Notes in Computer Science, 915:807–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adam:1995:SKI

[818] N. R. Adam, B. S. Fordham, and Y. Yesha. Some key issues in database systems in a digital library setting. Lecture Notes in Computer Science, 916: 9–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adam:1995:PRD

[819] N. R. Adam, M. Halem, and S. Naqvi. Promising research direction in digital libraries. Lecture Notes in Computer Science, 916:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lesk:1995:WWF

[820] M. Lesk. Which way to the future? the control of scholarly publication. Lecture Notes in Computer Science, 916:33–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Slonim:1995:NIS

[821] J. Slonim and L. BaronLaurie. Networked information systems as digital libraries. Lecture Notes in Computer Science, 916:51–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Myka:1995:AHC

[822] A. Myka and U. Guentzer. Automatic hypertext conversion of paper document collections. Lecture Notes in Computer Science, 916:65–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boehm:1995:ASD

[823] K. Boehm, K. Aberer, and E. Neuhold. Administering structured documents in digital libraries. Lecture Notes in Computer Science, 916:91–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srihari:1995:DRD

[824] S. N. Srihari, S. W. Lam, and J. J. Hull. Document recognition for a digital library. Lecture Notes in Computer Science, 916:119–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rus:1995:UNT

[825] D. Rus and K. Summers. Using nontextual cues for electronic document browsing. Lecture Notes in Computer Science, 916:129–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Futrelle:1995:CLE

[826] R. P. Futrelle, X. Zhang, and Y. Sekiya. Corpus linguistics for establishing the natural language content of digital library documents. Lecture Notes in Computer Science, 916:165–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Witten:1995:CFT

[827] I. H. Witten, A. Moffat, and T. C. Bell. Compression and full-text indexing for digital libraries. Lecture Notes in Computer Science, 916:181–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sloane:1995:DLH

[828] A. Sloane. The digital library and the home-based user. Lecture Notes in Computer Science, 916:203–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smotroff:1995:INL

[829] I. Smotroff, L. Hirschman, and S. Bayer. Integrating natural language with large dataspace visualization. Lecture Notes in Computer Science, 916: 209–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fayyad:1995:AAC

[830] U. M. Fayyad and P. Smyth. The automated analysis, cataloging and searching of digital image libraries: A machine learning approach. Lecture Notes in Computer Science, 916:225– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1995:VDS

[831] H. J. Zhang, S. W. Smoliar, Jian Hua Wu, and Chien Yong Low. A video database system for digital libraries. Lecture Notes in Computer Science, 916:253–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stackpole:1995:DST

[832] L. E. Stackpole, R. D. Atkinson, and J. Yokley. Developing the scientifictechnical digital library at a national laboratory. Lecture Notes in Computer Science, 916:265–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bhargava:1995:DRE

[833] B. Bhargava, M. Annamalai, S. Goel, and S. Li. DL-Raid: An environment for supporting digital library services. Lecture Notes in Computer Science, 916:281–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurosawa:1995:LEA

[834] Kaoru Kurosawa, Koji Okada, and Shigeo Tsujii. Low exponent attack against elliptic curve RSA. Lecture Notes in Computer Science, 917:376– 383, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beth:1995:MST

[835] T. Beth. Multifeature security through homomorphic encryption. Lecture Notes in Computer Science, 917:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Desmedt:1995:MNA

[836] Y. Desmedt, G. Di Crescenzo, and M. Burmester. Multiplicative non-Abelian sharing schemes and their application to threshold cryptography. Lecture Notes in Computer Science, 917:21–32, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okada:1995:LBS

[837] K. Okada and K. Kurosawa. Lower bound on the size of shares of nonperfect secret sharing schemes. Lecture Notes in Computer Science, 917:33–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jackson:1995:SMS

[838] W.-A. Jackson, K. M. Martin, and C. M. O’Keefe. On sharing many secrets. Lecture Notes in Computer Science, 917:42–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurosawa:1995:CIS

[839] K. Kurosawa and K. Okada. Combinatorial interpretation of secret sharing schemes. Lecture Notes in Computer Science, 917:55–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mihaljevic:1995:CAB

[840] M. J. Mihaljevic. A correlation attack on the binary sequence generators with time-varying output function. Lecture Notes in Computer Science, 917:67–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fuster-Sabater:1995:LCN

[841] A. Fuster-Sabater and P. Caballero-Gil. On the linear complexity of nonlinearly filtered PN-sequences. Lecture Notes in Computer Science, 917:80–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golic:1995:ISW

[842] J. D. Golic. Intrinsic statistical weakness of keystream generators. Lecture Notes in Computer Science, 917:91–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chee:1995:SBF

[843] S. Chee, S. Lee, and K. Kim. Semi-bent functions. Lecture Notes in Computer Science, 917:107–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seberry:1995:SCF

[844] J. Seberry, X.-M. Zhang, and Y. Zheng. Structures of cryptographic functions with strong avalanche characteristics. Lecture Notes in Computer Science, 917:119–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meadows:1995:FVC

[845] C. A. Meadows. Formal verification of cryptographic protocols: a survey. Lecture Notes in Computer Science, 917:133–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yacobi:1995:EEM

[846] Y. Yacobi. Efficient electronic money. Lecture Notes in Computer Science, 917:153–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Niemi:1995:HPB

[847] V. Niemi and A. Renvall. How to prevent buying of votes in computer elections. Lecture Notes in Computer Science, 917:164–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyd:1995:DAK

[848] C. Boyd and W. Mao. Design and analysis of key exchange protocols via secure channel identification. Lecture Notes in Computer Science, 917:171– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeSantis:1995:ZKP

[849] A. De Santis, T. Okamoto, and G. Persiano. Zero-knowledge proofs of computational power in the shared string model. Lecture Notes in Computer Science, 917:182–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Imai:1995:ISA

[850] H. Imai. Information security aspects of spread spectrum systems. Lecture Notes in Computer Science, 917:193– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Safavi-Naini:1995:CSC

[851] R. Safavi-Naini and L. Tombak. Combinatorial structure of A-codes with r-fold security. Lecture Notes in Computer Science, 917:211–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Horster:1995:MMR

[852] P. Horster, M. Michels, and H. Petersen. Meta-message recovery and meta-blind signature schemes based on the discrete logarithm problem and their applications. Lecture Notes in Computer Science, 917:224–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alabbadi:1995:DSS

[853] M. Alabbadi and S. B. Wicker. A digital signature scheme based on linear error-correcting block codes. Lecture Notes in Computer Science, 917:238– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beguin:1995:SAD

[854] P. Beguin and J.-J. Quisquater. Secure acceleration of DSS signatures using insecure server. Lecture Notes in Computer Science, 917:249–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Atkins:1995:MWS

[855] D. Atkins, M. Graff, A. K. Lenstra, and P. C. Leyland. The magic words are squeamish ossifrage. Lecture Notes in Computer Science, 917:263–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biham:1995:CMM

[856] E. Biham. Cryptanalysis of multiple modes of operation. Lecture Notes in Computer Science, 917:278–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tokita:1995:LCL

[857] T. Tokita, T. Sorimachi, and M. Matsui. Linear cryptanalysis of LOKI and s02DES. Lecture Notes in Computer Science, 917:293–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patarin:1995:CID

[858] J. Patarin. Collisions and inversions for Damgaard’s whole hash function. Lecture Notes in Computer Science, 917:307–321, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Charnes:1995:ASH

[859] C. Charnes and J. Pieprzyk. Attacking the SL2 hashing scheme. Lecture Notes in Computer Science, 917:322– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurosawa:1995:SCK

[860] K. Kurosawa, K. Okada, and K. Sakano. Security of the center in key distribution schemes. Lecture Notes in Computer Science, 917:333–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matsumoto:1995:ISK

[861] T. Matsumoto. Incidence structures for key sharing. Lecture Notes in Computer Science, 917:342–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1995:PKC

[862] P. Smith and C. Skinner. A publickey cryptosystem and a digital signature system based on the Lucas function analogue to discrete logarithms. Lecture Notes in Computer Science, 917:357–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abe:1995:HRN

[863] M. Abe and H. Morita. Higher radix nonrestoring modular multiplication algorithm and public-key LSI architecture with limited hardware resources. Lecture Notes in Computer Science, 917:365–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

OConnor:1995:UMA

[864] L. O’Connor and J. D. Golic. A unified Markov approach to differential and linear cryptanalysis. Lecture Notes in Computer Science, 917:387–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biham:1995:HSU

[865] E. Biham and A. Biryukov. How to strengthen DES using existing hardware. Lecture Notes in Computer Science, 917:398–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chang:1995:SCP

[866] X. Chang, Z.-D. Dai, and G. Gong. Some cryptographic properties of exponential functions. Lecture Notes in Computer Science, 917:415–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beaver:1995:FDS

[867] D. Beaver. Factoring: the DNA solution. Lecture Notes in Computer Science, 917:419–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stern:1995:COD

[868] J. Stern. Can one design a signature scheme based on error-correcting codes? Lecture Notes in Computer Science, 917:424–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carter:1995:DVD

[869] G. Carter, A. Clark, and L. Nielsen. DESV-1: a variation of the data encryption standard (DES). Lecture Notes in Computer Science, 917:427– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bibel:1995:ITP

[870] W. Bibel, S. Bruening, U. Egly, and D. Korn. Issues in theorem proving based on the connection method. Lecture Notes in Computer Science, 918: 1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeKogel:1995:RUS

[871] E. De Kogel. Rigid E-unification simplified. Lecture Notes in Computer Science, 918:17–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klingenbeck:1995:GFC

[872] S. Klingenbeck. Generating finite counter examples with semantic tableaux. Lecture Notes in Computer Science, 918:31–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neumann:1995:STI

[873] I. Neumann. Semantic tableaus for inheritance nets. Lecture Notes in Computer Science, 918:47–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Demri:1995:UCM

[874] S. Demri. Using connection method in modal logics: Some advantages. Lecture Notes in Computer Science, 918: 63–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Governatori:1995:LTM

[875] G. Governatori. Labelled tableaux for multi-modal logics. Lecture Notes in Computer Science, 918:79–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miglioli:1995:RSP

[876] P. Miglioli, U. Moscato, and M. Ornaghi. Refutation systems for propositional modal logics. Lecture Notes in Computer Science, 918:95–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmitt:1995:TIM

[877] S. Schmitt and C. Kreitz. On transforming intuitionistic matrix proofs into standard-sequent proofs. Lecture Notes in Computer Science, 918:106– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otten:1995:CBP

[878] J. Otten and C. Kreitz. A connection based proof method for intuitionistic logic. Lecture Notes in Computer Science, 918:122–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Underwood:1995:TIP

[879] J. Underwood. Tableau for intuitionistic predicate logic as metatheory. Lecture Notes in Computer Science, 918: 138–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caferra:1995:MBI

[880] R. Caferra and N. Peltier. Model building and interactive theory discovery. Lecture Notes in Computer Science, 918:154–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayr:1995:LDM

[881] K. Mayr. Link deletion in model elimination. Lecture Notes in Computer Science, 918:169–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Neugebauer:1995:SIR

[882] G. Neugebauer and U. Petermann. Specifications of inference rules and their automatic translation. Lecture Notes in Computer Science, 918:185– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baumgartner:1995:CME

[883] P. Baumgartner and F. Stolzenburg. Constraint model elimination and a PTTP-Implementation. Lecture Notes in Computer Science, 918:201–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaz:1995:NES

[884] M. Baaz and C. G. Fermueller. Nonelementary speedups between different versions of tableaux. Lecture Notes in Computer Science, 918:217–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Komara:1995:SRP

[885] J. Komara and P. J. Voda. Syntactic reduction of predicate tableaux to propositional tableaux. Lecture Notes in Computer Science, 918:231–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hudelmaier:1995:CLL

[886] J. Hudelmaier and P. Schroeder-Heister. Classical lambek logic. Lecture Notes in Computer Science, 918: 247–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeGroote:1995:LLI

[887] P. De Groote. Linear logic with isabelle: Pruning the proof search tree. Lecture Notes in Computer Science, 918:263–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meyer:1995:LAT

[888] R. K. Meyer, M. A. McRobbie, and N. D. Belnap. Linear analytic tableaux. Lecture Notes in Computer Science, 918:278–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohlhase:1995:HOT

[889] M. Kohlhase. Higher-order tableaux. Lecture Notes in Computer Science, 918:294–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heuerding:1995:PLC

[890] A. Heuerding, G. Jaeger, S. Schwendimann, and M. Seyfried. Propositional logics on the computer. Lecture Notes in Computer Science, 918:310– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pitt:1995:MYA

[891] J. Pitt. MacKE: Yet another proof assistant and automated pedagogic tool. Lecture Notes in Computer Science, 918:324–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schumann:1995:UTP

[892] J. Schumann. Using the theorem prover SETHEO for verifying the development of a communication protocol in FOCUS. Lecture Notes in Computer Science, 918:338–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delsanto:1995:EMP

[893] P. P. Delsanto, S. Biancotto, M. Scalerandi, and M. Rebaudengo. Exploiting massively parallel architectures for the solution of diffusion and propagation problems. Lecture Notes in Computer Science, 919:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bacci:1995:SEP

[894] B. Bacci, M. Danelutto, S. Pelagatti, and M. Vanneschi. Summarising an experiment in parallel programming language design. Lecture Notes in Computer Science, 919:7–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brezany:1995:LCP

[895] P. Brezany, T. A. Mueck, and E. Schikuta. Language, compiler and parallel database support for I/O intensive applications. Lecture Notes in Computer Science, 919:14–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Back:1995:UOE

[896] A. Back and S. Turner. Using optimistic execution techniques as a parallelisation tool for general purpose computing. Lecture Notes in Computer Science, 919:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guenther:1995:SMT

[897] H. Guenther, F. Brockners, and T. Bemmerl. System management tools for SHPC systems — partition management. Lecture Notes in Computer Science, 919:27–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mao:1995:MOM

[898] W. Mao. Multi-operation multimachine scheduling. Lecture Notes in Computer Science, 919:33–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corno:1995:PTA

[899] F. Corno, P. Prinetto, M. Rebaudeng, and M. Sonza Reorda. A PVM tool for automatic test generation on parallel and distributed systems. Lecture Notes in Computer Science, 919:39–44, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tolmie:1995:GLI

[900] D. Tolmie. Gigabit LAN issues — HIPPI, fibre channel, or ATM? Lecture Notes in Computer Science, 919:45–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:EMC

[901] Anonymous. The European meta computing utilizing integrated broadband communications (E=MC02) project. Lecture Notes in Computer Science, 919:54–59, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Manzoni:1995:OHP

[902] P. Manzoni. Obtaining high performance data transmission in the Internet. Lecture Notes in Computer Science, 919:60–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Clematis:1995:PPH

[903] A. Clematis, B. Falcidieno, D. Fernandez Prieto, and M. Spagnuolo. Parallel processing on heterogeneous networks for GIS applications. Lecture Notes in Computer Science, 919:67–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srinivasan:1995:BLR

[904] V. Srinivasan, A. Ghanwani, and E. Gelenbe. Block loss reduction in ATM networks. Lecture Notes in Computer Science, 919:73–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pozzetti:1995:CRD

[905] E. Pozzetti, G. Serazzi, V. Vetland, and J. Rolia. Characterizing the resource demands of TCP/IP. Lecture Notes in Computer Science, 919:79–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strumpen:1995:ICL

[906] V. Strumpen and T. L. Casavant. Implementing communication latency hiding in high-latency computer networks. Lecture Notes in Computer Science, 919:86–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smirni:1995:MSS

[907] E. Smirni and E. Rosti. Modeling speedup of SPMD applications on the Intel Paragon. Lecture Notes in Computer Science, 919:94–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calzarossa:1995:HAW

[908] M. Calzarossa, A. Merlo, D. Tessera, and G. Haring. A hierarchical approach to workload characterization for parallel systems. Lecture Notes in Computer Science, 919:102–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marenzoni:1995:PAC

[909] P. Marenzoni. Performance analysis of Cray T3D and Connection Machine CM-5: a comparison. Lecture Notes in Computer Science, 919:110–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boryczko:1995:NIC

[910] K. Boryczko, J. Kitowski, J. Moscinski, and A. Leszczynski. Numerically intensive computing as a benchmark for parallel computer architectures. Lecture Notes in Computer Science, 919:118–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piccardi:1995:PPE

[911] M. Piccardi, T. Salmon Cinotti, and S. Zuffi. A preliminary performance evaluation of the quadrics architecture with the DARPA image understanding benchmark. Lecture Notes in Computer Science, 919:124–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brachini:1995:IAP

[912] A. Brachini, A. Marconi, M. R. Nazzarelli, and S. Sabina. An integrated approach to performance and testing analysis for parallel systems. Lecture Notes in Computer Science, 919:130– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roest:1995:SPF

[913] M. R. Roest, E. A. H. Vollebregt, E. Ten Cate, and H.-X. Lin. Structured parallelisation of the flow simulation package TRIWAQ. Lecture Notes in Computer Science, 919:136–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mechoso:1995:HPC

[914] C. R. Mechoso. High-performance computing and networking for climate research. Lecture Notes in Computer Science, 919:142–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sunderland:1995:PSS

[915] A. Sunderland. Parallel solution strategies for triangular systems arising from oil reservoir simulations. Lecture Notes in Computer Science, 919:148– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanWaveren:1995:PSR

[916] G. M. Van Waveren and I. M. Godfrey. Parallelisation of surface-related multiple elimination. Lecture Notes in Computer Science, 919:156–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radi:1995:PRP

[917] B. Radi and J.-F. Estrade. Preliminary results on the parallelization of ARPEGE/IFS with the implementation of the full Meteo-France physics. Lecture Notes in Computer Science, 919:164–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meijer:1995:PSI

[918] P. M. Meijer, S. Poedts, J. P. Goedbloed, and A. Jakoby. A parallel semiimplicit method for 3D nonlinear magnetohydrodynamics. Lecture Notes in Computer Science, 919:170–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andre:1995:PDP

[919] F. Andre, M. Le Fur, Y. Maheo, and J.-L. Pazat. The Pandore dataparallel compiler and its portable runtime. Lecture Notes in Computer Science, 919:176–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dion:1995:MAL

[920] M. Dion and Y. Robert. Mapping affine loop nests: new results. Lecture Notes in Computer Science, 919: 184–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anand:1995:ETP

[921] S. S. Anand, D. A. Bell, J. G. Hughes, and C. M. Shapcott. Evidential techniques in parallel database mining. Lecture Notes in Computer Science, 919:190–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stamatelopoulos:1995:PPP

[922] F. Stamatelopoulos, G. Manis, and G. Papakonstantinou. A portable platform for parallel databases. Lecture Notes in Computer Science, 919:196– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Decker:1995:SAU

[923] K. M. Decker, J. J. Dvorak, R. M. Rehmann, and R. Ruehl. Satisfying application user requirements: A nextgeneration tool environment for parallel systems. Lecture Notes in Computer Science, 919:206–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeDoncker:1995:DPD

[924] E. De Doncker, A. Gupta, P. Ealy, and K. Rathbun. Development of a parallel and distributed integration package — PART I. Lecture Notes in Computer Science, 919:229–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grabner:1995:DPP

[925] S. Grabner, D. Kranzlmueller, J. Volkert, and J. Kepler. Debugging parallel programs using ATEMPT. Lecture Notes in Computer Science, 919:235– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saletore:1995:MDP

[926] V. A. Saletore and T. F. Neff. Messagedriven parallel computations on the MEIKO CS-2 parallel supercomputer. Lecture Notes in Computer Science, 919:241–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Veronese:1995:CSW

[927] S. Veronese and H. Othmer. A computational study of wave propagation in a model for anisotropic cardiac ventricular tissue. Lecture Notes in Computer Science, 919:248–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fatin:1995:PPS

[928] F. Fatin and R. J. Paul. Programming parallel simulations. Lecture Notes in Computer Science, 919:254–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carboni:1995:XIP

[929] P. Carboni, M. Fruscione, F. Guindani, and S. Punzi. XHIVE: interactive parallel application development using the PCF methodology. Lecture Notes in Computer Science, 919:257–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Loechner:1995:TAR

[930] V. Loechner and C. Mongenet. A toolbox for affine recurrence equations parallelization. Lecture Notes in Computer Science, 919:263–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anguita:1995:LLN

[931] D. Anguita, F. Passaggio, and R. Zunino. Learning in large neural networks. Lecture Notes in Computer Science, 919:269–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mattes:1995:NNP

[932] J. Mattes, D. Trystram, and J. Demongeot. Neural networks for parallel contrast enhancement in medical images. Lecture Notes in Computer Science, 919:275–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kihl:1995:NNB

[933] H. Kihl, J. P. Urban, J. Gresser, and S. Hagmann. Neural network based hand-eye positioning with a Transputer-based system. Lecture Notes in Computer Science, 919:281– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Willien:1995:PPM

[934] F. Willien, E. Piault, and F.-X. Roux. Parallel preconditioners on MIMD computers applied to petroleum industry. Lecture Notes in Computer Science, 919:287–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hortmann:1995:CHT

[935] M. Hortmann, M. Pophal, M. Schaefer, and K. Wechsler. Computation of heat transfer with methods of high performance scientific computing. Lecture Notes in Computer Science, 919:293–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marsh:1995:SLE

[936] A. Marsh and D. I. Kaklamani. Solution of large electromagnetic problems made feasible by HPC-Reducing execution times from months to hours. Lecture Notes in Computer Science, 919: 300–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoessel:1995:ETS

[937] A. Stoessel. An efficient tool for the study of 3D turbulent combustion phenomena on MPP computers. Lecture Notes in Computer Science, 919:306– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seyl:1995:PCS

[938] F. Seyl, R. Schinke, and O. Haan. A parallel code for simulating the ionization of hydrogen with short intense laser pulses. Lecture Notes in Computer Science, 919:312–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Erbacci:1995:PCM

[939] G. Erbacci, R. Ansaloni, B. Montagnini, and R. Scardovelli. Porting a coarse-mesh neutron diffusion code on a Cray T3D massively parallel computer. Lecture Notes in Computer Science, 919:318–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bubak:1995:LBL

[940] M. Bubak, J. Moschinski, M. Pogoda, and R. Slota. Load balancing for lattice gas and molecular dynamics simulations on networked workstations. Lecture Notes in Computer Science, 919: 329–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pimentel:1995:MME

[941] A. D. Pimentel, J. Van Brummen, T. Papathanassiadis, and P. M. A.

Sloot. Mermaid: modelling and evaluation research in MIMD architecture design. Lecture Notes in Computer Science, 919:335–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emmen:1995:FAE

[942] A. Emmen. A framework for analysis of European HPCN centres. Lecture Notes in Computer Science, 919: 341–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stueben:1995:CTP

[943] H. Stueben. The Cray T3D as a production machine at Konrad-Zuse-Zentrum Berlin. Lecture Notes in Computer Science, 919:347–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kallrath:1995:DOR

[944] J. Kallrath and A. Schreieck. Discrete optimisation and real-world problems. Lecture Notes in Computer Science, 919:351–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fox:1995:EIH

[945] G. C. Fox, K. A. Hawick, M. Podgorny, and K. Mills. The electronic InfoMall — HPCN enabling industry and commerce. Lecture Notes in Computer Science, 919:360–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conforti:1995:OME

[946] D. Conforti and L. De Luca. Optimal management of electric power systems via high performance computing. Lecture Notes in Computer Science, 919:366–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vounckx:1995:RMP

[947] J. Vounckx, G. Deconinck, and R. Lauwereins. Reconfiguration of massively parallel systems. Lecture Notes in Computer Science, 919:372– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thevenin:1995:SRF

[948] D. Thevenin, F. Behrendt, U. Maas, and J. Warnatz. Simulation of reacting flows with a portable parallel code using dynamic load balancing. Lecture Notes in Computer Science, 919:378–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stathopoulos:1995:DLB

[949] A. Stathopoulos and A. Ynnerman. Dynamic load balancing of atomic structure programs on a PVM cluster. Lecture Notes in Computer Science, 919:384–391, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanDriessche:1995:DLB

[950] R. Van Driessche and D. Roose. Dynamic load balancing with a spectral bisection algorithm for the constrained graph partitioning problem. Lecture Notes in Computer Science, 919:392– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joosen:1995:FLB

[951] W. Joosen, S. Bijnens, B. Robben, and J. Van Oeyen. Flexible load balancing software for parallel applications in a time-sharing environment. Lecture Notes in Computer Science, 919:398– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becker:1995:DBC

[952] W. Becker. Dynamic balancing complex workload in workstation networks — challenge, concepts and experience. Lecture Notes in Computer Science, 919:407–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hawick:1995:EHP

[953] K. A. Hawick and G. C. Fox. Exploiting high performance Fortran for computational fluid dynamics. Lecture Notes in Computer Science, 919:413–419, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kruizinga:1995:PIB

[954] P. Kruizinga and N. Petkov. Person identification based on multiscale matching of cortical images. Lecture Notes in Computer Science, 919:420– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fruscione:1995:PSP

[955] M. Fruscione, P. Stofella, and H.-J. Denuell. PARCS: a scalable parallel multimedia server for fast archiving and communication of medical images. Lecture Notes in Computer Science, 919:428–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carter:1995:CTC

[956] J. G. Carter, D. Cokljat, and R. J. Blake. Computation of turbulent coaxial jet flow on parallel systems. Lecture Notes in Computer Science, 919: 434–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fraaije:1995:IDD

[957] J. G. E. M. Fraaije and O. A. Evers. Implementation of dynamic density functional theory for self-organizing complex fluids on parallel computers. Lecture Notes in Computer Science, 919:441–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogels:1995:NSC

[958] M. E. S. Vogels, H. Van der Ven, and G. J. Hameetman. The need for super computers in aerospace research and industry. Lecture Notes in Computer Science, 919:448–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Song:1995:PHM

[959] Z. W. Song, D. Roose, C. S. Yu, and J. Berlamont. Parallelisation of a hydrodynamic model for the Northwest European continental shelf. Lecture Notes in Computer Science, 919:455– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernaschi:1995:PEI

[960] M. Bernaschi and G. Richelli. PVMe: an enhanced implementation of PVM for the IBM 9076 SP2. Lecture Notes in Computer Science, 919:461–471, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Topham:1995:PDA

[961] N. Topham and K. McDougall. Performance of the decoupled ACRI-1 architecture: the perfect club. Lecture Notes in Computer Science, 919:472– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tiemann:1995:AIS

[962] B. Tiemann, H. V. Muehll, I. Hasler, and E. Hiltebrand. Architecture and implementation of a single-board desktop supercomputer. Lecture Notes in Computer Science, 919:481–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ansaloni:1995:POQ

[963] R. Ansaloni, E. Rossi, and S. Evangelisti. Porting and optimizing a quantum-chemistry FCI algorithm on the Cray T3D. Lecture Notes in Computer Science, 919:488–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boehm:1995:HPC

[964] K. Boehm. High performance computing for one of the grand challenges. Lecture Notes in Computer Science, 919:496–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ellingworth:1995:RSA

[965] H. Ellingworth and S. R. Eddy. RNA structure alignment on a massively parallel computer. Lecture Notes in Computer Science, 919:502–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dzwinel:1995:PRM

[966] W. Dzwinel and J. Blasiak. Pattern recognition via molecular dynamics on vector supercomputers and networked workstations. Lecture Notes in Computer Science, 919:508–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schnekenburger:1995:PIP

[967] T. Schnekenburger and H. Spruth. Parallelization of an iterative placement algorithm using ParMod-C. Lecture Notes in Computer Science, 919:514– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vajtersic:1995:HPV

[968] M. Vajtersic. High-performance VLSI model elliptic solvers. Lecture Notes in Computer Science, 919:520–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brezany:1995:PIC

[969] P. Brezany, K. Sanjari, O. Cheron, and E. Van Konijnenburg. Processing irregular codes containing arrays with multi-dimensional distributions by the PREPARE HPF compiler. Lecture Notes in Computer Science, 919:526– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kremien:1995:SDS

[970] O. Kremien. Scalability in distributed systems, parallel systems and supercomputers. Lecture Notes in Computer Science, 919:532–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wolf:1995:HPC

[971] K. Wolf, A. Holtz, and M. Lang. High performance C++. Lecture Notes in Computer Science, 919:542–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dunlop:1995:TOP

[972] A. Dunlop, E. Hernandez, O. Naim, and T. Hey. A toolkit for optimising parallel performance. Lecture Notes in Computer Science, 919:548–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baraglia:1995:PSR

[973] R. Baraglia, D. Laforenza, and A. Lagana. Parallelization strategies for a reduced dimensionality calculation of quantum reactive scattering cross sections on a hypercube machine. Lecture Notes in Computer Science, 919: 554–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hausner:1995:EIP

[974] M. Hausner, M. Burrows, and C. A. Thekkath. Efficient implementation of PVM on the AN2 ATM network. Lecture Notes in Computer Science, 919:562–569, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aloisio:1995:UPW

[975] G. Aloisio and M. A. Bochicchio. The use of PVM with workstation clusters for distributed SAR data processing. Lecture Notes in Computer Science, 919:570–581, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alves:1995:WPC

[976] A. Alves, L. Silva, J. Carreira, and J. G. Silva. WPVM: parallel computing for the people. Lecture Notes in Computer Science, 919:582–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Appiani:1995:PSM

[977] E. Appiani, M. Bologna, M. Corvi, and M. Iardella. PVM in a sharedmemory industrial multiprocessor. Lecture Notes in Computer Science, 919: 588–593, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitschele-Thiel:1995:DPO

[978] A. Mitschele-Thiel. The DSPL project — an overview. Lecture Notes in Computer Science, 919:594–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calvin:1995:OTC

[979] C. Calvin, L. Colombet, and P. Michallon. Overlapping techniques of communications. Lecture Notes in Computer Science, 919:600–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blaszczyk:1995:PCE

[980] A. Blaszczyk, Z. Andjelic, P. Levin, and A. Ustundag. Parallel computation of electric fields in a heterogeneous workstation cluster. Lecture Notes in Computer Science, 919:606–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Huang:1995:MDT

[981] X. Huang and C. Steigner. A modeldriven tool for performance measurement and analysis of parallel programs. Lecture Notes in Computer Science, 919:612–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulz-Ziemer:1995:HIP

[982] G. Schulz-Ziemer and A. Geiger. HPF on Intel Paragon and CRAFT on CRAY T3D: basic performance measurements and experiments with a block-sparse CG-algorithm. Lecture Notes in Computer Science, 919:618– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simon:1995:SAN

[983] H. D. Simon and E. Strohmaier. Statistical analysis of NAS parallel benchmarks and LINPACK results. Lecture Notes in Computer Science, 919:626– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Overeinder:1995:PPE

[984] B. J. Overeinder and P. M. A. Sloot. Parallel performance evaluation through critical path analysis. Lecture Notes in Computer Science, 919:634– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bachler:1995:BPF

[985] G. Bachler, R. Greimel, H. Schiffermuller, and M. Bernaschi. Benchmarking the parallel FIRE code on IBM SP1-2 scalable parallel platforms. Lecture Notes in Computer Science, 919: 640–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cozzi:1995:HPD

[986] M. D. Cozzi. High performance disk systems for workstation environments. Lecture Notes in Computer Science, 919:646–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boulter:1995:PEH

[987] B. Boulter. Performance evaluation of HPF for scientific computing. Lecture Notes in Computer Science, 919:652– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vezolle:1995:PCE

[988] P. Vezolle. Parallel computational electromagnetics on the CRAY T3D using boundary element method. Lecture Notes in Computer Science, 919: 658–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taraglio:1995:EIB

[989] S. Taraglio, C. R. Casaccia, and F. Massaieli. An efficient implementation of a backpropagation learning algorithm on a quadrics parallel supercomputer. Lecture Notes in Computer Science, 919:664–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hakulinen:1995:ERP

[990] T. Hakulinen and F. Rademakers. Experience of running PIAF on the CS-2 at CERN. Lecture Notes in Computer Science, 919:672–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alda:1995:CDS

[991] W. Alda, W. Dzwinel, J. Kitowski, and J. Moscinski. Convection driven by sedimentation using molecular dynamics approach. Lecture Notes in Computer Science, 919:678–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hauser:1995:RTA

[992] R. Hauser and I. Legrand. A realtime application for the CS-2. Lecture Notes in Computer Science, 919:684– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asenjo:1995:SLF

[993] R. Asenjo and E. L. Zapata. Sparse LU factorization on the Cray T3D. Lecture Notes in Computer Science, 919: 690–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauber:1995:PSS

[994] T. Rauber and G. Ruenger. Parallel solution of a Schrödinger–Poisson system. Lecture Notes in Computer Science, 919:697–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birken:1995:DDM

[995] K. Birken and C. Helf. A dynamic data model for parallel adaptive PDE solvers. Lecture Notes in Computer Science, 919:703–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caselli:1995:ASG

[996] S. Caselli, G. Conte, and P. Marenzoni. Analysis and solution of generalized stochastic Petri nets: a data parallel CM-5 approach. Lecture Notes in Computer Science, 919:709–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pavani:1995:DDC

[997] R. Pavani and U. De Ros. A distributed divide-and-conquer approach to the parallel tridiagonal symmetric eigenvalue problem. Lecture Notes in Computer Science, 919:717–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanGijzen:1995:PIS

[998] M. Van Gijzen. Parallel iterative solution methods for linear finite element computations on the Cray T3D. Lecture Notes in Computer Science, 919: 723–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruno:1995:MAA

[999] G. Bruno and A. D’Acierno. The multialgorithmic approach to optimisation problems. Lecture Notes in Computer Science, 919:729–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauber:1995:ASP

[1000] T. Rauber, G. Ruenger, and R. Wilhelm. An application specific parallel programming paradigm. Lecture Notes in Computer Science, 919:735– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ronchi:1995:FDA

[1001] C. Ronchi, R. Iacono, and P. S.

Paolucci. Finite difference approximation to the shallow water equations on a quasi-uniform spherical grid. Lecture Notes in Computer Science, 919:741– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corana:1995:PAC

[1002] A. Corana and C. Rolando. A parallel approach to compute the Kolmogorov entropy from a time series. Lecture Notes in Computer Science, 919:748– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apolloni:1995:ILE

[1003] E. Apolloni and M. Talamo. ITU-Land ESPRIT project: a parallel DSS for environmental impact assessment. Lecture Notes in Computer Science, 919:754–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Letteron:1995:HPS

[1004] J.-M. Letteron and J. Bancroft. HOOD and parallelism in the softpar project. Lecture Notes in Computer Science, 919:762–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mager:1995:HHT

[1005] J. W. L. J. Mager and A. Ten Dam. HAMLET: HPCN technology for realtime, embedded applications. Lecture Notes in Computer Science, 919:772– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roettger:1995:IPD

[1006] M. Roettger, U.-P. Schroeder, and J. Simon. Implementation of a parallel and distributed mapping kernel for PARIX. Lecture Notes in Computer Science, 919:781–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codani:1995:PLL

[1007] J.-J. Codani and E. Glemet. Parallelism in LASSAP, a LArge scale sequence compArisons package. Lecture Notes in Computer Science, 919:787– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gerndt:1995:PAS

[1008] M. Gerndt and R. Berrendorf. Parallelizing applications with SVM-Fortran. Lecture Notes in Computer Science, 919:793–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patron:1995:ECT

[1009] M. Patron, T. Porcher, and F. Robin.

An evaluation of the Cray T3D at CEA/CEL-V. Lecture Notes in Computer Science, 919:799–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mierendorff:1995:EPI

[1010] H. Mierendorff, K. Stueben, C.-A.

Thole, and O. Thomas. Europort-1: Porting industrial codes to parallel architectures. Lecture Notes in Computer Science, 919:806–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Robinson:1995:PSP

[1011] D. Robinson, J. Demirdzic, and R. Krishna. Parallel STAR promises bright future. Lecture Notes in Computer Science, 919:813–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degrez:1995:PIC

[1012] G. Degrez, L. Giraud, M. Loriot, and A. Micelotta. Parallel industrial CFD calculations with N3S. Lecture Notes in Computer Science, 919:820–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Debae:1995:PVP

[1013] F. Debae, Y. Jiang, J. M. Marchal, and P. Henriksen. A parallel version of polyflow. Lecture Notes in Computer Science, 919:826–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ierotheou:1995:PNH

[1014] C. S. Ierotheou, C. R. Forsey, and U. Block. Parallelisation of a novel 3D hybrid structured/unstructured grid CFD production code. Lecture Notes in Computer Science, 919:831–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanKemenade:1995:PNS

[1015] V. Van Kemenade, M. J. Dayde, and J. B. Vos. Parallel Navier-stokes multiblock code to solve industrial aerodynamic design problems on high performance computers. Lecture Notes in Computer Science, 919:837–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ast:1995:GAA

[1016] M. Ast, H. Manz, J. Labarta, and A. Perez. A general approach for an automatic parallelization applied to the finite element code PERMAS. Lecture Notes in Computer Science, 919:844– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayer:1995:PMN

[1017] S. Mayer, P. Poschmann, S. Shamsian, and L. Komzsik. Parallel MSC/ NASTRAN on distributed memory computers. Lecture Notes in Computer Science, 919:850–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lonsdale:1995:EIC

[1018] G. Lonsdale, B. Elsner, J. Clinckemaillie, and S. Vlachoutsis. Experiences with industrial crashworthiness simulation using the portable, messagepassing PAM-CRASH code. Lecture Notes in Computer Science, 919:856– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Delsemme:1995:SMC

[1019] J.-P. Delsemme. SAMMI: S.A.M.C.E.F.

migration onto parallel platforms. Lecture Notes in Computer Science, 919: 863–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colbrook:1995:EEE

[1020] A. Colbrook, J. A. Elliott, M. Lemke, and F. Wray. EUROPORT2 — ES-PRIT European porting action no 2. Lecture Notes in Computer Science, 919:869–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frieze:1995:IAA

[1021] A. Frieze and M. Jerrum. Improved approximation algorithms for MAX k-CUT and MAX BISECTION. Lecture Notes in Computer Science, 920:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Even:1995:AMF

[1022] G. Even, J. Naor, B. Schieber, and M. Sudan. Approximating minimum feedback sets and multi-cuts in directed graphs. Lecture Notes in Computer Science, 920:14–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertsimas:1995:NFI

[1023] D. Bertsimas, C. Teo, and R. Vohra. Nonlinear formulations and improved randomized approximation algorithms for multicut problems. Lecture Notes in Computer Science, 920:29–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carr:1995:SCT

[1024] R. D. Carr. Separating clique tree and bipartition inequalities in polynomial time. Lecture Notes in Computer Science, 920:40–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mueller:1995:IOP

[1025] R. Mueller and A. S. Schulz. The interval order polytope of a digraph. Lecture Notes in Computer Science, 920:50–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1995:SPS

[1026] E. Cheng and W. H. Cunningham. Separation problems for the stable set polytope. Lecture Notes in Computer Science, 920:65–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bienstock:1995:CSF

[1027] D. Bienstock. Computational study of a family of mixed-integer quadratic programming problems. Lecture Notes in Computer Science, 920:80–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pisinger:1995:MAB

[1028] D. Pisinger. A minimal algorithm for the bounded knapsack problem. Lecture Notes in Computer Science, 920: 95–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Escudero:1995:FTP

[1029] L. F. Escudero, S. Martello, and P. Toth. A framework for tightening 0-1 programs based on extensions of pure 0-1 KP and SS problems. Lecture Notes in Computer Science, 920:110–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Helmberg:1995:CSP

[1030] C. Helmberg, S. Poljak, F. Rendl, and H. Wolkowicz. Combining semidefinite and polyhedral relaxations for integer programs. Lecture Notes in Computer Science, 920:124–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Deng:1995:DNO

[1031] X. Deng. Distributed near-optimal matching. Lecture Notes in Computer Science, 920:135–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pferschy:1995:RLB

[1032] U. Pferschy. The random linear bottleneck assignment problem. Lecture Notes in Computer Science, 920:145– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cherkassky:1995:IPR

[1033] B. V. Cherkassky and A. V. Goldberg. On implementing push-relabel method for the maximum flow problem. Lecture Notes in Computer Science, 920:157–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ali:1995:UHN

[1034] A. I. Ali, H.-S. Han, and J. L. Kennington. Use of hidden network structure in the set partitioning problem. Lecture Notes in Computer Science, 920: 172–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hartvigsen:1995:GMF

[1035] D. Hartvigsen. Generalized max flows and augmenting paths. Lecture Notes in Computer Science, 920:185– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kleinschmidt:1995:OMP

[1036] P. Kleinschmidt and S. Onn. Oriented matroid polytopes and polyhedral fans are signable. Lecture Notes in Computer Science, 920:198–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Novick:1995:CPB

[1037] B. Novick and A. Seboe. On combinatorial properties of binary spaces. Lecture Notes in Computer Science, 920: 212–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouchet:1995:CDC

[1038] A. Bouchet. Coverings and deltacoverings. Lecture Notes in Computer Science, 920:228–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barany:1995:TSM

[1039] I. Barany, H. E. Scarf, and D. Shallcross. The topological structure of maximal lattice free convex bodies: The general case. Lecture Notes in Computer Science, 920:244–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laburthe:1995:HBC

[1040] F. Laburthe. The Hilbert basis of the cut cone over the complete graph K6. Lecture Notes in Computer Science, 920:252–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hosten:1995:GIG

[1041] S. Hosten and B. Sturmfels. GRIN: An implementation of Gröbner bases for integer programming. Lecture Notes in Computer Science, 920:267–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crama:1995:SJE

[1042] Y. Crama and F. C. R. Spieksma.

Scheduling jobs of equal length: Complexity, facets and computational results. Lecture Notes in Computer Science, 920:277–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoogeveen:1995:FSP

[1043] H. Hoogeveen and S. Van de Velde. Formulating a scheduling problem with almost identical jobs by using positional completion times. Lecture Notes in Computer Science, 920:292–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Queyranne:1995:SUJ

[1044] M. Queyranne and A. S. Schulz. Scheduling unit jobs with compatible release dates on parallel machines with nonstationary speeds. Lecture Notes in Computer Science, 920:307–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conforti:1995:MMD

[1045] M. Conforti, G. Cornuejols, A. Kapoor, and K. Vuskovic. A Mickey-Mouse decomposition theorem. Lecture Notes in Computer Science, 920:321–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klinz:1995:MCD

[1046] B. Klinz and G. J. Woeginger. Minimum cost dynamic flows: The seriesparallel case. Lecture Notes in Computer Science, 920:329–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nobili:1995:IM

[1047] P. Nobili and A. Sassano. (0, 1) ideal matrices. Lecture Notes in Computer Science, 920:344–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juvan:1995:EGT

[1048] M. Juvan, J. Marincek, and B. Mohar.

Embedding graphs in the torus in linear time. Lecture Notes in Computer Science, 920:360–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ageev:1995:CSG

[1049] A. A. Ageev, A. V. Kostochka, and Z. Szigeti. A characterization of Seymour graphs. Lecture Notes in Computer Science, 920:364–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bartels:1995:MCC

[1050] J. E. Bartels and D. J. A. Welsh. The Markov chain of colourings. Lecture Notes in Computer Science, 920:373– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabow:1995:PAA

[1051] H. N. Gabow and K. S. Manu. Packing algorithms for arborescences (and spanning trees) in capacitated graphs. Lecture Notes in Computer Science, 920:388–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nagamochi:1995:FES

[1052] H. Nagamochi and T. Ibaraki. A faster edge splitting algorithm in multigraphs and its application to the edge-connectivity augmentation problem. Lecture Notes in Computer Science, 920:403–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frank:1995:HMS

[1053] A. Frank and T. Jordan. How to make a strongly connected digraph twoconnected. Lecture Notes in Computer Science, 920:414–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dahl:1995:POC

[1054] G. Dahl. Polyhedra and optimization in connection with a weak majorization ordering. Lecture Notes in Computer Science, 920:426–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ceria:1995:CSG

[1055] S. Ceria, G. Cornuejols, and M. Dawande. Combining and strengthening Gomory cuts. Lecture Notes in Computer Science, 920:438–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gu:1995:SIL

[1056] Z. Gu, G. L. Nemhauser, and M. W. P. Savelsbergh. Sequence independent lifting of cover inequalities. Lecture Notes in Computer Science, 920:452– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koyama:1995:FRT

[1057] Kenji Koyama. Fast RSA-type schemes based on singular cubic curves y^{2} + axy ≡ x^{3} (mod n). Lecture Notes in Computer Science, 921:329–340, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/0921/09210329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210329.
pdf.

Schnorr:1995:ACR

[1058] Claus P. Schnorr and H. H. Hörner. Attacking the Chor–Rivest cryptosystem by improved lattice reduction. Lecture Notes in Computer Science, 921: 1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210001.
pdf.

OConnor:1995:CDD

[1059] Luke O’Connor. Convergence in differential distributions. Lecture Notes in Computer Science, 921:13– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210013.
pdf.

Harpes:1995:GLC

[1060] Carlo Harpes, Gerhard G. Kramer, and James L. Massey. A generalization of linear cryptanalysis and the applicability of Matsui’s piling-up lemma. Lecture Notes in Computer Science, 921: 24–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210024.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0921/09210024.
                                        pdf.

Chen:1995:EGS

[1061] Lidong Chen and Torben Pryds Pedersen. On the efficiency of group signatures providing informationtheoretic anonymity. Lecture Notes in Computer Science, 921:39–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0921/09210039.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0921/09210039.
                                        pdf.

Franklin:1995:VSS

[1062] Matthew K. Franklin and Michael K. Reiter. Verifiable signature sharing. Lecture Notes in Computer Science, 921:50–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0921/09210050.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0921/09210050.
                                        pdf.

Lim:1995:SPS

[1063] Chae Hoon Lim and Pil Joong Lee. Server (prover/signer)-aided verification of identity proofs and signatures. Lecture Notes in Computer Science, 921:64–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0921/09210064.htm;   http:
                                        //link.springer-ny.com/link/service/

series/0558/papers/0921/09210064.
pdf.

Lercier:1995:CNP

[1064] Reynald Lercier and François Morain.

Counting the number of points on elliptic curves over finite fields: strategies and performances. Lecture Notes in Computer Science, 921:79– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210079.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210079.
pdf.

Weber:1995:IGN

[1065] Damian Weber. An implementation of the general number field sieve to compute discrete logarithms mod p. Lecture Notes in Computer Science, 921: 95–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210095.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210095.
pdf.

Montgomery:1995:BLA

[1066] Peter L. Montgomery. A block Lanczos algorithm for finding dependencies over GF (2). Lecture Notes in Computer Science, 921:106–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210106.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/0921/09210106.
                                        pdf.

Pfitzmann:1995:HBA

[1067] Birgit Pfitzmann, Matthias Schunter, and Michael Waidner. How to break another “provably secure” payment system. Lecture Notes in Computer Science, 921:121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0921/09210121.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0921/09210121.
                                        pdf.

Crepeau:1995:QOM

[1068] Claude Crépeau and Louis Salvail. Quantum oblivious mutual identification. Lecture Notes in Computer Science, 921:133–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0921/09210133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0921/09210133.
                                        pdf.

Desmedt:1995:STC

[1069] Yvo Desmedt. Securing traceability of ciphertexts — towards a secure software key escrow system. Lecture Notes in Computer Science, 921:147– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0921/09210147.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/0921/09210147.
pdf.

Gehrmann:1995:SMA

[1070] Christian Gehrmann. Secure multiround authentication protocols. Lecture Notes in Computer Science, 921: 158–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210158.
pdf.

Gennaro:1995:VSS

[1071] Rosario Gennaro and Silvio Micali. Verifiable secret sharing as secure computation. Lecture Notes in Computer Science, 921:168–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210168.
pdf.

Jackson:1995:ESS

[1072] Wen-Ai Jackson, Keith M. Martin, and Christine M. O’Keefe. Efficient secret sharing without a mutually trusted authority. Lecture Notes in Computer Science, 921:183– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210183.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/0921/09210183.
pdf.

Beguin:1995:GSC

[1073] Philippe Béguin and Antonella Cresti. General short computational secret sharing schemes. Lecture Notes in Computer Science, 921:194–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210194.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210194.
pdf.

Stadler:1995:FBS

[1074] Markus A. Stadler, Jean-Marc Piveteau, and Jan L. Camenisch. Fair blind signatures. Lecture Notes in Computer Science, 921:209–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210209.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210209.
pdf.

Jakobsson:1995:RCF

[1075] Markus Jakobsson. Ripping coins for a fair exchange. Lecture Notes in Computer Science, 921:220–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210220.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210220.
pdf.

Brands:1995:RBS

[1076] Stefan Brands. Restrictive blinding of secret-key certificates. Lecture Notes in Computer Science, 921:231– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210231.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210231.
pdf.

Golic:1995:TFC

[1077] Jovan Dj. Golic. Towards fast correlation attacks on irregularly clocked shift registers. Lecture Notes in Computer Science, 921:248–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210248.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210248.
pdf.

Klapper:1995:LPN

[1078] Andrew Klapper and Mark Goresky.

Large periods nearly de Bruijn FCSR sequences. Lecture Notes in Computer Science, 921:263–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210263.
pdf.

Zhang:1995:NRF

[1079] Xian-Mo Zhang and Yuliang Zheng. On nonlinear resilient functions. Lecture Notes in Computer Science, 921:274– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0921/09210274.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0921/09210274.
                                        pdf.

Kurosawa:1995:CBA

[1080] Kaoru Kurosawa and Satoshi Obana. Combinatorial bounds for authentication codes with arbitration. Lecture Notes in Computer Science, 921:289– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0921/09210289.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0921/09210289.
                                        pdf.

Krawczyk:1995:NHF

[1081] Hugo Krawczyk. New hash functions for message authentication. Lecture Notes in Computer Science, 921:301– 310, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0921/09210301.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0921/09210301.
                                        pdf.

Bierbrauer:1995:CUH

[1082] Jürgen Bierbrauer. A^{2}-codes from universal hash classes. Lecture Notes in Computer Science, 921:311–318, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210311.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210311.
pdf.

Pointcheval:1995:NIS

[1083] David Pointcheval. A new identification scheme based on the perceptrons problem. Lecture Notes in Computer Science, 921:319–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210319.
pdf.

Sakurai:1995:RAC

[1084] Kouichi Sakurai and Hiroki Shizuya.

Relationships among the computational powers of breaking discrete log cryptosystems. Lecture Notes in Computer Science, 921:341–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210341.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210341.
pdf.

Naeslund:1995:UHF

[1085] M. Naeslund. Universal hash functions and hard core bits. Lecture Notes in Computer Science, 921:356–366, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naslund:1995:UHF

[1086] Mats Näslund. Universal hash functions & hard core bits. Lecture Notes in Computer Science, 921:356– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0921/09210356.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0921/09210356.
                                        pdf.

Crescenzo:1995:RRB

[1087] Giovanni Di Crescenzo. Recycling random bits in composed perfect zero-knowledge. Lecture Notes in Computer Science, 921:367–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0921/09210367.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0921/09210367.
                                        pdf.

DiCrescenzo:1995:RRB

[1088] G. Di Crescenzo. Recycling random bits in composed perfect zeroknowledge. Lecture Notes in Computer Science, 921:367–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1995:MIH

[1089] Chih-Hung Wang, Tzonelih Hwang, and Jiun-Jang Tsai. On the Matsumoto and Imai’s human identification scheme. Lecture Notes in Computer Science, 921:382–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210382.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210382.
pdf.

Sako:1995:RFM

[1090] Kazue Sako and Joe Kilian. Receiptfree mix-type voting scheme — A practical solution to the implementation of a voting booth. Lecture Notes in Computer Science, 921:393– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210393.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210393.
pdf.

Naccache:1995:CAR

[1091] David Naccache, David M’Raihi, William Wolfowicz, and Adina Di Porto. Are crypto-accelerators really inevitable? 20 bit zero-knowledge in less than a second on simple 8bit microcontrollers. Lecture Notes in Computer Science, 921:404–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0921/09210404.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0921/09210404.
pdf.

DAmiano:1995:ANP

[1092] Stefano D’Amiano and Giovanni Di Crescenzo. Anonymous NIZK proofs of knowledge with preprocessing. Lecture Notes in Computer Science, 921: 413–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0921/09210413.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0921/09210413.
                                        pdf.

Anonymous:1995:AIa

[1093] Anonymous. Author index. Lecture Notes in Computer Science, 921:417– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/0921/0921indx.pdf.

Anonymous:1995:GRS

[1094] Anonymous. Graph rewriting systems — the basic concepts. Lecture Notes in Computer Science, 922:10–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:UGR

[1095] Anonymous. UBS-Graph rewriting systems — matching subgraphs constant time. Lecture Notes in Computer Science, 922:38–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:PAG

[1096] Anonymous. Programmed attributed graph rewrite systems — an advanced modelling formalism. Lecture Notes in Computer Science, 922:92–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:AMG

[1097] Anonymous. The abstract machine for graph rewriting — supporting a fast implementation. Lecture Notes in Computer Science, 922:125–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:GIF

[1098] Anonymous. A graphical implementation of functional languages — A case study in UBS-Graph rewriting systems. Lecture Notes in Computer Science, 922:165–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Podelski:1995:DAT

[1099] A. Podelski and P. Van Roy. A detailed algorithm testing guards over feature trees. Lecture Notes in Computer Science, 923:11–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Codognet:1995:GSC

[1100] C. Codognet and P. Codognet. A generalized semantics for concurrent constraint languages and their abstract interpretation. Lecture Notes in Computer Science, 923:39–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ertl:1995:HLC

[1101] M. Anton Ertl and A. Krall. High-level constraints over finite domains. Lecture Notes in Computer Science, 923:51–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gensel:1995:ICO

[1102] J. Gensel. Integrating constraints in an object-based knowledge representation system. Lecture Notes in Computer Science, 923:67–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boizumault:1995:CAE

[1103] P. Boizumault, Y. Delon, and L. Peridy. A CLP approach for examination planning. Lecture Notes in Computer Science, 923:85–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freuder:1995:MPS

[1104] E. C. Freuder. The many paths to satisfaction. Lecture Notes in Computer Science, 923:103–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wallace:1995:DAC

[1105] R. J. Wallace. Directed arc consistency preprocessing. Lecture Notes in Computer Science, 923:121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smith:1995:SED

[1106] B. M. Smith. In search of exceptionally difficult constraint satisfaction problems. Lecture Notes in Computer Science, 923:139–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bessiere:1995:UBS

[1107] C. Bessiere and J.-C. Regin. Using bidirectionality to speed up arc-consistency processing. Lecture Notes in Computer Science, 923:157–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freuder:1995:UMC

[1108] E. C. Freuder. Using metalevel constraint knowledge to reduce constraint checking. Lecture Notes in Computer Science, 923:171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prosser:1995:FCB

[1109] P. Prosser. Forward checking with backmarking. Lecture Notes in Computer Science, 923:185–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rossi:1995:RHV

[1110] F. Rossi. Redundant hidden variables in finite domain constraint problems. Lecture Notes in Computer Science, 923:205–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

File:1995:SPC

[1111] G. File, G. Nardiello, and A. Tirabosco. Semantic properties of CHIP (FD). Lecture Notes in Computer Science, 923:225–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vlasie:1995:CHC

[1112] D. R. Vlasie. Combining hill climbing and forward checking. Lecture Notes in Computer Science, 923:247–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiben:1995:GEG

[1113] A. Eiben, P.-E. Raue, and Z. Ruttkay. GA-easy and GA-hard constraint satisfaction problems. Lecture Notes in Computer Science, 923:267–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreoli:1995:RBO

[1114] J.-M. Andreoli, H. Gallaire, and R. Pareschi. Rule-based object coordination. Lecture Notes in Computer Science, 924:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bijnens:1995:SIR

[1115] S. Bijnens, W. Joosen, and P. Verbaeten. Sender-initiated and receiverinitiated coordination in a global object space. Lecture Notes in Computer Science, 924:14–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1995:CPT

[1116] N. Brown. Correctness-preserving transformations for the design of parallel programs. Lecture Notes in Computer Science, 924:29–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Busi:1995:DCC

[1117] N. Busi and R. Gorrieri. Distributed conflicts in communicating systems. Lecture Notes in Computer Science, 924:49–??, 1995. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Carriero:1995:BL

[1118] N. Carriero, D. Gelernter, and L. Zuck.

Bauhaus linda. Lecture Notes in Computer Science, 924:66–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ciancarini:1995:OSC

[1119] P. Ciancarini, K. K. Jensen, and D. Yankelevich. On the operational semantics of a coordination language. Lecture Notes in Computer Science, 924:77–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Froelund:1995:AIB

[1120] S. Froelund and G. Agha. Abstracting interactions based on message sets. Lecture Notes in Computer Science, 924:107–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Minsky:1995:LGL

[1121] N. H. Minsky and J. Leichter. Lawgoverned Linda as a coordination model. Lecture Notes in Computer Science, 924:125–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nierstrasz:1995:RCL

[1122] O. Nierstrasz and T. D. Meijler. Requirements for a composition language. Lecture Notes in Computer Science, 924:147–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papathomas:1995:MAO

[1123] M. Papathomas, G. S. Blair, and G. Coulson. A model for active object coordination and for distributed multimedia applications. Lecture Notes in Computer Science, 924:162–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tolksdorf:1995:MUC

[1124] R. Tolksdorf. A machine for uncoupled coordination and its concurrent behavior. Lecture Notes in Computer Science, 924:176–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fokker:1995:FP

[1125] J. Fokker. Functional parsers. Lecture Notes in Computer Science, 925:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wadler:1995:MFP

[1126] P. Wadler. Monads for functional programming. Lecture Notes in Computer Science, 925:24–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hughes:1995:DPP

[1127] J. Hughes. The design of a prettyprinting library. Lecture Notes in Computer Science, 925:53–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jones:1995:FPO

[1128] M. P. Jones. Functional programming with overloading and higherorder polymorphism. Lecture Notes in Computer Science, 925:97–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hallgren:1995:PF

[1129] T. Hallgren and M. Carlsson. Programming with fudgets. Lecture Notes in Computer Science, 925:137–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanEekelen:1995:CMS

[1130] M. C. J. D. Van Eekelen and R. M. J.

Plasmeijer. Constructing medium sized efficient functional programs in clean. Lecture Notes in Computer Science, 925:183–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Meijer:1995:MMF

[1131] E. Meijer and J. Jeuring. Merging monads and folds for functional programming. Lecture Notes in Computer Science, 925:228–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kieburtz:1995:PA

[1132] R. B. Kieburtz and J. Lewis. Programming with algebras. Lecture Notes in Computer Science, 925:267–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Launchbury:1995:GAF

[1133] J. Launchbury. Graph algorithms with a functional flavour. Lecture Notes in Computer Science, 925:308–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bucci:1995:ISQ

[1134] G. Bucci. The impact of software quality. Lecture Notes in Computer Science, 926:10–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuntzmann-Combelles:1995:QAS

[1135] A. Kuntzmann-Combelles. Quantitative approach to software process improvement. Lecture Notes in Computer Science, 926:16–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuvaja:1995:BSP

[1136] P. Kuvaja. BOOTSTRAP: A software process assessment and improvement methodology. Lecture Notes in Computer Science, 926:31–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coletta:1995:SPI

[1137] A. Coletta. The SPICE project: An international standard for software process assessment, improvement and capability determination. Lecture Notes in Computer Science, 926:49–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Campanai:1995:QES

[1138] M. Campanai, E. Ferretti, and V. Valori. Quality estimation of software applications for banking. Lecture Notes in Computer Science, 926:64–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruno:1995:VSR

[1139] G. Bruno and R. Agarwal. Validating software requirements using operational models. Lecture Notes in Computer Science, 926:78–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lofrumento:1995:OOD

[1140] G. Lofrumento and S. Pileri. Objectoriented design of the UPT service: Evaluation of the design quality and experience learned. Lecture Notes in Computer Science, 926:94–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Staalhane:1995:DMR

[1141] T. Staalhane. Development of a model for reusability assessment. Lecture Notes in Computer Science, 926:111– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karlsson:1995:IRS

[1142] E.-A. Karlsson and J.-M. Morel. The impact of reuse on software quality. Lecture Notes in Computer Science, 926:124–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sommerville:1995:PGA

[1143] I. Sommerville, L. Masera, and C. Demaria. Practical guidelines for Ada reuse in an industrial environment. Lecture Notes in Computer Science, 926:138–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cicu:1995:PQP

[1144] A. Cicu. Process quality problems from the point of view of the user and of the purchaser of software. Lecture Notes in Computer Science, 926:148–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Binucci:1995:SQU

[1145] L. Binucci. Software quality of use: Evaluation by MUSiC. Lecture Notes in Computer Science, 926:165– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bakker:1995:RIE

[1146] G. Bakker and F. P. Hirdes. Recent industrial experiences with software product metrics. Lecture Notes in Computer Science, 926:179–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spinelli:1995:QMS

[1147] A. Spinelli, D. Pina, P. Salvaneschi, and E. Crivelli. Quality measurement of software products: an experience about a large automation system. Lecture Notes in Computer Science, 926:192–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bazzana:1995:QBS

[1148] G. Bazzana, R. Delmiglio, A. Lora, and O. Balestrini. Quantifying the benefits of software testing: an experience report from the GSM application domain. Lecture Notes in Computer Science, 926:207–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruno:1995:BUM

[1149] G. Bruno, M. Varani, V. Vico, and C. Offerman. Benefits of using modelbased testing tools. Lecture Notes in Computer Science, 926:224–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertolino:1995:STD

[1150] A. Bertolino. Software testing for dependability assessment. Lecture Notes in Computer Science, 926:236– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alferes:1995:ATS

[1151] J. J. Alferes and L. M. Pereira. An argumentation theoretic semantics based on non-refutable falsity. Lecture Notes in Computer Science, 927:3–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lifschitz:1995:DPA

[1152] V. Lifschitz and H. Turner. From disjunctive programs to abduction. Lecture Notes in Computer Science, 927: 23–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Przymusinski:1995:SND

[1153] T. C. Przymusinski. Semantics of normal and disjunctive logic programs: A unifying framework. Lecture Notes in Computer Science, 927:43–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Witteveen:1995:ENP

[1154] C. Witteveen. Every normal program has a nearly-stable model. Lecture Notes in Computer Science, 927:68–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

You:1995:LPA

[1155] J.-H. You and L. Y. Yuan. Logic programming with assumption denial. Lecture Notes in Computer Science, 927: 85–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barback:1995:RBP

[1156] M. D. Barback and J. Lobo. A resolution-based procedure for default theories with extensions. Lecture Notes in Computer Science, 927:101– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brass:1995:GAB

[1157] S. Brass and J. Diz. A general approach to bottom-up computation of disjunctive semantics. Lecture Notes in Computer Science, 927:127–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Costantini:1995:SSP

[1158] S. Costantini and G. A. Lanzarone. Static semantics as program transformation and well-founded computation. Lecture Notes in Computer Science, 927:156–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degerstedt:1995:MCW

[1159] L. Degerstedt and U. Nilsson. Magic computation for well-founded semantics. Lecture Notes in Computer Science, 927:181–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruiz:1995:CSP

[1160] C. Ruiz and J. Minker. Computing stable and partial stable models of extended disjunctive logic programs. Lecture Notes in Computer Science, 927:205–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiter:1995:CRA

[1161] T. Eiter, G. Gottlob, and N. Leone.

Complexity results for abductive logic programming. Lecture Notes in Computer Science, 928:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Denecker:1995:TIA

[1162] M. Denecker. A terminological interpretation of (abductive) logic programming. Lecture Notes in Computer Science, 928:15–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damasio:1995:AVE

[1163] C. V. Damasio and L. M. Pereira. Abduction over 3-valued extended logic programs. Lecture Notes in Computer Science, 928:29–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marek:1995:LCL

[1164] V. W. Marek, A. Nerode, and J. B. Remmel. On logical constraints in logic programming. Lecture Notes in Computer Science, 928:43–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aquilino:1995:OCD

[1165] D. Aquilino, P. Asirelli, C. Renso, and F. Turini. An operator for composing deductive data bases with theories of constraints. Lecture Notes in Computer Science, 928:57–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

HalfeldFerrariAlves:1995:URD

[1166] M. Halfeld Ferrari Alves, D. Laurent, and N. Spyratos. Update rules in datalog programs. Lecture Notes in Computer Science, 928:71–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brass:1995:CSS

[1167] S. Brass and J. Dix. Characterizations of the stable semantics by partial evaluation. Lecture Notes in Computer Science, 928:85–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blair:1995:GCL

[1168] H. A. Blair. Game characterizations of logic program properties. Lecture Notes in Computer Science, 928:99–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berman:1995:CWF

[1169] K. A. Berman, J. S. Schlipf, and J. V.

Franco. Computing the well-founded semantics faster. Lecture Notes in Computer Science, 928:113–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lifschitz:1995:LCW

[1170] V. Lifschitz, N. McCain, T. C. Przymusinski, and R. F. Staerk. Loop checking and the well-founded semantics. Lecture Notes in Computer Science, 928:127–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fitting:1995:ARS

[1171] M. Fitting. Annotated revision specification programs. Lecture Notes in Computer Science, 928:143–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Przymusinski:1995:UMI

[1172] T. C. Przymusinski and H. Turner. Update by means of inference rules. Lecture Notes in Computer Science, 928: 156–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DaSilva:1995:SWS

[1173] J. C. P. Da Silva and S. R. M. Veloso.

A sphere world semantics for default reasoning. Lecture Notes in Computer Science, 928:175–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Witteveen:1995:RC

[1174] C. Witteveen and W. Van der Hoek. Revision by communication. Lecture Notes in Computer Science, 928:189– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gabbay:1995:HUP

[1175] D. Gabbay, L. Giordano, A. Martelli, and N. Olivetti. Hypothetical updates, priority and inconsistency in a logic programming language. Lecture Notes in Computer Science, 928:203– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miller:1995:SCS

[1176] R. Miller. Situation calculus specifications for event calculus logic programs. Lecture Notes in Computer Science, 928:217–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yuan:1995:ELP

[1177] L.-Y. Yuan and J.-H. You. On the extension of logic programming with negation through uniform proofs. Lecture Notes in Computer Science, 928: 231–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bochman:1995:DCR

[1178] A. Bochman. Default consequence relations as a logical framework for logic programs. Lecture Notes in Computer Science, 928:245–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mikitiuk:1995:SRE

[1179] A. Mikitiuk and M. Truszczynski.

Skeptical rational extensions. Lecture Notes in Computer Science, 928:259– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cholewinski:1995:RSD

[1180] P. Cholewinski. Reasoning with stratified default theories. Lecture Notes in Computer Science, 928:273–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ng:1995:IMO

[1181] R. T. Ng and X. Tian. Incremental methods for optimizing partial instantiation. Lecture Notes in Computer Science, 928:287–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Staerk:1995:TPP

[1182] R. F. Staerk. A transformation of propositional Prolog programs into classical logic. Lecture Notes in Computer Science, 928:302–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dung:1995:NIA

[1183] P. M. Dung and T. C. Son. Nonmonotonic inheritance, argumentation and logic programming. Lecture Notes in Computer Science, 928:316–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brogi:1995:AFE

[1184] A. Brogi, E. Lamma, P. Mancarella, and P. Mello. An abductive framework for extended logic programming. Lecture Notes in Computer Science, 928:330–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakama:1995:ECT

[1185] C. Sakama and K. Inoue. Embedding circumscriptive theories in general disjunctive programs. Lecture Notes in Computer Science, 928:344–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kalinski:1995:SCO

[1186] J. Kalinski. Stable classes and operator pairs for disjunctive programs. Lecture Notes in Computer Science, 928: 358–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pearce:1995:NAS

[1187] D. Pearce. Nonmonotonicity and answer set inference. Lecture Notes in Computer Science, 928:372–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rajasekar:1995:TES

[1188] A. Rajasekar. Trans-epistemic semantics for logic programs. Lecture Notes in Computer Science, 928:388– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Toni:1995:CAS

[1189] F. Toni and A. C. Kakas. Computing the acceptability semantics. Lecture Notes in Computer Science, 928: 401–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schuster:1995:ALM

[1190] P. Schuster. Artificial life and molecular evolutionary biology. Lecture Notes in Computer Science, 929:3–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pattee:1995:ALN

[1191] H. H. Pattee. Artificial life needs a real epistemology. Lecture Notes in Computer Science, 929:23–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prem:1995:GES

[1192] E. Prem. Grounding and the entailment structure in robots and artificial life. Lecture Notes in Computer Science, 929:39–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gutowitz:1995:MFT

[1193] H. Gutowitz and C. Langton. Mean field theory of the edge of chaos. Lecture Notes in Computer Science, 929: 52–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wheeler:1995:ECM

[1194] M. Wheeler. Escaping from the Cartesian mind-set: Heidegger and artificial life. Lecture Notes in Computer Science, 929:65–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Umerez:1995:SCG

[1195] J. Umerez. Semantic closure: A guiding notion to ground artificial life. Lecture Notes in Computer Science, 929:77–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kampis:1995:IOV

[1196] G. Kampis. The inside and outside views of life. Lecture Notes in Computer Science, 929:95–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lazcano:1995:PCA

[1197] A. Lazcano. Prebiotic chemistry, artificial life, and complexity theory: What do they tell us about the origin of biological systems? Lecture Notes in Computer Science, 929:105–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nutho:1995:CRM

[1198] J. C. Nutho, P. Chacon, A. Moreno, and F. Moran. Compartimentation in replicator models. Lecture Notes in Computer Science, 929:116–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Forst:1995:EDO

[1199] C. V. Forst, C. Reidys, and J. Weber. Evolutionary dynamics and optimization: Neutral networks as model-landscapes for RNA secondarystructure folding-landscapes. Lecture Notes in Computer Science, 929:128– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Woodcock:1995:PES

[1200] G. Woodcock and P. G. Higgs. Population evolution in a single peak fitness landscape — how high are the clouds? Lecture Notes in Computer Science, 929:148–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McMullin:1995:RD

[1201] B. McMullin. Replicators don’t! Lecture Notes in Computer Science, 929: 158–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moya:1995:RVB

[1202] A. Moya, E. Domingo, and J. J. Holland. RNA viruses: a bridge between life and artificial life. Lecture Notes in Computer Science, 929:170–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yagil:1995:CAS

[1203] G. Yagil. Complexity analysis of a self-organizing vs. a template-directed system. Lecture Notes in Computer Science, 929:179–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yamamoto:1995:TAE

[1204] T. Yamamoto and K. Kaneko. Tile automaton for evolution of metabolism. Lecture Notes in Computer Science, 929:188–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cliff:1995:TRQ

[1205] D. Cliff and G. F. Miller. Tracking the red queen: Measurements of adaptive progress in co-evolutionary simulations. Lecture Notes in Computer Science, 929:200–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maley:1995:CMR

[1206] C. Maley. The coevolution of mutation rates. Lecture Notes in Computer Science, 929:219–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ikegami:1995:CMT

[1207] T. Ikegami and T. Hashimoto. Coevolution of machines and tapes. Lecture Notes in Computer Science, 929:234– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juille:1995:ICE

[1208] H. Juille. Incremental co-evolution of organisms: A new approach for optimization and discovery of strategies. Lecture Notes in Computer Science, 929:246–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Numaoka:1995:SCE

[1209] C. Numaoka. Symbiosis and coevolution in animals. Lecture Notes in Computer Science, 929:261–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bull:1995:AE

[1210] L. Bull, T. C. Fogarty, and A. G.

Pipe. Artificial endosymbiosis. Lecture Notes in Computer Science, 929: 273–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maeshiro:1995:MAE

[1211] T. Maeshiro and M. Kimura. Mathematical analysis of evolutionary process. Lecture Notes in Computer Science, 929:290–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oppacher:1995:EHR

[1212] F. Oppacher and D. Deugo. The evolution of hierarchical representations. Lecture Notes in Computer Science, 929:302–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wagner:1995:AMD

[1213] G. P. Wagner. Adaptation and the modular design of organisms. Lecture Notes in Computer Science, 929:317– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaneko:1995:TDD

[1214] K. Kaneko and T. Yomo. A theory of differentiation with dynamic clustering. Lecture Notes in Computer Science, 929:329–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nolfi:1995:EAN

[1215] S. Nolfi and D. Parisi. Evolving artificial neural networks that develop in time. Lecture Notes in Computer Science, 929:353–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rocha:1995:CGA

[1216] L. Mateus Rocha. Contextual genetic algorithms: Evolving developmental rules. Lecture Notes in Computer Science, 929:368–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rutkowska:1995:CDD

[1217] J. C. Rutkowska. Can development be designed? what we may learn from the cog project. Lecture Notes in Computer Science, 929:383–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ono:1995:EOI

[1218] N. Ono, T. Ohira, and A. T. Rahmani. Emergent organization of interspecies communication in Q-learning artificial organisms. Lecture Notes in Computer Science, 929:396–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carneiro:1995:SNR

[1219] J. Carneiro and J. Stewart. Self and nonself revisited: Lessons from modelling the immune network. Lecture Notes in Computer Science, 929:406– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vaario:1995:FS

[1220] J. Vaario and K. Shimohara. On formation of structures. Lecture Notes in Computer Science, 929:421–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parisi:1995:LAM

[1221] D. Parisi and F. Cecconi. Learning in the active mode. Lecture Notes in Computer Science, 929:439–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pipe:1995:LSC

[1222] A. G. Pipe, B. Carse, T. C. Fogarty, and A. Winfield. Learning subjective “cognitive maps” in the presence of sensory-motor errors. Lecture Notes in Computer Science, 929:463–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

HautopLund:1995:SUS

[1223] H. Hautop Lund. Specialization under social conditions in shared environments. Lecture Notes in Computer Science, 929:477–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stanley:1995:IPD

[1224] E. A. Stanley, D. Ashlock, and M. D.

Smucker. Iterated Prisoner’s dilemma with choice and refusal of partners: Evolutionary results. Lecture Notes in Computer Science, 929:490–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adami:1995:ADA

[1225] C. Adami, C. Titus Brown, and M. R. Haggerty. Abundance-distributions in artificial life and stochastic models: “age and area” revisited. Lecture Notes in Computer Science, 929:503– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rasmussen:1995:ETS

[1226] S. Rasmussen and C. L. Barrett. Elements of a theory of simulation. Lecture Notes in Computer Science, 929: 515–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mulhauser:1995:SSP

[1227] G. R. Mulhauser. To simulate or not to simulate: A problem of minimising functional logic depth. Lecture Notes in Computer Science, 929:530–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sipper:1995:QUC

[1228] M. Sipper. Quasi-uniform computationuniversal cellular automata. Lecture Notes in Computer Science, 929:544– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tempesti:1995:NSR

[1229] G. Tempesti. A new self-reproducing cellular automaton capable of construction and computation. Lecture Notes in Computer Science, 929:555–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibathez:1995:SIB

[1230] J. Ibathez, D. Anabitarte, I. Azpeitia, and O. Barrera. Self-inspection based reproduction in cellular automata. Lecture Notes in Computer Science, 929: 564–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mataric:1995:ELP

[1231] M. J. Mataric. Evaluation of learning performance of situated embodied agents. Lecture Notes in Computer Science, 929:579–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Teimoorzadeh:1995:SDA

[1232] K. Teimoorzadeh. Seeing in the dark with artificial bats. Lecture Notes in Computer Science, 929:590–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lambrinos:1995:NAL

[1233] D. Lambrinos. Navigating with an adaptive light compass. Lecture Notes in Computer Science, 929:602– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zapata:1995:CAU

[1234] R. Zapata, P. Lepinay, and P. Deplanques. Collision avoidance using an egocentric memory of proximity. Lecture Notes in Computer Science, 929:614–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Correia:1995:UAV

[1235] L. Correia and A. Steiger-Garcao. A useful autonomous vehicle with a hierarchical behavior control. Lecture Notes in Computer Science, 929:625– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thompson:1995:EER

[1236] A. Thompson. Evolving electronic robot controllers that exploit hardware resources. Lecture Notes in Computer Science, 929:640–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scheier:1995:CSM

[1237] C. Scheier and R. Pfeifer. Classification as sensory-motor coordination: A case study on autonomous agents. Lecture Notes in Computer Science, 929:657–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cruse:1995:HPF

[1238] H. Cruse, C. Bartling, and T. Kindermann. High-pass filtered positive feedback. decentralized control of cooperation. Lecture Notes in Computer Science, 929:668–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pebody:1995:LAE

[1239] M. Pebody. Learning and adaptivity: Enhancing reactive behaviour architectures in real-world interaction systems. Lecture Notes in Computer Science, 929:679–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cherian:1995:IFM

[1240] S. Cherian and W. Troxell. Interactivistm: a functional model of representation for behavior-based systems. Lecture Notes in Computer Science, 929:691–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jakobi:1995:NRG

[1241] N. Jakobi, P. Husbands, and I. Harvey. Noise and the reality gap: The use of simulation in evolutionary robotics. Lecture Notes in Computer Science, 929:704–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tani:1995:EDS

[1242] J. Tani. Essential dynamical structure in learnable autonomous robots. Lecture Notes in Computer Science, 929: 721–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vainio:1995:OPR

[1243] M. Vainio, T. Schoenberg, A. Halmest, and P. Jakubik. Optimizing the performance of a robot society in structured environment through genetic algorithms. Lecture Notes in Computer Science, 929:733–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

May:1995:SGE

[1244] R. M. May, S. Bohoeffer, and M. A.

Nowak. Spatial games and evolution of cooperation. Lecture Notes in Computer Science, 929:749–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBourcier:1995:ASM

[1245] P. De Bourcier and M. Wheeler. Aggressive signaling meets adaptive receiving: Further experiments in synthetic behavioural ecology. Lecture Notes in Computer Science, 929:760– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fletcher:1995:MFB

[1246] R. P. Fletcher, C. Cannings, and P. G. Blackwell. Modelling foraging behaviour of ant colonies. Lecture Notes in Computer Science, 929:772– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lachmann:1995:CCA

[1247] M. Lachmann and G. Sella. The computationally complete ant colony: Global coordination in a system with no hierarchy. Lecture Notes in Computer Science, 929:784–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bourgine:1995:MCH

[1248] P. Bourgine and D. Snyers. Mimicry and coevolution of hedonic agents. Lecture Notes in Computer Science, 929: 801–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hashimoto:1995:ESG

[1249] T. Hashimoto and T. Ikegami. Evolution of symbolic grammar systems. Lecture Notes in Computer Science, 929: 812–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sulis:1995:DCA

[1250] W. Sulis. Driven cellular automata, adaptation and the binding problem. Lecture Notes in Computer Science, 929:824–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ouzounis:1995:FCL

[1251] C. Ouzounis, A. Valencia, J. Tamames, and P. Bork. The functional composition of living machines as a design principle for artificial organisms. Lecture Notes in Computer Science, 929: 843–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Higgs:1995:TRF

[1252] P. G. Higgs and S. R. Morgan. Thermodynamics of RNA folding: When is an RNA molecule in equilibrium. Lecture Notes in Computer Science, 929: 852–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Calabretta:1995:ALM

[1253] R. Calabretta, S. Nolfi, and D. Parisi.

An artificial life model for predicting the tertiary structure of unknown proteins that emulates the folding process. Lecture Notes in Computer Science, 929:862–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lahoz-Beltra:1995:ECE

[1254] R. Lahoz-Beltra and S. R. Hameroff.

Energy cost evaluation of computing capabilities in biomolecular and artificial matter. Lecture Notes in Computer Science, 929:876–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwefel:1995:CES

[1255] H.-P. Schwefel and G. Rudolph. Contemporary evolution strategies. Lecture Notes in Computer Science, 929:893– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hordijk:1995:UR

[1256] W. Hordijk and B. Manderick. The usefulness of recombination. Lecture Notes in Computer Science, 929:908– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bartscht:1995:ILI

[1257] E. Bartscht, J. Engel, and C. Mueller-Schloer. The investigation of lamarckian inheritance with classifier systems in a massively parallel simulation environment. Lecture Notes in Computer Science, 929:920–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiben:1995:OCM

[1258] A. E. Eiben, C. H. M. Van Kemenade, and J. N. Kok. Orgy in the computer: Multi-parent reproduction in genetic algorithms. Lecture Notes in Computer Science, 929:934–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lobo:1995:STN

[1259] S. Lobo, A. J. Garcia-Tejedor, R. Rodriguez-Galan, and L. Lopez. A simplification of the theory of neural groups selection for adaptive control. Lecture Notes in Computer Science, 929:946–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grossberg:1995:TUP

[1260] S. Grossberg. Are there universal principles of brain computation? Lecture Notes in Computer Science, 930:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MenendezdelaPrida:1995:MCN

[1261] L. Menendez de la Prida. Modeling cortical networks. Lecture Notes in Computer Science, 930:7–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mira:1995:COC

[1262] J. Mira, A. Manjarres, S. Ros, and A. E. Delgado. Cooperative organization of connectivity patterns and receptive fields in the visual pathway: Application to adaptive thresholding. Lecture Notes in Computer Science, 930:15–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alexandre:1995:NIA

[1263] F. Alexandre and F. Guyot. Neurobiological inspiration for the architecture and functioning of cooperating neural networks. Lecture Notes in Computer Science, 930:24–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duro:1995:SMB

[1264] R. J. Duro, J. Santos, and A. Gomez.

Synaptic modulation based artificial neural networks. Lecture Notes in Computer Science, 930:31–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stetter:1995:SOC

[1265] M. Stetter, M. Kussinger, A. Schels, and E. Seeger. Self-organization of cortical receptive fields and columnar structures in a hebb trained neural network. Lecture Notes in Computer Science, 930:37–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hoekstra:1995:ASC

[1266] J. Hoekstra and M. Maouli. An analytical solution of the compartmental model for use in local learning in artificial neural networks. Lecture Notes in Computer Science, 930:45–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wallace:1995:SAA

[1267] J. G. Wallace and K. Bluff. Should ANN be ANGN? Lecture Notes in Computer Science, 930:53–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lourens:1995:MRH

[1268] T. Lourens. Modeling retinal high and low contrast sensitivity filters. Lecture Notes in Computer Science, 930:61–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miro:1995:NMS

[1269] J. Miro. Neural micro-structures. three simple models. Lecture Notes in Computer Science, 930:69–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daffertshofer:1995:AHS

[1270] A. Daffertshofer and H. Haken. Adaptive hierarchical structures. Lecture Notes in Computer Science, 930:76–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andreu:1995:ORI

[1271] E. Andreu, B. Soria, S. Bolea, and J. V.

Sanchez-Andres. Optimal range of input resistance in the oscillatory behavior of the pancreatic β-cell. Lecture Notes in Computer Science, 930:85–89, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kruizinga:1995:CMP

[1272] P. Kruizinga and N. Petkov. A computational model of periodic-pattern-selective cells. Lecture Notes in Computer Science, 930:90–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rybak:1995:MAS

[1273] I. A. Rybak, J. F. R. Paton, and J. S. Schwaber. Modeling and analysis of some neural mechanisms for the genesis and control of respiratory pattern. Lecture Notes in Computer Science, 930:100–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moran:1995:NNM

[1274] F. Moran and M. A. Andrade. A neural network model for plasticity in adult striate cortex. Lecture Notes in Computer Science, 930:108–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valderrama:1995:RTN

[1275] E. Valderrama, R. Villa, E. Cabruja, and P. Garrido. Regenerative-type neural interface. Lecture Notes in Computer Science, 930:114–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Villa:1995:NPA

[1276] R. Villa and J. Aguilo. New perspectives in auditory coding: Bases for a new cochlear behavioural model. Lecture Notes in Computer Science, 930: 121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mpodozis:1995:NSC

[1277] J. Mpodozis, J.-C. Letelier, and H. Maturana. Nervous system as a closed neural network: Behavioral and cognitive consequences. Lecture Notes in Computer Science, 930:130– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fernandez:1995:LAP

[1278] M. A. Fernandez, J. Mira, M. T. Lopez, and J. R. Alvarez. Local accumulation of persistent activity at synaptic level: Application to motion analysis. Lecture Notes in Computer Science, 930:137–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gratha:1995:HOB

[1279] M. Gratha, A. D’Anjou, F. X. Albizuri, and A. De la Hera. High order Boltzmann machines with continuous units: Some experimental results. Lecture Notes in Computer Science, 930:144–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nishii:1995:ACM

[1280] J. Nishii. An adaptive control model of a locomotion by the central pattern generator. Lecture Notes in Computer Science, 930:151–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grabec:1995:SCN

[1281] I. Grabec. Self-consistent neural receptive fields. Lecture Notes in Computer Science, 930:158–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofman:1995:SPN

[1282] P. M. Hofman, F. B. Rodriguez, J. A.

Sigueenza, and V. Lopez. A simple probabilistic neural model producing multimodal ISHs. Lecture Notes in Computer Science, 930:166–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

SignesPont:1995:ANN

[1283] M. T. Signes Pont and J. V. Sanchez-Andres. An associative neural network to model the developing mammalian hippocampus. Lecture Notes in Computer Science, 930:174–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weitze:1995:IPL

[1284] M.-D. Weitze and G. L. Hofacker. The implementation of propositional logic in random neural networks. Lecture Notes in Computer Science, 930:180– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stone:1995:LRE

[1285] J. V. Stone and A. Bray. A learning rule for extracting temporal invariances. Lecture Notes in Computer Science, 930:189–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Han:1995:ISF

[1286] J. Han and C. Moraga. The influence of the sigmoid function parameters on the speed of backpropagation learning. Lecture Notes in Computer Science, 930:195–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ho:1995:GTL

[1287] A. M. C.-L. Ho and P. De Wilde. General transient length upper bound for recurrent neural networks. Lecture Notes in Computer Science, 930:202– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreno-Diaz:1995:SMN

[1288] R. Moreno-Diaz and K. N. Leibovic. On some methods in neuromathematics. Lecture Notes in Computer Science, 930:209–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Farhat:1995:LND

[1289] N. H. Farhat and E. Del Moral Hernandez. Logistic networks with DNA-Like encoding and interactions. Lecture Notes in Computer Science, 930: 216–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christodoulou:1995:RSF

[1290] C. Christodoulou and T. Clarkson.

A review on the stochastic firing behaviour of real neurons and how it can be modelled. Lecture Notes in Computer Science, 930:223–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Augereau:1995:BGA

[1291] B. Augereau, T. Simon, J. Bernard, and B. Heit. The BP-gammaL1 algorithm: Non-chaotic and accelerated learning in a MLP network. Lecture Notes in Computer Science, 930:231– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jasic:1995:APB

[1292] T. Jasic and H. L. Poh. Analysis of pruning in backpropagation networks for artificial and real world mapping problems. Lecture Notes in Computer Science, 930:239–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuzmina:1995:ONH

[1293] M. G. Kuzmina, E. A. Manykin, and I. I. Surina. Oscillatory networks with Hebbian matrix of connections. Lecture Notes in Computer Science, 930: 246–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gimenez:1995:NAI

[1294] V. Gimenez, P. Gomez-Vilda, E. Torrano, and M. Perez-Castellanos. A new algorithm for implementing a recursive neural network. Lecture Notes in Computer Science, 930:252–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BarahonadaFonseca:1995:VIP

[1295] J. Barahona da Fonseca, I. Barahona da Fonseca, and J. Simoes da Fonseca. Visual information processing from the viewpoint of symbolic operations. Lecture Notes in Computer Science, 930:260–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

BarahonadaFonseca:1995:DCB

[1296] I. Barahona da Fonseca, J. Barahona da Fonseca, and J. Simoes da Fonseca. Dentritic computation in the brain. Lecture Notes in Computer Science, 930:268–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hruby:1995:SNM

[1297] P. Hruby. Stochastic neuronal models with realistic synaptic inputs and oscillatory inputs. Lecture Notes in Computer Science, 930:276–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joya:1995:NPC

[1298] G. Joya and F. Sandoval. A neural paradigm for controlling autonomous systems with reflex behaviour and learning capability. Lecture Notes in Computer Science, 930:283–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez:1995:FAA

[1299] A. M. Gonzalez, C. Santa Cruz, V. Lopez, and J. R. Dorronsoro. Fast automatic architecture selection in RBF networks. Lecture Notes in Computer Science, 930:291–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viana:1995:CBC

[1300] L. Viana. Collective behaviour of a chain of Hopfield subnetworks interconnected unidirectionally. Lecture Notes in Computer Science, 930:298– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Butchart:1995:CST

[1301] K. Butchart, N. Davey, and R. Adams.

A comparative study of three neural networks that use soft competition. Lecture Notes in Computer Science, 930:308–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dracopoulos:1995:NNG

[1302] D. C. Dracopoulos and A. J. Jones.

Neural networks and genetic algorithms for the attitude control problem. Lecture Notes in Computer Science, 930:315–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flanagan:1995:SOA

[1303] J. A. Flanagan and M. Hasler. Selforganising artificial neural networks. Lecture Notes in Computer Science, 930:322–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stocker:1995:DCB

[1304] E. Stocker, Y. Lecourtier, and A. Ennaji. A distributed classifier based on Yprel networks cooperation. Lecture Notes in Computer Science, 930:330– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kosak:1995:FSM

[1305] A. Kosak and K. Goser. A fractal selforganizing map with partially chaotic neurons. Lecture Notes in Computer Science, 930:338–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cervera:1995:MSO

[1306] E. Cervera and A. P. Del Pobil. Multiple self-organizing maps for supervised learning. Lecture Notes in Computer Science, 930:345–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feng:1995:ASA

[1307] J. Feng and B. Tirozzi. An application of the saturated attractor analysis to three typical models. Lecture Notes in Computer Science, 930:353–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jutten:1995:LEN

[1308] C. Jutten. Learning in evolute neural architectures: An ill-posed problem? Lecture Notes in Computer Science, 930:361–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelbrecht:1995:ASU

[1309] A. P. Engelbrecht, I. Cloete, J. Geldenhuys, and J. M. Zurada. Automatic scaling using gamma learning for feedforward neural networks. Lecture Notes in Computer Science, 930:374– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelbrecht:1995:DSI

[1310] A. P. Engelbrecht, I. Cloete, and J. M.

Zurada. Determining the significance of input parameters using sensitivity analysis. Lecture Notes in Computer Science, 930:382–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aizenberg:1995:MVN

[1311] N. N. Aizenberg, I. N. Aizenberg, and G. A. Krivosheev. Multi-valued neurons: Learning, networks, application to image recognition and extrapolation of temporal series. Lecture Notes in Computer Science, 930:389–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhao:1995:IEA

[1312] Q. Zhao and T. Higuchi. Individual evolutionary algorithm and its application to learning of nearest neighbor based MLP. Lecture Notes in Computer Science, 930:396–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voz:1995:PVS

[1313] J.-L. Voz, M. Verleysen, P. Thissen, and J.-D. Legat. A practical view of suboptimal Bayesian classification with radial Gaussian kernels. Lecture Notes in Computer Science, 930:404– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corbacho:1995:SBL

[1314] F. J. Corbacho and M. A. Arbib. Schema based learning and learning to detour. Lecture Notes in Computer Science, 930:412–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goeppert:1995:NCV

[1315] J. Goeppert and W. Rosenstiel. Neurons with continuous varying activation in self-organizing maps. Lecture Notes in Computer Science, 930:419– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trejo:1995:IBP

[1316] L. A. Trejo and C. Sandoval. Improving back-propagation: Epsilonback-propagation. Lecture Notes in Computer Science, 930:427–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castatho:1995:FSA

[1317] M. A. Castatho, E. Vidal, and F. Casacuberta. Finite state automata and connectionist machines. Lecture Notes in Computer Science, 930:433– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guan:1995:LTP

[1318] Y. Guan, T. Clarkson, and J. G.

Taylor. Learning transformed prototypes (LTP) — A statistical pattern classification technique of neural networks. Lecture Notes in Computer Science, 930:441–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

FerreiroGarcia:1995:FFE

[1319] R. Ferreiro Garcia and F. J. Perez Castelo. Fuzzy function estimators as basis on learning from experience. Lecture Notes in Computer Science, 930: 448–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Flexer:1995:CSF

[1320] A. Flexer. Connectionists and statisticians, friends or foes? Lecture Notes in Computer Science, 930:454–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leisenberg:1995:UNN

[1321] M. Leisenberg. Unsupervised neural networks for speech perception with cochlear implant systems for the profoundly deaf. Lecture Notes in Computer Science, 930:462–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zalama:1995:OAM

[1322] E. Zalama, P. Gaudiano, and J. L.

Lopez Coronado. Obstacle avoidance by means of an operant conditioning model. Lecture Notes in Computer Science, 930:471–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MorcegoSeix:1995:QAG

[1323] B. Morcego Seix, A. Catala Mallofre, and N. Piera Carrete. Qualitative approach to gradient based learning algorithms. Lecture Notes in Computer Science, 930:478–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vico:1995:CRN

[1324] F. J. Vico, F. Ortega, J. Almaraz, and F. Sandoval. Character recognition with neural assemblies. Lecture Notes in Computer Science, 930:486– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreter:1995:LAS

[1325] Z. Schreter. Learning by attentional scanning. Lecture Notes in Computer Science, 930:492–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lis:1995:SRN

[1326] J. Lis. The synthesis of the ranked neural networks applying genetic algorithm with the dynamic probability of mutation. Lecture Notes in Computer Science, 930:498–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marin:1995:GVL

[1327] F. J. Marin, F. Garcia, and F. Sandoval. Global versus local heuristic terminal attractor. Lecture Notes in Computer Science, 930:505–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanstein:1995:DLR

[1328] A. Kanstein and K. Goser. Dynamic learning of radial basis functions for fuzzy clustering. Lecture Notes in Computer Science, 930:513–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chentouf:1995:ILS

[1329] R. Chentouf and C. Jutten. Incremental learning with a stopping criterion — experimental results. Lecture Notes in Computer Science, 930:519–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benitez-Diaz:1995:LAG

[1330] D. Benitez-Diaz and J. Garcia-Quesada. Learning algorithm with Gaussian membership function for fuzzy RBF neural networks. Lecture Notes in Computer Science, 930:527– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thimm:1995:NNI

[1331] G. Thimm and E. Fiesler. Neural network initialization. Lecture Notes in Computer Science, 930:535–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nejad:1995:BNN

[1332] A. F. Nejad and T. D. Gedeon. Bidirectional neural networks reduce generalisation error. Lecture Notes in Computer Science, 930:543–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gedeon:1995:BBV

[1333] T. D. Gedeon, P. M. Wong, and D. Harris. Balancing bias and variance: Network topology and pattern set reduction techniques. Lecture Notes in Computer Science, 930:551–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puzenat:1995:PAN

[1334] D. Puzenat. Priming an artificial neural classifier. Lecture Notes in Computer Science, 930:559–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aleksander:1995:ANU

[1335] I. Aleksander. Artificial neuroconsciousness an update. Lecture Notes in Computer Science, 930:566–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeWilde:1995:PLP

[1336] P. De Wilde. Physical and linguistic problems in the modelling of consciousness by neural networks. Lecture Notes in Computer Science, 930:584– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miura:1995:NNM

[1337] K.-I. Miura and T. Nagano. A neural network model for the velocity vector of an object and its consistency with psychological phenomena. Lecture Notes in Computer Science, 930:589–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crestani:1995:IER

[1338] F. Crestani. Implementation and evaluation of a relevance feedback device based on neural networks. Lecture Notes in Computer Science, 930:597– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carrasco:1995:SOR

[1339] R. C. Carrasco and M. L. Forcada.

Second-order recurrent neural networks can learn regular grammars from noisy strings. Lecture Notes in Computer Science, 930:605–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viktor:1995:EDR

[1340] H. L. Viktor and I. Cloete. Extracting DNF rules from artificial neural networks. Lecture Notes in Computer Science, 930:611–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prem:1995:DSG

[1341] E. Prem. Dynamic symbol grounding, state construction and the problem of teleology. Lecture Notes in Computer Science, 930:619–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monte:1995:AIE

[1342] E. Monte, J. M. Calvet, and S. Vilarrubla. Analysis of industrial economics by means of neural nets. Lecture Notes in Computer Science, 930: 627–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diaz-Otero:1995:ESF

[1343] F. Diaz-Otero, A. Caballero, A. Lorenzo, and J. A. Sigueenza. Effects of spatial frequency and stimulus size on the orientation sensitivity of humans. Lecture Notes in Computer Science, 930:634–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Adamo:1995:OOD

[1344] J. M. Adamo and D. Anguita. Object oriented design of a simulator for large BP neural networks. Lecture Notes in Computer Science, 930:642–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Varona:1995:IXN

[1345] P. Varona and J. A. Sigueenza. Introducing XSim: A neural network simulator that incorporates biological parameters. Lecture Notes in Computer Science, 930:650–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Santos:1995:NHC

[1346] J. Santos, R. P. Otero, and J. Mira. NETTOOL: A hybrid connectionistsymbolic development environment. Lecture Notes in Computer Science, 930:658–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chiaberge:1995:SDE

[1347] M. Chiaberge, G. Di Bene, S. Di Pascoli, and R. Lambert. EL-SIM: a development environment for neuro-fuzzy intelligent controllers. Lecture Notes in Computer Science, 930:666–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheneval:1995:PIE

[1348] Y. Cheneval. Packlib, an interactive environment to develop modular software for data processing. Lecture Notes in Computer Science, 930:673–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weitzenfeld:1995:NNS

[1349] A. Weitzenfeld. NSL-Neural simulation language. Lecture Notes in Computer Science, 930:683–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Torralba:1995:LCA

[1350] A. Torralba. Low-cost accelerator for the simulation of cellular neural networks. Lecture Notes in Computer Science, 930:689–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thissen:1995:VSN

[1351] P. Thissen, M. Verleysen, J.-D. Legat, and J. Madrenas. A VLSI system for neural Bayesian and LVQ classification. Lecture Notes in Computer Science, 930:696–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thissen:1995:APD

[1352] P. Thissen, M. Verleysen, and J.- D. Legat. An associative processor dedicated to classification by neural methods. Lecture Notes in Computer Science, 930:704–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Martin-del-Brio:1995:HOM

[1353] B. Martin del Brio and J. Blasco-Alberto. Hardware-oriented models for VLSI implementation of self-organizing maps. Lecture Notes in Computer Science, 930:712–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roth:1995:HRS

[1354] U. Roth, A. Jahnke, and H. Klar. Hardware requirements for spike-processing neural networks. Lecture Notes in Computer Science, 930:720–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pelayo:1995:VAI

[1355] F. J. Pelayo, E. Ros, P. Martin-Smith, and F. J. Fernandez. A VLSI approach to the implementation of additive and shunting neural networks. Lecture Notes in Computer Science, 930: 728–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anguita:1995:LPA

[1356] M. Anguita, F. J. Pelayo, F. J. Fernandez, and A. Prieto. A low-power analog implementation of cellular neural networks. Lecture Notes in Computer Science, 930:736–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benitez:1995:APB

[1357] J. M. Benitez, J. Ortega, and I. Requena. Asynchronously parallel Boltzmann machines mapped onto distributed-memory multiprocessors. Lecture Notes in Computer Science, 930:744–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castillo:1995:CCF

[1358] F. Castillo, J. A. Garcia, J. M. Moreno, and J. Cabestany. A coprocessor card for fast neural network emulation. Lecture Notes in Computer Science, 930:752–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moreno:1995:DHI

[1359] J. M. Moreno, J. Madrenas, S. San Anselmo, and F. Castillo. Digital hardware implementation of ROI incremental algorithms. Lecture Notes in Computer Science, 930:761–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maria:1995:CIR

[1360] N. Maria, A. Guerin-Dugue, J. M.

Moreno, and F. Blayo. Comparing implementations of radial basis function neural networks on three parallel machines. Lecture Notes in Computer Science, 930:771–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blayo:1995:IRB

[1361] F. Blayo, A. Guerin-Dugue, and N. Maria. Implementing radial basis functions neural networks on the systolic MANTRA machine. Lecture Notes in Computer Science, 930:781– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Torralba:1995:MPS

[1362] A. Torralba, F. Colodro, and L. G. Franquelo. A mixed parallel-sequential SHNN for large networks. Lecture Notes in Computer Science, 930:789– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vermesan:1995:MVA

[1363] O. Vermesan. A modular VLSI architecture for neural networks implementation. Lecture Notes in Computer Science, 930:794–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Strey:1995:MPN

[1364] A. Strey, N. Avellana, R. Holgado, and J. A. Fernandez. A massively parallel neurocomputer with a reconfigurable arithmetical unit. Lecture Notes in Computer Science, 930:800–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saxena:1995:AOF

[1365] I. Saxena and E. Fiesler. An all-optical forward propagation multilayer neural network. Lecture Notes in Computer Science, 930:807–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mayes:1995:VCM

[1366] D. J. Mayes and A. Hamilton. A VLSI current mode synapse chip. Lecture Notes in Computer Science, 930:815– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beiu:1995:OMN

[1367] V. Beiu and J. G. Taylor. Optimal mapping of neural networks onto FPGA’s — A new constructive algorithm. Lecture Notes in Computer Science, 930:822–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miranda:1995:CSW

[1368] E. Miranda and L. M. Reyneri. A CPWM synapsis for weighted radial basis functions. Lecture Notes in Computer Science, 930:830–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernier:1995:TPG

[1369] J. L. Bernier, J. J. Merelo, J. Ortega, and A. Prieto. Test pattern generation for analog circuits using neural networks and evolutive algorithms. Lecture Notes in Computer Science, 930:838–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herault:1995:ASP

[1370] J. Herault. About some perception problems in neural networks. Lecture Notes in Computer Science, 930:845– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vilaritho:1995:ONN

[1371] D. L. Vilaritho, D. Cabello, and A. Mosquera. Optimization neural networks for image segmentation. Lecture Notes in Computer Science, 930:860– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheung:1995:SRI

[1372] W. P. Cheung, C. K. Lee, and K. C. Li. Segmentation of range images: A neural network approach. Lecture Notes in Computer Science, 930:868–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ruiz-del-Solar:1995:NAP

[1373] J. Ruiz del Solar and M. Koeppen. A neural architecture for preattentive segmentation of sewage pipes video images. Lecture Notes in Computer Science, 930:875–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jaramillo-Moran:1995:CMG

[1374] M. A. Jaramillo-Moran, F. J. Lopez-Aligue, and M. Macias-Macias. A CNN model for grey scale image processing. Lecture Notes in Computer Science, 930:882–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Natowicz:1995:KSO

[1375] R. Natowicz. Kohonen’s self organizing maps for contour segmentation of gray level and color images. Lecture Notes in Computer Science, 930:890–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puntonet:1995:GBP

[1376] C. G. Puntonet, M. Rodriguez-Alvarez, and A. Prieto. A geometrical based procedure for source separation mapped to a neural network. Lecture Notes in Computer Science, 930:898– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

RuizGarcia:1995:QOC

[1377] A. Ruiz Garcia and F. J. Arcas Tunez.

Quasi-optimum combination of multilayer perceptrons for adaptive multiclass pattern recognition. Lecture Notes in Computer Science, 930:906– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Echanobe:1995:TRS

[1378] J. Echanobe, J. R. Gonzalez De Mendivil, and J. R. Garitagoitia. A text recognition system based on a neural network and on a deformed system. Lecture Notes in Computer Science, 930:913–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shams:1995:SRM

[1379] S. Shams. Simultaneous recognition of multiple objects using the MEM model. Lecture Notes in Computer Science, 930:919–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cho:1995:LHC

[1380] J.-W. Cho, S.-Y. Lee, and C. H.

Park. On-line handwritten character recognition by a hybrid method based on neural networks and pattern matching. Lecture Notes in Computer Science, 930:926–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

PerezMaroto:1995:AAS

[1381] J. Perez Maroto, Y. A. Dimitriadis, J. M. Cano Izquierdo, and J. L. Lopez Coronado. Analysis and application of the STORE neural model in recognizing handwritten symbols. Lecture Notes in Computer Science, 930:934– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ibarra-Pico:1995:AOA

[1382] F. Ibarra-Pico, J. M. Garcia-Chamizo, R. Rizo-Aldeguer, and D. Corredor Lacha. An adaptative orthogonal asociative memory and its application to character recognition. Lecture Notes in Computer Science, 930:942–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

AsensiMuthoz:1995:TCR

[1383] D. Asensi Muthoz, A. Almagro Leon, and F. Ibarra Pico. Texture classification on real time using semi-cover vector and an orthogonal neural network. Lecture Notes in Computer Science, 930:948–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Palagi:1995:ATS

[1384] P. M. Palagi and A. Guerin-Dugue. An architecture for texture segmentation: from energy features to region detection. Lecture Notes in Computer Science, 930:956–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomez-Vilda:1995:LBT

[1385] P. Gomez-Vilda, V. Rodellar, V. Nieto, and M. A. Hombrados. A latticebased time-delay neural network for speech processing. Lecture Notes in Computer Science, 930:963–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Noda:1995:AIR

[1386] I. Noda. Acquisition of internal representation by learning of identitymapping using overload learning. Lecture Notes in Computer Science, 930: 971–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bandera:1995:MCM

[1387] C. Bandera, I. M. Conde, J. Jerez, and M. Gonzalez. A multiacuity connectionist model for local speed estimation. Lecture Notes in Computer Science, 930:979–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arroyo:1995:UAN

[1388] F. Arroyo, A. Gonzalo, and J. R.

Hilera. Using artificial neural networks for ultrasonic signals processing from simple geometric shapes. Lecture Notes in Computer Science, 930:987– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braspenning:1995:NNA

[1389] P. J. Braspenning. Neural networks as associative devices. Lecture Notes in Computer Science, 931:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weijters:1995:BNG

[1390] A. J. M. M. Weijters and G. A. J. Hoppenbrouwers. Backpropagation networks for grapheme-phoneme conversion. Lecture Notes in Computer Science, 931:11–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henseler:1995:BP

[1391] J. Henseler. Back propagation. Lecture Notes in Computer Science, 931:37–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peters:1995:P

[1392] H. J. M. Peters. Perceptrons. Lecture Notes in Computer Science, 931:67–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vrieze:1995:KN

[1393] O. J. Vrieze. Kohonen network. Lecture Notes in Computer Science, 931:83–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Postma:1995:ART

[1394] E. O. Postma and P. T. W. Hudson. Adaptive resonance theory. Lecture Notes in Computer Science, 931:101– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spieksma:1995:BM

[1395] F. C. R. Spieksma. Boltzmann machines. Lecture Notes in Computer Science, 931:119–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lenting:1995:RIB

[1396] J. H. J. Lenting. Representation issues in Boltzmann machines. Lecture Notes in Computer Science, 931:131– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Postma:1995:ON

[1397] E. O. Postma. Optimization networks. Lecture Notes in Computer Science, 931:145–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crama:1995:LSC

[1398] Y. Crama, A. W. J. Kolen, and E. J.

Pesch. Local search in combinatorial optimization. Lecture Notes in Computer Science, 931:157–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boekhoudt:1995:PIC

[1399] P. Boekhoudt. Process identification and control. Lecture Notes in Computer Science, 931:175–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanLuenen:1995:LCU

[1400] W. T. C. Van Luenen. Learning controllers using neural networks. Lecture Notes in Computer Science, 931:205– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cardon:1995:KIS

[1401] H. R. A. Cardon and R. Hoogstraten.

Key issues for successful industrial neural-network applications: an application in geology. Lecture Notes in Computer Science, 931:235–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Braspenning:1995:NC

[1402] P. J. Braspenning. Neural cognodynamics. Lecture Notes in Computer Science, 931:247–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hudson:1995:CUN

[1403] P. T. W. Hudson and E. O. Postma. Choosing and using a neural net. Lecture Notes in Computer Science, 931: 273–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lehman:1995:PIW

[1404] M. M. Lehman. Process improvement: The way forward. Lecture Notes in Computer Science, 932:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olive:1995:MEB

[1405] A. Olive and M.-R. Sancho. A method for explaining the behaviour of conceptual models. Lecture Notes in Computer Science, 932:12–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Burg:1995:CXL

[1406] J. F. M. Burg and R. P. Van de Riet. COLOR-X: Linguistically-based event modeling: A general approach to dynamic modeling. Lecture Notes in Computer Science, 932:26–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pastor-Collado:1995:STD

[1407] J. A. Pastor-Collado and A. Olive. Supporting transaction design in conceptual modelling of information systems. Lecture Notes in Computer Science, 932:40–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Opdahl:1995:FMP

[1408] A. L. Opdahl and G. Sindre. Facet models for problem analysis. Lecture Notes in Computer Science, 932:54–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duffy:1995:FRA

[1409] D. Duffy, C. MacNish, J. McDermid, and P. Morris. A framework for requirements analysis using automated reasoning. Lecture Notes in Computer Science, 932:68–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krogstie:1995:TDU

[1410] J. Krogstie, O. I. Lindland, and G. Sindre. Towards a deeper understanding of quality in requirements engineering. Lecture Notes in Computer Science, 932:82–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andonoff:1995:MIC

[1411] E. Andonoff, G. Hubert, A. Le Parc, and G. Zurfluh. Modelling inheritance, composition and relationship links between objects, object versions and class versions. Lecture Notes in Computer Science, 932:96–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dattolo:1995:HVM

[1412] A. Dattolo and V. Loia. Hypertext version management in an actor-based framework. Lecture Notes in Computer Science, 932:112–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plihon:1995:MWW

[1413] V. Plihon and C. Rolland. Modelling ways-of-working. Lecture Notes in Computer Science, 932:126–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dignum:1995:MCB

[1414] F. Dignum and H. Weigand. Modelling communication between cooperative systems. Lecture Notes in Computer Science, 932:140–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aalto:1995:CAO

[1415] J.-M. Aalto. Challenges in applying objects large systems. Lecture Notes in Computer Science, 932:154–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Verhoef:1995:FFI

[1416] T. F. Verhoef and A. H. M. Ter Hofstede. Feasibility of flexible information modelling support. Lecture Notes in Computer Science, 932:168–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herbst:1995:MMB

[1417] H. Herbst. A meta-model for business rules in systems analysis. Lecture Notes in Computer Science, 932:186– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rossi:1995:MME

[1418] M. Rossi and S. Brinkkemper. Metrics in method engineering. Lecture Notes in Computer Science, 932:200– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shklar:1995:IUA

[1419] L. Shklar, A. Sheth, V. Kashyap, and K. Shah. InfoHarness: Use of automatically generated metadata for search and retrieval of heterogeneous information. Lecture Notes in Computer Science, 932:217–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pettersson:1995:DUI

[1420] M. Pettersson. Designing the user interface on top of a conceptual model. Lecture Notes in Computer Science, 932:231–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Viehstaedt:1995:GRM

[1421] G. Viehstaedt and M. Minas. Graphical representation and manipulation of complex structures based on a formal model. Lecture Notes in Computer Science, 932:243–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grundy:1995:PIS

[1422] J. C. Grundy and J. R. Venable. Providing integrated support for multiple development notations. Lecture Notes in Computer Science, 932:255– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bounab:1995:FAT

[1423] M. Bounab and C. Godart. A federated approach to tool integration. Lecture Notes in Computer Science, 932: 269–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gibson:1995:DKR

[1424] M. D. Gibson and K. Conheeney. Domain knowledge reuse during requirements engineering. Lecture Notes in Computer Science, 932:283–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geppert:1995:STR

[1425] A. Geppert and K. R. Dittrich. Strategies and techniques: Reusable artifacts for the construction of database management systems. Lecture Notes in Computer Science, 932:297–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bracchi:1995:INI

[1426] G. Bracchi. The impact of new information architectures on industry and government transformation. Lecture Notes in Computer Science, 932:311– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauh:1995:STN

[1427] O. Rauh and E. Stickel. Standard transformations for the normalization of ER schemata. Lecture Notes in Computer Science, 932:313–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Albrecht:1995:RAD

[1428] M. Albrecht, E. Buchholz, A. Duesterhoeft, and M. Altus. The rapid application and database development (RADD) workbench — A comfortable database design tool. Lecture Notes in Computer Science, 932:327–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Siau:1995:PSU

[1429] K. Siau, Y. Wand, and I. Benbasat. A psychological study on the use of relationship concept — some preliminary findings. Lecture Notes in Computer Science, 932:341–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parmakson:1995:ASQ

[1430] P. Parmakson. Alignment of software quality and service quality. Lecture Notes in Computer Science, 932:355– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Narat:1995:GSM

[1431] V. Narat and A. Vila. A guide for software maintenance evaluation: Experience report. Lecture Notes in Computer Science, 932:366–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laitinen:1995:NNS

[1432] K. Laitinen. Natural naming in software development: Feedback from practitioners. Lecture Notes in Computer Science, 932:375–??, 1995. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aspinall:1995:SST

[1433] D. Aspinall. Subtyping with singleton types. Lecture Notes in Computer Science, 933:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bono:1995:SFH

[1434] V. Bono and L. Liquori. A subtyping for the Fisher-Honsell-Mitchell lambda calculus of objects. Lecture Notes in Computer Science, 933:16–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brauener:1995:GTE

[1435] T. Brauener. The girard translation extended with recursion. Lecture Notes in Computer Science, 933:31–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Compagnoni:1995:DHO

[1436] A. B. Compagnoni. Decidability of higher-order subtyping with intersection types. Lecture Notes in Computer Science, 933:46–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Herbelin:1995:LCS

[1437] H. Herbelin. A lambda-calculus structure isomorphic to Gentzen-style sequent calculus structure. Lecture Notes in Computer Science, 933:61–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuper:1995:UFD

[1438] J. Kuper. Usability: Formalising (un)definedness in typed lambda calculus. Lecture Notes in Computer Science, 933:76–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zaionc:1995:LRO

[1439] M. Zaionc. Lambda representation of operations between different term algebras. Lecture Notes in Computer Science, 933:91–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baaz:1995:SUG

[1440] M. Baaz and G. Salzer. Semiunification and generalizations of a particularly simple form. Lecture Notes in Computer Science, 933:106–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benton:1995:MLN

[1441] N. Benton. A mixed linear and non-linear logic: Proofs, terms and models. Lecture Notes in Computer Science, 933:121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gordeev:1995:CFF

[1442] L. Gordeev. Cut free formalization of logic with finitely many variables part I. Lecture Notes in Computer Science, 933:136–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pudlak:1995:HLB

[1443] P. Pudlak and S. R. Buss. How to lie without being (easily) convicted and the lengths of proofs in propositional calculus. Lecture Notes in Computer Science, 933:151–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courcelle:1995:MSO

[1444] B. Courcelle. Monadic second-order logic and linear orderings of finite structures. Lecture Notes in Computer Science, 933:163–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Durand:1995:FOS

[1445] A. Durand and S. Ranaivoson. Firstorder spectra with one binary predicate. Lecture Notes in Computer Science, 933:177–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Grandjean:1995:MLD

[1446] E. Grandjean and F. Olive. Monadic logical definability of NP-Complete problems. Lecture Notes in Computer Science, 933:190–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lautemann:1995:LCF

[1447] C. Lautemann, T. Schwentick, and D. Therien. Logics for context-free languages. Lecture Notes in Computer Science, 933:205–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Malmstroem:1995:LAM

[1448] A. Malmstroem. Log-approximable minimization problems on random inputs. Lecture Notes in Computer Science, 933:217–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McArthur:1995:CLL

[1449] M. McArthur. Convergence and 0-1 laws for L0k,omega under arbitrary measures. Lecture Notes in Computer Science, 933:228–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stolboushkin:1995:FOC

[1450] A. P. Stolboushkin and M. A. Taitslin. Is first order contained in an initial segment of PTIME? Lecture Notes in Computer Science, 933:242–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finkelstein:1995:LPT

[1451] S. E. Finkelstein, P. Freyd, and J. Lipton. Logic programming in tau categories. Lecture Notes in Computer Science, 933:249–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kriauciukas:1995:RRS

[1452] V. Kriauciukas and M. Walicki. Reasoning and rewriting with set-relations I: Ground completeness. Lecture Notes in Computer Science, 933:264– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeNivelle:1995:RGN

[1453] H. De Nivelle. Resolution games and non-liftable resolution orderings. Lecture Notes in Computer Science, 933: 279–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schulz:1995:ETL

[1454] K. U. Schulz. On existential theories of list concatenation. Lecture Notes in Computer Science, 933:294–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tammet:1995:CRD

[1455] T. Tammet. Completeness of resolution for definite answers with case analysis. Lecture Notes in Computer Science, 933:309–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voda:1995:SBF

[1456] P. J. Voda. Subrecursion as a basis for a feasible programming language. Lecture Notes in Computer Science, 933:324–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crole:1995:SMS

[1457] R. L. Crole and A. D. Gordon. A sound metalogical semantics for input/output effects. Lecture Notes in Computer Science, 933:339–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fairtlough:1995:IML

[1458] M. Fairtlough and M. Mendler. An intuitionistic modal logic with applications to the formal verification of hardware. Lecture Notes in Computer Science, 933:354–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lester:1995:TMC

[1459] D. Lester and S. Mintchev. Towards machine-checked compiler correctness for higher-order pure functional languages. Lecture Notes in Computer Science, 933:369–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moschovakis:1995:PPF

[1460] Y. N. Moschovakis and G. T. Whitney.

Powerdomains, powerstructures and fairness. Lecture Notes in Computer Science, 933:382–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piessens:1995:CFD

[1461] F. Piessens and E. Steegmans. Canonical forms for data-specifications. Lecture Notes in Computer Science, 933: 397–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hermida:1995:AVS

[1462] C. Hermida and B. Jacobs. An algebraic view of structural induction. Lecture Notes in Computer Science, 933: 412–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hofmann:1995:ITT

[1463] M. Hofmann. On the interpretation of type theory in locally Cartesian closed categories. Lecture Notes in Computer Science, 933:427–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chagrov:1995:AAP

[1464] A. V. Chagrov and V. B. Shehtman. Algorithmic aspects of propositional tense logics. Lecture Notes in Computer Science, 933:442–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cholewinski:1995:SDT

[1465] P. Cholewinski. Stratified default theories. Lecture Notes in Computer Science, 933:456–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Klarlund:1995:HCO

[1466] N. Klarlund. A homomorphism concept for omega-regularity. Lecture Notes in Computer Science, 933:471–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leivant:1995:RRC

[1467] D. Leivant and J.-Y. Marion. Ramified recurrence and computational complexity II: Substitution and polyspace. Lecture Notes in Computer Science, 933:486–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marandjian:1995:GFR

[1468] H. B. Marandjian. General form recursive equations I. Lecture Notes in Computer Science, 933:501–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rybakov:1995:MLP

[1469] V. V. Rybakov. Modal logics preserving admissible for S4 inference rules. Lecture Notes in Computer Science, 933:512–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sazonov:1995:BST

[1470] V. Y. Sazonov. A bounded set theory with anti-foundation axiom and inductive definability. Lecture Notes in Computer Science, 933:527–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Musen:1995:CBA

[1471] M. A. Musen, S. W. Tu, A. K. Das, and Y. Shahar. A component-based architecture for automation of protocoldirected therapy. Lecture Notes in Computer Science, 934:3–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rector:1995:CTK

[1472] A. L. Rector. Coordinating taxonomies: Key to re-usable concept representations. Lecture Notes in Computer Science, 934:17–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Binsted:1995:GPP

[1473] K. Binsted, A. Cawsey, and R. Jones.

Generating personalised patient information using the medical record. Lecture Notes in Computer Science, 934: 29–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rassinoux:1995:AMJ

[1474] A.-M. Rassinoux, C. Juge, P.-A. Michel, and R. H. Baud. Analysis of medical jargon: The RECIT system. Lecture Notes in Computer Science, 934:42–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smart:1995:MKR

[1475] J. F. Smart and M. Roux. Medical knowledge representation for medical report analysis. Lecture Notes in Computer Science, 934:53–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keravnou:1995:MMC

[1476] E. T. Keravnou. Modelling medical concepts as time-objects. Lecture Notes in Computer Science, 934:67–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chittaro:1995:MMR

[1477] L. Chittaro, M. Del Rosso, and M. Dojat. Modeling medical reasoning with the event calculus: An application to the management of mechanical ventilation. Lecture Notes in Computer Science, 934:79–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Larizza:1995:GFB

[1478] C. Larizza, G. Bernuzzi, and M. Stefanelli. A general framework for building patient monitoring systems. Lecture Notes in Computer Science, 934:91–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barahona:1995:SQM

[1479] P. Barahona. Semi-qualitative models and simulation for biomedical applications. Lecture Notes in Computer Science, 934:103–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ironi:1995:GEP

[1480] L. Ironi and M. Stefanelli. Generating explanations of pathophysiological systems behaviors from qualitative simulation of compartmental models. Lecture Notes in Computer Science, 934:115–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Carenini:1995:IBB

[1481] G. Carenini, S. Monti, and G. Banks.

An information-based Bayesian approach to history taking. Lecture Notes in Computer Science, 934:129– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ramoni:1995:MDM

[1482] M. Ramoni, A. Riva, M. Stefanelli, and V. Patel. Medical decision making using ignorant influence diagrams. Lecture Notes in Computer Science, 934: 139–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hejlesen:1995:DPC

[1483] O. K. Hejlesen, S. Andreassen, and S. K. Andersen. Dynamic propagation in causal probabilistic networks with instantiated variables. Lecture Notes in Computer Science, 934:151–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Safran:1995:ASP

[1484] E. Safran, D. Pittet, F. Borst, and G. Thurler. Alerts as starting point for hospital infection surveillance and control. Lecture Notes in Computer Science, 934:165–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lanzola:1995:CSA

[1485] G. Lanzola, S. Falasconi, and M. Stefanelli. Cooperative software agents for patient management. Lecture Notes in Computer Science, 934:173–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Riva:1995:HLC

[1486] A. Riva and R. Bellazzi. High level control strategies for diabetes therapy. Lecture Notes in Computer Science, 934:185–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Miksch:1995:TPU

[1487] S. Miksch, W. Horn, C. Popow, and F. Paky. Therapy planning using qualitative trend descriptions. Lecture Notes in Computer Science, 934:197– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1995:AAC

[1488] R. Schmidt, L. Boscher, B. Heindl, and G. Schmid. Adaptation and abstraction in a case-based antibiotics therapy adviser. Lecture Notes in Computer Science, 934:209–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Diamond:1995:FEK

[1489] L. W. Diamond, D. T. Nguyen, P. Ralph, and B. Sheridan. Field evaluations of a knowledge-based system for peripheral blood interpretation. Lecture Notes in Computer Science, 934:221–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Darmoni:1995:FES

[1490] S. J. Darmoni, P. Massari, J.-M. Droy, and T. Blanc. Functional evaluation of SETH: An expert system in clinical toxicology. Lecture Notes in Computer Science, 934:231–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Downs:1995:ENN

[1491] J. Downs, R. F. Harrison, and S. S.

Cross. Evaluating a neural network decision-support tool for the diagnosis of breast cancer. Lecture Notes in Computer Science, 934:239–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1995:KBS

[1492] D. T. Nguyen, I. A. Park, P. Cherubino, and P. B. Tamino. Knowledgebased systems for lymph node pathology: A comparison of two approaches. Lecture Notes in Computer Science, 934:251–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kindler:1995:MLM

[1493] H. Kindler, D. Densow, B. Fischer, and T. M. Fliedner. Mapping laboratory medicine onto the select and test model to facilitate knowledgebased report generation in laboratory medicine. Lecture Notes in Computer Science, 934:265–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ohmann:1995:MLT

[1494] C. Ohmann, Q. Yang, V. Moustakis, and K. Lang. Machine learning techniques applied to the diagnosis of acute abdominal pain. Lecture Notes in Computer Science, 934:276–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puppe:1995:RBM

[1495] B. Puppe. Reflections on building medical decision support systems and corresponding implementation in diagnostics shell D3. Lecture Notes in Computer Science, 934:282–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Quaglini:1995:DMC

[1496] S. Quaglini, M. Stefanelli, and F. Locatelli. Decision models for costeffectiveness analysis: A means for knowledge sharing and quality control in health care multidisciplinary tasks. Lecture Notes in Computer Science, 934:295–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alpay:1995:MBA

[1497] L. Alpay, A. Nowlan, D. Solomon, and C. Lovis. Model-based application:

The galen structured clinical user interface. Lecture Notes in Computer Science, 934:307–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kanoui:1995:KBM

[1498] H. Kanoui, M. Joubert, and R. Favard.

A knowledge-based modeling of hospital information systems components. Lecture Notes in Computer Science, 934:319–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lovis:1995:UCS

[1499] C. Lovis, P.-A. Michel, R. H. Baud, and J.-R. Scherrer. Use of a conceptual semi-automatic ICD-9 encoding system in an hospital environment. Lecture Notes in Computer Science, 934:331–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Waschulzik:1995:QAI

[1500] T. Waschulzik, W. Brauer, M. Foerster, and K. Kirchner. Quality assurance and increased efficiency in medical projects with neural networks by using a structured development method for feedforward neural networks (SENN). Lecture Notes in Computer Science, 934:343–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Downs:1995:PNN

[1501] J. Downs, R. F. Harrison, and R. L.

Kennedy. A prototype neural network decision-support tool for the early diagnosis of acute myocardial infarction. Lecture Notes in Computer Science, 934:355–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bianchi:1995:INN

[1502] N. Bianchi and C. Diamantini. Integration of neural networks and rule based systems in the interpretation of liver biopsy images. Lecture Notes in Computer Science, 934:367–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Spinu:1995:CAA

[1503] C. Spinu, C. Garbay, and J. M. Chassery. A cooperative and adaptive approach to medical image segmentation. Lecture Notes in Computer Science, 934:379–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsumoto:1995:CIK

[1504] S. Tsumoto, H. Tanaka, H. Amano, and K. Ohyama. COBRA: Integration of knowledge-bases with case-databases in the domain of congenital malformation. Lecture Notes in Computer Science, 934:393–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bichindaritz:1995:CBM

[1505] I. Bichindaritz. Case-based medical multi-expertise: An example in psychiatry. Lecture Notes in Computer Science, 934:395–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Combi:1995:TND

[1506] C. Combi, F. Pinciroli, and G. Pozzi. TIME-NESIS: A data model in managing time granularity of naturallanguage clinical information. Lecture Notes in Computer Science, 934:397– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tsumoto:1995:IES

[1507] S. Tsumoto and H. Tanaka. Induction of expert system rules from clinical databases based on rough set theory and resampling methods. Lecture Notes in Computer Science, 934:399– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brigl:1995:SKA

[1508] B. Brigl, A. Grau, P. Ringleb, and T. Steiner. Sequential knowledge acquisition: Combining models and cases. Lecture Notes in Computer Science, 934:401–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hajek:1995:MFE

[1509] P. Hajek and D. Harmancova. Medical fuzzy expert systems and reasoning about beliefs. Lecture Notes in Computer Science, 934:403–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurzynski:1995:DHA

[1510] M. Kurzynski, M. Wozniak, and A. Blinowska. Diagnosis of human acid-base balance states via combined pattern recognition of Markov chains. Lecture Notes in Computer Science, 934:405– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Olkhovsky:1995:IFP

[1511] I. F. Olkhovsky and S. I. Blokhina. Intelligence formation problems in children at an early age applying new computer technologies under conditions of rehabilitation center. Lecture Notes in Computer Science, 934:407–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McClelland:1995:T

[1512] I. McClelland, K. Adamson, and N. Black. Telecardiology. Lecture Notes in Computer Science, 934:409– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Steve:1995:MSM

[1513] G. Steve, A. Gangemi, and A. Rossi Mori. Modelling a sharable medical concept system: Ontological foundation in GALEN. Lecture Notes in Computer Science, 934:411–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DellaMea:1995:GBA

[1514] V. Della Mea, N. Finato, and C. A. Beltrami. A graph-based approach to the structural analysis of proliferative breast lesions. Lecture Notes in Computer Science, 934:413–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Taddei:1995:WCD

[1515] A. Taddei, M. Niccolai, M. Raciti, and C. Michelassi. A workstation for clinical decision support in a local area network for cardiology. Lecture Notes in Computer Science, 934:415–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Salzsieder:1995:KBE

[1516] E. Salzsieder, U. Fischer, A. Hierle, and U. Oppel. Knowledge-based education tool to improve quality in diabetes care. Lecture Notes in Computer Science, 934:417–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Keller:1995:NPD

[1517] F. Keller, R. Arnold, T. Frankewitsch, and D. Zellner. NEPHARM: A pharmacokinetic database for adjusting drug dosage to impaired renal function. Lecture Notes in Computer Science, 934:419–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christodoulou:1995:HAK

[1518] E. Christodoulou. A hybrid architecture for knowledge-based systems. Lecture Notes in Computer Science, 934: 421–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dojat:1995:RMC

[1519] M. Dojat and F. Pachet. Representing medical context using rulebased object-oriented programming techniques. Lecture Notes in Computer Science, 934:423–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ultsch:1995:INN

[1520] A. Ultsch, D. Korus, and T. O.

Kleine. Integration of neural networks and knowledge-based systems in medicine. Lecture Notes in Computer Science, 934:425–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puppe:1995:GCK

[1521] F. Puppe, B. Reinhardt, and K. Poeck. Generated critic in the knowledge based neurology trainer. Lecture Notes in Computer Science, 934:427– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Polikarpova:1995:AAQ

[1522] N. Polikarpova. An approach to analysis of qualitative data with insufficient number of quantization levels. Lecture Notes in Computer Science, 934:429– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gamberger:1995:ILR

[1523] D. Gamberger. Inductively learned rule for breast cancer domain with improved interobserver reproducibility. Lecture Notes in Computer Science, 934:431–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shahsavar:1995:DEK

[1524] N. Shahsavar and O. Wigertz. Development and evaluation of a knowledgebased system to support ventilator therapy management. Lecture Notes in Computer Science, 934:433–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiocchi:1995:NSP

[1525] F. Fiocchi, A. Gamba, R. Pizzi, and F. Sicurello. A neural support to the prognostic evaluation of cardiac surgery. Lecture Notes in Computer Science, 934:435–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zeilinger:1995:DSS

[1526] G. Zeilinger, J. De Mey, G. Gell, and G. Vrisk. DECISion-Support system for radiological diagnostic. Lecture Notes in Computer Science, 934:437– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caleb:1995:PIA

[1527] P. Caleb, P. K. Sharpe, and R. Jones. A preliminary investigation into the analysis of electromyographic activity using a system of multiple neural networks. Lecture Notes in Computer Science, 934:439–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sell:1995:KBS

[1528] C. Sell, S. Herfurth, A. Rutscher, and E. Salzsieder. Knowledge-based system to predict the effect of pregnancy on progression of diabetic retinopathy. Lecture Notes in Computer Science, 934:441–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anselmi:1995:SEM

[1529] R. Anselmi, G. Paoli, G. Ghiso, and F. Foppiano. A software to evaluate multislices radiotherapic treatment planning. Lecture Notes in Computer Science, 934:443–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heras:1995:TTE

[1530] J. Heras and R. P. Otero. TKRtool: An expert system for total knee replacement management. Lecture Notes in Computer Science, 934: 444–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Best:1995:RVB

[1531] E. Best and M. Koutny. A refined view of the box algebra. Lecture Notes in Computer Science, 935:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mattern:1995:DAC

[1532] F. Mattern. Distributed algorithms and causally consistent observations. Lecture Notes in Computer Science, 935:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Voss:1995:PNI

[1533] K. Voss. Petri nets for information systems modelling. Lecture Notes in Computer Science, 935:23–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barkaoui:1995:LEN

[1534] K. Barkaoui, J. M. Couvreur, and C. Dutheillet. On liveness in extended non self-controlling nets. Lecture Notes in Computer Science, 935:25–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basten:1995:ASH

[1535] T. Basten and M. Voorhoeve. An algebraic semantics for hierarchical P/ T nets. Lecture Notes in Computer Science, 935:45–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bastide:1995:PNB

[1536] R. Bastide and P. Palanque. A Petri net based environment for the design of event-driven interfaces. Lecture Notes in Computer Science, 935:66–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Battiston:1995:ISH

[1537] E. Battiston, O. Botti, E. Crivelli, and F. De Cindio. An incremental specification of a hydroelectric power plant control systems using a class of modular algebraic nets. Lecture Notes in Computer Science, 935:84–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Best:1995:CCH

[1538] E. Best, H. Fleischack, W. Fraczak, and R. P. Hopkins. A class of composable high level Petri nets with an application to the semantics of B(PN)02. Lecture Notes in Computer Science, 935:103–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brgan:1995:EAC

[1539] R. Brgan and D. Poitrenand. An efficient algorithm for the computation of stubborn sets of well formed Petri nets. Lecture Notes in Computer Science, 935:121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruno:1995:CED

[1540] G. Bruno, A. Castella, R. Agarwal, and M. P. Pescarmona. CAB: an environment for developing concurrent application. Lecture Notes in Computer Science, 935:141–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bucholz:1995:NES

[1541] P. Bucholz. A notion of equivalence for stochastic Petri nets. Lecture Notes in Computer Science, 935:161–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Caselli:1995:PSS

[1542] S. Caselli, G. Conte, and P. Marenzoni. Parallel state space exploration for GSPN models. Lecture Notes in Computer Science, 935:181–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Christensen:1995:MSS

[1543] S. Christensen and L. Petrucci. Modular state space analysis of coloured Petri nets. Lecture Notes in Computer Science, 935:201–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gomes:1995:PCD

[1544] L. Gomes and A. Steiger-Garcao. Programmable controller design based on a synchronized colored Petri net model and integrating fuzzy reasoning. Lecture Notes in Computer Science, 935: 218–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haddad:1995:SRG

[1545] S. Haddad, J. M. Ilie, M. Taghelit, and B. Zouari. Symbolic reachability graph and partial symmetries. Lecture Notes in Computer Science, 935:238– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katoen:1995:CBN

[1546] J.-P Katoen. Causal behaviours and nets. Lecture Notes in Computer Science, 935:258–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lakos:1995:CPN

[1547] C. Lakos. From coloured Petri nets to object Petri nets. Lecture Notes in Computer Science, 935:278–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lee:1995:HRA

[1548] D.-I. Lee, S. Kumagai, and S. Kodama. Handles and reachability analysis of free choice nets. Lecture Notes in Computer Science, 935:298–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liu:1995:PBS

[1549] Z. Liu. Performance bounds for stochastic timed Petri nets. Lecture Notes in Computer Science, 935:316– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Molla-Hosseini:1995:IMM

[1550] M. Molla-Hosseini, R. M. Kerr, R. B. Randall, and R. B. Platfoot. An inspection model with minimal and major maintenance for a flexible manufacturing cell using generalized stochastic Petri nets. Lecture Notes in Computer Science, 935:335–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reineke:1995:SRT

[1551] H. Reineke. Subclasses of recognizable trace languages. Lecture Notes in Computer Science, 935:357–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roig:1995:VAC

[1552] O. Roig, J. Cortadella, and E. Pastor. Verification of asynchronous circuits by BDD-based model checking of Petri nets. Lecture Notes in Computer Science, 935:374–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schmidt:1995:PRT

[1553] K. Schmidt. Parameterized reachability trees for algebraic Petri nets. Lecture Notes in Computer Science, 935: 392–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoef:1995:EST

[1554] S. Schoef, M. Sonnenschein, and R. Wieting. Efficient simulation of THOR nets. Lecture Notes in Computer Science, 935:412–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schreiber:1995:FEP

[1555] G. Schreiber. Functional equivalences of Petri nets. Lecture Notes in Computer Science, 935:432–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Senac:1995:HTS

[1556] P. Senac, P. De Saqui-Sannes, and R. Willrich. Hierarchical time stream Petri net: A model for hypermedia systems. Lecture Notes in Computer Science, 935:451–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Telek:1995:MRS

[1557] M. Telek and A. Bobbio. Markov regenerative stochastic Petri nets with age type general transitions. Lecture Notes in Computer Science, 935:471– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valero:1995:TPT

[1558] V. Valero, D. De Frutos, and F. Cuartero. Timed processes of timed Petri nets. Lecture Notes in Computer Science, 935:490–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gries:1995:ELT

[1559] D. Gries. Equational logic as a tool. Lecture Notes in Computer Science, 936:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wing:1995:TMS

[1560] J. M. Wing. Teaching mathematics to software engineers. Lecture Notes in Computer Science, 936:18–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ralston:1995:RET

[1561] T. Ralston, S. Gerhart, and D. Craigen. The role of education and training in the industrial application of formal methods. Lecture Notes in Computer Science, 936:41–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Orlowska:1995:IA

[1562] E. Orlowska. Information algebras. Lecture Notes in Computer Science, 936:50–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apt:1995:VLP

[1563] K. R. Apt and I. Luitjes. Verification of logic programs with delay declarations. Lecture Notes in Computer Science, 936:66–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goguen:1995:ICB

[1564] J. A. Goguen and R. Diaconescu. An introduction to category-based equational logic. Lecture Notes in Computer Science, 936:91–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parikh:1995:KBC

[1565] R. Parikh. Knowledge-based computation. Lecture Notes in Computer Science, 936:127–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haxthausen:1995:OSA

[1566] A. E. Haxthausen. Order-sorted algebraic specifications with higher-order functions. Lecture Notes in Computer Science, 936:133–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bidoit:1995:PCB

[1567] M. Bidoit and R. Hennicker. Proving the correctness of behavioural implementations. Lecture Notes in Computer Science, 936:152–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dam:1995:DPE

[1568] M. Dam. On the decidability of process equivalences for the π-calculus. Lecture Notes in Computer Science, 936:169– 183, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oriat:1995:DIM

[1569] C. Oriat. Detecting isomorphisms of modular specifications with diagrams. Lecture Notes in Computer Science, 936:184–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prehofer:1995:HON

[1570] C. Prehofer. Higher-order narrowing with convergent systems. Lecture Notes in Computer Science, 936:199– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badouel:1995:CFE

[1571] E. Badouel, P. Darondeau, and J.-C. Raoult. Context-free event domains are recognizable. Lecture Notes in Computer Science, 936:214–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Terrasse:1995:ENS

[1572] D. Terrasse. Encoding natural semantics in coq. Lecture Notes in Computer Science, 936:230–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jacobs:1995:MCC

[1573] B. Jacobs. Mongruences and cofree coalgebras. Lecture Notes in Computer Science, 936:245–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cengarle:1995:STP

[1574] M. V. Cengarle. Semantic typing for parametric algebraic specifications. Lecture Notes in Computer Science, 936:261–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jagadeesan:1995:CTC

[1575] L. J. Jagadeesan and R. Jagadeesan. Causality and true concurrency: A data-flow analysis of the pi-calculus. Lecture Notes in Computer Science, 936:277–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeAlfaro:1995:VCT

[1576] L. De Alfaro and Z. Manna. Verification in continuous time by discrete reasoning. Lecture Notes in Computer Science, 936:292–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ferrari:1995:DMC

[1577] G. Ferrari and U. Montanari. Dynamic matrices and the cost analysis of concurrent programs. Lecture Notes in Computer Science, 936:307–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:1995:PNT

[1578] A. Cheng. Petri nets, traces, and local model checking. Lecture Notes in Computer Science, 936:322–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leonard:1995:AFD

[1579] E. I. Leonard and A. E. Zwarico. An algebraic framework for developing and maintaining real-time systems. Lecture Notes in Computer Science, 936:338– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Collette:1995:LFC

[1580] P. Collette and E. Knapp. Logical foundations for compositional verification and development of concurrent programs in UNITY. Lecture Notes in Computer Science, 936:353–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corradini:1995:CMI

[1581] A. Corradini and F. Gadducci. CPO models for infinite term rewriting. Lecture Notes in Computer Science, 936: 368–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Montanari:1995:CRT

[1582] A. Montanari and M. De Rijke. Completeness results for two-sorted metric temporal logics. Lecture Notes in Computer Science, 936:385–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brown:1995:MPW

[1583] N. Brown and A. Mokkedem. On mechanizing proofs within a complete proof system for unity. Lecture Notes in Computer Science, 936:400–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kapur:1995:ARA

[1584] D. Kapur and M. Subramaniam. Automated reasoning about parallel algorithms using powerlists. Lecture Notes in Computer Science, 936:416– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dold:1995:RVA

[1585] A. Dold. Representing, verifying and applying software development steps using the PVS system. Lecture Notes in Computer Science, 936:431–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gogolla:1995:ADT

[1586] M. Gogolla and R. Herzig. An algebraic development technique for information systems. Lecture Notes in Computer Science, 936:446–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bumbulis:1995:FMA

[1587] P. Bumbulis, P. S. C. Alencar, D. D.

Cowan, and C. J. P. Lucena. A framework for machine-assisted user interface verification. Lecture Notes in Computer Science, 936:461–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Heisel:1995:SUF

[1588] M. Heisel. Specification of the Unix file system: A comparative case study. Lecture Notes in Computer Science, 936: 475–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Isobe:1995:CCB

[1589] Y. Isobe, Y. Sato, and K. Ohmaki.

A calculus of countable broadcasting systems. Lecture Notes in Computer Science, 936:489–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bergeron:1995:STD

[1590] A. Bergeron. Symbolic timing devices. Lecture Notes in Computer Science, 936:504–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bagai:1995:ACW

[1591] R. Bagai and R. Sunderraman. An algebraic construction of the wellfounded model. Lecture Notes in Computer Science, 936:518–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Falaschi:1995:CCC

[1592] M. Falaschi, M. Gabbrielli, K. Marriott, and C. Palamidessi. Confluence in concurrent constraint programming. Lecture Notes in Computer Science, 936:531–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rajagopalan:1995:GAD

[1593] P. Rajagopalan and C. P. Tsang. A generic algebra for data collections based on constructive logic. Lecture Notes in Computer Science, 936:546– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jayaraman:1995:POP

[1594] B. Jayaraman, M. Osorio, and K. Moon. Partial order programming. Lecture Notes in Computer Science, 936:561–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouhoula:1995:SSA

[1595] A. Bouhoula and M. Rusinowitch. SPIKE: A system for automatic inductive proofs. Lecture Notes in Computer Science, 936:576–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eusterbrock:1995:SKB

[1596] J. Eusterbrock. SEAMLESS: Knowledge based evolutionary system synthesis. Lecture Notes in Computer Science, 936:578–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jamil:1995:OOF

[1597] H. M. Jamil and L. V. S. Lakshmanan.

An object-oriented front-end for deductive databases. Lecture Notes in Computer Science, 936:581–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jayaraman:1995:SPF

[1598] B. Jayaraman and K. Moon. The SuRE programming framework. Lecture Notes in Computer Science, 936: 585–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lakshmanan:1995:DSM

[1599] L. V. S. Lakshmanan, I. N. Subramanian, D. Papoulis, and N. Shiri. A declarative system for multi-database interoperability. Lecture Notes in Computer Science, 936:586–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sahraoui:1995:MS

[1600] H. A. Sahraoui. The METAGEN system. Lecture Notes in Computer Science, 936:590–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bafna:1995:CSB

[1601] V. Bafna, S. Muthukrishnan, and R. Ravi. Computing similarity between RNA strings. Lecture Notes in Computer Science, 937:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benham:1995:CTM

[1602] C. Benham, S. Kannan, and T. Warnow. Of chicken teeth and mouse eyes, or generalized character compatibility. Lecture Notes in Computer Science, 937:17–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Breslauer:1995:ESM

[1603] D. Breslauer and L. Gasieniec. Efficient string matching on coded texts. Lecture Notes in Computer Science, 937:27–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cobbs:1995:FAM

[1604] A. L. Cobbs. Fast approximate matching using suffix trees. Lecture Notes in Computer Science, 937:41–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dancik:1995:CSS

[1605] V. Dancik. Common subsequences and supersequences and their expected length. Lecture Notes in Computer Science, 937:55–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fu:1995:PMD

[1606] J. Fu. Pattern matching in directed graphs. Lecture Notes in Computer Science, 937:64–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasieniec:1995:CSS

[1607] L. Gasieniec, W. Plandowski, and W. Rytter. Constant-space string matching with smaller number of comparisons: Sequential sampling. Lecture Notes in Computer Science, 937:78–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giancarlo:1995:MDP

[1608] R. Giancarlo and R. Grossi. Multidimensional pattern matching with dimensional wildcards. Lecture Notes in Computer Science, 937:90–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goldberg:1995:MPN

[1609] L. A. Goldberg, P. W. Goldberg, C. A.

Phillips, and E. Sweedyk. Minimizing phylogenetic number to find good evolutionary trees. Lecture Notes in Computer Science, 937:102–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:1995:MSP

[1610] S. K. Gupta, J. D. Kececioglu, and A. A. Schaeffer. Making the shortestpaths approach to sum-of-pairs multiple sequence alignment more space efficient in practice. Lecture Notes in Computer Science, 937:128–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hanks:1995:EAD

[1611] L. Hanks, R. K. Cytron, and W. Gillett. An efficient algorithm for developing topologically valid matchings. Lecture Notes in Computer Science, 937:144–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hannenhalli:1995:PTA

[1612] S. Hannenhalli. Polynomial-time algorithm for computing translocation distance between genomes. Lecture Notes in Computer Science, 937:162– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hein:1995:CCE

[1613] J. Hein, T. Jiang, L. Wang, and K. Zhang. On the complexity of comparing evolutionary trees. Lecture Notes in Computer Science, 937:177– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaerkkaeinen:1995:SCC

[1614] J. Kaerkkaeinen. Suffix cactus: A cross between suffix tree and suffix array. Lecture Notes in Computer Science, 937:191–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Karpinski:1995:PMS

[1615] M. Karpinski, W. Rytter, and A. Shinohara. Pattern-matching for strings with short descriptions. Lecture Notes in Computer Science, 937:205– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Knecht:1995:PAS

[1616] L. Knecht. Pairwise alignment with scoring on tuples. Lecture Notes in Computer Science, 937:215–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kucherov:1995:MSS

[1617] G. Kucherov and M. Rusinowitch.

Matching a set of strings with variable length don’t cares. Lecture Notes in Computer Science, 937:230–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lesk:1995:TDP

[1618] A. M. Lesk. Three-dimensional pattern matching in protein structure analysis. Lecture Notes in Computer Science, 937:248–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mewes:1995:GAP

[1619] H. W. Mewes and K. Heumann.

Genome analysis: Pattern search in biological macromolecules (invited lecture). Lecture Notes in Computer Science, 937:261–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohri:1995:MPA

[1620] M. Mohri. Matching patterns of an automaton. Lecture Notes in Computer Science, 937:286–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Muthukrishnan:1995:NRO

[1621] S. Muthukrishnan. New results and open problems related to non-standard stringology. Lecture Notes in Computer Science, 937:298–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1995:SMH

[1622] K. Park and Dong Kyue Kim. String matching in hypertext. Lecture Notes in Computer Science, 937:318–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ravi:1995:AAM

[1623] R. Ravi and J. D. Kececioglu. Approximation algorithms for multiple sequence alignment under a fixed evolutionary tree. Lecture Notes in Computer Science, 937:330–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rick:1995:NFA

[1624] C. Rick. A new flexible algorithm for the longest common subsequence problem. Lecture Notes in Computer Science, 937:340–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roche:1995:SRF

[1625] E. Roche. Smaller representations for finite-state transducers and finite-state automata. Lecture Notes in Computer Science, 937:352–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sagot:1995:MSC

[1626] M.-F. Sagot, A. Viari, and H. Soldano. Multiple sequence comparison — A peptide matching approach. Lecture Notes in Computer Science, 937: 366–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vishkin:1995:TPS

[1627] U. Vishkin. On a technique for parsing a string (invited lecture). Lecture Notes in Computer Science, 937:386– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yao:1995:DLS

[1628] A. C. Yao and F. F. Yao. Dictionary look-up with small errors. Lecture Notes in Computer Science, 937: 387–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1995:EDB

[1629] K. Zhang, J. T. L. Wang, and D. Shasha. On the editing distance between undirected acyclic graphs and related problems. Lecture Notes in Computer Science, 937:395–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Johansen:1995:AIS

[1630] D. Johansen and G. Hartvigsen. Architectural issues in the StormCast system. Lecture Notes in Computer Science, 938:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shrivastava:1995:LLB

[1631] S. K. Shrivastava. Lessons learned from building and using the Arjuna distributed programming system. Lecture Notes in Computer Science, 938:17–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whetten:1995:HPT

[1632] B. Whetten, T. Montgomery, and S. Kaplan. A high performance totally ordered multicast protocol. Lecture Notes in Computer Science, 938:33–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanRenesse:1995:NAG

[1633] R. Van Renesse, K. P. Birman, T. Von Eicken, and K. Marzullo. New applications for group computing. Lecture Notes in Computer Science, 938:58–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolland:1995:SIS

[1634] M. Kolland. Support for information sharing in CSCW based on causally and totally ordered group communication. Lecture Notes in Computer Science, 938:64–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dolev:1995:DTS

[1635] D. Dolev and D. Malki. The design of the Transis system. Lecture Notes in Computer Science, 938:83–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reiter:1995:RTB

[1636] M. K. Reiter. The rampart toolkit for building high-integrity services. Lecture Notes in Computer Science, 938:99–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alvisi:1995:DOC

[1637] L. Alvisi and K. Marzullo. Deriving optimal checkpoint protocols for distributed shared memory architectures. Lecture Notes in Computer Science, 938:111–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guerraoui:1995:TMV

[1638] R. Guerraoui and A. Schiper. Transaction model vs. virtual synchrony models: bridging the gap. Lecture Notes in Computer Science, 938:121–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Magee:1995:MDC

[1639] J. Magee, S. Eisenbach, and J. Kramer.

Modelling Darwin in the λ-calculus. Lecture Notes in Computer Science, 938:133–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geihs:1995:TOS

[1640] K. Geihs, B. Baer, and A. Puder. Towards open service environments. Lecture Notes in Computer Science, 938: 153–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Reisig:1995:CPD

[1641] W. Reisig. Correctness proofs of distributed algorithms. Lecture Notes in Computer Science, 938:164–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dawson:1995:DFI

[1642] S. Dawson and F. Jahanian. Deterministic fault injection of distributed systems. Lecture Notes in Computer Science, 938:178–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mattern:1995:NBL

[1643] F. Mattern and S. Fuenfrocken. A non-blocking lightweight implementation of causal order message delivery. Lecture Notes in Computer Science, 938:197–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Birman:1995:MPA

[1644] K. P. Birman. Merits of a probabilistic approach to properties in process group systems. Lecture Notes in Computer Science, 938:214–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mizuno:1995:SCD

[1645] M. Mizuno, M. Raynal, and J. Z. Zhou. Sequential consistency in distributed systems. Lecture Notes in Computer Science, 938:224–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwarz:1995:SLT

[1646] R. Schwarz. Speedup limits for tightlycoupled parallel computations. Lecture Notes in Computer Science, 938:242– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bryant:1995:MDI

[1647] R. E. Bryant. Multipliers and dividers: Insights on arithmetic circuit verification. Lecture Notes in Computer Science, 939:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bern:1995:GRO

[1648] J. Bern, C. Meinel, and A. Slobodova.

Global rebuilding of OBDD’s — avoiding memory requirement maxima. Lecture Notes in Computer Science, 939: 4–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dsouza:1995:GBM

[1649] A. Dsouza and B. Bloom. Generating BDD models for process algebra terms. Lecture Notes in Computer Science, 939:16–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Basin:1995:HVU

[1650] D. A. Basin and N. Klarlund. Hardware verification using monadic secondorder logic. Lecture Notes in Computer Science, 939:31–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jonsson:1995:VSP

[1651] B. Jonsson and L. Kempe. Verifying safety properties of a class of infinitestate distributed algorithms. Lecture Notes in Computer Science, 939:42–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dingel:1995:MCI

[1652] J. Dingel and T. Filkorn. Model checking for infinite state systems using data abstraction, assumption-commitment style reasoning and theorem proving. Lecture Notes in Computer Science, 939:54–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gribomont:1995:CTT

[1653] E. P. Gribomont and D. Rossetto. CAVEAT: Technique and tool for computer aided verification and transformation. Lecture Notes in Computer Science, 939:70–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rajan:1995:IMC

[1654] S. Rajan, N. Shankar, and M. K. Srivas. An integration of model checking with automated proof checking. Lecture Notes in Computer Science, 939:84–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hojati:1995:ADA

[1655] R. Hojati and R. K. Brayton. Automatic datapath abstraction in hardware systems. Lecture Notes in Computer Science, 939:98–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rauzy:1995:TCC

[1656] A. Rauzy. Toupie = μ-calculus + constraints. Lecture Notes in Computer Science, 939:114–126, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jagadeesan:1995:SPV

[1657] L. J. Jagadeesan, C. Puchol, and J. E. Von Olnhausen. Safety property verification of Esterel programs and applications to telecommunications software. Lecture Notes in Computer Science, 939:127–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emerson:1995:MCM

[1658] E. A. Emerson. Methods for mucalculus model checking. Lecture Notes in Computer Science, 939:141– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1995:ECB

[1659] H. R. Andersen and B. Vergauwen. Efficient checking of behavioural relations and modal assertions using fixed-point inversion. Lecture Notes in Computer Science, 939:142–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aziz:1995:IUW

[1660] A. Aziz, V. Singhal, F. Balarin, and R. K. Brayton. It usually works: The temporal logic of stochastic systems. Lecture Notes in Computer Science, 939:155–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alur:1995:LLC

[1661] R. Alur and T. A. Henzinger. Local liveness for compositional modeling of fair reactive systems. Lecture Notes in Computer Science, 939:166–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

McMillan:1995:TTV

[1662] K. L. McMillan. Trace theoretic verification of asynchronous circuits using unfoldings. Lecture Notes in Computer Science, 939:180–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouajjani:1995:DCL

[1663] A. Bouajjani, Y. Lakhnech, and R. Robbana. From duration calculus to linear hybrid automata. Lecture Notes in Computer Science, 939:196– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sokolsky:1995:LMC

[1664] O. V. Sokolsky and S. A. Smolka. Local model checking for real-time systems. Lecture Notes in Computer Science, 939:211–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1995:AAN

[1665] T. A. Henzinger and P.-H. Ho. Algorithmic analysis of nonlinear hybrid systems. Lecture Notes in Computer Science, 939:225–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lescow:1995:PSP

[1666] H. Lescow. On polynomial-size programs winning finite-state games. Lecture Notes in Computer Science, 939: 239–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Krishnan:1995:RIC

[1667] S. C. Krishnan, A. Puri, R. K. Brayton, and P. P. Varaiya. The Rabin index and chain automata, with applications to automata and games. Lecture Notes in Computer Science, 939:253–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vardi:1995:ATA

[1668] M. Y. Vardi. An automata-theoretic approach to fair realizability and synthesis. Lecture Notes in Computer Science, 939:267–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Aziz:1995:SCF

[1669] A. Aziz, F. Balarin, R. K. Brayton, and M. D. DiBenedetto. Supervisory control of finite state machines. Lecture Notes in Computer Science, 939:279–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cousot:1995:CIS

[1670] P. Cousot and R. Cousot. Compositional and inductive semantic definitions in fixpoint, equational, constraint, closure-condition, rule-based and game-theoretic form. Lecture Notes in Computer Science, 939:293– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Emerson:1995:USW

[1671] E. A. Emerson and A. P. Sistla. Utilizing symmetry when model checking under fairness assumptions: an automatatheoretic approach. Lecture Notes in Computer Science, 939:309–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kupferman:1995:ABT

[1672] O. Kupferman. Augmenting branching temporal logics with existential quantification over atomic propositions. Lecture Notes in Computer Science, 939:325–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurshan:1995:MAS

[1673] R. P. Kurshan, M. Merritt, A. Orda, and S. R. Sachs. Modelling asynchrony with a synchronous model. Lecture Notes in Computer Science, 939:339– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esparza:1995:MCP

[1674] J. Esparza and A. Kiehn. On the model checking problem for branching time logics and basic parallel processes. Lecture Notes in Computer Science, 939:353–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eiriksson:1995:UFV

[1675] A. T. Eiriksson and K. L. McMillan. Using formal verification/analysis methods on the critical path in system design. Lecture Notes in Computer Science, 939:367–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ho:1995:AAA

[1676] P.-H. Ho and H. Wong-Toi. Automated analysis of an audio control protocol. Lecture Notes in Computer Science, 939:381–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fidge:1995:IVS

[1677] C. Fidge, P. Kearney, and M. Utting. Interactively verifying a simple real-time scheduler. Lecture Notes in Computer Science, 939:395–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dill:1995:VRT

[1678] D. Dill and H. Wong-Toi. Verification of real-time systems by successive over and under approximation. Lecture Notes in Computer Science, 939: 409–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hulgaard:1995:ETA

[1679] H. Hulgaard and S. M. Burns. Efficient timing analysis of a class of Petri nets. Lecture Notes in Computer Science, 939:423–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouajjani:1995:VOR

[1680] A. Bouajjani and R. Robbana. Verifying omega-regular properties for a subclass of linear hybrid systems. Lecture Notes in Computer Science, 939:437– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cattell:1995:ODS

[1681] R. G. G. Cattell. Object databases and standards. Lecture Notes in Computer Science, 940:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ghafoor:1995:MDM

[1682] A. Ghafoor. Multimedia database management: Perspectives and challenges. Lecture Notes in Computer Science, 940:12–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courtenage:1995:CIP

[1683] S. Courtenage and A. Poulovassilis. Combining inheritance and parametric polymorphism in a functional database language. Lecture Notes in Computer Science, 940:24–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sutton:1995:EFD

[1684] D. Sutton and C. Small. Extending functional database languages to update completeness. Lecture Notes in Computer Science, 940:47–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nordbotten:1995:RGD

[1685] J. C. Nordbotten and M. E. Crosby. Recognising graphic detail — an experiment in user interpretation of data models. Lecture Notes in Computer Science, 940:64–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Haw:1995:PND

[1686] D. Haw, C. A. Goble, and A. L. Rector. The pragmatics of naive database enquiry. Lecture Notes in Computer Science, 940:79–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mitchell:1995:UCD

[1687] K. J. Mitchell, J. B. Kennedy, and P. J. Barclay. Using a conceptual data language to describe a database and its interface. Lecture Notes in Computer Science, 940:101–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sieg:1995:SQP

[1688] J. C. Sieg, D. Pinkney, and J. Lamoureaux. Scheduling query plans with buffer-requirement estimates. Lecture Notes in Computer Science, 940:120– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Veenhof:1995:OSJ

[1689] H. M. Veenhof, P. M. G. Apers, and A. W. Houtsmathe. Optimisation of spatial joins using filters. Lecture Notes in Computer Science, 940:136– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gukal:1995:ETV

[1690] S. Gukal, E. Omiecinski, and U. Ramachandran. An efficient transient versioning method. Lecture Notes in Computer Science, 940:155–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hussak:1995:CCT

[1691] W. Hussak and J. A. Keane. Concurrency control of tiered flat transactions. Lecture Notes in Computer Science, 940:172–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kim:1995:TSP

[1692] H. Kim, S. Lee, and H.-J. Kim. Two step pruning: A distributed query optimisation algorithm. Lecture Notes in Computer Science, 940:183–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kaufmann:1995:TSU

[1693] H. Kaufmann and H.-J. Schek. Text search using database systems revisited — some experiments. Lecture Notes in Computer Science, 940:204–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alzahrani:1995:IMO

[1694] R. M. Alzahrani, M. A. Qutaisha, N. J. Fiddian, and W. A. Gray. Integrity management in an object-oriented federated database environment. Lecture Notes in Computer Science, 940:226– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Watson:1995:AIG

[1695] P. Watson and G. Catlow. The architecture of the ICL GOLDRUSH MegaSERVER. Lecture Notes in Computer Science, 940:249–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kerridge:1995:WSI

[1696] J. Kerridge, D. Walter, and R. Guiton. W-SQL: An interface for scalable, highly parallel database machines. Lecture Notes in Computer Science, 940:263–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:LRC

[1697] Anonymous. Language restriction: Complexity of minimal reasoning. Lecture Notes in Computer Science, 941: 27–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:ALT

[1698] Anonymous. Approximation of a logical theory. Lecture Notes in Computer Science, 941:67–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:UCR

[1699] Anonymous. Using complexity results for evaluating approximation techniques. Lecture Notes in Computer Science, 941:123–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bockle:1995:EFG

[1700] Günter Böckle. Exploitation of finegrain parallelism. Lecture Notes in Computer Science, 942:ix + 188, 1995. CODEN LNCSD9. ISBN 3-540-60054-X, 0-387-60054-X. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:FPO

[1701] Anonymous. Fundamental principles of object-oriented systems. Lecture Notes in Computer Science, 943:8–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:BCM

[1702] Anonymous. Basic concepts for metaclasses in the object-oriented data model. Lecture Notes in Computer Science, 943:14–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:SDM

[1703] Anonymous. Semantic data modelling. Lecture Notes in Computer Science, 943:71–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:MSD

[1704] Anonymous. Metaclasses for semantic data modelling. Lecture Notes in Computer Science, 943:82–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:OCD

[1705] Anonymous. Object class definition by generalization. Lecture Notes in Computer Science, 943:123–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:MOC

[1706] Anonymous. Metaclasses for object class definition by generalization. Lecture Notes in Computer Science, 943: 150–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:1995:MOO

[1707] Anonymous. Metaclasses in other object-oriented systems. Lecture Notes in Computer Science, 943:172–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ehrenfeucht:1995:TS

[1708] A. Ehrenfeucht, T. Harju, and G. Rozenberg. Theory of 2-Structures. Lecture Notes in Computer Science, 944:1–??, 1995. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic).

Diekert:1995:DCT

[1709] V. Diekert and P. Gastin. A domain for concurrent termination: A generalization of Mazurkiewicz traces. Lecture Notes in Computer Science, 944:15–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Esik:1995:NAE

[1710] Z. Esik and M. Bertol. Nonfinite axiomatizability of the equational theory of shuffle. Lecture Notes in Computer Science, 944:27–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kuich:1995:AEA

[1711] W. Kuich. The algebraic equivalent of AFL theory. Lecture Notes in Computer Science, 944:39–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Culik:1995:FST

[1712] K. Culik and J. Kari. Finite state transformations of images. Lecture Notes in Computer Science, 944:51–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lipponen:1995:PCP

[1713] M. Lipponen. Post correspondence problem: Words possible as primitive solutions. Lecture Notes in Computer Science, 944:63–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Metivier:1995:CCS

[1714] Y. Metivier, G. Richomme, and P.- A. Wacrenier. Computing the closure of sets of words under partial commutations. Lecture Notes in Computer Science, 944:75–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1995:ICG

[1715] H. L. Bodlaender and B. De Fluiter. Intervalizing k-colored graphs. Lecture Notes in Computer Science, 944:87–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chen:1995:NAF

[1716] Z.-Z. Chen. NC algorithms for finding a maximal set of paths with application to compressing strings. Lecture Notes in Computer Science, 944:99–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giancarlo:1995:CCS

[1717] R. Giancarlo and R. Grossi. On the construction of classes of suffix trees for square matrices: Algorithms and applications. Lecture Notes in Computer Science, 944:111–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parra:1995:HUM

[1718] A. Parra and P. Scheffler. How to use the minimal separators of a graph for its chordal triangulation. Lecture Notes in Computer Science, 944:123– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bermond:1995:FGS

[1719] J.-C. Bermond, L. Gargano, A. A.

Rescigno, and U. Vaccaro. Fast gossiping by short messages. Lecture Notes in Computer Science, 944:135–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bao:1995:BFA

[1720] F. Bao and Y. Igarashi. Break finite automata public key cryptosystem. Lecture Notes in Computer Science, 944:147–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nikoletseas:1995:SMS

[1721] S. Nikoletseas, J. Reif, P. Spirakis, and M. Yung. Short memory in stochastic graphs: Fully dynamic connectivity in poly-log expected time. Lecture Notes in Computer Science, 944:159– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blundo:1995:NRB

[1722] C. Blundo, A. De Santis, G. Persiano, and U. Vaccaro. On the number of random bits in totally private computation. Lecture Notes in Computer Science, 944:171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Freivalds:1995:LTB

[1723] R. Freivalds and M. Karpinski. Lower time bounds for randomized computation. Lecture Notes in Computer Science, 944:183–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koebler:1995:NCC

[1724] J. Koebler and O. Watanabe. New collapse consequences of NP having small circuits. Lecture Notes in Computer Science, 944:196–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balcazar:1995:CSS

[1725] J. L. Balcazar. The complexity of searching succinctly represented graphs. Lecture Notes in Computer Science, 944:208–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bauernoeppel:1995:OSC

[1726] F. Bauernoeppel, E. Kranakis, D. Krizanc, and A. Maheshwari. Optimal shooting: Characterizations and applications. Lecture Notes in Computer Science, 944:220–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Auletta:1995:PRT

[1727] V. Auletta, D. Parente, and G. Persiano. Placing resources in a tree: Dynamic and static algorithms. Lecture Notes in Computer Science, 944:232– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chaudhuri:1995:SPQ

[1728] S. Chaudhuri and C. D. Zaroliagis. Shortest path queries in digraphs of small treewidth. Lecture Notes in Computer Science, 944:244–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Golin:1995:DPA

[1729] M. J. Golin and G. Rote. A dynamic programming algorithm for constructing optimal prefix-free codes for unequal letter costs. Lecture Notes in Computer Science, 944:256–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bodlaender:1995:PAO

[1730] H. L. Bodlaender and T. Hagerup. Parallel algorithms with optimal speedup for bounded treewidth. Lecture Notes in Computer Science, 944:268–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

RauchHenzinger:1995:AMC

[1731] M. Rauch Henzinger. Approximating minimum cuts under insertions. Lecture Notes in Computer Science, 944: 280–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Corneil:1995:LTA

[1732] D. G. Corneil, S. Olariu, and L. Stewart. Linear time algorithms for dominating pairs in asteroidal triple-free graphs. Lecture Notes in Computer Science, 944:292–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leonardi:1995:LRM

[1733] S. Leonardi and A. Marchetti-Spaccamela. On-line resource management with applications to routing and scheduling. Lecture Notes in Computer Science, 944:303–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Petersen:1995:ASD

[1734] H. Petersen. Alternation in simple devices. Lecture Notes in Computer Science, 944:315–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1995:HAF

[1735] T. A. Henzinger. Hybrid automata with finite bisimulations. Lecture Notes in Computer Science, 944:324– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez:1995:GSL

[1736] L.-M. Lopez and P. Narbel. Generalized Sturmian languages. Lecture Notes in Computer Science, 944:336– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pin:1995:PCU

[1737] J.-E. Pin and P. Weil. Polynomial closure and unambiguous product. Lecture Notes in Computer Science, 944: 348–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ben-Amram:1995:LBA

[1738] A. M. Ben-Amram and Z. Galil. Lower bounds on algebraic random access machines. Lecture Notes in Computer Science, 944:360–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pietracaprina:1995:IDP

[1739] A. Pietracaprina and G. Pucci. Improved deterministic PRAM simulation on the mesh. Lecture Notes in Computer Science, 944:372–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ambos-Spies:1995:OPT

[1740] K. Ambos-Spies. On optimal polynomial time approximations: Plevelability vs. cap delta-levelability. Lecture Notes in Computer Science, 944:384–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fenner:1995:WUS

[1741] S. A. Fenner, J. H. Lutz, and E. Mayordomo. Weakly useful sequences. Lecture Notes in Computer Science, 944: 393–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schwentick:1995:GCM

[1742] T. Schwentick. Graph connectivity, monadic NP and built-in relations of moderate degree. Lecture Notes in Computer Science, 944:405–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henzinger:1995:EPC

[1743] T. A. Henzinger, P. W. Kopke, and H. Wong-Toi. The expressive power of clocks. Lecture Notes in Computer Science, 944:417–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Paun:1995:GSG

[1744] G. Paun. Grammar systems: A grammatical approach to distribution and cooperation. Lecture Notes in Computer Science, 944:429–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harju:1995:CSE

[1745] T. Harju, J. Karhumaeki, and W. Plandowski. Compactness of systems of equations in semigroups. Lecture Notes in Computer Science, 944:444– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Duris:1995:SVN

[1746] P. Duris and Z. Galil. Sensing versus nonsensing automata. Lecture Notes in Computer Science, 944:455–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bozanis:1995:NUB

[1747] P. Bozanis, N. Kitsios, C. Makris, and A. Tsakalidis. New upper bounds for generalized intersection searching problems. Lecture Notes in Computer Science, 944:464–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Becker:1995:OVO

[1748] B. Becker, R. Drechsler, and M. Theobald. OKFDDs versus OBDDs and OFDDs. Lecture Notes in Computer Science, 944:475–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Marathe:1995:BND

[1749] M. V. Marathe, R. Ravi, R. Sundaram, and S. S. Ravi. Bicriteria network design problems. Lecture Notes in Computer Science, 944:487–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Le:1995:DOS

[1750] N.-M. Le. On determining optimal strategies in pursuit games in the plane. Lecture Notes in Computer Science, 944:499–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rubio:1995:EO

[1751] A. Rubio. Extension orderings. Lecture Notes in Computer Science, 944: 511–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Greco:1995:PMO

[1752] S. Greco, D. Sacca, and C. Zaniolo.

The PushDown method to optimize chain logic programs. Lecture Notes in Computer Science, 944:523–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Andersen:1995:ASR

[1753] J. H. Andersen, K. J. Kristoffersen, K. G. Larsen, and J. Niedermann. Automatic synthesis of real time systems. Lecture Notes in Computer Science, 944:535–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blum:1995:SCF

[1754] M. Blum, B. Codenotti, P. Gemmell, and T. Shahoumian. Self-correcting for function fields of finite transcendental degree. Lecture Notes in Computer Science, 944:547–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fortnow:1995:MCL

[1755] L. Fortnow, R. Freivalds, W. I. Gasarch, and M. Kummer. Measure, category and learning theory. Lecture Notes in Computer Science, 944:558– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cosnard:1995:CEE

[1756] M. Cosnard and E. Goles. A characterization of the existence of energies for neural networks. Lecture Notes in Computer Science, 944:570–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Juergensen:1995:VLC

[1757] H. Juergensen and S. Konstantinidis. Variable-length codes for error correction. Lecture Notes in Computer Science, 944:581–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Khuller:1995:GMD

[1758] S. Khuller, E. Rivlin, and A. Rosenfeld. Graphbots: Mobility in discrete spaces. Lecture Notes in Computer Science, 944:593–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Best:1995:SRN

[1759] E. Best and M. Koutny. Solving recursive net equations. Lecture Notes in Computer Science, 944:605–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dawar:1995:IDI

[1760] A. Dawar, L. Hella, and P. G. Kolaitis.

Implicit definability and infinitary logic in finite model theory. Lecture Notes in Computer Science, 944:624–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vogler:1995:LSL

[1761] W. Vogler. The limit of splitnlanguage equivalence. Lecture Notes in Computer Science, 944:636–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Natarajan:1995:DFT

[1762] V. Natarajan and R. Cleaveland. Divergence and fair testing. Lecture Notes in Computer Science, 944:648– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Degano:1995:CMP

[1763] P. Degano and C. Priami. Causality for mobile processes. Lecture Notes in Computer Science, 944:660–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sangiorgi:1995:IMA

[1764] D. Sangiorgi. Internal mobility and agent-passing calculi. Lecture Notes in Computer Science, 944:672–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shafer:1995:PFC

[1765] G. Shafer. Philosophical foundations for causal networks. Lecture Notes in Computer Science, 945:3–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Smets:1995:NSP

[1766] P. Smets. Non-standard probabilistic and non probabilistic representations of uncertainty. Lecture Notes in Computer Science, 945:13–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohlas:1995:ASE

[1767] J. Kohlas and H. W. Brachinger. Argumentation systems and evidence theory. Lecture Notes in Computer Science, 945:41–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Daniel:1995:ASR

[1768] M. Daniel. Algebraic structures related to Dempster-Shafer theory. Lecture Notes in Computer Science, 945:51–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:1995:ELP

[1769] Z. Wang and W. Wang. Extension of lower probabilities and coherence of belief measures. Lecture Notes in Computer Science, 945:62–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padet:1995:NMO

[1770] C. Padet, A. Ramer, and R. Yager. A note on multi-objective information measures. Lecture Notes in Computer Science, 945:70–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mellouli:1995:DMU

[1771] K. Mellouli. Decision making using belief functions: Evaluation of information. Lecture Notes in Computer Science, 945:77–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

David:1995:HPJ

[1772] A. P. David, U. Kjoerulff, and S. L.

Lauritzen. Hybrid propagation in junction trees. Lecture Notes in Computer Science, 945:87–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cano:1995:HAT

[1773] A. Cano and S. Moral. Heuristic algorithms for the triangulation of graphs. Lecture Notes in Computer Science, 945:98–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Xu:1995:CMM

[1774] H. Xu. Computing marginals from the marginal representation in Markov trees. Lecture Notes in Computer Science, 945:108–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wong:1995:RBN

[1775] S. K. M. Wong, Y. Xiang, and X. Nie.

Representation of Bayesian networks as relational databases. Lecture Notes in Computer Science, 945:117–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lopez:1995:DID

[1776] M. Lopez. Decision influence diagrams with fuzzy utilities. Lecture Notes in Computer Science, 945:131–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Liang:1995:CNT

[1777] J. Liang, Q. Ren, Z. Xu, and J. Fang. Causal networks and their toolkit in KSE. Lecture Notes in Computer Science, 945:139–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Acid:1995:ACN

[1778] S. Acid and L. M. De Campos. Approximations of causal networks by polytrees: an empirical study. Lecture Notes in Computer Science, 945:149– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ezawa:1995:EPI

[1779] K. J. Ezawa. Evidence propagation on influence diagrams and value of evidence. Lecture Notes in Computer Science, 945:159–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wellman:1995:SVQ

[1780] M. P. Wellman. Some varieties of qualitative probability. Lecture Notes in Computer Science, 945:171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parsons:1995:QVQ

[1781] S. Parsons and A. Saffiotti. The qualitative verification of quantitative uncertainty. Lecture Notes in Computer Science, 945:180–189, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeCampos:1995:UMU

[1782] L. M. De Campos, J. F. Huete, and S. Moral. Uncertainty management using probability intervals. Lecture Notes in Computer Science, 945:190– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gilio:1995:PCC

[1783] A. Gilio. Probabilistic consistency of conditional probability bounds. Lecture Notes in Computer Science, 945: 200–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alonso:1995:TCV

[1784] M. C. Alonso and M. A. Gil. Testing the convenience of a variate for stratification in estimating the Gini-Simpson diversity. Lecture Notes in Computer Science, 945:210–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ralescu:1995:FPT

[1785] D. Ralescu. Fuzzy probabilities and their applications to statistical inference. Lecture Notes in Computer Science, 945:217–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Georgescu:1995:BFA

[1786] C. Georgescu, A. Afshari, and G. Bornard. A Bayesian functional approach to fuzzy system representation. Lecture Notes in Computer Science, 945:223–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Morales:1995:IMA

[1787] D. Morales, L. Pardo, M. Salicru, and M. L. Menendez. Information matrices associated to (h,0)-divergence measures: Applications to testing hypotheses. Lecture Notes in Computer Science, 945:233–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pardo:1995:IET

[1788] J. A. Pardo, M. C. Pardo, M. L. Vicente, and M. D. Esteban. Informational energy test for equality of variances. Lecture Notes in Computer Science, 945:241–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bertoluzza:1995:CIM

[1789] C. Bertoluzza, T. Brezmes, and G. Naval. Compositive information measure of a fuzzy set. Lecture Notes in Computer Science, 945:248–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vallee:1995:IT

[1790] R. Vallee. Informational time. Lecture Notes in Computer Science, 945: 255–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1995:UTC

[1791] D. Dubois, F. Dupin de Saint Cyr, and H. Prade. Updating, transition constraints and possibilistic Markov chains. Lecture Notes in Computer Science, 945:263–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gebhardt:1995:NFP

[1792] J. Gebhardt and R. Kruse. A numerical framework for possibilistic abduction. Lecture Notes in Computer Science, 945:281–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

FarithasdelCerro:1995:PTI

[1793] L. Farithas del Cerro and A. Herzig.

Possibility theory and independence. Lecture Notes in Computer Science, 945:292–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cappelle:1995:GPF

[1794] B. Cappelle and E. E. Kerre. A general possibilistic framework for reliability theory. Lecture Notes in Computer Science, 945:311–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sandri:1995:PSN

[1795] S. Sandri and G. Bittencourt. Possibilistic semantic nets. Lecture Notes in Computer Science, 945:318–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Trillas:1995:ENE

[1796] E. Trillas, S. Cubillo, and A. R. De Soto. An essay on name and extension of rule-given properties. Lecture Notes in Computer Science, 945:329– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abar:1995:NAL

[1797] C. A. A. P. Abar and M. Yamashita. On non-alethic logic. Lecture Notes in Computer Science, 945:339–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Studeny:1995:MPD

[1798] M. Studeny. Marginal problem in different calculi of AI. Lecture Notes in Computer Science, 945:348–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Besnard:1995:ICL

[1799] P. Besnard and Y. Moinard. Implication connectives for logics with right weakening. Lecture Notes in Computer Science, 945:360–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guiasu:1995:SL

[1800] S. Guiasu. Stochastic logic. Lecture Notes in Computer Science, 945:370– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gottwald:1995:AHP

[1801] S. Gottwald. An approach to handle partially sound rules of inference. Lecture Notes in Computer Science, 945: 380–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bruening:1995:UCT

[1802] S. Bruening and T. Schaub. Using classical theorem-proving techniques for approximate reasoning: Revised report. Lecture Notes in Computer Science, 945:389–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gasquet:1995:TBD

[1803] O. Gasquet and A. Herzig. Translationbased deduction methods for modal logics. Lecture Notes in Computer Science, 945:399–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DaSilva:1995:SKR

[1804] R. M. Da Silva, A. E. C. Pereira, and M. A. Netto. A system of knowledge representation based in formulae of predicate calculus whose variables are annotated by expressions of A “fuzzy” terminological logic. Lecture Notes in Computer Science, 945:409–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bigham:1995:UPB

[1805] J. Bigham. Using preference based heuristics to control abductive reasoning. Lecture Notes in Computer Science, 945:418–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Escalada-Imaz:1995:EIP

[1806] G. Escalada-Imaz and F. Manya. Efficient interpretation of propositional multiple-valued logic programs. Lecture Notes in Computer Science, 945: 428–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Godo:1995:MVE

[1807] L. Godo, W. Van der Hoek, J.-J. C. Meyer, and C. Sierra. Many-valued epistemic states: An application to a reflexive architecture: Milord-II. Lecture Notes in Computer Science, 945: 440–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gan:1995:ACF

[1808] W. S. Gan. Acoustical chaotic fractal images for medical imaging. Lecture Notes in Computer Science, 945: 455–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakai:1995:CCP

[1809] K. Sakai, T. Katayama, S. Wada, and H. Oiwa. Chaos causes perspective reversals for ambiguous patterns. Lecture Notes in Computer Science, 945:463–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cohen:1995:ICA

[1810] M. E. Cohen, D. L. Hudson, M. F. Anderson, and P. C. Deedwania. Implications of a continuous approach to chaotic modeling. Lecture Notes in Computer Science, 945:473–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bernard-Weil:1995:HCI

[1811] E. Bernard-Weil. “homeostatic” control of imbalanced strange attractors with or without asking to a change in chaotic dynamics. Lecture Notes in Computer Science, 945:483–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Park:1995:FBR

[1812] S. Park and J. D. Palmer. A featurebased reuse library. Lecture Notes in Computer Science, 945:495–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gall:1995:MUO

[1813] H. Gall and R. Kloesch. Managing uncertainty in an object recovery process. Lecture Notes in Computer Science, 945:501–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simos:1995:JFF

[1814] M. A. Simos. Juggling in free fall:

Uncertainty management aspects of domain analysis methods. Lecture Notes in Computer Science, 945:512– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Balentine:1995:HFM

[1815] B. J. Balentine, M. K. Zand, and M. H.

Samadzadeh. Hybrid fuzzy metrics for software reusability. Lecture Notes in Computer Science, 945:522–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bakhouch:1995:PTS

[1816] M. Bakhouch. A program transformation system for software reuse. Lecture Notes in Computer Science, 945: 534–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mittermeir:1995:AVS

[1817] R. T. Mittermeir and L. Wuerfl. Abstract visualization of software. A basis for a complex hash-key? Lecture Notes in Computer Science, 945:545– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mukunoki:1995:RIU

[1818] M. Mukunoki, M. Minoh, and K. Ikeda. Retrieval of images using pixel-based objects models. Lecture Notes in Computer Science, 945:557–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhang:1995:ORB

[1819] W. Zhang and A. Ralescu. Object recognition based on pattern features. Lecture Notes in Computer Science, 945:567–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bothe:1995:KPS

[1820] H. H. Bothe and N. Von Boetticher. Key-picture selection for the analysis of visual speech with fuzzy methods. Lecture Notes in Computer Science, 945:577–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anger:1995:EST

[1821] F. D. Anger and R. V. Rodriguez. Effective scheduling of tasks under weak temporal interval constraints. Lecture Notes in Computer Science, 945:584– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nimier:1995:ICI

[1822] V. Nimier. Introducing contextual information in multisensor tracking algorithms. Lecture Notes in Computer Science, 945:595–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

FerreiroGarcia:1995:FGS

[1823] R. Ferreiro Garcia. FAM on gain scheduling control of highly nonlinear and disturbed processes. Lecture Notes in Computer Science, 945:605– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guenoche:1995:HCA

[1824] A. Guenoche. How to choose according to partial evaluations? Lecture Notes in Computer Science, 945:611– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moulet:1995:RMU

[1825] M. Moulet. The role of measurement uncertainty in numeric law induction. Lecture Notes in Computer Science, 945:619–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alechina:1995:AT

[1826] N. Alechina. For all typical. Lecture Notes in Computer Science, 946:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alvarado:1995:PKA

[1827] M. Alvarado. A posteriori knowledge:

From ambiguous knowledge and undefined information to knowledge. Lecture Notes in Computer Science, 946:9–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Baroni:1995:MUR

[1828] P. Baroni, G. Guida, and S. Mussi. Modeling uncertain relational knowledge: The AV-Qualified production rules approach. Lecture Notes in Computer Science, 946:18–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benevides:1995:MDL

[1829] M. R. F. Benevides. Multiple database logic. Lecture Notes in Computer Science, 946:28–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benferhat:1995:LAR

[1830] S. Benferhat, D. Dubois, and H. Prade. A local approach to reasoning under inconsistency in stratified knowledge bases. Lecture Notes in Computer Science, 946:36–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Besnard:1995:QCL

[1831] P. Besnard and A. Hunter. Quasiclassical logic: Non-trivializable classical reasoning from inconsistent information. Lecture Notes in Computer Science, 946:44–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bigham:1995:CBP

[1832] J. Bigham, Z. Luo, and D. Banerjee. A cost bounded possibilities ATMS. Lecture Notes in Computer Science, 946:52–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boldrin:1995:SCP

[1833] L. Boldrin. A substructural connective for possibilistic logic. Lecture Notes in Computer Science, 946:60–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bouckaert:1995:CGS

[1834] R. R. Bouckaert and M. Studeny.

Chain graphs: Semantics and expressiveness. Lecture Notes in Computer Science, 946:69–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeCampos:1995:ATP

[1835] L. M. De Campos, J. Gebhardt, and R. Kruse. Axiomatic treatment of possibilistic independence. Lecture Notes in Computer Science, 946:77–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Castillo:1995:PSP

[1836] E. Castillo, J. M. Gutierrez, and A. S.

Hadi. Parametric structure of probabilities in Bayesian networks. Lecture Notes in Computer Science, 946:89–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cayrol:1995:NMS

[1837] C. Cayrol and M.-C. Lagasquie-Schiex.

Non-monotonic syntax-based entailment: A classification of consequence relations. Lecture Notes in Computer Science, 946:107–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Chateauneuf:1995:LMT

[1838] A. Chateauneuf and J.-Y. Jaffray. Local Möbius transforms of monotone capacities. Lecture Notes in Computer Science, 946:115–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cholvy:1995:ARM

[1839] L. Cholvy. Automated reasoning with merged contradictory information whose reliability depends on topics. Lecture Notes in Computer Science, 946:125–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cordier:1995:LTB

[1840] M.-O. Cordier and J. Lang. Linking transition-based update and base revision. Lecture Notes in Computer Science, 946:133–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cordier:1995:PTU

[1841] M.-O. Cordier and P. Siegel. Prioritized transitions for updates. Lecture Notes in Computer Science, 946:142– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DouanyaNguetse:1995:PSD

[1842] G.-B. Douanya Nguetse, P. Hansen, and B. Jaumard. Probabilistic satisfiability and decomposition. Lecture Notes in Computer Science, 946:151– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1995:UPI

[1843] D. Dubois, F. Dupin de St.-Cyr, and H. Prade. Update postulates without inertia. Lecture Notes in Computer Science, 946:162–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dubois:1995:SBC

[1844] D. Dubois, F. Esteva, P. Garcia, and L. Godo. Similarity-based consequence relations. Lecture Notes in Computer Science, 946:171–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dunin-Keplicz:1995:EAA

[1845] B. Dunin-Keplicz and A. Radzikowska.

Epistemic approach to actions with typical effects. Lecture Notes in Computer Science, 946:180–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Engelfriet:1995:NBS

[1846] J. Engelfriet, H. Herre, and J. Treur.

Nonmonotonic belief state frames and reasoning frames. Lecture Notes in Computer Science, 946:189–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ezawa:1995:BNB

[1847] K. J. Ezawa and T. Schuermann. A Bayesian network based learning system. Lecture Notes in Computer Science, 946:197–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Geerts:1995:SD

[1848] P. Geerts and D. Vermeir. Specificity by default. Lecture Notes in Computer Science, 946:207–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Glesner:1995:CFD

[1849] S. Glesner and D. Koller. Constructing flexible dynamic belief networks from first-order probabilistic knowledge bases. Lecture Notes in Computer Science, 946:217–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Guiasu:1995:FSL

[1850] S. Guiasu. On the formalism of stochastic logic. Lecture Notes in Computer Science, 946:227–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hunter:1995:UDL

[1851] A. Hunter. Using default logic in information retrieval. Lecture Notes in Computer Science, 946:235–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jensen:1995:SAB

[1852] F. V. Jensen, S. H. Aldenryd, and K. B. Jensen. Sensitivity analysis in Bayesian networks. Lecture Notes in Computer Science, 946:243–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joshi:1995:BAI

[1853] A. V. Joshi, S. C. Sahasrabudhe, and K. Shankar. Bayesian approximation and invariance of Bayesian belief functions. Lecture Notes in Computer Science, 946:251–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kohlas:1995:MBD

[1854] J. Kohlas, P. J. A. Monney, R. Haenni, and N. Lehmann. Model-based diagnostics using hints. Lecture Notes in Computer Science, 946:259–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kramosil:1995:AAE

[1855] I. Kramosil. An axiomatic approach to extensional probability measures. Lecture Notes in Computer Science, 946: 267–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

VanLinder:1995:DDR

[1856] B. Van Linder, W. Van der Hoek, and J.-J. C. Meyer. The dynamics of default reasoning. Lecture Notes in Computer Science, 946:277–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Linke:1995:LHD

[1857] T. Linke and T. Schaub. Lemma handling in default logic theorem provers. Lecture Notes in Computer Science, 946:285–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lukasiewicz:1995:URC

[1858] T. Lukasiewicz. Uncertain reasoning in concept lattices. Lecture Notes in Computer Science, 946:293–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mengin:1995:TPD

[1859] J. Mengin. A theorem prover for default logic based on prioritized conflict resolution and an extended resolution principle. Lecture Notes in Computer Science, 946:301–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nicolas:1995:TPL

[1860] P. Nicolas and B. Duval. A theorem prover for Lukaszewicz’ open default theory. Lecture Notes in Computer Science, 946:311–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oppel:1995:TDT

[1861] U. G. Oppel. Two different types of discontinuity of Bayesian learning in causal probabilistic networks. Lecture Notes in Computer Science, 946:320– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Papini:1995:REP

[1862] O. Papini and A. Rauzy. Revision in extended propositional calculus. Lecture Notes in Computer Science, 946: 328–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parsons:1995:UQU

[1863] S. Parsons. Using qualitative uncertainty in protein topology prediction. Lecture Notes in Computer Science, 946:336–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Radzikowska:1995:CFF

[1864] A. Radzikowska. Circumscribing features and fluents: Reasoning about action with default effects. Lecture Notes in Computer Science, 946:344– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rhodes:1995:UME

[1865] P. C. Rhodes and G. R. Garside.

Using maximum entropy to compute marginal probabilities in a causal binary tree need not take exponential time. Lecture Notes in Computer Science, 946:352–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schlechta:1995:TSA

[1866] K. Schlechta. A two-stage approach to first order default reasoning. Lecture Notes in Computer Science, 946: 379–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schlechta:1995:RTC

[1867] K. Schlechta. A reduction of the theory of confirmation to the notions of distance and measure. Lecture Notes in Computer Science, 946:387–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schubert:1995:CBS

[1868] J. Schubert. Cluster-based specification techniques in Dempster-Shafer theory. Lecture Notes in Computer Science, 946:395–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

State:1995:TKR

[1869] L. State and R. State. AND/OR trees for knowledge representation. Lecture Notes in Computer Science, 946:405– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tan:1995:WDD

[1870] Y.-H. Tan and L. W. N. Van der Torre. Why defeasible Deontic logic needs a multi preference semantics. Lecture Notes in Computer Science, 946:412– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weydert:1995:NDA

[1871] E. Weydert. Numeric defaults: About an expressive first-order framework for reasoning with infinitesimal probabilities. Lecture Notes in Computer Science, 946:420–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apt:1995:TAP

[1872] K. R. Apt. Towards automatic parallelization of logic programs. Lecture Notes in Computer Science, 947:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bird:1995:FAD

[1873] R. S. Bird. Functional algorithm design. Lecture Notes in Computer Science, 947:2–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Broy:1995:MSE

[1874] M. Broy. Mathematics of software engineering. Lecture Notes in Computer Science, 947:18–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Coquand:1995:PCI

[1875] T. Coquand. Program construction in intuitionistic type theory. Lecture Notes in Computer Science, 947:49–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shankar:1995:CAC

[1876] N. Shankar. Computer-aided computing. Lecture Notes in Computer Science, 947:50–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abdallah:1995:DPA

[1877] A. E. Abdallah. Derivation of parallel algorithms from functional specifications to CSP processes. Lecture Notes in Computer Science, 947:67–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Achatz:1995:AIM

[1878] K. Achatz and W. Schulte. Architecture independent massive parallelization of divide-and-conquer algorithms. Lecture Notes in Computer Science, 947:97–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Back:1995:ESP

[1879] R.-J. R. Back and M. Butler. Exploring summation and product operators in the refinement calculus. Lecture Notes in Computer Science, 947: 128–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Back:1995:ASS

[1880] R.-J. R. Back, A. J. Martin, and K. Sere. An action system specification of the Caltech asynchronous microprocessor. Lecture Notes in Computer Science, 947:159–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Berghammer:1995:FDC

[1881] R. Berghammer and B. Von Karger. Formal derivation of CSP programs from temporal specifications. Lecture Notes in Computer Science, 947:180– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

DeBoer:1995:CPS

[1882] F. S. De Boer and M. Van Hulst. A compositional proof system for asynchronously communicating processes. Lecture Notes in Computer Science, 947:197–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Curtis:1995:GC

[1883] S. Curtis and G. Lowe. A graphical calculus. Lecture Notes in Computer Science, 947:214–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dielissen:1995:SEF

[1884] V. J. Dielissen and A. Kaldewaij. A simple, efficient, and flexible implementation of flexible arrays. Lecture Notes in Computer Science, 947:232– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Doornbos:1995:IRD

[1885] H. Doornbos and R. Backhouse. Induction and recursion on datatypes. Lecture Notes in Computer Science, 947: 242–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frappier:1995:PCP

[1886] M. Frappier, A. Mili, and J. Desharnais. Program construction by parts. Lecture Notes in Computer Science, 947:257–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gibbons:1995:IAA

[1887] J. Gibbons. An initial-algebra approach to directed acyclic graphs. Lecture Notes in Computer Science, 947: 282–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Groenboom:1995:APA

[1888] R. Groenboom, C. Hendriks, I. Polak, and J. Terlouw. Algebraic proof assistants in HOL. Lecture Notes in Computer Science, 947:304–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hesselink:1995:ATD

[1889] W. H. Hesselink. Angelic termination in Dijkstra’s calculus. Lecture Notes in Computer Science, 947:322–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Monin:1995:EPE

[1890] J.-F. Monin. Extracting programs with exceptions in an impredicative type system. Lecture Notes in Computer Science, 947:335–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parent:1995:SPP

[1891] C. Parent. Synthesizing proofs from programs in the calculus of inductive constructions. Lecture Notes in Computer Science, 947:351–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Russling:1995:GSB

[1892] M. Russling. A general scheme for breadth-first graph transversal. Lecture Notes in Computer Science, 947: 380–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srinivas:1995:SFS

[1893] Y. V. Srinivas and R. Juellig. Specware: Formal support for composing software. Lecture Notes in Computer Science, 947:399–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoelen:1995:RRS

[1894] K. Stoelen. A refinement relation supporting the transition from unbounded to bounded communication buffers. Lecture Notes in Computer Science, 947:423–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Udink:1995:IUP

[1895] R. T. Udink and J. N. Kok. ImpUNITY: UNITY with procedures and local variables. Lecture Notes in Computer Science, 947:452–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bassalygo:1995:SC

[1896] L. A. Bassalygo. Supports of a code. Lecture Notes in Computer Science, 948:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Benecke:1995:CIC

[1897] C. Benecke, R. Grund, R. Hohberger, and A. Kerber. Chemical isomerism, a challenge for algebraic combinatorics and for computer science. Lecture Notes in Computer Science, 948:4–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Honkala:1995:AMC

[1898] I. Honkala, S. Litsyn, and A. Tietaevaeinen. On algebraic methods in covering radius problems. Lecture Notes in Computer Science, 948:21–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pardo:1995:HLU

[1899] L. M. Pardo. How lower and upper complexity bounds meet in elimination theory. Lecture Notes in Computer Science, 948:33–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schoenhage:1995:BPM

[1900] A. Schoenhage. Bivariate polynomial multiplication patterns. Lecture Notes in Computer Science, 948:70–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Apel:1995:DEF

[1901] J. Apel. Division of entire functions by polynomial ideals. Lecture Notes in Computer Science, 948:82–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ashikhmin:1995:VMC

[1902] A. Ashikhmin, A. Barg, G. Cohen, and L. Huguet. Variations on minimal codewords in linear codes. Lecture Notes in Computer Science, 948:96–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Armendariz:1995:CRP

[1903] I. Armendariz and P. Solerno. On the computation of the radical of polynomial complete intersection ideals. Lecture Notes in Computer Science, 948: 106–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beth:1995:WFL

[1904] T. Beth, H. Kalouti, and D. E. Lazic. Which families of long binary linear codes have a binomial weight distribution? Lecture Notes in Computer Science, 948:120–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Betten:1995:DSD

[1905] A. Betten, A. Kerber, A. Kohnert, and R. Laue. The discovery of simple 7-designs with automorphism group PL(2,32). Lecture Notes in Computer Science, 948:131–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bocharova:1995:FEC

[1906] I. E. Bocharova and B. D. Kudryashov.

Fast exponentiation in cryptography. Lecture Notes in Computer Science, 948:146–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boyvalenkov:1995:MSC

[1907] P. Boyvalenkov and I. Landgev. On maximal spherical codes I. Lecture Notes in Computer Science, 948:158– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Colin:1995:FCG

[1908] A. Colin. Formal computation of Galois groups with relative resolvents. Lecture Notes in Computer Science, 948: 169–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Conti:1995:CAT

[1909] P. Conti and C. Traverso. A case of automatic theorem proving in Euclidean geometry: the maclane 83 theorem. Lecture Notes in Computer Science, 948:183–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fripertinger:1995:ICI

[1910] H. Fripertinger and A. Kerber. Isometry classes of indecomposable linear codes. Lecture Notes in Computer Science, 948:194–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Giusti:1995:WPE

[1911] M. Giusti, J. Heintz, J. E. Morais, and L. M. Pardo. When polynomial equation systems can be “solved” fast? Lecture Notes in Computer Science, 948:205–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gonzalez-Vega:1995:USF

[1912] L. Gonzalez-Vega and G. Trujillo. Using symmetric functions to describe the solution set of a zero dimensional ideal. Lecture Notes in Computer Science, 948:232–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Graebe:1995:TSF

[1913] H.-G. Graebe. Triangular systems and factorized Gröbner bases. Lecture Notes in Computer Science, 948:248– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hache:1995:CAF

[1914] G. Hache. Computation in algebraic function fields for effective construction of algebraic-geometric codes. Lecture Notes in Computer Science, 948:262– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Harada:1995:SES

[1915] M. Harada and V. D. Tonchev. Singlyeven self-dual codes and Hadamard matrices. Lecture Notes in Computer Science, 948:279–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hong:1995:ICP

[1916] H. Hong. Implicitization of curves parameterized by generalized trigonometric polynomials. Lecture Notes in Computer Science, 948:285–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kamiya:1995:RPP

[1917] Y. Kamiya, T. Miki, and Y. Iwadare.

Randomness properties of partial gamma-planes as LSI test inputs and their implementations. Lecture Notes in Computer Science, 948:297–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kolev:1995:MCC

[1918] E. Kolev. Mixed covering codes with two binary and four ternary coordinates. Lecture Notes in Computer Science, 948:312–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koseleff:1995:AAE

[1919] P.-V. Koseleff. About approximations of exponentials. Lecture Notes in Computer Science, 948:323–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lacan:1995:NNB

[1920] J. Lacan and E. Delpeyroux. A note on normal bases. Lecture Notes in Computer Science, 948:334–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Levy-dit-Vehel:1995:CRL

[1921] F. Levy dit Vehel and S. Litsyn. On the covering radius of long Goppa codes. Lecture Notes in Computer Science, 948:341–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Matera:1995:IMR

[1922] G. Matera. Integration of multivariate rational functions given by straight-line programs. Lecture Notes in Computer Science, 948:347–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

MorenoMaza:1995:PGC

[1923] M. Moreno Maza and R. Rioboo. Polynomial GCD computations over towers of algebraic extensions. Lecture Notes in Computer Science, 948:365– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nguyen:1995:DMP

[1924] Q. L. Nguyen and D. H. Wood. Displacements of matrix products. Lecture Notes in Computer Science, 948: 383–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Peladan-Germa:1995:TIS

[1925] A. Peladan-Germa. Testing identities of series defined by algebraic partial differential equations. Lecture Notes in Computer Science, 948:393–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sadik:1995:CFR

[1926] B. Sadik. The complexity of formal resolution of linear partial differential equations. Lecture Notes in Computer Science, 948:408–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sakata:1995:FPI

[1927] S. Sakata and M. Kurihara. A fast parallel implementation of the Berlekamp– Massey algorithm with a one-D systolic array architecture. Lecture Notes in Computer Science, 948:415–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schapira:1995:TCF

[1928] P. Schapira. Tomography of constructible functions. Lecture Notes in Computer Science, 948:427–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shih:1995:DFS

[1929] J. A. Shih. On the determination of formal solutions of a system of partial differential equations. Lecture Notes in Computer Science, 948:436–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sole:1995:DEE

[1930] P. Sole. D4, E6, E8 and the AGM. Lecture Notes in Computer Science, 948: 448–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Valibouze:1995:CGG

[1931] A. Valibouze. Computation of the Galois groups of the resolvent factors for the direct and inverse Galois problems. Lecture Notes in Computer Science, 948:456–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Weil:1995:FID

[1932] J.-A. Weil. First integrals and Darboux polynomials of homogeneous linear differential systems. Lecture Notes in Computer Science, 948:469–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feitelson:1995:PJS

[1933] D. G. Feitelson and L. Rudolph. Parallel job scheduling: Issues and approaches. Lecture Notes in Computer Science, 949:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Alverson:1995:STM

[1934] G. Alverson, S. Kahan, R. Korry, and C. McCann. Scheduling on the Tera MTA. Lecture Notes in Computer Science, 949:19–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Barton:1995:SMD

[1935] J. M. Barton and N. Bitar. A scalable multi-discipline, multipleprocessor scheduling framework for IRIX. Lecture Notes in Computer Science, 949:45–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hori:1995:TSS

[1936] A. Hori, T. Yokota, Y. Ishikawa, and S. Sakai. Time space sharing scheduling and architectural support. Lecture Notes in Computer Science, 949:92–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sobalvarro:1995:DBC

[1937] P. G. Sobalvarro and W. E. Weihl.

Demand-based coscheduling of parallel jobs on multiprogrammed multiprocessors. Lecture Notes in Computer Science, 949:106–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Parsons:1995:MSH

[1938] E. W. Parsons and K. C. Sevcik.

Multiprocessor scheduling for highvariability service time distributions. Lecture Notes in Computer Science, 949:127–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Setia:1995:IBM

[1939] S. K. Setia. The interaction between memory allocation and adaptive partitioning in message-passing multicomputers. Lecture Notes in Computer Science, 949:146–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosti:1995:ANW

[1940] E. Rosti, E. Smirni, G. Serazzi, and L. W. Dowdy. Analysis of nonwork-conserving processor partitioning policies. Lecture Notes in Computer Science, 949:165–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Stoica:1995:MSP

[1941] I. Stoica, H. Abdel-Wahab, and A. Pothen. A microeconomic scheduler for parallel computers. Lecture Notes in Computer Science, 949:200–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Squillante:1995:BLD

[1942] M. S. Squillante. On the benefits and limitations of dynamic partitioning in parallel computer systems. Lecture Notes in Computer Science, 949:219– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kipersztok:1995:IFC

[1943] O. Kipersztok and J. C. Patterson. Intelligent fuzzy control to augment scheduling capabilities of network queuing systems. Lecture Notes in Computer Science, 949:239–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pruyne:1995:PPD

[1944] J. Pruyne and M. Livny. Parallel processing on dynamic resources with CARMI. Lecture Notes in Computer Science, 949:259–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henderson:1995:JSU

[1945] R. L. Henderson. Job scheduling under the portable batch system. Lecture Notes in Computer Science, 949: 279–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lifka:1995:AIS

[1946] D. A. Lifka. The ANL/IBM SP scheduling system. Lecture Notes in Computer Science, 949:295–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rosenkrantz:1995:RCT

[1947] M. E. Rosenkrantz, D. J. Schneider, R. Leibensperger, and M. Shore. Requirements of the Cornell Theory Center for resource management and process scheduling. Lecture Notes in Computer Science, 949:304–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saphir:1995:JMR

[1948] W. Saphir, L. A. Tanner, and B. Traversat. Job management requirements for NAS parallel systems and clusters. Lecture Notes in Computer Science, 949:319–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Feitelson:1995:JCP

[1949] D. G. Feitelson and B. Nitzberg. Job characteristics of a production parallel scientific workload on the NASA Ames iPSC/860. Lecture Notes in Computer Science, 949:337–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naor:1995:VC

[1950] Moni Naor and Adi Shamir. Visual cryptography. Lecture Notes in Computer Science, 950:1–12, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500001.
pdf.

Csirmaz:1995:SSM

[1951] László Csirmaz. The size of a share must be large. Lecture Notes in Computer Science, 950:13–22, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500013.
pdf.

vanDijk:1995:LCP

[1952] Marten van Dijk. A linear construction of perfect secret sharing schemes. Lecture Notes in Computer Science, 950: 23–34, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500023.
pdf.

Dijk:1995:LCP

[1953] Marten van Dijk. A linear construction of perfect secret sharing schemes. Lecture Notes in Computer Science, 950:

23–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500023.
pdf.

Blundo:1995:DRR

[1954] Carlo Blundo, Antonio Giorgio Gaggia, and Douglas R. Stinson. On the dealer’s randomness required in secret sharing schemes. Lecture Notes in Computer Science, 950:35– 46, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500035.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500035.
pdf.

Schnorr:1995:BBC

[1955] Claus P. Schnorr and Serge Vaudenay. Black box cryptanalysis of hash networks based on multipermutations. Lecture Notes in Computer Science, 950:47–57, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500047.
pdf.

Joux:1995:PAA

[1956] Antoine Joux and Louis Granboulan.

A practical attack against knapsack based hash functions. Lecture Notes in Computer Science, 950:58– 66, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500058.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500058.
                                        pdf.

Franklin:1995:BWS

[1957] Matthew Franklin and Moti Yung. The blinding of weak signatures. Lecture Notes in Computer Science, 950:67– 76, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500067.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500067.
                                        pdf.

Naccache:1995:CDI

[1958] David Naccache, David M’Raïhi, Serge Vaudenay, and Dan Raphaeli. Can D.S.A. be improved? complexity trade-offs with the Digital Signature Standard. Lecture Notes in Computer Science, 950:77– 85, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500077.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500077.
                                        pdf.

Chaum:1995:DCS

[1959] David Chaum. Designated confirmer signatures. Lecture Notes in Computer Science, 950:86–91, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500086.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500086.
pdf.

Bellare:1995:OAE

[1960] Mihir Bellare and Phillip Rogaway. Optimal asymmetric encryption. Lecture Notes in Computer Science, 950:92– 111, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500092.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500092.
pdf.

Orton:1995:MIT

[1961] Glenn A. Orton. A multipleiterated trapdoor for dense compact knapsacks. Lecture Notes in Computer Science, 950:112–130, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500112.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500112.
pdf.

Chabaud:1995:SSC

[1962] Florent Chabaud. On the security of some cryptosystems based on error-correcting codes. Lecture Notes in Computer Science, 950:131–139, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500131.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500131.
                                        pdf.

Chen:1995:PDP

[1963] Lidong Chen, Ivan Bjerre Damgård, and Torben Pryds Pedersen. Parallel divertibility of proofs of knowledge. Lecture Notes in Computer Science, 950:140–155, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0950/09500140.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500140.
                                        pdf.

DAmiano:1995:MDM

[1964] Stefano D’Amiano and Giovanni Di Crescenzo. Methodology for digital money based on general cryptographic tools. Lecture Notes in Computer Science, 950:156–170, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500156.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500156.
                                        pdf.

Chen:1995:NGS

[1965] Lidong Chen and Torben Pryds Pedersen. New group signature schemes. Lecture Notes in Computer Science, 950:171–181, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0950/09500171.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0950/09500171.
pdf.

Nyberg:1995:MRS

[1966] Kaisa Nyberg and Rainer A. Rueppel. Message recovery for signature schemes based on the discrete logarithm problem. Lecture Notes in Computer Science, 950:182–193, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500182.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500182.
pdf.

Li:1995:TMS

[1967] Chuan-Ming Li, Tzonelih Hwang, and Narn-Yih Lee. Thresholdmultisignature schemes where suspected forgery implies traceability of adversarial shareholders. Lecture Notes in Computer Science, 950:194–204, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500194.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500194.
pdf.

Meier:1995:SSG

[1968] Willi Meier and Othmar Staffelbach.

The self-shrinking generator. Lecture Notes in Computer Science, 950:205– 214, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500205.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500205.
                                        pdf.

Goresky:1995:FRB

[1969] Mark Goresky and Andrew Klapper. Feedback registers based on ramified extensions of the 2-adic numbers. Lecture Notes in Computer Science, 950: 215–222, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0950/09500215.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500215.
                                        pdf.

Goettfert:1995:GLB

[1970] R. Goettfert and H. Niederreiter. A general lower bound for the linear complexity of the product of shift-register sequences. Lecture Notes in Computer Science, 950:223–229, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gottfert:1995:GLB

[1971] Rainer Göttfert and Harald Niederreiter. A general lower bound for the linear complexity of the product of shift-register sequences. Lecture Notes in Computer Science, 950:223– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500223.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/0950/09500223.
pdf.

Golic:1995:EPC

[1972] Jovan Dj. Golic and Luke O’Connor.

Embedding and probabilistic correlation attacks on clock-controlled shift registers. Lecture Notes in Computer Science, 950:230–243, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0950/09500230.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0950/09500230.pdf.

Taylor:1995:NOU

[1973] Richard Taylor. Near optimal unconditionally secure authentication. Lecture Notes in Computer Science, 950:244– 253, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500244.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500244.
pdf.

Safavi-Naini:1995:ACP

[1974] Reihaneh Safavi-Naini and L. Tombak.

Authentication codes in plaintext and chosen-content attacks. Lecture Notes in Computer Science, 950:254–265, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500254.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500254.
pdf.

Cachin:1995:LIR

[1975] Christian Cachin and Ueli M. Maurer. Linking information reconciliation and privacy amplification. Lecture Notes in Computer Science, 950:266– 274, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500266.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500266.
                                        pdf.

Burmester:1995:SEC

[1976] Mike V. D. Burmester and Yvo Desmedt. A secure and efficient conference key distribution system. Lecture Notes in Computer Science, 950:275– 286, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500275.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500275.
                                        pdf.

Blundo:1995:SRB

[1977] Carlo Blundo and Antonella Cresti. Space requirements for broadcast encryption. Lecture Notes in Computer Science, 950:287–298, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/0950/09500287.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        0950/09500287.pdf.

Zheng:1995:HBR

[1978] Yuliang Zheng. How to break and repair Leighton and Micali’s key agreement protocol. Lecture Notes in Computer Science, 950:299–305, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500299.
pdf.

Eng:1995:STD

[1979] Tony Eng and Tatsuaki Okamoto.

Single-term divisible electronic coins. Lecture Notes in Computer Science, 950:306–319, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0950/09500306.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0950/09500306.
pdf.

Syverson:1995:FRK

[1980] Paul Syverson and Catherine Meadows. Formal requirements for key distribution protocols. Lecture Notes in Computer Science, 950:320–331, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500320.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500320.
pdf.

Pfitzmann:1995:BEA

[1981] Birgit Pfitzmann. Breaking an efficient anonymous channel. Lecture Notes in Computer Science, 950:332– 340, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500332.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500332.
                                        pdf.

Biham:1995:MLC

[1982] Eli Biham. On Matsui’s linear cryptanalysis. Lecture Notes in Computer Science, 950:341–355, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500341.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500341.
                                        pdf.

Chabaud:1995:LBD

[1983] Florent Chabaud and Serge Vaudenay. Links between differential and linear cryptanalysis. Lecture Notes in Computer Science, 950:356–365, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500356.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500356.
                                        pdf.

Matsui:1995:CBO

[1984] Mitsuru Matsui. On correlation between the order of S-boxes and the strength of DES. Lecture Notes in Computer Science, 950:366–375, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500366.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500366.
pdf.

Seberry:1995:RAN

[1985] Jennifer Seberry, Xian-Mo Zhang, and Yuliang Zheng. Relationships among nonlinearity criteria. Lecture Notes in Computer Science, 950:376–388, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500376.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500376.
pdf.

deRooij:1995:EEU

[1986] Peter de Rooij. Efficient exponentiation using precomputation and vector addition chains. Lecture Notes in Computer Science, 950:389–399, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500389.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500389.
pdf.

Rooij:1995:EEU

[1987] Peter de Rooij. Efficient exponentiation using precomputation and vector addition chains. Lecture Notes in Computer Science, 950:389–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500389.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500389.
                                        pdf.

Damm:1995:MFH

[1988] F. Damm, F.-P. Heider, and G. Wambach. MIMD-factorisation on hypercubes. Lecture Notes in Computer Science, 950:400–409, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/0950/09500400.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500400.
                                        pdf.

Knudsen:1995:NAA

[1989] Lars Ramkilde Knudsen and Xuejia Lai. New attacks on all double block length hash functions of hash rate 1, including the parallel-DM. Lecture Notes in Computer Science, 950:410– 418, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500410.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500410.
                                        pdf.

Knudsen:1995:NPW

[1990] Lars Ramkilde Knudsen. New potentially “weak” keys for DES and LOKI. Lecture Notes in Computer Science, 950:419–424, 1995. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/0950/09500419.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/0950/09500419.
pdf.

Jakobsson:1995:BUU

[1991] Markus Jakobsson. Blackmailing using undeniable signatures. Lecture Notes in Computer Science, 950:425– 427, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500425.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500425.
pdf.

Camenisch:1995:BSB

[1992] Jan L. Camenisch, Jean-Marc Piveteau, and Markus A. Stadler. Blind signatures based on the discrete logarithm problem. Lecture Notes in Computer Science, 950:428–432, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500428.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500428.
pdf.

Charnes:1995:CSE

[1993] C. Charnes, Luke O’Connor, Józef P.

Pieprzyk, Reihaneh Safavi-Naini, and Yuliang Zheng. Comments on Soviet encryption algorithm. Lecture Notes in Computer Science, 950:433– 438, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500433.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500433.
                                        pdf.

Nyberg:1995:LAB

[1994] Kaisa Nyberg. Linear approximation of block ciphers. Lecture Notes in Computer Science, 950:439–444, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500439.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500439.
                                        pdf.

Shamir:1995:MEV

[1995] Adi Shamir. Memory efficient variants of public-key schemes for smart card applications. Lecture Notes in Computer Science, 950:445–449, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0950/09500445.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0950/09500445.
                                        pdf.

Menicocci:1995:SAC

[1996] Renato Menicocci. A systematic attack on clock controlled cascades. Lecture Notes in Computer Science, 950:450– 455, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/0950/09500450.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500450.
pdf.

Johansson:1995:CIA

[1997] Thomas Johansson and Bernard J. M.

Smeets. On A^{2}-codes including arbiter’s attacks. Lecture Notes in Computer Science, 950:456–460, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500456.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500456.
pdf.

Biham:1995:IDA

[1998] Eli Biham and Alex Biryukov.

An improvement of Davies’ attack on DES. Lecture Notes in Computer Science, 950:461–467, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0950/09500461.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0950/09500461.
pdf.

Hruby:1995:QDQ

[1999] J. Hruby. Q-deformed quantum cryptography. Lecture Notes in Computer Science, 950:468–472, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0950/09500468.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        0950/09500468.pdf.

Anonymous:1995:AIb

[2000] Anonymous. Author index. Lecture Notes in Computer Science, 950:473– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/0950/0950indx.pdf.

Kuijpers:1995:LRT

[2001] B. Kuijpers, J. Paredaens, and J. Van den Bussche. Lossless representation of topological spatial data. Lecture Notes in Computer Science, 951:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Vandeurzen:1995:DLL

[2002] L. Vandeurzen, M. Gyssens, and D. Van Gucht. On the desirability and limitations of linear spatial database models. Lecture Notes in Computer Science, 951:14–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Finke:1995:QVD

[2003] U. Finke and K. H. Hinrichs. The quad view data structure — A representation for planar subdivisions. Lecture Notes in Computer Science, 951:29–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Koperski:1995:DSA

[2004] K. Koperski and J. Han. Discovery of spatial association rules in geographic information databases. Lecture Notes in Computer Science, 951:47–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ester:1995:KDL

[2005] M. Ester, H.-P. Kriegel, and X. Xu.

Knowledge discovery in large spatial databases: Focusing techniques for efficient class identification. Lecture Notes in Computer Science, 951:67–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hjaltason:1995:RSD

[2006] G. R. Hjaltason and H. Samet. Ranking in spatial databases. Lecture Notes in Computer Science, 951:83–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gaede:1995:ORS

[2007] V. Gaede. Optimal redundancy in spatial database systems. Lecture Notes in Computer Science, 951:96–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Blott:1995:AGM

[2008] S. Blott and A. Vckovski. Accessing geographical metafiles through a database storage system. Lecture Notes in Computer Science, 951:117– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Henrich:1995:ESA

[2009] A. Henrich and J. Moeller. Extending a spatial access structure to support additional standard attributes. Lecture Notes in Computer Science, 951:132– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Puppo:1995:TFM

[2010] E. Puppo and G. Dettori. Towards a formal model for multiresolution spatial maps. Lecture Notes in Computer Science, 951:152–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Rigaux:1995:MSP

[2011] P. Rigaux and M. Scholl. Multi-scale partitions: Application to spatial and statistical databases. Lecture Notes in Computer Science, 951:170–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Frank:1995:SOG

[2012] A. U. Frank and W. Kuhn. Specifying open GIS with functional languages. Lecture Notes in Computer Science, 951:184–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Shekhar:1995:LBH

[2013] S. Shekhar, S. Ravada, V. Kumar, and D. Chubb. Load-balancing in high performance GIS: Declustering polygonal maps. Lecture Notes in Computer Science, 951:196–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gueting:1995:IRA

[2014] R. Hartmut Gueting, T. De Ridder, and M. Schneider. Implementation of the ROSE algebra: Efficient algorithms for realm-based spatial data types. Lecture Notes in Computer Science, 951:216–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Seidl:1995:MSR

[2015] T. Seidl and H.-P. Kriegel. A 3D molecular surface representation supporting neighborhood queries. Lecture Notes in Computer Science, 951:240–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scarponcini:1995:ILA

[2016] P. Scarponcini, D. C. St. Clair, and G. W. Zobrist. An inferencing language for automated spatial reasoning about graphic entities. Lecture Notes in Computer Science, 951:259–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sharma:1995:ICK

[2017] J. Sharma and D. M. Flewelling.

Inferences from combined knowledge about topology and directions. Lecture Notes in Computer Science, 951: 279–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nabil:1995:PIR

[2018] M. Nabil, J. Shepherd, and A. H. H.

Ngu. 2D projection interval relationships: A symbolic representation of spatial relationships. Lecture Notes in Computer Science, 951:292–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Winter:1995:TRB

[2019] S. Winter. Topological relations between discrete regions. Lecture Notes in Computer Science, 951:310–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lo:1995:GST

[2020] M.-L. Lo and C. V. Ravishankar. Generating seeded trees from data sets. Lecture Notes in Computer Science, 951:328–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abel:1995:SJS

[2021] D. J. Abel, Beng Chin Ooi, K.-L. Tan, and R. Power. Spatial join strategies in distributed spatial DBMS. Lecture Notes in Computer Science, 951:348– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Arctur:1995:CBI

[2022] D. Arctur, E. Anwar, J. Alexander, and S. Chakravarthy. Comparison and benchmarks for import of VPF geographic data from object-oriented and relational database files. Lecture Notes in Computer Science, 951:368– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Franklin:1995:CED

[2023] W. R. Franklin. Compressing elevation data. Lecture Notes in Computer Science, 951:385–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Thomas:1995:ERO

[2024] Dave Thomas. Experiences on the road to object utopia (extended abstract). Lecture Notes in Computer Science, 952:1, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/0952/09520001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0952/09520001.
pdf.

Agesen:1995:CPA

[2025] Ole Agesen. The Cartesian product algorithm. Lecture Notes in Computer Science, 952:2–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0952/09520002.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0952/09520002.pdf.

Bruce:1995:PTS

[2026] Kim B. Bruce, Angela Schuett, and Robert van Gent. PolyTOIL: A type-safe polymorphic objectoriented language. Lecture Notes in Computer Science, 952:27–51, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0952/09520027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0952/09520027.
pdf.

Thatte:1995:OI

[2027] Satish R. Thatté. Object imaging. Lecture Notes in Computer Science, 952: 52–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0952/09520052.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/0952/09520052.
pdf.

Dean:1995:OOO

[2028] Jeffrey Dean, David Grove, and Craig Chambers. Optimization of object-oriented programs using static class hierarchy analysis. Lecture Notes in Computer Science, 952:77– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0952/09520077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0952/09520077.
pdf.

Bertino:1995:OMM

[2029] Elisa Bertino and Giovanna Guerrini. Objects with multiple most specific classes. Lecture Notes in Computer Science, 952:102–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0952/09520102.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0952/09520102.
pdf.

Steyaert:1995:MCO

[2030] Patrick Steyaert and Wolfgang De Meuter. A marriage of class-and object-based inheritance without unwanted children. Lecture Notes in Computer Science, 952:127–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0952/09520127.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/0952/09520127.
pdf.

Abadi:1995:SM

[2031] Martin Abadi and Luca Cardelli. On subtyping and matching. Lecture Notes in Computer Science, 952:145– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0952/09520145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0952/09520145.
pdf.

Stroud:1995:UMP

[2032] R. J. Stroud and Z. Wu. Using metaobject protocols to implement atomic data types. Lecture Notes in Computer Science, 952:168– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0952/09520168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0952/09520168.
pdf.

McAffer:1995:MLP

[2033] Jeff McAffer. Meta-level programming with CodA. Lecture Notes in Computer Science, 952:190–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0952/09520190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0952/09520190.
pdf.

Arditi:1995:OOF

[2034] Laurent Arditi and Hélère Collavizza. An object-oriented framework for the formal verification of processors. Lecture Notes in Computer Science, 952: 215–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0952/09520215.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0952/09520215.
                                        pdf.

Seligmann:1995:IMG

[2035] Jacob Seligmann and Steffen Grarup. Incremental mature garbage collection using the train algorithm. Lecture Notes in Computer Science, 952:235– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0952/09520235.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0952/09520235.
                                        pdf.

Driesen:1995:MDP

[2036] Karel Driesen, Urs Hölzle, and Jan Vitek. Message dispatch on pipelined processors. Lecture Notes in Computer Science, 952:253–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0952/09520253.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/0952/09520253.
                                        pdf.

Hoelzle:1995:DOO

[2037] U. Hoelzle and D. Ungar. Do objectoriented languages need special hardware support? Lecture Notes in Computer Science, 952:283–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holzle:1995:DOO

[2038] Urs Hölzle and David Ungar. Do object-oriented languages need special hardware support? Lecture Notes in Computer Science, 952:283– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0952/09520283.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0952/09520283.
pdf.

Smith:1995:PEI

[2039] Randall B. Smith and David Ungar. Programming as an experience: The inspiration for self. Lecture Notes in Computer Science, 952:303– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0952/09520303.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0952/09520303.
pdf.

Satoh:1995:TAI

[2040] Ichiro Satoh and Mario Tokoro. Time and asynchrony in interactions among distributed real-time objects. Lecture Notes in Computer Science, 952:331– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0952/09520331.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0952/09520331.
pdf.

Florijn:1995:OPF

[2041] Gert Florijn. Object protocols as functional parsers. Lecture Notes in Computer Science, 952:351–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0952/09520351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0952/09520351.
pdf.

Lea:1995:IBP

[2042] Doug Lea and Jos Marlowe. Interfacebased protocol specification of open systems using PSL. Lecture Notes in Computer Science, 952:374–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0952/09520374.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0952/09520374.
pdf.

Schmidt:1995:EUD

[2043] Doug Schmidt and Paul Stephenson. Experience using design patterns to evolve communication software across diverse OS platforms. Lecture Notes in Computer Science, 952:399– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0952/09520399.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0952/09520399.
pdf.

Kiessling:1995:SPU

[2044] Heiko Kießling and Uwe Krüger. Sharing properties in a uniform object space. Lecture Notes in Computer Science, 952:424–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/0952/09520424.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0952/09520424.pdf.

Amsellem:1995:CHP

[2045] Maurice Amsellem. ChyPro: A hypermedia programming environment for Smalltalk-80. Lecture Notes in Computer Science, 952:449–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0952/09520449.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/0952/09520449.
pdf.

Anonymous:1995:AIc

[2046] Anonymous. Author index. Lecture Notes in Computer Science, 952:471– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/0952/0952ind.pdf.

Milner:1995:CSM

[2047] R. Milner. Control structures: A model of interaction. Lecture Notes in Computer Science, 953:1–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Pavlovic:1995:CCP

[2048] D. Pavlovic. Convenient category of processes and simulations I: Modulo strong bisimilarity. Lecture Notes in Computer Science, 953:3–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Badouel:1995:DBN

[2049] E. Badouel and P. Darondeau. Dualities between nets and automata induced by schizophrenic objects. Lecture Notes in Computer Science, 953:24–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kawahara:1995:RST

[2050] Y. Kawahara. Relational set theory. Lecture Notes in Computer Science, 953:44–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Soloviev:1995:PMC

[2051] S. Soloviev. Proof of a S. MacLane conjecture. Lecture Notes in Computer Science, 953:59–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Asperti:1995:EAS

[2052] A. Asperti and A. Ciabattoni. Effective applicative structures. Lecture Notes in Computer Science, 953:81–??, 1995.

CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hyland:1995:RC

[2053] J. M. E. Hyland and E. Moggi. The S-replete construction. Lecture Notes in Computer Science, 953:96–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Simpson:1995:CPC

[2054] A. K. Simpson. The convex powerdomain in a category of posets realized by CPOs. Lecture Notes in Computer Science, 953:117–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fiore:1995:LKD

[2055] M. P. Fiore. Lifting as a KZ-Doctrine. Lecture Notes in Computer Science, 953:146–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Backhouse:1995:CFP

[2056] R. Backhouse, M. Bijsterveld, R. Van Geldrop, and J. Van der Woude. Categorical fixed point calculus. Lecture Notes in Computer Science, 953:159– ??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Brookes:1995:CTT

[2057] S. Brookes. A category-theoretic treatment of a parallel Algol-like language. Lecture Notes in Computer Science, 953:180–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Altenkirch:1995:CRR

[2058] T. Altenkirch, M. Hofmann, and T. Streicher. Categorical reconstruction of a reduction free normalization proof. Lecture Notes in Computer Science, 953:182–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hasegawa:1995:DTL

[2059] M. Hasegawa. Decomposing typed lambda calculus into a couple of categorical programming languages. Lecture Notes in Computer Science, 953: 200–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Otto:1995:VCP

[2060] J. R. Otto. V-comprehensions and P space. Lecture Notes in Computer Science, 953:220–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fourman:1995:PCS

[2061] M. Fourman and H. Thielecke. A proposed categorical semantics for ML modules. Lecture Notes in Computer Science, 953:240–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).